Gnus 5.7 Manual

by Lars Magne Ingebrigtsen

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1995,96,97 Free Software Foundation, Inc. Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one. Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions.

The Gnus Newsreader

Gnus is the advanced, self-documenting, customizable, extensible unreal-time newsreader for GNU Emacs.

Oops. That sounds oddly familiar, so let’s start over again to avoid being accused of plagiarism:

Gnus is a message-reading laboratory. It will let you look at just about anything as if it were a newsgroup. You can read mail with it, you can browse directories with it, you can ftp with it—you can even read news with it!

Gnus tries to empower people who read news the same way Emacs empowers people who edit text. Gnus sets no limits to what the user should be allowed to do. Users are encouraged to extend Gnus to make it behave like they want it to behave. A program should not control people; people should be empowered to do what they want by using (or abusing) the program.

2 Gnus Manual

1 Starting Gnus

If your system administrator has set things up properly, starting Gnus and reading news is extremely easy—you just type M-x gnus in your Emacs.

If you want to start Gnus in a different frame, you can use the command M-x gnus-

other-frame instead.

If things do not go smoothly at startup, you have to twiddle some variables in your ‘~/.gnus’ file. This file is similar to ‘~/.emacs’, but is read when gnus starts.

If you puzzle at any terms used in this manual, please refer to the terminology section (see Section 10.3 [Terminology], page 204).

1.1 Finding the News

The gnus-select-method variable says where Gnus should look for news. This variable should be a list where the first element says how and the second element says where. This method is your native method. All groups not fetched with this method are foreign groups.

For instance, if the ‘news.somewhere.edu’ nntp server is where you want to get your daily dosage of news from, you’d say:

     (setq gnus-select-method '(nntp "news.somewhere.edu"))

If you want to read directly from the local spool, say:

     (setq gnus-select-method '(nnspool ""))

If you can use a local spool, you probably should, as it will almost certainly be much faster.

If this variable is not set, Gnus will take a look at the NNTPSERVER environment variable. If that variable isn’t set, Gnus will see whether gnus-nntpserver-file (‘/etc/nntpserver’ by default) has any opinions on the matter. If that fails as well, Gnus will try to use the machine running Emacs as an nntp server. That’s a long shot, though.

If gnus-nntp-server is set, this variable will override gnus-select-method. You should therefore set gnus-nntp-server to nil, which is what it is by default.

You can also make Gnus prompt you interactively for the name of an nntp server. If you give a non-numerical prefix to gnus (i.e., C-u M-x gnus), Gnus will let you choose between the servers in the gnus-secondary-servers list (if any). You can also just type in the name of any server you feel like visiting.

However, if you use one nntp server regularly and are just interested in a couple of groups from a different server, you would be better served by using the B command in the group buffer. It will let you have a look at what groups are available, and you can subscribe to any of the groups you want to. This also makes ‘.newsrc’ maintenance much tidier. See Section 2.9 [Foreign Groups], page 18.

A slightly different approach to foreign groups is to set the gnus-secondary-select-methods variable. The select methods listed in this variable are in many ways just as native as the gnus-select-method server. They will also be queried for active files during startup (if that’s required), and new newsgroups that appear on these servers will be subscribed (or not) just as native groups are.

For instance, if you use the nnmbox backend to read your mail, you would typically set this variable to

     (setq gnus-secondary-select-methods '((nnmbox "")))

1.2 The First Time

If no startup files exist, Gnus will try to determine what groups should be subscribed by default.

If the variable gnus-default-subscribed-newsgroups is set, Gnus will subscribe you to just those groups in that list, leaving the rest killed. Your system administrator should have set this variable to something useful.

Since she hasn’t, Gnus will just subscribe you to a few arbitrarily picked groups (i.e., ‘*.newusers’). (Arbitrary is defined here as whatever Lars thinks you should read.)

You’ll also be subscribed to the Gnus documentation group, which should help you with most common problems.

If gnus-default-subscribed-newsgroups is t, Gnus will just use the normal functions for handling new groups, and not do anything special.

1.3 The Server is Down

If the default server is down, Gnus will understandably have some problems starting. However, if you have some mail groups in addition to the news groups, you may want to start Gnus anyway.

Gnus, being the trusting sort of program, will ask whether to proceed without a native select method if that server can’t be contacted. This will happen whether the server doesn’t actually exist (i.e., you have given the wrong address) or the server has just momentarily taken ill for some reason or other. If you decide to continue and have no foreign groups, you’ll find it difficult to actually do anything in the group buffer. But, hey, that’s your problem. Blllrph!

If you know that the server is definitely down, or you just want to read your mail without bothering with the server at all, you can use the gnus-no-server command to start Gnus. That might come in handy if you’re in a hurry as well. This command will not attempt to contact your primary server—instead, it will just activate all groups on level 1 and 2. (You should preferably keep no native groups on those two levels.)

1.4 Slave Gnusae

You might want to run more than one Emacs with more than one Gnus at the same time. If you are using different ‘.newsrc’ files (e.g., if you are using the two different Gnusae to read from two different servers), that is no problem whatsoever. You just do it.

The problem appears when you want to run two Gnusae that use the same .newsrc file. To work around that problem some, we here at the Think-Tank at the Gnus Towers

have come up with a new concept: Masters and slaves. (We have applied for a patent on this concept, and have taken out a copyright on those words. If you wish to use those words in conjunction with each other, you have to send $1 per usage instance to me. Usage of the patent (Master/Slave Relationships In Computer Applications) will be much more expensive, of course.)

Anyways, you start one Gnus up the normal way with M-x gnus (or however you do it). Each subsequent slave Gnusae should be started with M-x gnus-slave. These slaves won’t save normal ‘.newsrc’ files, but instead save slave files that contain information only on what groups have been read in the slave session. When a master Gnus starts, it will read (and delete) these slave files, incorporating all information from them. (The slave files will be read in the sequence they were created, so the latest changes will have precedence.)

Information from the slave files has, of course, precedence over the information in the normal (i.e., master) .newsrc file.

1.5 Fetching a Group

It is sometimes convenient to be able to just say “I want to read this group and I don’t care whether Gnus has been started or not”. This is perhaps more useful for people who write code than for users, but the command gnus-fetch-group provides this functionality in any case. It takes the group name as a parameter.

1.6 New Groups

If you are satisfied that you really never want to see any new groups, you can set gnuscheck-new-newsgroups to nil. This will also save you some time at startup. Even if this variable is nil, you can always subscribe to the new groups just by pressing U in the group buffer (see Section 2.13 [Group Maintenance], page 24). This variable is ask-server by default. If you set this variable to always, then Gnus will query the backends for new groups even when you do the g command (see Section 2.17.1 [Scanning New Messages], page 30).

1.6.1 Checking New Groups

Gnus normally determines whether a group is new or not by comparing the list of groups from the active file(s) with the lists of subscribed and dead groups. This isn’t a particularly fast method. If gnus-check-new-newsgroups is ask-server, Gnus will ask the server for new groups since the last time. This is both faster and cheaper. This also means that you can get rid of the list of killed groups altogether, so you may set gnus-save-killed-list to nil, which will save time both at startup, at exit, and all over. Saves disk space, too. Why isn’t this the default, then? Unfortunately, not all servers support this command.

I bet I know what you’re thinking now: How do I find out whether my server supports ask-server? No? Good, because I don’t have a fail-safe answer. I would suggest just setting this variable to ask-server and see whether any new groups appear within the next few days. If any do, then it works. If none do, then it doesn’t work. I could write a function to make Gnus guess whether the server supports ask-server, but it would just be a guess. So I won’t. You could telnet to the server and say HELP and see whether it lists ‘NEWGROUPS’ among the commands it understands. If it does, then it might work. (But there are servers that lists ‘NEWGROUPS’ without supporting the function properly.)

This variable can also be a list of select methods. If so, Gnus will issue an ask-server command to each of the select methods, and subscribe them (or not) using the normal methods. This might be handy if you are monitoring a few servers for new groups. A side effect is that startup will take much longer, so you can meditate while waiting. Use the mantra “dingnusdingnusdingnus” to achieve permanent bliss.

1.6.2 Subscription Methods

What Gnus does when it encounters a new group is determined by the gnus-subscribe-

newsgroup-method variable.

This variable should contain a function. This function will be called with the name of the new group as the only parameter.

Some handy pre-fab functions are:

gnus-subscribe-zombies

Make all new groups zombies. This is the default. You can browse the zombies later (with A z) and either kill them all off properly (with S z), or subscribe to them (with u).

gnus-subscribe-randomly

Subscribe all new groups in arbitrary order. This really means that all new groups will be added at “the top” of the group buffer.

gnus-subscribe-alphabetically

Subscribe all new groups in alphabetical order.

gnus-subscribe-hierarchically

Subscribe all new groups hierarchically. The difference between this

          function and gnus-subscribe-alphabetically is slight. gnus-subscribe-

alphabetically will subscribe new groups in a strictly alphabetical fashion, while this function will enter groups into it’s hierarchy. So if you want to have the ‘rec’ hierarchy before the ‘comp’ hierarchy, this function will not mess that configuration up. Or something like that.

gnus-subscribe-interactively

Subscribe new groups interactively. This means that Gnus will ask you about all new groups. The groups you choose to subscribe to will be subscribed hierarchically.

gnus-subscribe-killed

Kill all new groups.

   A closely related variable is gnus-subscribe-hierarchical-interactive. (That’s

quite a mouthful.) If this variable is non-nil, Gnus will ask you in a hierarchical fashion whether to subscribe to new groups or not. Gnus will ask you for each sub-hierarchy whether you want to descend the hierarchy or not.

One common mistake is to set the variable a few paragraphs above (gnus-subscribe-

newsgroup-method) to gnus-subscribe-hierarchical-interactive. This is an error.

This will not work. This is ga-ga. So don’t do it.

1.6.3 Filtering New Groups

A nice and portable way to control which new newsgroups should be subscribed (or ignored) is to put an options line at the start of the ‘.newsrc’ file. Here’s an example:

     options -n !alt.all !rec.all sci.all

This line obviously belongs to a serious-minded intellectual scientific person (or she may just be plain old boring), because it says that all groups that have names beginning with ‘alt’ and ‘rec’ should be ignored, and all groups with names beginning with ‘sci’ should be subscribed. Gnus will not use the normal subscription method for subscribing these groups.

gnus-subscribe-options-newsgroup-method is used instead. This variable defaults to
gnus-subscribe-alphabetically.

If you don’t want to mess with your ‘.newsrc’ file, you can just set the two variables

gnus-options-subscribe and gnus-options-not-subscribe. These two variables do ex-

actly the same as the ‘.newsrc’ ‘options -n’ trick. Both are regexps, and if the new group matches the former, it will be unconditionally subscribed, and if it matches the latter, it will be ignored.

Yet another variable that meddles here is gnus-auto-subscribed-groups. It works exactly like gnus-options-subscribe, and is therefore really superfluous, but I thought it would be nice to have two of these. This variable is more meant for setting some ground rules, while the other variable is used more for user fiddling. By default this variable makes all new groups that come from mail backends (nnml, nnbabyl, nnfolder, nnmbox, and nnmh) subscribed. If you don’t like that, just set this variable to nil.

New groups that match this regexp are subscribed using gnus-subscribe-options-

newsgroup-method.

1.7 Changing Servers

Sometimes it is necessary to move from one nntp server to another. This happens very rarely, but perhaps you change jobs, or one server is very flaky and you want to use another.

Changing the server is pretty easy, right? You just change gnus-select-method to point to the new server?

Wrong!

Article numbers are not (in any way) kept synchronized between different nntp servers, and the only way Gnus keeps track of what articles you have read is by keeping track of article numbers. So when you change gnus-select-method, your ‘.newsrc’ file becomes worthless.

Gnus provides a few functions to attempt to translate a ‘.newsrc’ file from one server to another. They all have one thing in common—they take a looong time to run. You don’t want to use these functions more than absolutely necessary.

If you have access to both servers, Gnus can request the headers for all the articles you have read and compare Message-IDs and map the article numbers of the read articles and article marks. The M-x gnus-change-server command will do this for all your native groups. It will prompt for the method you want to move to.

You can also move individual groups with the M-x gnus-group-move-group-to-server command. This is useful if you want to move a (foreign) group from one server to another.

If you don’t have access to both the old and new server, all your marks and read ranges

have become worthless. You can use the M-x gnus-group-clear-data-on-native-groups

command to clear out all data that you have on your native groups. Use with caution.

1.8 Startup Files

Now, you all know about the ‘.newsrc’ file. All subscription information is traditionally stored in this file.

Things got a bit more complicated with gnus. In addition to keeping the ‘.newsrc’ file updated, it also used a file called ‘.newsrc.el’ for storing all the information that didn’t fit into the ‘.newsrc’ file. (Actually, it also duplicated everything in the ‘.newsrc’ file.) gnus would read whichever one of these files was the most recently saved, which enabled people to swap between gnus and other newsreaders.

That was kinda silly, so Gnus went one better: In addition to the ‘.newsrc’ and ‘.newsrc.el’ files, Gnus also has a file called ‘.newsrc.eld’. It will read whichever of these files that are most recent, but it will never write a ‘.newsrc.el’ file. You should never delete the ‘.newsrc.eld’ file—it contains much information not stored in the ‘.newsrc’ file.

You can turn off writing the ‘.newsrc’ file by setting gnus-save-newsrc-file to nil, which means you can delete the file and save some space, as well as exiting from Gnus faster. However, this will make it impossible to use other newsreaders than Gnus. But hey, who would want to, right?

If gnus-save-killed-list (default t) is nil, Gnus will not save the list of killed groups to the startup file. This will save both time (when starting and quitting) and space (on disk). It will also mean that Gnus has no record of what groups are new or old, so the automatic new groups subscription methods become meaningless. You should always set

gnus-check-new-newsgroups to nil or ask-server if you set this variable to nil (see

Section 1.6 [New Groups], page 5). This variable can also be a regular expression. If that’s the case, remove all groups that do not match this regexp before saving. This can be useful in certain obscure situations that involve several servers where not all servers support

ask-server.

The gnus-startup-file variable says where the startup files are. The default value is ‘~/.newsrc’, with the Gnus (El Dingo) startup file being whatever that one is, with a ‘.eld’ appended.

gnus-save-newsrc-hook is called before saving any of the newsrc files, while gnus-

save-quick-newsrc-hook is called just before saving the ‘.newsrc.eld’ file, and gnus-

save-standard-newsrc-hook is called just before saving the ‘.newsrc’ file. The latter two are commonly used to turn version control on or off. Version control is on by default when saving the startup files. If you want to turn backup creation off, say something like:

     (defun turn-off-backup ()
       (set (make-local-variable 'backup-inhibited) t))
     (add-hook 'gnus-save-quick-newsrc-hook 'turn-off-backup)
     (add-hook 'gnus-save-standard-newsrc-hook 'turn-off-backup)
   When Gnus starts, it will read the gnus-site-init-file (‘.../site-lisp/gnus’ by

default) and gnus-init-file (‘~/.gnus’ by default) files. These are normal Emacs Lisp files and can be used to avoid cluttering your ‘~/.emacs’ and ‘site-init’ files with Gnus stuff. Gnus will also check for files with the same names as these, but with ‘.elc’ and ‘.el’ suffixes. In other words, if you have set gnus-init-file to ‘~/.gnus’, it will look for ‘~/.gnus.elc’, ‘~/.gnus.el’, and finally ‘~/.gnus’ (in this order).

1.9 Auto Save

Whenever you do something that changes the Gnus data (reading articles, catching up, killing/subscribing groups), the change is added to a special dribble buffer. This buffer is auto-saved the normal Emacs way. If your Emacs should crash before you have saved the ‘.newsrc’ files, all changes you have made can be recovered from this file.

If Gnus detects this file at startup, it will ask the user whether to read it. The auto save file is deleted whenever the real startup file is saved.

If gnus-use-dribble-file is nil, Gnus won’t create and maintain a dribble buffer. The default is t.

Gnus will put the dribble file(s) in gnus-dribble-directory. If this variable is nil, which it is by default, Gnus will dribble into the directory where the ‘.newsrc’ file is located. (This is normally the user’s home directory.) The dribble file will get the same file permissions as the .newsrc file.

If gnus-always-read-dribble-file is non-nil, Gnus will read the dribble file on startup without querying the user.

1.10 The Active File

When Gnus starts, or indeed whenever it tries to determine whether new articles have arrived, it reads the active file. This is a very large file that lists all the active groups and articles on the server.

Before examining the active file, Gnus deletes all lines that match the regexp gnusignored-newsgroups. This is done primarily to reject any groups with bogus names, but you can use this variable to make Gnus ignore hierarchies you aren’t ever interested in. However, this is not recommended. In fact, it’s highly discouraged. Instead, see Section 1.6 [New Groups], page 5 for an overview of other variables that can be used instead.

The active file can be rather Huge, so if you have a slow network, you can set gnusread-active-file to nil to prevent Gnus from reading the active file. This variable is some by default.

Gnus will try to make do by getting information just on the groups that you actually subscribe to.

Note that if you subscribe to lots and lots of groups, setting this variable to nil will probably make Gnus slower, not faster. At present, having this variable nil will slow Gnus down considerably, unless you read news over a 2400 baud modem.

This variable can also have the value some. Gnus will then attempt to read active info only on the subscribed groups. On some servers this is quite fast (on sparkling, brand new INN servers that support the LIST ACTIVE group command), on others this isn’t fast at all. In any case, some should be faster than nil, and is certainly faster than t over slow lines.

If this variable is nil, Gnus will ask for group info in total lock-step, which isn’t very fast. If it is some and you use an nntp server, Gnus will pump out commands as fast as it can, and read all the replies in one swoop. This will normally result in better performance, but if the server does not support the aforementioned LIST ACTIVE group command, this isn’t very nice to the server.

In any case, if you use some or nil, you should definitely kill all groups that you aren’t interested in to speed things up.

Note that this variable also affects active file retrieval from secondary select methods.

1.11 Startup Variables

gnus-load-hook

A hook run while Gnus is being loaded. Note that this hook will normally be run just once in each Emacs session, no matter how many times you start Gnus.

gnus-before-startup-hook

A hook run after starting up Gnus successfully.

gnus-startup-hook

A hook run as the very last thing after starting up Gnus

gnus-started-hook

A hook that is run as the very last thing after starting up Gnus successfully.

gnus-started-hook

A hook that is run after reading the ‘.newsrc’ file(s), but before generating the group buffer.

gnus-check-bogus-newsgroups

If non-nil, Gnus will check for and delete all bogus groups at startup. A bogus group is a group that you have in your ‘.newsrc’ file, but doesn’t exist on the news server. Checking for bogus groups can take quite a while, so to save time and resources it’s best to leave this option off, and do the checking for bogus groups once in a while from the group buffer instead (see Section 2.13 [Group Maintenance], page 24).

gnus-inhibit-startup-message

If non-nil, the startup message won’t be displayed. That way, your boss might not notice as easily that you are reading news instead of doing your job. Note that this variable is used before ‘.gnus.el’ is loaded, so it should be set in .emacs instead.

gnus-no-groups-message

Message displayed by Gnus when no groups are available.

gnus-play-startup-jingle

If non-nil, play the Gnus jingle at startup.

gnus-startup-jingle

Jingle to be played if the above variable is non-nil. The default is

‘Tuxedomoon.Jingle4.au’.

2 The Group Buffer

The group buffer lists all (or parts) of the available groups. It is the first buffer shown when Gnus starts, and will never be killed as long as Gnus is active.

2.1 Group Buffer Format

2.1.1 Group Line Specification

The default format of the group buffer is nice and dull, but you can make it as exciting and ugly as you feel like.

Here’s a couple of example group lines:

          25: news.announce.newusers
      *    0: alt.fan.andrea-dworkin

Quite simple, huh?

You can see that there are 25 unread articles in ‘news.announce.newusers’. There are no unread articles, but some ticked articles, in ‘alt.fan.andrea-dworkin’ (see that little asterisk at the beginning of the line?).

You can change that format to whatever you want by fiddling with the gnus-group-line-format variable. This variable works along the lines of a format specification, which is pretty much the same as a printf specifications, for those of you who use (feh!) C. See Section 8.4 [Formatting Variables], page 154.

‘%M%S%5y: %(%g%)\n’ is the value that produced those lines above.

There should always be a colon on the line; the cursor always moves to the colon after performing an operation. Nothing else is required—not even the group name. All displayed text is just window dressing, and is never examined by Gnus. Gnus stores all real information it needs using text properties.

(Note that if you make a really strange, wonderful, spreadsheet-like layout, everybody will believe you are hard at work with the accounting instead of wasting time reading news.)

Here’s a list of all available format characters:

‘M’ An asterisk if the group only has marked articles.

‘S’ Whether the group is subscribed.

‘L’ Level of subscribedness.

‘N’ Number of unread articles.

‘I’ Number of dormant articles.

‘T’ Number of ticked articles.

‘R’ Number of read articles.

‘t’ Estimated total number of articles. (This is really max-number minus minnumber plus 1.)

‘y’ Number of unread, unticked, non-dormant articles.

‘i’ Number of ticked and dormant articles. ‘g’ Full group name. ‘G’ Group name. ‘D’ Newsgroup description. ‘o’ ‘m’ if moderated. ‘O’ ‘(m)’ if moderated. ‘s’ Select method. ‘n’ Select from where. ‘z’ A string that looks like ‘<%s:%n>’ if a foreign select method is used. ‘P’ Indentation based on the level of the topic (see Section 2.16 [Group Topics],

page 25).

‘c’ Short (collapsed) group name. The gnus-group-uncollapsed-levels variable says how many levels to leave at the end of the group name. The default is 1—this will mean that group names like ‘gnu.emacs.gnus’ will be shortened

          to ‘g.emacs.gnus’.

‘m’ ‘%’ (gnus-new-mail-mark) if there has arrived new mail to the group lately. ‘d’ A string that says when you last read the group (see Section 2.17.3 [Group

Timestamp], page 31).

‘u’ User defined specifier. The next character in the format string should be a letter. Gnus will call the function gnus-user-format-function-‘X’, where ‘X’ is the letter following ‘%u’. The function will be passed a single dummy parameter as argument. The function should return a string, which will be inserted into the buffer just like information from any other specifier.

All the “number-of” specs will be filled with an asterisk (‘*’) if no info is available—for instance, if it is a non-activated foreign group, or a bogus native group.

2.1.2 Group Modeline Specification

The mode line can be changed by setting gnus-group-mode-line-format (see Section 8.4.2 [Mode Line Formatting], page 155). It doesn’t understand that many format specifiers: ‘S’ The native news server. ‘M’ The native select method.

2.1.3 Group Highlighting

Highlighting in the group buffer is controlled by the gnus-group-highlight variable. This is an alist with elements that look like (form . face). If form evaluates to something non-nil, the face will be used on the line.

Here’s an example value for this variable that might look nice if the background is dark:

     (face-spec-set 'my-group-face-1
                    '((t (:foreground "Red" :bold t))))
     (face-spec-set 'my-group-face-2
                    '((t (:foreground "SeaGreen" :bold t))))
     (face-spec-set 'my-group-face-3
                    '((t (:foreground "SpringGreen" :bold t))))
     (face-spec-set 'my-group-face-4
                    '((t (:foreground "SteelBlue" :bold t))))
     (face-spec-set 'my-group-face-5
                    '((t (:foreground "SkyBlue" :bold t))))
     (setq gnus-group-highlight
           '(((> unread 200) . my-group-face-1)
             ((and (< level 3) (zerop unread)) . my-group-face-2)
             ((< level 3) . my-group-face-3)
             ((zerop unread) . my-group-face-4)
             (t . my-group-face-5)))

Also see Section 8.6 [Faces and Fonts], page 160.

Variables that are dynamically bound when the forms are evaluated include:

group The group name.

unread The number of unread articles in the group.

method The select method.

mailp Whether the group is a mail group.

level The level of the group.

score The score of the group.

ticked The number of ticked articles in the group.

total The total number of articles in the group. Or rather, MAX-NUMBER minus MIN-NUMBER plus one.

topic When using the topic minor mode, this variable is bound to the current topic being inserted.

When the forms are evaled, point is at the beginning of the line of the group in question, so you can use many of the normal Gnus functions for snarfing info on the group.

gnus-group-update-hook is called when a group line is changed. It will not be called

when gnus-visual is nil. This hook calls gnus-group-highlight-line by default.

2.2 Group Maneuvering

All movement commands understand the numeric prefix and will behave as expected, hopefully.

n Go to the next group that has unread articles (gnus-group-next-unread-

          group).

p

DEL Go to the previous group that has unread articles (gnus-group-prev-unread-

group).
N  Go to the next group (gnus-group-next-group).

P Go to the previous group (gnus-group-prev-group).

M-n Go to the next unread group on the same (or lower) level (gnus-group-next-

unread-group-same-level).

M-p Go to the previous unread group on the same (or lower) level (gnus-group-

prev-unread-group-same-level).

Three commands for jumping to groups:

j Jump to a group (and make it visible if it isn’t already) (gnus-group-jump-to-group). Killed groups can be jumped to, just like living groups.

, Jump to the unread group with the lowest level (gnus-group-best-unread-

group).

. Jump to the first group with unread articles (gnus-group-first-unread-

group).

If gnus-group-goto-unread is nil, all the movement commands will move to the next group, not the next unread group. Even the commands that say they move to the next unread group. The default is t.

2.3 Selecting a Group

SPACE Select the current group, switch to the summary buffer and display the first unread article (gnus-group-read-group). If there are no unread articles in the group, or if you give a non-numerical prefix to this command, Gnus will offer to fetch all the old articles in this group from the server. If you give a numerical prefix N, N determines the number of articles Gnus will fetch. If N is positive, Gnus fetches the N newest articles, if N is negative, Gnus fetches the abs(N) oldest articles.

RET Select the current group and switch to the summary buffer (gnus-group-

select-group). Takes the same arguments as gnus-group-read-group—the

only difference is that this command does not display the first unread article automatically upon group entry.

M-RET This does the same as the command above, but tries to do it with the minimum amount of fuzz (gnus-group-quick-select-group). No scoring/killing will be performed, there will be no highlights and no expunging. This might be useful if you’re in a real hurry and have to enter some humongous group. If you give a 0 prefix to this command (i.e., 0 M-RET), Gnus won’t even generate the summary buffer, which is useful if you want to toggle threading before generating the summary buffer (see Section 3.24.3 [Summary Generation Commands], page 77).

M-SPACE This is yet one more command that does the same as the RET command, but this one does it without expunging and hiding dormants (gnus-group-visible-

          select-group).

M-C-RET Finally, this command selects the current group ephemerally without doing any

          processing of its contents (gnus-group-select-group-ephemerally). Even

threading has been turned off. Everything you do in the group after selecting it in this manner will have no permanent effects.

The gnus-large-newsgroup variable says what Gnus should consider to be a big group. This is 200 by default. If the group has more (unread and/or ticked) articles than this, Gnus will query the user before entering the group. The user can then specify how many articles should be fetched from the server. If the user specifies a negative number (-n), the n oldest articles will be fetched. If it is positive, the n articles that have arrived most recently will be fetched.

gnus-auto-select-first control whether any articles are selected automatically when entering a group with the SPACE command.

nil Don’t select any articles when entering the group. Just display the full summary buffer.

t Select the first unread article when entering the group.

best Select the highest scored article in the group when entering the group.

If you want to prevent automatic selection in some group (say, in a binary group with Huge articles) you can set this variable to nil in gnus-select-group-hook, which is called when a group is selected.

2.4 Subscription Commands

S t

u Toggle subscription to the current group (gnus-group-unsubscribe-current-

          group).
S s

U Prompt for a group to subscribe, and then subscribe it. If it was subscribed already, unsubscribe it instead (gnus-group-unsubscribe-group).

S k
C-k  Kill the current group (gnus-group-kill-group).
S y

C-y Yank the last killed group (gnus-group-yank-group).

C-x C-t Transpose two groups (gnus-group-transpose-groups). This isn’t really a subscription command, but you can use it instead of a kill-and-yank sequence sometimes.

S w

C-w Kill all groups in the region (gnus-group-kill-region).

S z  Kill all zombie groups (gnus-group-kill-all-zombies).

S C-k Kill all groups on a certain level (gnus-group-kill-level). These groups can’t be yanked back after killing, so this command should be used with some caution. The only time where this command comes in really handy is when you have a ‘.newsrc’ with lots of unsubscribed groups that you want to get rid off. S C-k on level 7 will kill off all unsubscribed groups that do not have message numbers in the ‘.newsrc’ file.

Also see Section 2.6 [Group Levels], page 16.

2.5 Group Data

c Mark all unticked articles in this group as read (gnus-group-catchup-

          current). gnus-group-catchup-group-hook is called when catching up a

group from the group buffer.

C Mark all articles in this group, even the ticked ones, as read (gnus-group-

          catchup-current-all).

M-c Clear the data from the current group—nix out marks and the list of read

          articles (gnus-group-clear-data).
M-x gnus-group-clear-data-on-native-groups

If you have switched from one nntp server to another, all your marks and read ranges have become worthless. You can use this command to clear out all data that you have on your native groups. Use with caution.

2.6 Group Levels

All groups have a level of subscribedness. For instance, if a group is on level 2, it is more subscribed than a group on level 5. You can ask Gnus to just list groups on a given level or lower (see Section 2.11 [Listing Groups], page 22), or to just check for new articles in groups on a given level or lower (see Section 2.17.1 [Scanning New Messages], page 30).

Remember: The higher the level of the group, the less important it is.

S l Set the level of the current group. If a numeric prefix is given, the next n groups will have their levels set. The user will be prompted for a level.

Gnus considers groups from levels 1 to gnus-level-subscribed (inclusive) (default

5) to be subscribed, gnus-level-subscribed (exclusive) and gnus-level-unsubscribed

(inclusive) (default 7) to be unsubscribed, gnus-level-zombie to be zombies (walking dead) (default 8) and gnus-level-killed to be killed (completely dead) (default 9). Gnus treats subscribed and unsubscribed groups exactly the same, but zombie and killed groups have no information on what articles you have read, etc, stored. This distinction between dead and living groups isn’t done because it is nice or clever, it is done purely for reasons of efficiency.

It is recommended that you keep all your mail groups (if any) on quite low levels (e.g. 1 or 2).

If you want to play with the level variables, you should show some care. Set them once, and don’t touch them ever again. Better yet, don’t touch them at all unless you know exactly what you’re doing.

Two closely related variables are gnus-level-default-subscribed (default 3) and gnus-level-default-unsubscribed (default 6), which are the levels that new groups will be put on if they are (un)subscribed. These two variables should, of course, be inside the relevant valid ranges.

If gnus-keep-same-level is non-nil, some movement commands will only move to groups of the same level (or lower). In particular, going from the last article in one group to the next group will go to the next group of the same level (or lower). This might be handy if you want to read the most important groups before you read the rest.

All groups with a level less than or equal to gnus-group-default-list-level will be listed in the group buffer by default.

If gnus-group-list-inactive-groups is non-nil, non-active groups will be listed along with the unread groups. This variable is t by default. If it is nil, inactive groups won’t be listed.

   If gnus-group-use-permanent-levels is non-nil, once you give a level prefix to g or

l, all subsequent commands will use this level as the “work” level.

Gnus will normally just activate (i. e., query the server about) groups on level gnusactivate-level or less. If you don’t want to activate unsubscribed groups, for instance, you might set this variable to 5. The default is 6.

2.7 Group Score

You would normally keep important groups on high levels, but that scheme is somewhat restrictive. Don’t you wish you could have Gnus sort the group buffer according to how often you read groups, perhaps? Within reason?

This is what group score is for. You can assign a score to each group. You can then sort the group buffer based on this score. Alternatively, you can sort on score and then level. (Taken together, the level and the score is called the rank of the group. A group that is on level 4 and has a score of 1 has a higher rank than a group on level 5 that has a score of 300. (The level is the most significant part and the score is the least significant part.))

If you want groups you read often to get higher scores than groups you read seldom

you can add the gnus-summary-bubble-group function to the gnus-summary-exit-hook

hook. This will result (after sorting) in a bubbling sort of action. If you want to see that in action after each summary exit, you can add gnus-group-sort-groups-by-rank or gnus-group-sort-groups-by-score to the same hook, but that will slow things down somewhat.

2.8 Marking Groups

If you want to perform some command on several groups, and they appear subsequently in the group buffer, you would normally just give a numerical prefix to the command. Most group commands will then do your bidding on those groups.

However, if the groups are not in sequential order, you can still perform a command on several groups. You simply mark the groups first with the process mark and then execute the command.

#

M m Set the mark on the current group (gnus-group-mark-group).

M-#

M u Remove the mark from the current group (gnus-group-unmark-group).

M U  Remove the mark from all groups (gnus-group-unmark-all-groups).

M w Mark all groups between point and mark (gnus-group-mark-region).

M b Mark all groups in the buffer (gnus-group-mark-buffer).

M r Mark all groups that match some regular expression (gnus-group-mark-

          regexp).

Also see Section 8.1 [Process/Prefix], page 153.

If you want to execute some command on all groups that have been marked with the process mark, you can use the M-& (gnus-group-universal-argument) command. It will prompt you for the command to be executed.

2.9 Foreign Groups

Below are some group mode commands for making and editing general foreign groups, as well as commands to ease the creation of a few special-purpose groups. All these commands insert the newly created groups under point—gnus-subscribe-newsgroup-method is not consulted.

G m Make a new group (gnus-group-make-group). Gnus will prompt you for a name, a method and possibly an address. For an easier way to subscribe to nntp groups, see Section 2.14 [Browse Foreign Server], page 24.

G r Rename the current group to something else (gnus-group-rename-group). This is valid only on some groups—mail groups mostly. This command might very well be quite slow on some backends.

G c Customize the group parameters (gnus-group-customize).

G e Enter a buffer where you can edit the select method of the current group (gnus-

          group-edit-group-method).

G p Enter a buffer where you can edit the group parameters (gnus-group-edit-

          group-parameters).

G E Enter a buffer where you can edit the group info (gnus-group-edit-group).

G d Make a directory group (see Section 6.4.1 [Directory Groups], page 113). You will be prompted for a directory name (gnus-group-make-directory-group).

G h  Make the Gnus help group (gnus-group-make-help-group).

G a Make a Gnus archive group (gnus-group-make-archive-group). By default a group pointing to the most recent articles will be created (gnus-group-recent-archive-directory), but given a prefix, a full group will be created from

          gnus-group-archive-directory.

G k Make a kiboze group. You will be prompted for a name, for a regexp to match groups to be “included” in the kiboze group, and a series of strings to match on headers (gnus-group-make-kiboze-group). See Section 6.5.2 [Kibozed Groups], page 123.

G D Read an arbitrary directory as if it were a newsgroup with the nneething backend (gnus-group-enter-directory). See Section 6.4.2 [Anything Groups], page 113.

G f Make a group based on some file or other (gnus-group-make-doc-group). If you give a prefix to this command, you will be prompted for a file name and a file type. Currently supported types are babyl, mbox, digest, mmdf, news,

rnews, clari-briefs, rfc934, rfc822-forward, and forward. If you run this

command without a prefix, Gnus will guess at the file type. See Section 6.4.3 [Document Groups], page 114.

G u Create one of the groups mentioned in gnus-useful-groups (gnus-group-

make-useful-group).

G w Make an ephemeral group based on a web search (gnus-group-make-web-group). If you give a prefix to this command, make a solid group instead. You will be prompted for the search engine type and the search string. Valid search engine types include dejanews, altavista and reference. See Section 6.4.5 [Web Searches], page 120. If you use the dejanews search engine, you can limit the search to a particular group by using a match string like ‘~g alt.sysadmin.recovery shaving’.

G DEL This function will delete the current group (gnus-group-delete-group). If given a prefix, this function will actually delete all the articles in the group, and forcibly remove the group itself from the face of the Earth. Use a prefix only if you are absolutely sure of what you are doing. This command can’t be used on read-only groups (like nntp group), though.

G V  Make a new, fresh, empty nnvirtual group (gnus-group-make-empty-

virtual). See Section 6.5.1 [Virtual Groups], page 122.

G v  Add the current group to an nnvirtual group (gnus-group-add-to-virtual).

Uses the process/prefix convention.

See Chapter 6 [Select Methods], page 91, for more information on the various select methods.

If gnus-activate-foreign-newsgroups is a positive number, Gnus will check all foreign groups with this level or lower at startup. This might take quite a while, especially if you subscribe to lots of groups from different nntp servers. Also see Section 2.6 [Group Levels], page 16; gnus-activate-level also affects activation of foreign newsgroups.

2.10 Group Parameters

The group parameters store information local to a particular group. Here’s an example group parameter list:

((to-address . "ding@gnus.org")
(auto-expire . t))

We see that each element consists of a "dotted pair"—the thing before the dot is the key, while the thing after the dot is the value. All the parameters have this form except local variable specs, which are not dotted pairs, but proper lists.

The following group parameters can be used:

to-address

Address used by when doing followups and new posts.

     (to-address .  "some@where.com")

This is primarily useful in mail groups that represent closed mailing lists— mailing lists where it’s expected that everybody that writes to the mailing list is subscribed to it. Since using this parameter ensures that the mail only goes to the mailing list itself, it means that members won’t receive two copies of your followups. Using to-address will actually work whether the group is foreign or not. Let’s say there’s a group on the server that is called ‘fa.4ad-l’. This is a real newsgroup, but the server has gotten the articles from a mail-to-news gateway. Posting directly to this group is therefore impossible—you have to send mail to the mailing list address instead.

to-list Address used when doing a a in that group.

     (to-list . "some@where.com")

It is totally ignored when doing a followup—except that if it is present in a news group, you’ll get mail group semantics when doing f. If you do an a command in a mail group and you have neither a to-list group parameter nor a to-address group parameter, then a to-list group parameter will be added automatically upon sending the message if gnus-add-to-list is set to t. If you do an a command in a mail group and you don’t have a to-list group parameter, one will be added automatically upon sending the message.

visible If the group parameter list has the element (visible . t), that group will always be visible in the Group buffer, regardless of whether it has any unread articles.

broken-reply-to

Elements like (broken-reply-to . t) signals that Reply-To headers in this group are to be ignored. This can be useful if you’re reading a mailing list group where the listserv has inserted Reply-To headers that point back to the listserv itself. This is broken behavior. So there!

to-group Elements like (to-group . "some.group.name") means that all posts in that

group will be sent to some.group.name.

newsgroup

If you have (newsgroup . t) in the group parameter list, Gnus will treat all responses as if they were responses to news articles. This can be useful if you have a mail group that’s really a mirror of a news group. gcc-self If (gcc-self . t) is present in the group parameter list, newly composed mes-

sages will be Gcc’d to the current group. If (gcc-self . none) is present, no Gcc: header will be generated, if (gcc-self . "string") is present, this string will be inserted literally as a gcc header. This parameter takes precedence over any default Gcc rules as described later (see Section 5.5 [Archived Messages], page 86).

auto-expire

If the group parameter has an element that looks like (auto-expire . t), all articles read will be marked as expirable. For an alternative approach, see Section 6.3.7 [Expiring Mail], page 106.

total-expire

If the group parameter has an element that looks like (total-expire . t), all read articles will be put through the expiry process, even if they are not marked as expirable. Use with caution. Unread, ticked and dormant articles are not eligible for expiry.

expiry-wait

If the group parameter has an element that looks like (expiry-wait . 10),

this value will override any nnmail-expiry-wait and nnmail-expiry-wait-

function when expiring expirable messages. The value can either be a number of days (not necessarily an integer) or the symbols never or immediate.

score-file

Elements that look like (score-file . "file") will make ‘file’ into the current score file for the group in question. All interactive score entries will be put into this file.

adapt-file

Elements that look like (adapt-file . "file") will make ‘file’ into the current adaptive file for the group in question. All adaptive score entries will be put into this file.

admin-address

When unsubscribing from a mailing list you should never send the unsubscription notice to the mailing list itself. Instead, you’d send messages to the administrative address. This parameter allows you to put the admin address somewhere convenient.

display Elements that look like (display . MODE) say which articles to display on entering the group. Valid values are:

all Display all articles, both read and unread.

default Display the default visible articles, which normally includes unread and ticked articles.

comment Elements that look like (comment . "This is a comment") are arbitrary comments on the group. They are currently ignored by Gnus, but provide a place for you to store information on particular groups.

(variable form)

You can use the group parameters to set variables local to the group you are entering. If you want to turn threading off in ‘news.answers’, you could put

(gnus-show-threads nil) in the group parameters of that group. gnus-show-

threads will be made into a local variable in the summary buffer you enter, and the form nil will be evaled there. This can also be used as a group-specific hook function, if you’d like. If you want to hear a beep when you enter a group, you could put something like

(dummy-variable (ding)) in the parameters of that group. dummy-variable

will be set to the result of the (ding) form, but who cares?

Use the G p command to edit group parameters of a group. You might also be interested in reading about topic parameters (see Section 2.16.5 [Topic Parameters], page 29).

2.11 Listing Groups

These commands all list various slices of the groups available.

l

A s List all groups that have unread articles (gnus-group-list-groups). If the numeric prefix is used, this command will list only groups of level ARG and lower. By default, it only lists groups of level five (i. e., gnus-group-default-list-level) or lower (i.e., just subscribed groups).

L

A u List all groups, whether they have unread articles or not (gnus-group-list-all-groups). If the numeric prefix is used, this command will list only groups of level ARG and lower. By default, it lists groups of level seven or lower (i.e., just subscribed and unsubscribed groups).

A l List all unread groups on a specific level (gnus-group-list-level). If given a prefix, also list the groups with no unread articles.

A k List all killed groups (gnus-group-list-killed). If given a prefix argument, really list all groups that are available, but aren’t currently (un)subscribed. This could entail reading the active file from the server.

A z  List all zombie groups (gnus-group-list-zombies).

A m List all unread, subscribed groups with names that match a regexp (gnus-

group-list-matching).
A M  List groups that match a regexp (gnus-group-list-all-matching).

A A List absolutely all groups in the active file(s) of the server(s) you are connected to (gnus-group-list-active). This might very well take quite a while. It might actually be a better idea to do a A M to list all matching, and just give ‘.’ as the thing to match on. Also note that this command may list groups that don’t exist (yet)—these will be listed as if they were killed groups. Take the output with some grains of salt.

A a List all groups that have names that match a regexp (gnus-group-apropos).

A d List all groups that have names or descriptions that match a regexp (gnus-

          group-description-apropos).

Groups that match the gnus-permanently-visible-groups regexp will always be shown, whether they have unread articles or not. You can also add the visible element to the group parameters in question to get the same effect.

Groups that have just ticked articles in it are normally listed in the group buffer. If

gnus-list-groups-with-ticked-articles is nil, these groups will be treated just like

totally empty groups. It is t by default.

2.12 Sorting Groups

The C-c C-s (gnus-group-sort-groups) command sorts the group buffer according to the function(s) given by the gnus-group-sort-function variable. Available sorting functions include:

gnus-group-sort-by-alphabet

Sort the group names alphabetically. This is the default.

gnus-group-sort-by-real-name

Sort the group alphabetically on the real (unprefixed) group names.

gnus-group-sort-by-level

Sort by group level.

gnus-group-sort-by-score

Sort by group score. See Section 2.7 [Group Score], page 17.

gnus-group-sort-by-rank

Sort by group score and then the group level. The level and the score are, when taken together, the group’s rank. See Section 2.7 [Group Score], page 17.

gnus-group-sort-by-unread

Sort by number of unread articles.

gnus-group-sort-by-method

Sort alphabetically on the select method.

gnus-group-sort-function can also be a list of sorting functions. In that case, the most significant sort key function must be the last one.

There are also a number of commands for sorting directly according to some sorting criteria:

G S a Sort the group buffer alphabetically by group name (gnus-group-sort-

          groups-by-alphabet).

G S u Sort the group buffer by the number of unread articles (gnus-group-sort-

          groups-by-unread).
G S l  Sort the group buffer by group level (gnus-group-sort-groups-by-level).
G S v  Sort the group buffer by group score (gnus-group-sort-groups-by-score).

See Section 2.7 [Group Score], page 17.

G S r Sort the group buffer by group rank (gnus-group-sort-groups-by-rank). See Section 2.7 [Group Score], page 17.

G S m Sort the group buffer alphabetically by backend name (gnus-group-sort-

groups-by-method).

When given a prefix, all these commands will sort in reverse order. You can also sort a subset of the groups:

G P a Sort the process/prefixed groups in the group buffer alphabetically by group

name (gnus-group-sort-selected-groups-by-alphabet).

G P u Sort the process/prefixed groups in the group buffer by the number of unread

articles (gnus-group-sort-selected-groups-by-unread).

G P l Sort the process/prefixed groups in the group buffer by group level (gnus-

group-sort-selected-groups-by-level).

G P v Sort the process/prefixed groups in the group buffer by group score

(gnus-group-sort-selected-groups-by-score). See Section 2.7 [Group

Score], page 17.

G P r Sort the process/prefixed groups in the group buffer by group rank

(gnus-group-sort-selected-groups-by-rank). See Section 2.7 [Group

Score], page 17.

G P m Sort the process/prefixed groups in the group buffer alphabetically by backend

name (gnus-group-sort-selected-groups-by-method).

2.13 Group Maintenance

b Find bogus groups and delete them (gnus-group-check-bogus-groups).

F Find new groups and process them (gnus-group-find-new-groups). With 1 C-u, use the ask-server method to query the server for new groups. With 2 C-u’s, use most complete method possible to query the server for new groups, and subscribe the new groups as zombies.

C-c C-x Run all expirable articles in the current group through the expiry process (if

any) (gnus-group-expire-articles).

C-c M-C-x Run all articles in all groups through the expiry process (gnus-group-expire-

all-groups).

2.14 Browse Foreign Server

B You will be queried for a select method and a server name. Gnus will then attempt to contact this server and let you browse the groups there (gnus-

group-browse-foreign-server).

A new buffer with a list of available groups will appear. This buffer will use the gnusbrowse-mode. This buffer looks a bit (well, a lot) like a normal group buffer.

Here’s a list of keystrokes available in the browse mode:

n  Go to the next group (gnus-group-next-group).

p Go to the previous group (gnus-group-prev-group).

SPACE Enter the current group and display the first article (gnus-browse-read-

          group).
RET  Enter the current group (gnus-browse-select-group).

u Unsubscribe to the current group, or, as will be the case here, subscribe to it

          (gnus-browse-unsubscribe-current-group).
l
q  Exit browse mode (gnus-browse-exit).

? Describe browse mode briefly (well, there’s not much to describe, is there)

          (gnus-browse-describe-briefly).

2.15 Exiting Gnus

Yes, Gnus is ex(c)iting.

z Suspend Gnus (gnus-group-suspend). This doesn’t really exit Gnus, but it kills all buffers except the Group buffer. I’m not sure why this is a gain, but then who am I to judge?

q  Quit Gnus (gnus-group-exit).

Q Quit Gnus without saving the ‘.newsrc’ files (gnus-group-quit). The dribble file will be saved, though (see Section 1.9 [Auto Save], page 9).

   gnus-suspend-gnus-hook is called when you suspend Gnus and gnus-exit-gnus-hook

is called when you quit Gnus, while gnus-after-exiting-gnus-hook is called as the final item when exiting Gnus.

If you wish to completely unload Gnus and all its adherents, you can use the gnus-unload command. This command is also very handy when trying to customize meta-variables.

Note:

Miss Lisa Cannifax, while sitting in English class, felt her feet go numbly heavy and herself fall into a hazy trance as the boy sitting behind her drew repeated lines with his pencil across the back of her plastic chair.

2.16 Group Topics

If you read lots and lots of groups, it might be convenient to group them hierarchically according to topics. You put your Emacs groups over here, your sex groups over there, and the rest (what, two groups or so?) you put in some misc section that you never bother with anyway. You can even group the Emacs sex groups as a sub-topic to either the Emacs groups or the sex groups—or both! Go wild!

Here’s an example:

     Gnus
       Emacs -- I wuw it!
          3: comp.emacs
          2: alt.religion.emacs
         Naughty Emacs
          452: alt.sex.emacs
            0: comp.talk.emacs.recovery
       Misc
          8: comp.binaries.fractals
         13: comp.sources.unix

To get this fab functionality you simply turn on (ooh!) the gnus-topic minor mode— type t in the group buffer. (This is a toggling command.)

Go ahead, just try it. I’ll still be here when you get back. La de dum... Nice tune, that... la la la... What, you’re back? Yes, and now press l. There. All your groups are now listed under ‘misc’. Doesn’t that make you feel all warm and fuzzy? Hot and bothered?

If you want this permanently enabled, you should add that minor mode to the hook for the group mode:

     (add-hook 'gnus-group-mode-hook 'gnus-topic-mode)

2.16.1 Topic Variables

Now, if you select a topic, it will fold/unfold that topic, which is really neat, I think. The topic lines themselves are created according to the gnus-topic-line-format vari-

able (see Section 8.4 [Formatting Variables], page 154). Valid elements are:

‘i’ Indentation.

‘n’ Topic name.

‘v’ Visibility.

‘l’ Level.

‘g’ Number of groups in the topic.

‘a’ Number of unread articles in the topic.

‘A’ Number of unread articles in the topic and all its subtopics.

Each sub-topic (and the groups in the sub-topics) will be indented with gnus-topic-indent-level times the topic level number of spaces. The default is 2.

gnus-topic-mode-hook is called in topic minor mode buffers.

The gnus-topic-display-empty-topics says whether to display even topics that have no unread articles in them. The default is t.

2.16.2 Topic Commands

When the topic minor mode is turned on, a new T submap will be available. In addition, a few of the standard keys change their definitions slightly.

T n Prompt for a new topic name and create it (gnus-topic-create-topic).

T m Move the current group to some other topic (gnus-topic-move-group). This command uses the process/prefix convention (see Section 8.1 [Process/Prefix], page 153).

T c Copy the current group to some other topic (gnus-topic-copy-group). This command uses the process/prefix convention (see Section 8.1 [Process/Prefix], page 153).

T D Remove a group from the current topic (gnus-topic-remove-group). This command is mainly useful if you have the same group in several topics and wish to remove it from one of the topics. You may also remove a group from all topics, but in that case, Gnus will add it to the root topic the next time you start Gnus. In fact, all new groups (which, naturally, don’t belong to any topic) will show up in the root topic. This command uses the process/prefix convention (see Section 8.1 [Process/Prefix], page 153).

T M Move all groups that match some regular expression to a topic (gnus-topic-

move-matching).

T C Copy all groups that match some regular expression to a topic (gnus-topic-

copy-matching).
T H  Toggle hiding empty topics (gnus-topic-toggle-display-empty-topics).

T # Mark all groups in the current topic with the process mark (gnus-topic-mark-

topic).

T M-# Remove the process mark from all groups in the current topic (gnus-topic-

unmark-topic).
RET

SPACE Either select a group or fold a topic (gnus-topic-select-group). When you perform this command on a group, you’ll enter the group, as usual. When done on a topic line, the topic will be folded (if it was visible) or unfolded (if it was folded already). So it’s basically a toggling command on topics. In addition, if you give a numerical prefix, group on that level (and lower) will be displayed.

T TAB

TAB “Indent” the current topic so that it becomes a sub-topic of the previous topic (gnus-topic-indent). If given a prefix, “un-indent” the topic instead.

M-TAB “Un-indent” the current topic so that it becomes a sub-topic of the parent of

its current parent (gnus-topic-unindent).

C-k Kill a group or topic (gnus-topic-kill-group). All groups in the topic will be removed along with the topic.

C-y Yank the previously killed group or topic (gnus-topic-yank-group). Note that all topics will be yanked before all groups.

T r  Rename a topic (gnus-topic-rename).
T DEL  Delete an empty topic (gnus-topic-delete).

A T List all groups that Gnus knows about in a topics-ified way (gnus-topic-list-

          active).

G p Edit the topic parameters (gnus-topic-edit-parameters). See Section 2.16.5 [Topic Parameters], page 29.

2.16.3 Topic Sorting

You can sort the groups in each topic individually with the following commands:

T S a Sort the current topic alphabetically by group name (gnus-topic-sort-

          groups-by-alphabet).

T S u Sort the current topic by the number of unread articles (gnus-topic-sort-

          groups-by-unread).
T S l  Sort the current topic by group level (gnus-topic-sort-groups-by-level).
T S v  Sort the current topic by group score (gnus-topic-sort-groups-by-score).

See Section 2.7 [Group Score], page 17.

T S r Sort the current topic by group rank (gnus-topic-sort-groups-by-rank). See Section 2.7 [Group Score], page 17.

T S m Sort the current topic alphabetically by backend name (gnus-topic-sort-

          groups-by-method).

See Section 2.12 [Sorting Groups], page 23, for more information about group sorting.

2.16.4 Topic Topology

So, let’s have a look at an example group buffer:

     Gnus
       Emacs -- I wuw it!
          3: comp.emacs
          2: alt.religion.emacs
         Naughty Emacs
          452: alt.sex.emacs
            0: comp.talk.emacs.recovery
       Misc
          8: comp.binaries.fractals
         13: comp.sources.unix

So, here we have one top-level topic (‘Gnus’), two topics under that, and one sub-topic under one of the sub-topics. (There is always just one (1) top-level topic). This topology can be expressed as follows:

     (("Gnus" visible)
      (("Emacs -- I wuw it!" visible)
       (("Naughty Emacs" visible)))
      (("Misc" visible)))

This is in fact how the variable gnus-topic-topology would look for the display above. That variable is saved in the ‘.newsrc.eld’ file, and shouldn’t be messed with manually— unless you really want to. Since this variable is read from the ‘.newsrc.eld’ file, setting it in any other startup files will have no effect.

This topology shows what topics are sub-topics of what topics (right), and which topics are visible. Two settings are currently allowed—visible and invisible.

2.16.5 Topic Parameters

All groups in a topic will inherit group parameters from the parent (and ancestor) topic parameters. All valid group parameters are valid topic parameters (see Section 2.10 [Group Parameters], page 19).

Group parameters (of course) override topic parameters, and topic parameters in subtopics override topic parameters in super-topics. You know. Normal inheritance rules. (Rules is here a noun, not a verb, although you may feel free to disagree with me here.)

     Gnus
       Emacs
          3: comp.emacs
          2: alt.religion.emacs
        452: alt.sex.emacs
         Relief
          452: alt.sex.emacs
            0: comp.talk.emacs.recovery
       Misc
          8: comp.binaries.fractals
         13: comp.sources.unix
        452: alt.sex.emacs

The ‘Emacs’ topic has the topic parameter (score-file . "emacs.SCORE"); the

‘Relief’ topic has the topic parameter (score-file . "relief.SCORE"); and the ‘Misc’

topic has the topic parameter (score-file . "emacs.SCORE"). In addition,

‘alt.religion.emacs’ has the group parameter (score-file . "religion.SCORE").

Now, when you enter ‘alt.sex.emacs’ in the ‘Relief’ topic, you will get the ‘relief.SCORE’ home score file. If you enter the same group in the ‘Emacs’ topic, you’ll get the ‘emacs.SCORE’ home score file. If you enter the group ‘alt.religion.emacs’, you’ll get the ‘religion.SCORE’ home score file.

This seems rather simple and self-evident, doesn’t it? Well, yes. But there are some problems, especially with the total-expiry parameter. Say you have a mail group in two topics; one with total-expiry and one without. What happens when you do M-x gnusexpire-all-expirable-groups? Gnus has no way of telling which one of these topics you mean to expire articles from, so anything may happen. In fact, I hereby declare that it is undefined what happens. You just have to be careful if you do stuff like that.

2.17 Misc Group Stuff

^Enter the server buffer (gnus-group-enter-server-mode). See Section 6.1 [The Server Buffer], page 91.

a Post an article to a group (gnus-group-post-news). If given a prefix, the current group name will be used as the default.

m Mail a message somewhere (gnus-group-mail).

Variables for the group buffer:

gnus-group-mode-hook

is called after the group buffer has been created.

gnus-group-prepare-hook

is called after the group buffer is generated. It may be used to modify the buffer in some strange, unnatural way.

gnus-group-prepared-hook

is called as the very last thing after the group buffer has been generated. It may be used to move point around, for instance.

gnus-permanently-visible-groups

Groups matching this regexp will always be listed in the group buffer, whether they are empty or not.

2.17.1 Scanning New Messages

g Check the server(s) for new articles. If the numerical prefix is used, this command will check only groups of level arg and lower (gnus-group-get-new-news). If given a non-numerical prefix, this command will force a total rereading of the active file(s) from the backend(s).

M-g Check whether new articles have arrived in the current group (gnus-group-

get-new-news-this-group). gnus-goto-next-group-when-activating says

whether this command is to move point to the next group or not. It is t by default.

C-c M-g Activate absolutely all groups (gnus-activate-all-groups).

R Restart Gnus (gnus-group-restart). This saves the ‘.newsrc’ file(s), closes the connection to all servers, clears up all run-time Gnus variables, and then starts Gnus all over again.

gnus-get-new-news-hook is run just before checking for new news.

gnus-after-getting-new-news-hook is run after checking for new news.

2.17.2 Group Information

H f Try to fetch the FAQ for the current group (gnus-group-fetch-faq). Gnus will try to get the FAQ from gnus-group-faq-directory, which is usually a directory on a remote machine. This variable can also be a list of directories. In that case, giving a prefix to this command will allow you to choose between the various sites. ange-ftp (or efs) will be used for fetching the file. If fetching from the first site is unsuccessful, Gnus will attempt to go through gnus-group-faq-directory and try to open them one by one.

H d

C-c C-d Describe the current group (gnus-group-describe-group). If given a prefix, force Gnus to re-read the description from the server.

M-d Describe all groups (gnus-group-describe-all-groups). If given a prefix, force Gnus to re-read the description file from the server.

H v

V Display current Gnus version numbers (gnus-version).

?  Give a very short help message (gnus-group-describe-briefly).
C-c C-i Go to the Gnus info node (gnus-info-find-node).

2.17.3 Group Timestamp

It can be convenient to let Gnus keep track of when you last read a group. To set the

ball rolling, you should add gnus-group-set-timestamp to gnus-select-group-hook:
     (add-hook 'gnus-select-group-hook 'gnus-group-set-timestamp)

After doing this, each time you enter a group, it’ll be recorded.

This information can be displayed in various ways—the easiest is to use the ‘%d’ spec in the group line format:

     (setq gnus-group-line-format
           "%M\%S\%p\%P\%5y: %(%-40,40g%) %d\n")

This will result in lines looking like:

     *        0: mail.ding                                19961002T012943
              0: custom                                   19961002T012713

As you can see, the date is displayed in compact ISO 8601 format. This may be a bit too much, so to just display the date, you could say something like:

     (setq gnus-group-line-format
           "%M\%S\%p\%P\%5y: %(%-40,40g%) %6,6~(cut 2)d\n")

2.17.4 File Commands

r Re-read the init file (gnus-init-file, which defaults to ‘~/.gnus’) (gnus-

          group-read-init-file).
s  Save the ‘.newsrc.eld’ file (and ‘.newsrc’ if wanted) (gnus-group-save-

newsrc). If given a prefix, force saving the file(s) whether Gnus thinks it is necessary or not.

32 Gnus Manual

3 The Summary Buffer

A line for each article is displayed in the summary buffer. You can move around, read articles, post articles and reply to articles.

The most common way to a summary buffer is to select a group from the group buffer (see Section 2.3 [Selecting a Group], page 14).

You can have as many summary buffers open as you wish.

3.1 Summary Buffer Format

Gnus will use the value of the gnus-extract-address-components variable as a function for getting the name and address parts of a From header. Two pre-defined functions exist: gnus-extract-address-components, which is the default, quite fast, and too simplistic solution; and mail-extract-address-components, which works very nicely, but is slower. The default function will return the wrong answer in 5% of the cases. If this is unacceptable to you, use the other function instead.

gnus-summary-same-subject is a string indicating that the current article has the same subject as the previous. This string will be used with those specs that require it. The default is "".

3.1.1 Summary Buffer Lines

You can change the format of the lines in the summary buffer by changing the gnussummary-line-format variable. It works along the same lines as a normal format string, with some extensions (see Section 8.4 [Formatting Variables], page 154).

   The default string is ‘%U%R%z%I%(%[%4L: %-20,20n%]%) %s\n’.

The following format specification characters are understood:

‘N’ Article number.

‘S’ Subject string.

‘s’ Subject if the article is the root of the thread or the previous article had a differ-

          ent subject, gnus-summary-same-subject otherwise. (gnus-summary-same-

subject defaults to "".)

‘F’ Full From header.

‘n’ The name (from the From header).

‘a’ The name (from the From header). This differs from the n spec in that it uses the function designated by the gnus-extract-address-components variable, which is slower, but may be more thorough.

‘A’ The address (from the From header). This works the same way as the a spec.

‘L’ Number of lines in the article.

‘c’ Number of characters in the article.

‘I’ Indentation based on thread level (see Section 3.9.1 [Customizing Threading], page 47).

‘T’ Nothing if the article is a root and lots of spaces if it isn’t (it pushes everything

after it off the screen).

‘[’ Opening bracket, which is normally ‘[’, but can also be ‘<’ for adopted articles

(see Section 3.9.1 [Customizing Threading], page 47).

‘]’ Closing bracket, which is normally ‘]’, but can also be ‘>’ for adopted articles. ‘>’ One space for each thread level. ‘<’ Twenty minus thread level spaces. ‘U’ Unread. ‘R’ This misleadingly named specifier is the secondary mark. This mark will say

whether the article has been replied to, has been cached, or has been saved.

‘i’ Score as a number (see Chapter 7 [Scoring], page 131). ‘z’ Zcore, ‘+’ if above the default level and ‘-’ if below the default level. If the

difference between gnus-summary-default-score and the score is less than

          gnus-summary-zcore-fuzz, this spec will not be used.

‘V’ Total thread score.

‘x’  Xref.
‘D’  Date.

‘d’ The Date in DD-MMM format. ‘o’ The Date in YYYYMMDDTHHMMSS format.

‘M’  Message-ID.
‘r’  References.

‘t’ Number of articles in the current sub-thread. Using this spec will slow down

summary buffer generation somewhat.

‘e’ An ‘=’ (gnus-not-empty-thread-mark) will be displayed if the article has any

children.

‘P’ The line number. ‘O’ Download mark. ‘u’ User defined specifier. The next character in the format string should be a

letter. Gnus will call the function gnus-user-format-function-‘X’, where ‘X’ is the letter following ‘%u’. The function will be passed the current header as argument. The function should return a string, which will be inserted into the summary just like information from any other summary specifier.

The ‘%U’ (status), ‘%R’ (replied) and ‘%z’ (zcore) specs have to be handled with care. For reasons of efficiency, Gnus will compute what column these characters will end up in, and “hard-code” that. This means that it is invalid to have these specs after a variable-length spec. Well, you might not be arrested, but your summary buffer will look strange, which is bad enough.

The smart choice is to have these specs as far to the left as possible. (Isn’t that the case with everything, though? But I digress.)

This restriction may disappear in later versions of Gnus.

3.1.2 Summary Buffer Mode Line

You can also change the format of the summary mode bar (see Section 8.4.2 [Mode Line Formatting], page 155). Set gnus-summary-mode-line-format to whatever you like. The

default is ‘Gnus: %%b [%A] %Z’.

Here are the elements you can play with: ‘G’ Group name. ‘p’ Unprefixed group name. ‘A’ Current article number. ‘z’ Current article score. ‘V’ Gnus version. ‘U’ Number of unread articles in this group. ‘e’ Number of unread articles in this group that aren’t displayed in the summary

buffer.

‘Z’ A string with the number of unread and unselected articles represented either as ‘<%U(+%e) more>’ if there are both unread and unselected articles, and just as ‘<%U more>’ if there are just unread articles and no unselected ones.

‘g’ Shortish group name. For instance, ‘rec.arts.anime’ will be shortened to

          ‘r.a.anime’.

‘S’ Subject of the current article. ‘u’ User-defined spec (see Section 8.4.4 [User-Defined Specs], page 156). ‘s’ Name of the current score file (see Chapter 7 [Scoring], page 131). ‘d’ Number of dormant articles (see Section 3.7.1 [Unread Articles], page 42). ‘t’ Number of ticked articles (see Section 3.7.1 [Unread Articles], page 42). ‘r’ Number of articles that have been marked as read in this session. ‘E’ Number of articles expunged by the score files.

3.1.3 Summary Highlighting

gnus-visual-mark-article-hook

This hook is run after selecting an article. It is meant to be used for highlighting the article in some way. It is not run if gnus-visual is nil.

gnus-summary-update-hook

This hook is called when a summary line is changed. It is not run if gnus-

          visual is nil.
gnus-summary-selected-face

This is the face (or font as some people call it) used to highlight the current article in the summary buffer.

gnus-summary-highlight

Summary lines are highlighted according to this variable, which is a list where the elements are of the format (FORM . FACE). If you would, for instance, like ticked articles to be italic and high-scored articles to be bold, you could set this variable to something like

     (((eq mark gnus-ticked-mark) . italic)
      ((> score default) . bold))

As you may have guessed, if FORM returns a non-nil value, FACE will be applied to the line.

3.2 Summary Maneuvering

All the straight movement commands understand the numeric prefix and behave pretty much as you’d expect.

None of these commands select articles.

G M-n

M-n Go to the next summary line of an unread article (gnus-summary-next-

unread-subject).
G M-p

M-p Go to the previous summary line of an unread article (gnus-summary-prev-

unread-subject).
G j

j Ask for an article number or Message-ID, and then go to that article (gnus-

summary-goto-article).

G g Ask for an article number and then go to the summary line of that article without displaying the article (gnus-summary-goto-subject).

If Gnus asks you to press a key to confirm going to the next group, you can use the C-n and C-p keys to move around the group buffer, searching for the next group to read without actually returning to the group buffer.

Variables related to summary movement:

gnus-auto-select-next

If you issue one of the movement commands (like n) and there are no more unread articles after the current one, Gnus will offer to go to the next group. If this variable is t and the next group is empty, Gnus will exit summary mode and return to the group buffer. If this variable is neither t nor nil, Gnus will select the next group, no matter whether it has any unread articles or not. As a special case, if this variable is quietly, Gnus will select the next group without asking for confirmation. If this variable is almost-quietly, the same will happen only if you are located on the last article in the group. Finally, if this variable is slightly-quietly, the Z n command will go to the next group without confirmation. Also see Section 2.6 [Group Levels], page 16.

gnus-auto-select-same

If non-nil, all the movement commands will try to go to the next article with the same subject as the current. (Same here might mean roughly equal. See gnus-summary-gather-subject-limit for details (see Section 3.9.1 [Customizing Threading], page 47).) If there are no more articles with the same subject, go to the first unread article. This variable is not particularly useful if you use a threaded display.

gnus-summary-check-current

If non-nil, all the “unread” movement commands will not proceed to the next (or previous) article if the current article is unread. Instead, they will choose the current article.

gnus-auto-center-summary

If non-nil, Gnus will keep the point in the summary buffer centered at all times. This makes things quite tidy, but if you have a slow network connection, or simply do not like this un-Emacsism, you can set this variable to nil to get the normal Emacs scrolling action. This will also inhibit horizontal recentering of the summary buffer, which might make it more inconvenient to read extremely long threads.

3.3 Choosing Articles

3.3.1 Choosing Commands

None of the following movement commands understand the numeric prefix, and they all select and display an article.

SPACE Select the current article, or, if that one’s read already, the next unread article

          (gnus-summary-next-page).
G n
n  Go to next unread article (gnus-summary-next-unread-article).
G p
p  Go to previous unread article (gnus-summary-prev-unread-article).
G N
N  Go to the next article (gnus-summary-next-article).
G P
P  Go to the previous article (gnus-summary-prev-article).

G C-n Go to the next article with the same subject (gnus-summary-next-same-

          subject).

G C-p Go to the previous article with the same subject (gnus-summary-prev-same-

          subject).
G f
.  Go to the first unread article (gnus-summary-first-unread-article).

G b

, Go to the article with the highest score (gnus-summary-best-unread-

article).
G l
l  Go to the previous article read (gnus-summary-goto-last-article).

G o Pop an article off the summary history and go to this article (gnus-summary-pop-article). This command differs from the command above in that you can pop as many previous articles off the history as you like, while l toggles the two last read articles. For a somewhat related issue (if you use these commands a lot), see Section 3.14 [Article Backlog], page 55.

3.3.2 Choosing Variables

Some variables relevant for moving and selecting articles:

gnus-auto-extend-newsgroup

All the movement commands will try to go to the previous (or next) article, even if that article isn’t displayed in the Summary buffer if this variable is nonnil. Gnus will then fetch the article from the server and display it in the article buffer.

gnus-select-article-hook

This hook is called whenever an article is selected. By default it exposes any threads hidden under the selected article.

gnus-mark-article-hook

This hook is called whenever an article is selected. It is intended to be used for marking articles as read. The default value is gnus-summary-mark-read-and-unread-as-read, and will change the mark of almost any article you read to gnus-unread-mark. The only articles not affected by this function are ticked, dormant, and expirable articles. If you’d instead like to just have unread articles

marked as read, you can use gnus-summary-mark-unread-as-read instead. It

will leave marks like gnus-low-score-mark, gnus-del-mark (and so on) alone.

3.4 Scrolling the Article

SPACE Pressing SPACE will scroll the current article forward one page, or, if you have come to the end of the current article, will choose the next article (gnus-

summary-next-page).

DEL Scroll the current article back one page (gnus-summary-prev-page).

RET Scroll the current article one line forward (gnus-summary-scroll-up).

M-RET Scroll the current article one line backward (gnus-summary-scroll-down).

A g

g (Re)fetch the current article (gnus-summary-show-article). If given a prefix, fetch the current article, but don’t run any of the article treatment functions. This will give you a “raw” article, just the way it came from the server.

A <

< Scroll to the beginning of the article (gnus-summary-beginning-of-article).

A >

> Scroll to the end of the article (gnus-summary-end-of-article).

A s

s Perform an isearch in the article buffer (gnus-summary-isearch-article).

h  Select the article buffer (gnus-summary-select-article-buffer).

3.5 Reply, Followup and Post

3.5.1 Summary Mail Commands

Commands for composing a mail message:

S r

r Mail a reply to the author of the current article (gnus-summary-reply).

S R

R Mail a reply to the author of the current article and include the original mes-

          sage (gnus-summary-reply-with-original). This command uses the pro-

cess/prefix convention.

S w Mail a wide reply to the author of the current article (gnus-summary-wide-reply). A wide reply is a reply that goes out to all people listed in the To, From (or Reply-to) and Cc headers.

S W Mail a wide reply to the current article and include the original message (gnussummary-reply-with-original). This command uses the process/prefix convention.

S o m Forward the current article to some other person (gnus-summary-mail-forward). If given a prefix, include the full headers of the forwarded article.

S m
m  Send a mail to some other person (gnus-summary-mail-other-window).

S D b If you have sent a mail, but the mail was bounced back to you for some reason (wrong address, transient failure), you can use this command to resend that

          bounced mail (gnus-summary-resend-bounced-mail). You will be popped

into a mail buffer where you can edit the headers before sending the mail off again. If you give a prefix to this command, and the bounced mail is a reply to some other mail, Gnus will try to fetch that mail and display it for easy perusal of its headers. This might very well fail, though.

S D r Not to be confused with the previous command, gnus-summary-resend-message will prompt you for an address to send the current message off to, and then send it to that place. The headers of the message won’t be altered—but lots of headers that say Resent-To, Resent-From and so on will be added. This means that you actually send a mail to someone that has a To header that (probably) points to yourself. This will confuse people. So, natcherly you’ll only do that if you’re really eVIl. This command is mainly used if you have several accounts and want to ship a mail to a different account of yours. (If you’re both root and postmaster and get a mail for postmaster to the root account, you may want to resend it to postmaster. Ordnung mu sein! This command understands the process/prefix convention (see Section 8.1 [Process/Prefix], page 153).

S O m Digest the current series (see Section 3.16 [Decoding Articles], page 59) and forward the result using mail (gnus-uu-digest-mail-forward). This command uses the process/prefix convention (see Section 8.1 [Process/Prefix], page 153).

S M-c Send a complaint about excessive crossposting to the author of the current

article (gnus-summary-mail-crosspost-complaint).

This command is provided as a way to fight back against the current crossposting pandemic that’s sweeping Usenet. It will compose a reply using the gnuscrosspost-complaint variable as a preamble. This command understands the process/prefix convention (see Section 8.1 [Process/Prefix], page 153) and will prompt you before sending each mail.

Also see ⟨undefined⟩ [(message)Header Commands], page ⟨undefined⟩ for more information.

3.5.2 Summary Post Commands

Commands for posting a news article:

S p

a Post an article to the current group (gnus-summary-post-news).

S f

f Post a followup to the current article (gnus-summary-followup).

S F

F Post a followup to the current article and include the original message (gnussummary-followup-with-original). This command uses the process/prefix convention.

S n Post a followup to the current article via news, even if you got the message

through mail (gnus-summary-followup-to-mail).

S N Post a followup to the current article via news, even if you got the message through mail and include the original message (gnus-summary-followup-to-mail-with-original). This command uses the process/prefix convention.

S o p Forward the current article to a newsgroup (gnus-summary-post-forward). If given a prefix, include the full headers of the forwarded article.

S O p Digest the current series and forward the result to a newsgroup (gnus-uu-digest-mail-forward). This command uses the process/prefix convention.

S u Uuencode a file, split it into parts, and post it as a series (gnus-uu-post-news). (see Section 3.16.5.3 [Uuencoding and Posting], page 62).

Also see ⟨undefined⟩ [(message)Header Commands], page ⟨undefined⟩ for more information.

3.6 Canceling Articles

Have you ever written something, and then decided that you really, really, really wish you hadn’t posted that?

Well, you can’t cancel mail, but you can cancel posts.

Find the article you wish to cancel (you can only cancel your own articles, so don’t try any funny stuff). Then press C or S c (gnus-summary-cancel-article). Your article will be canceled—machines all over the world will be deleting your article. This command uses the process/prefix convention (see Section 8.1 [Process/Prefix], page 153).

Be aware, however, that not all sites honor cancels, so your article may live on here and there, while most sites will delete the article in question.

Gnus will use the “current” select method when canceling. If you want to use the standard posting method, use the ‘a’ symbolic prefix (see Section 8.3 [Symbolic Prefixes], page 154).

If you discover that you have made some mistakes and want to do some corrections, you can post a superseding article that will replace your original article.

Go to the original article and press S s (gnus-summary-supersede-article). You will be put in a buffer where you can edit the article all you want before sending it off the usual way.

The same goes for superseding as for canceling, only more so: Some sites do not honor superseding. On those sites, it will appear that you have posted almost the same article twice.

If you have just posted the article, and change your mind right away, there is a trick you can use to cancel/supersede the article without waiting for the article to appear on your site first. You simply return to the post buffer (which is called *sent ...*). There you will find the article you just posted, with all the headers intact. Change the Message-ID header to a Cancel or Supersedes header by substituting one of those words for the word Message-ID. Then just press C-c C-c to send the article as you would do normally. The previous article will be canceled/superseded.

Just remember, kids: There is no ’c’ in ’supersede’.

3.7 Marking Articles

There are several marks you can set on an article.

You have marks that decide the readedness (whoo, neato-keano neologism ohoy!) of the article. Alphabetic marks generally mean read, while non-alphabetic characters generally mean unread.

In addition, you also have marks that do not affect readedness.

3.7.1 Unread Articles

The following marks mark articles as (kinda) unread, in one form or other. ‘!’ Marked as ticked (gnus-ticked-mark).

Ticked articles are articles that will remain visible always. If you see an article that you find interesting, or you want to put off reading it, or replying to it, until sometime later, you’d typically tick it. However, articles can be expired, so if you want to keep an article forever, you’ll have to make it persistent (see Section 3.13 [Persistent Articles], page 55).

‘?’ Marked as dormant (gnus-dormant-mark).

Dormant articles will only appear in the summary buffer if there are followups to it. If you want to see them even if they don’t have followups, you can use the / D command (see Section 3.8 [Limiting], page 45).

‘SPACE’  Marked as unread (gnus-unread-mark).

Unread articles are articles that haven’t been read at all yet.

3.7.2 Read Articles

All the following marks mark articles as read. ‘r’ These are articles that the user has marked as read with the d command man-

ually, more or less (gnus-del-mark).

‘R’ Articles that have actually been read (gnus-read-mark). ‘O’ Articles that were marked as read in previous sessions and are now old (gnus-

          ancient-mark).

‘K’ Marked as killed (gnus-killed-mark). ‘X’ Marked as killed by kill files (gnus-kill-file-mark). ‘Y’ Marked as read by having too low a score (gnus-low-score-mark). ‘C’ Marked as read by a catchup (gnus-catchup-mark).

‘G’  Canceled article (gnus-canceled-mark)

‘F’ SOUPed article (gnus-souped-mark). See Section 6.4.4 [SOUP], page 117. ‘Q’ Sparsely reffed article (gnus-sparse-mark). See Section 3.9.1 [Customizing

Threading], page 47.

‘M’ Article marked as read by duplicate suppression (gnus-duplicated-mark). See

Section 3.27 [Duplicate Suppression], page 80. All these marks just mean that the article is marked as read, really. They are interpreted

differently when doing adaptive scoring, though.

One more special mark, though: ‘E’ Marked as expirable (gnus-expirable-mark).

Marking articles as expirable (or have them marked as such automatically) doesn’t make much sense in normal groups—a user doesn’t control expiring of news articles, but in mail groups, for instance, articles marked as expirable can be deleted by Gnus at any time.

3.7.3 Other Marks

There are some marks that have nothing to do with whether the article is read or not.

• You can set a bookmark in the current article. Say you are reading a long thesis on cats’ urinary tracts, and have to go home for dinner before you’ve finished reading the thesis. You can then set a bookmark in the article, and Gnus will jump to this bookmark the next time it encounters the article. See Section 3.7.4 [Setting Marks], page 43.

• All articles that you have replied to or made a followup to (i.e., have answered) will be marked with an ‘A’ in the second column (gnus-replied-mark).

• Articles stored in the article cache will be marked with an ‘*’ in the second column (gnus-cached-mark). See Section 3.12 [Article Caching], page 54.

• Articles “saved” (in some manner or other; not necessarily religiously) are marked with an ‘S’ in the second column (gnus-saved-mark).

• If the ‘%e’ spec is used, the presence of threads or not will be marked with gnus-not-

    empty-thread-mark and gnus-empty-thread-mark in the third column, respectively.

• Finally we have the process mark (gnus-process-mark). A variety of commands react to the presence of the process mark. For instance, X u (gnus-uu-decode-uu) will uudecode and view all articles that have been marked with the process mark. Articles marked with the process mark have a ‘#’ in the second column.

You might have noticed that most of these “non-readedness” marks appear in the second column by default. So if you have a cached, saved, replied article that you have processmarked, what will that look like?

Nothing much. The precedence rules go as follows: process -> cache -> replied -> saved. So if the article is in the cache and is replied, you’ll only see the cache mark and not the replied mark.

3.7.4 Setting Marks

All the marking commands understand the numeric prefix.

M c

M-u Clear all readedness-marks from the current article (gnus-summary-clear-mark-forward). In other words, mark the article as unread.

M t
!  Tick the current article (gnus-summary-tick-article-forward). See Sec-

tion 3.12 [Article Caching], page 54.

M ?

? Mark the current article as dormant (gnus-summary-mark-as-dormant). See Section 3.12 [Article Caching], page 54.

M d
d  Mark the current article as read (gnus-summary-mark-as-read-forward).

D Mark the current article as read and move point to the previous line (gnus-

          summary-mark-as-read-backward).

M k

k Mark all articles that have the same subject as the current one as read, and

          then select the next unread article (gnus-summary-kill-same-subject-and-
          select).
M K

C-k Mark all articles that have the same subject as the current one as read (gnus-

          summary-kill-same-subject).

M C Mark all unread articles as read (gnus-summary-catchup).

M C-c Mark all articles in the group as read—even the ticked and dormant articles

          (gnus-summary-catchup-all).

M H Catchup the current group to point (gnus-summary-catchup-to-here).

C-w Mark all articles between point and mark as read (gnus-summary-mark-

          region-as-read).

M V k Kill all articles with scores below the default score (or below the numeric prefix)

          (gnus-summary-kill-below).
M e

E Mark the current article as expirable (gnus-summary-mark-as-expirable).

M b Set a bookmark in the current article (gnus-summary-set-bookmark).

M B Remove the bookmark from the current article (gnus-summary-remove-

          bookmark).

M V c Clear all marks from articles with scores over the default score (or over the

          numeric prefix) (gnus-summary-clear-above).

M V u Tick all articles with scores over the default score (or over the numeric prefix)

          (gnus-summary-tick-above).

M V m Prompt for a mark, and mark all articles with scores over the default score (or over the numeric prefix) with this mark (gnus-summary-clear-above).

The gnus-summary-goto-unread variable controls what action should be taken after setting a mark. If non-nil, point will move to the next/previous unread article. If nil, point will just move one line up or down. As a special case, if this variable is never, all the marking commands as well as other commands (like SPACE) will move to the next article, whether it is unread or not. The default is t.

3.7.5 Setting Process Marks

M P p

# Mark the current article with the process mark (gnus-summary-mark-as-

          processable).
M P u

M-# Remove the process mark, if any, from the current article (gnus-summary-

          unmark-as-processable).

M P U Remove the process mark from all articles (gnus-summary-unmark-all-

          processable).

M P i Invert the list of process marked articles (gnus-uu-invert-processable).

M P R Mark articles that have a Subject header that matches a regular expression

          (gnus-uu-mark-by-regexp).

M P r Mark articles in region (gnus-uu-mark-region).

M P t Mark all articles in the current (sub)thread (gnus-uu-mark-thread).

M P T Unmark all articles in the current (sub)thread (gnus-uu-unmark-thread).

M P v Mark all articles that have a score above the prefix argument (gnus-uu-mark-

          over).

M P s Mark all articles in the current series (gnus-uu-mark-series).

M P S Mark all series that have already had some articles marked (gnus-uu-mark-

          sparse).

M P a Mark all articles in series order (gnus-uu-mark-series).

M P b Mark all articles in the buffer in the order they appear (gnus-uu-mark-buffer).

M P k Push the current process mark set onto the stack and unmark all articles (gnus-

          summary-kill-process-mark).

M P y Pop the previous process mark set from the stack and restore it (gnus-summary-

          yank-process-mark).

M P w Push the current process mark set onto the stack (gnus-summary-save-

          process-mark).

3.8 Limiting

It can be convenient to limit the summary buffer to just show some subset of the articles currently in the group. The effect most limit commands have is to remove a few (or many) articles from the summary buffer.

All limiting commands work on subsets of the articles already fetched from the servers. None of these commands query the server for additional articles.

/ /

/ s Limit the summary buffer to articles that match some subject (gnus-summary-

          limit-to-subject).

/ a Limit the summary buffer to articles that match some author (gnus-summary-

          limit-to-author).
/ u

x Limit the summary buffer to articles not marked as read (gnus-summary-limit-to-unread). If given a prefix, limit the buffer to articles strictly unread. This means that ticked and dormant articles will also be excluded.

/ m Ask for a mark and then limit to all articles that have been marked with that

          mark (gnus-summary-limit-to-marks).

/ t Ask for a number and then limit the summary buffer to articles older than (or equal to) that number of days (gnus-summary-limit-to-marks). If given a prefix, limit to articles younger than that number of days.

/ n Limit the summary buffer to the current article (gnus-summary-limit-to-articles). Uses the process/prefix convention (see Section 8.1 [Process/Prefix], page 153).

/ w Pop the previous limit off the stack and restore it (gnus-summary-pop-limit). If given a prefix, pop all limits off the stack.

/ v Limit the summary buffer to articles that have a score at or above some score

          (gnus-summary-limit-to-score).
/ E

M S Include all expunged articles in the limit (gnus-summary-limit-include-

          expunged).

/ D Include all dormant articles in the limit (gnus-summary-limit-include-

          dormant).

/ * Include all cached articles in the limit (gnus-summary-limit-include-

          cached).

/ d Exclude all dormant articles from the limit (gnus-summary-limit-exclude-

          dormant).

/ T Include all the articles in the current thread in the limit.

/ c Exclude all dormant articles that have no children from the limit

          (gnus-summary-limit-exclude-childless-dormant).

/ C Mark all excluded unread articles as read (gnus-summary-limit-mark-excluded-as-read). If given a prefix, also mark excluded ticked and dormant articles as read.

3.9 Threading

Gnus threads articles by default. To thread is to put responses to articles directly after the articles they respond to—in a hierarchical fashion.

Threading is done by looking at the References headers of the articles. In a perfect world, this would be enough to build pretty trees, but unfortunately, the References header is often broken or simply missing. Weird news propagation excarcerbates the problem, so one has to employ other heuristics to get pleasing results. A plethora of approaches exists, as detailed in horrible detail in see Section 3.9.1 [Customizing Threading], page 47.

First, a quick overview of the concepts:

root The top-most article in a thread; the first article in the thread.

thread A tree-like article structure.

sub-thread A small(er) section of this tree-like structure. loose threads

Threads often lose their roots due to article expiry, or due to the root already having been read in a previous session, and not displayed in the summary buffer. We then typically have many sub-threads that really belong to one thread, but are without connecting roots. These are called loose threads.

thread gathering An attempt to gather loose threads into bigger threads.

sparse threads A thread where the missing articles have been “guessed” at, and are displayed as empty lines in the summary buffer.

3.9.1 Customizing Threading

3.9.1.1 Loose Threads

gnus-summary-make-false-root

If non-nil, Gnus will gather all loose subtrees into one big tree and create a dummy root at the top. (Wait a minute. Root at the top? Yup.) Loose subtrees occur when the real root has expired, or you’ve read or killed the root in a previous session. When there is no real root of a thread, Gnus will have to fudge something. This variable says what fudging method Gnus should use. There are four possible values:

adopt Gnus will make the first of the orphaned articles the parent. This parent will adopt all the other articles. The adopted articles will be marked as such by pointy brackets (‘<>’) instead of the standard square brackets (‘[]’). This is the default method.

dummy Gnus will create a dummy summary line that will pretend to be the parent. This dummy line does not correspond to any real article, so selecting it will just select the first real article after the dummy

          article. gnus-summary-dummy-line-format is used to specify the

format of the dummy roots. It accepts only one format spec: ‘S’, which is the subject of the article. See Section 8.4 [Formatting Variables], page 154.

empty Gnus won’t actually make any article the parent, but simply leave the subject field of all orphans except the first empty. (Actually, it will use gnus-summary-same-subject as the subject (see Section 3.1 [Summary Buffer Format], page 33).)

none Don’t make any article parent at all. Just gather the threads and display them after one another.

nil Don’t gather loose threads.

gnus-summary-gather-subject-limit

Loose threads are gathered by comparing subjects of articles. If this variable is nil, Gnus requires an exact match between the subjects of the loose threads before gathering them into one big super-thread. This might be too strict a requirement, what with the presence of stupid newsreaders that chop off long subject lines. If you think so, set this variable to, say, 20 to require that only the first 20 characters of the subjects have to match. If you set this variable to a really low number, you’ll find that Gnus will gather everything in sight into one thread, which isn’t very helpful. If you set this variable to the special value fuzzy, Gnus will use a fuzzy string comparison algorithm on the subjects (see Section 8.16 [Fuzzy Matching], page 171).

gnus-simplify-subject-fuzzy-regexp

This can either be a regular expression or list of regular expressions that match strings that will be removed from subjects if fuzzy subject simplification is used.

gnus-simplify-ignored-prefixes
If you set gnus-summary-gather-subject-limit to something as low as 10,

you might consider setting this variable to something sensible:

     (setq gnus-simplify-ignored-prefixes
           (concat
            "\\`\\[?\\("
            (mapconcat
             'identity
             '("looking"
               "wanted" "followup" "summary\\( of\\)?"
               "help" "query" "problem" "question"
               "answer" "reference" "announce"
               "How can I" "How to" "Comparison of"
               ;; ...
               )
             "\\|")
            "\\)\\s *\\("
            (mapconcat 'identity
                       '("for" "for reference" "with" "about")
                       "\\|")
            "\\)?\\]?:?[ \t]*"))

All words that match this regexp will be removed before comparing two subjects.

gnus-simplify-subject-functions
If non-nil, this variable overrides gnus-summary-gather-subject-limit.

This variable should be a list of functions to apply to the Subject string iteratively to arrive at the simplified version of the string. Useful functions to put in this list include:

gnus-simplify-subject-re

Strip the leading ‘Re:’.

gnus-simplify-subject-fuzzy

Simplify fuzzily.

gnus-simplify-whitespace

Remove excessive whitespace.

You may also write your own functions, of course.

gnus-summary-gather-exclude-subject

Since loose thread gathering is done on subjects only, that might lead to many false hits, especially with certain common subjects like ‘’ and ‘(none)’. To make the situation slightly better, you can use the regexp gnus-summary-gather-exclude-subject to say what subjects should be excluded from the gathering process. The default is ‘^*$\\|^(none)$’.

gnus-summary-thread-gathering-function

Gnus gathers threads by looking at Subject headers. This means that totally unrelated articles may end up in the same “thread”, which is confusing. An alternate approach is to look at all the Message-IDs in all the References headers to find matches. This will ensure that no gathered threads ever include unrelated articles, but it also means that people who have posted with broken newsreaders won’t be gathered properly. The choice is yours—plague or cholera:

gnus-gather-threads-by-subject

This function is the default gathering function and looks at Subjects exclusively.

gnus-gather-threads-by-references

This function looks at References headers exclusively.

If you want to test gathering by References, you could say something like:

     (setq gnus-summary-thread-gathering-function
           'gnus-gather-threads-by-references)

3.9.1.2 Filling In Threads

gnus-fetch-old-headers

If non-nil, Gnus will attempt to build old threads by fetching more old headers—headers to articles marked as read. If you would like to display as few summary lines as possible, but still connect as many loose threads as possible, you should set this variable to some or a number. If you set it to a number, no more than that number of extra old headers will be fetched. In either case, fetching old headers only works if the backend you are using carries overview files—this would normally be nntp, nnspool and nnml. Also remember that if the root of the thread has been expired by the server, there’s not much Gnus can do about that. This variable can also be set to invisible. This won’t have any visible effects, but is useful if you use the A T command a lot (see Section 3.20 [Finding the Parent], page 71).

gnus-build-sparse-threads

Fetching old headers can be slow. A low-rent similar effect can be gotten by setting this variable to some. Gnus will then look at the complete References headers of all articles and try to string together articles that belong in the same thread. This will leave gaps in the threading display where Gnus guesses that an article is missing from the thread. (These gaps appear like normal summary lines. If you select a gap, Gnus will try to fetch the article in question.) If this variable is t, Gnus will display all these “gaps” without regard for whether they are useful for completing the thread or not. Finally, if this variable is more, Gnus won’t cut off sparse leaf nodes that don’t lead anywhere. This variable is nil by default.

3.9.1.3 More Threading

gnus-show-threads

If this variable is nil, no threading will be done, and all of the rest of the variables here will have no effect. Turning threading off will speed group selection up a bit, but it is sure to make reading slower and more awkward.

gnus-thread-hide-subtree

If non-nil, all threads will be hidden when the summary buffer is generated.

gnus-thread-expunge-below

All threads that have a total score (as defined by gnus-thread-score-function) less than this number will be expunged. This variable is nil by default, which means that no threads are expunged.

gnus-thread-hide-killed

if you kill a thread and this variable is non-nil, the subtree will be hidden.

gnus-thread-ignore-subject

Sometimes somebody changes the subject in the middle of a thread. If this variable is non-nil, the subject change is ignored. If it is nil, which is the default, a change in the subject will result in a new thread.

gnus-thread-indent-level

This is a number that says how much each sub-thread should be indented. The default is 4.

3.9.1.4 Low-Level Threading

gnus-parse-headers-hook

Hook run before parsing any headers. The default value is (gnus-decode-rfc1522), which means that QPized headers will be slightly decoded in a hackish way. This is likely to change in the future when Gnus becomes MIMEified.

gnus-alter-header-function

If non-nil, this function will be called to allow alteration of article header structures. The function is called with one parameter, the article header vector, which it may alter in any way. For instance, if you have a mail-to-news gateway which alters the Message-IDs in systematic ways (by adding prefixes and such), you can use this variable to un-scramble the Message-IDs so that they are more meaningful. Here’s one example:

               (setq gnus-alter-header-function 'my-alter-message-id)
               (defun my-alter-message-id (header)
                 (let ((id (mail-header-id header)))
                   (when (string-match
                          "\\(<[^<>@]*\\)\\.?cygnus\\..*@\\([^<>@]*>\\)" id)
                     (mail-header-set-id
                      (concat (match-string 1 id) "@" (match-string 2 id))
                      header))))

3.9.2 Thread Commands

T k

M-C-k Mark all articles in the current (sub-)thread as read (gnus-summary-kill-thread). If the prefix argument is positive, remove all marks instead. If the prefix argument is negative, tick articles instead.

T l

M-C-l Lower the score of the current (sub-)thread (gnus-summary-lower-thread). T i Increase the score of the current (sub-)thread (gnus-summary-raise-thread). T # Set the process mark on the current (sub-)thread (gnus-uu-mark-thread). T M-# Remove the process mark from the current (sub-)thread (gnus-uu-unmark-

          thread).
T T  Toggle threading (gnus-summary-toggle-threads).

T s Expose the (sub-)thread hidden under the current article, if any

          (gnus-summary-show-thread).

T h Hide the current (sub-)thread (gnus-summary-hide-thread).

T S  Expose all hidden threads (gnus-summary-show-all-threads).
T H  Hide all threads (gnus-summary-hide-all-threads).

T t Re-thread the current article’s thread (gnus-summary-rethread-current).

This works even when the summary buffer is otherwise unthreaded.

T ^Make the current article the child of the marked (or previous) article (gnus-

          summary-reparent-thread).

The following commands are thread movement commands. They all understand the numeric prefix.

T n  Go to the next thread (gnus-summary-next-thread).
T p  Go to the previous thread (gnus-summary-prev-thread).
T d  Descend the thread (gnus-summary-down-thread).
T u  Ascend the thread (gnus-summary-up-thread).
T o  Go to the top of the thread (gnus-summary-top-thread).

If you ignore subject while threading, you’ll naturally end up with threads that have several different subjects in them. If you then issue a command like ‘T k’ (gnus-summary-kill-thread) you might not wish to kill the entire thread, but just those parts of the thread that have the same subject as the current article. If you like this idea, you can fiddle

with gnus-thread-operation-ignore-subject. If it is non-nil (which it is by default),

subjects will be ignored when doing thread commands. If this variable is nil, articles in the same thread with different subjects will not be included in the operation in question. If this variable is fuzzy, only articles that have subjects fuzzily equal will be included (see Section 8.16 [Fuzzy Matching], page 171).

3.10 Sorting

If you are using a threaded summary display, you can sort the threads by setting gnusthread-sort-functions, which is a list of functions. By default, sorting is done on article numbers. Ready-made sorting predicate functions include gnus-thread-sort-by-number,

gnus-thread-sort-by-author, gnus-thread-sort-by-subject, gnus-thread-sort-by-
date, gnus-thread-sort-by-score, and gnus-thread-sort-by-total-score.

Each function takes two threads and returns non-nil if the first thread should be sorted before the other. Note that sorting really is normally done by looking only at the roots of each thread. If you use more than one function, the primary sort key should be the last function in the list. You should probably always include gnus-thread-sort-by-number in the list of sorting functions—preferably first. This will ensure that threads that are equal with respect to the other sort criteria will be displayed in ascending article order.

If you would like to sort by score, then by subject, and finally by number, you could do something like:

     (setq gnus-thread-sort-functions
           '(gnus-thread-sort-by-number
             gnus-thread-sort-by-subject
             gnus-thread-sort-by-total-score))

The threads that have highest score will be displayed first in the summary buffer. When threads have the same score, they will be sorted alphabetically. The threads that have the same score and the same subject will be sorted by number, which is (normally) the sequence in which the articles arrived.

If you want to sort by score and then reverse arrival order, you could say something like:

     (setq gnus-thread-sort-functions
           '((lambda (t1 t2)
               (not (gnus-thread-sort-by-number t1 t2)))
             gnus-thread-sort-by-score))

The function in the gnus-thread-score-function variable (default +) is used for calculating the total score of a thread. Useful functions might be max, min, or squared means, or whatever tickles your fancy.

If you are using an unthreaded display for some strange reason or other, you have to fiddle with the gnus-article-sort-functions variable. It is very similar to the gnus-thread-sort-functions, except that it uses slightly different functions for article comparison. Available sorting predicate functions are gnus-article-

sort-by-number, gnus-article-sort-by-author, gnus-article-sort-by-subject,
gnus-article-sort-by-date, and gnus-article-sort-by-score.

If you want to sort an unthreaded summary display by subject, you could say something like:

     (setq gnus-article-sort-functions
           '(gnus-article-sort-by-number
             gnus-article-sort-by-subject))

3.11 Asynchronous Article Fetching

If you read your news from an nntp server that’s far away, the network latencies may make reading articles a chore. You have to wait for a while after pressing n to go to the next article before the article appears. Why can’t Gnus just go ahead and fetch the article while you are reading the previous one? Why not, indeed.

First, some caveats. There are some pitfalls to using asynchronous article fetching, especially the way Gnus does it.

Let’s say you are reading article 1, which is short, and article 2 is quite long, and you are not interested in reading that. Gnus does not know this, so it goes ahead and fetches article 2. You decide to read article 3, but since Gnus is in the process of fetching article 2, the connection is blocked.

To avoid these situations, Gnus will open two (count ’em two) connections to the server. Some people may think this isn’t a very nice thing to do, but I don’t see any real alternatives. Setting up that extra connection takes some time, so Gnus startup will be slower.

Gnus will fetch more articles than you will read. This will mean that the link between your machine and the nntp server will become more loaded than if you didn’t use article pre-fetch. The server itself will also become more loaded—both with the extra article requests, and the extra connection.

Ok, so now you know that you shouldn’t really use this thing... unless you really want to.

Here’s how: Set gnus-asynchronous to t. The rest should happen automatically.

You can control how many articles are to be pre-fetched by setting gnus-use-article-prefetch. This is 30 by default, which means that when you read an article in the group, the backend will pre-fetch the next 30 articles. If this variable is t, the backend will pre-fetch all the articles it can without bound. If it is nil, no pre-fetching will be done.

There are probably some articles that you don’t want to pre-fetch—read articles, for instance. The gnus-async-prefetch-article-p variable controls whether an article is to be pre-fetched. This function should return non-nil when the article in question is to be pre-fetched. The default is gnus-async-read-p, which returns nil on read articles. The function is called with an article data structure as the only parameter.

If, for instance, you wish to pre-fetch only unread articles shorter than 100 lines, you could say something like:

     (defun my-async-short-unread-p (data)
       "Return non-nil for short, unread articles."

       (and (gnus-data-unread-p data)
            (< (mail-header-lines (gnus-data-header data))
               100)))
     (setq gnus-async-prefetch-article-p 'my-async-short-unread-p)

These functions will be called many, many times, so they should preferably be short and sweet to avoid slowing down Gnus too much. It’s probably a good idea to byte-compile things like this.

Articles have to be removed from the asynch buffer sooner or later. The gnus-

prefetched-article-deletion-strategy says when to remove articles. This is a list

that may contain the following elements:

read Remove articles when they are read.

exit Remove articles when exiting the group.

The default value is (read exit).

3.12 Article Caching

If you have an extremely slow nntp connection, you may consider turning article caching on. Each article will then be stored locally under your home directory. As you may surmise, this could potentially use huge amounts of disk space, as well as eat up all your inodes so fast it will make your head swim. In vodka.

Used carefully, though, it could be just an easier way to save articles.

To turn caching on, set gnus-use-cache to t. By default, all articles ticked or marked as dormant will then be copied over to your local cache (gnus-cache-directory). Whether this cache is flat or hierarchal is controlled by the gnus-use-long-file-name variable, as usual.

When re-selecting a ticked or dormant article, it will be fetched from the cache instead of from the server. As articles in your cache will never expire, this might serve as a method of saving articles while still keeping them where they belong. Just mark all articles you want to save as dormant, and don’t worry.

When an article is marked as read, is it removed from the cache.

The entering/removal of articles from the cache is controlled by the gnus-cache-enter-articles and gnus-cache-remove-articles variables. Both are lists of symbols. The first is (ticked dormant) by default, meaning that ticked and dormant articles will be put in the cache. The latter is (read) by default, meaning that articles marked as read are removed from the cache. Possibly symbols in these two lists are ticked, dormant, unread and read.

So where does the massive article-fetching and storing come into the picture? The gnusjog-cache command will go through all subscribed newsgroups, request all unread articles, score them, and store them in the cache. You should only ever, ever ever ever, use this command if 1) your connection to the nntp server is really, really, really slow and 2) you have a really, really, really huge disk. Seriously. One way to cut down on the number of articles downloaded is to score unwanted articles down and have them marked as read. They will not then be downloaded by this command.

It is likely that you do not want caching on all groups. For instance, if your nnml mail is located under your home directory, it makes no sense to cache it somewhere else under your home directory. Unless you feel that it’s neat to use twice as much space.

To limit the caching, you could set gnus-cacheable-groups to a regexp of groups to cache, ‘^nntp’ for instance, or set the gnus-uncacheable-groups regexp to ‘^nnml’, for instance. Both variables are nil by default. If a group matches both variables, the group is not cached.

The cache stores information on what articles it contains in its active file (gnus-cache-active-file). If this file (or any other parts of the cache) becomes all messed up for some reason or other, Gnus offers two functions that will try to set things right. M-x

gnus-cache-generate-nov-databases will (re)build all the nov files, and gnus-cache-

generate-active will (re)generate the active file.

3.13 Persistent Articles

Closely related to article caching, we have persistent articles. In fact, it’s just a different way of looking at caching, and much more useful in my opinion.

Say you’re reading a newsgroup, and you happen on to some valuable gem that you want to keep and treasure forever. You’d normally just save it (using one of the many saving commands) in some file. The problem with that is that it’s just, well, yucky. Ideally you’d prefer just having the article remain in the group where you found it forever; untouched by the expiry going on at the news server.

This is what a persistent article is—an article that just won’t be deleted. It’s implemented using the normal cache functions, but you use two explicit commands for managing persistent articles:

* Make the current article persistent (gnus-cache-enter-article).

M-* Remove the current article from the persistent articles (gnus-cache-remove-article). This will normally delete the article.

Both these commands understand the process/prefix convention.

To avoid having all ticked articles (and stuff) entered into the cache, you should set gnus-use-cache to passive if you’re just interested in persistent articles:

     (setq gnus-use-cache 'passive)

3.14 Article Backlog

If you have a slow connection, but the idea of using caching seems unappealing to you (and it is, really), you can help the situation some by switching on the backlog. This is where Gnus will buffer already read articles so that it doesn’t have to re-fetch articles you’ve already read. This only helps if you are in the habit of re-selecting articles you’ve recently read, of course. If you never do that, turning the backlog on will slow Gnus down a little bit, and increase memory usage some.

If you set gnus-keep-backlog to a number n, Gnus will store at most n old articles in a buffer for later re-fetching. If this variable is non-nil and is not a number, Gnus will store all read articles, which means that your Emacs will grow without bound before exploding and taking your machine down with you. I put that in there just to keep y’all on your toes.

This variable is nil by default.

3.15 Saving Articles

Gnus can save articles in a number of ways. Below is the documentation for saving articles in a fairly straight-forward fashion (i.e., little processing of the article is done before it is saved). For a different approach (uudecoding, unsharing) you should use gnus-uu (see Section 3.16 [Decoding Articles], page 59).

If gnus-save-all-headers is non-nil, Gnus will not delete unwanted headers before saving the article.

If the preceding variable is nil, all headers that match the gnus-saved-headers regexp will be kept, while the rest will be deleted before saving.

O o

o Save the current article using the default article saver (gnus-summary-save-

          article).

O m Save the current article in mail format (gnus-summary-save-article-mail).

O r Save the current article in rmail format (gnus-summary-save-article-rmail).

O f Save the current article in plain file format (gnus-summary-save-article-

          file).

O F Write the current article in plain file format, overwriting any previous file con-

          tents (gnus-summary-write-article-file).

O b Save the current article body in plain file format (gnus-summary-save-

          article-body-file).

O h Save the current article in mh folder format (gnus-summary-save-article-

          folder).

O v Save the current article in a VM folder (gnus-summary-save-article-vm).

O p Save the current article in a pipe. Uhm, like, what I mean is—Pipe the current

          article to a process (gnus-summary-pipe-output).

All these commands use the process/prefix convention (see Section 8.1 [Process/Prefix], page 153). If you save bunches of articles using these functions, you might get tired of being prompted for files to save each and every article in. The prompting action is controlled by the gnus-prompt-before-saving variable, which is always by default, giving you that excessive prompting action you know and loathe. If you set this variable to t instead, you’ll be prompted just once for each series of articles you save. If you like to really have Gnus do all your thinking for you, you can even set this variable to nil, which means that you will never be prompted for files to save articles in. Gnus will simply save all the articles in the default files.

You can customize the gnus-default-article-saver variable to make Gnus do what you want it to. You can use any of the four ready-made functions below, or you can create your own.

gnus-summary-save-in-rmail

This is the default format, babyl. Uses the function in the gnus-rmail-save-name variable to get a file name to save the article in. The default is gnus-

plain-save-name.
gnus-summary-save-in-mail

Save in a Unix mail (mbox) file. Uses the function in the gnus-mail-save-name variable to get a file name to save the article in. The default is gnus-plain-

save-name.
gnus-summary-save-in-file

Append the article straight to an ordinary file. Uses the function in the gnusfile-save-name variable to get a file name to save the article in. The default

is gnus-numeric-save-name.
gnus-summary-save-body-in-file

Append the article body to an ordinary file. Uses the function in the gnusfile-save-name variable to get a file name to save the article in. The default

is gnus-numeric-save-name.
gnus-summary-save-in-folder

Save the article to an MH folder using rcvstore from the MH library. Uses the function in the gnus-folder-save-name variable to get a file name to save the article in. The default is gnus-folder-save-name, but you can also use gnus-Folder-save-name, which creates capitalized names.

gnus-summary-save-in-vm

Save the article in a VM folder. You have to have the VM mail reader to use this setting.

All of these functions, except for the last one, will save the article in the gnus-article-save-directory, which is initialized from the SAVEDIR environment variable. This is ‘~/News/’ by default.

As you can see above, the functions use different functions to find a suitable name of a file to save the article in. Below is a list of available functions that generate names:

gnus-Numeric-save-name
File names like ‘~/News/Alt.andrea-dworkin/45’.
gnus-numeric-save-name
File names like ‘~/News/alt.andrea-dworkin/45’.
gnus-Plain-save-name
File names like ‘~/News/Alt.andrea-dworkin’.
gnus-plain-save-name
File names like ‘~/News/alt.andrea-dworkin’.

You can have Gnus suggest where to save articles by plonking a regexp into the gnussplit-methods alist. For instance, if you would like to save articles related to Gnus in the file ‘gnus-stuff’, and articles related to VM in vm-stuff, you could set this variable to something like:

     (("^Subject:.*gnus\\|^Newsgroups:.*gnus" "gnus-stuff")
      ("^Subject:.*vm\\|^Xref:.*vm" "vm-stuff")
      (my-choosing-function "../other-dir/my-stuff")
      ((equal gnus-newsgroup-name "mail.misc") "mail-stuff"))

We see that this is a list where each element is a list that has two elements—the match and the file. The match can either be a string (in which case it is used as a regexp to match on the article head); it can be a symbol (which will be called as a function with the group name as a parameter); or it can be a list (which will be evaled). If any of these actions have a non-nil result, the file will be used as a default prompt. In addition, the result of the operation itself will be used if the function or form called returns a string or a list of strings.

You basically end up with a list of file names that might be used when saving the current article. (All “matches” will be used.) You will then be prompted for what you really want to use as a name, with file name completion over the results from applying this variable.

This variable is ((gnus-article-archive-name)) by default, which means that Gnus will look at the articles it saves for an Archive-name line and use that as a suggestion for the file name.

Here’s an example function to clean up file names somewhat. If you have lots of mail groups called things like ‘nnml:mail.whatever’, you may want to chop off the beginning of these group names before creating the file name to save to. The following will do just that:

     (defun my-save-name (group)
       (when (string-match "^nnml:mail." group)
         (substring group (match-end 0))))
     (setq gnus-split-methods
           '((gnus-article-archive-name)
             (my-save-name)))

Finally, you have the gnus-use-long-file-name variable. If it is nil, all the preceding functions will replace all periods (‘.’) in the group names with slashes (‘/’)—which means that the functions will generate hierarchies of directories instead of having all the files in the top level directory (‘~/News/alt/andrea-dworkin’ instead of ‘~/News/alt.andrea-dworkin’.) This variable is t by default on most systems. However, for historical reasons, this is nil on Xenix and usg-unix-v machines by default.

This function also affects kill and score file names. If this variable is a list, and the list contains the element not-score, long file names will not be used for score files, if it contains the element not-save, long file names will not be used for saving, and if it contains the element not-kill, long file names will not be used for kill files.

If you’d like to save articles in a hierarchy that looks something like a spool, you could

     (setq gnus-use-long-file-name '(not-save)) ; to get a hierarchy
     (setq gnus-default-article-saver 'gnus-summary-save-in-file) ; no encoding

Then just save with o. You’d then read this hierarchy with ephemeral nneething groups—G D in the group buffer, and the top level directory as the argument (‘~/News/’). Then just walk around to the groups/directories with nneething.

3.16 Decoding Articles

Sometime users post articles (or series of articles) that have been encoded in some way or other. Gnus can decode them for you.

All these functions use the process/prefix convention (see Section 8.1 [Process/Prefix], page 153) for finding out what articles to work on, with the extension that a “single article” means “a single series”. Gnus can find out by itself what articles belong to a series, decode all the articles and unpack/view/save the resulting file(s).

Gnus guesses what articles are in the series according to the following simplish rule: The subjects must be (nearly) identical, except for the last two numbers of the line. (Spaces are largely ignored, however.)

For example: If you choose a subject called ‘cat.gif (2/3)’, Gnus will find all the articles that match the regexp ‘^cat.gif ([0-9]+/[0-9]+).*$’.

Subjects that are non-standard, like ‘cat.gif (2/3) Part 6 of a series’, will not be properly recognized by any of the automatic viewing commands, and you have to mark the articles manually with #.

3.16.1 Uuencoded Articles

X u Uudecodes the current series (gnus-uu-decode-uu).

X U Uudecodes and saves the current series (gnus-uu-decode-uu-and-save).

X v u Uudecodes and views the current series (gnus-uu-decode-uu-view).

X v U Uudecodes, views and saves the current series (gnus-uu-decode-uu-and-save-

          view).

Remember that these all react to the presence of articles marked with the process mark. If, for instance, you’d like to decode and save an entire newsgroup, you’d typically do M P

a (gnus-uu-mark-all) and then X U (gnus-uu-decode-uu-and-save).

All this is very much different from how gnus-uu worked with gnus 4.1, where you had explicit keystrokes for everything under the sun. This version of gnus-uu generally assumes that you mark articles in some way (see Section 3.7.5 [Setting Process Marks], page 44) and then press X u.

Note: When trying to decode articles that have names matching gnus-uu-notify-

files, which is hard-coded to ‘[Cc][Ii][Nn][Dd][Yy][0-9]+.\\(gif\\|jpg\\)’, gnus-

uu will automatically post an article on ‘comp.unix.wizards’ saying that you have just viewed the file in question. This feature can’t be turned off.

3.16.2 Shell Archives

Shell archives (“shar files”) used to be a popular way to distribute sources, but it isn’t used all that much today. In any case, we have some commands to deal with these:

X s Unshars the current series (gnus-uu-decode-unshar).

X S Unshars and saves the current series (gnus-uu-decode-unshar-and-save).

X v s Unshars and views the current series (gnus-uu-decode-unshar-view).

X v S Unshars, views and saves the current series (gnus-uu-decode-unshar-and-

          save-view).

3.16.3 PostScript Files

X p Unpack the current PostScript series (gnus-uu-decode-postscript).

X P Unpack and save the current PostScript series (gnus-uu-decode-postscript-

          and-save).
X v p  View the current PostScript series (gnus-uu-decode-postscript-view).

X v P View and save the current PostScript series (gnus-uu-decode-postscript-

          and-save-view).

3.16.4 Other Files

X o Save the current series (gnus-uu-decode-save).

X b Unbinhex the current series (gnus-uu-decode-binhex). This doesn’t really work yet.

3.16.5 Decoding Variables

Adjective, not verb.

3.16.5.1 Rule Variables

Gnus uses rule variables to decide how to view a file. All these variables are of the form

           (list '(regexp1 command2)
                 '(regexp2 command2)
                 ...)
gnus-uu-user-view-rules

This variable is consulted first when viewing files. If you wish to use, for instance, sox to convert an ‘.au’ sound file, you could say something like:

               (setq gnus-uu-user-view-rules
                     (list '(\"\\\\.au$\" \"sox %s -t .aiff > /dev/audio\")))
gnus-uu-user-view-rules-end

This variable is consulted if Gnus couldn’t make any matches from the user and default view rules.

gnus-uu-user-archive-rules

This variable can be used to say what commands should be used to unpack archives.

3.16.5.2 Other Decode Variables

gnus-uu-grabbed-file-functions

All functions in this list will be called right after each file has been successfully decoded—so that you can move or view files right away, and don’t have to wait for all files to be decoded before you can do anything. Ready-made functions you can put in this list are:

gnus-uu-grab-view

View the file.

gnus-uu-grab-move

Move the file (if you’re using a saving function.)

gnus-uu-be-dangerous

Specifies what to do if unusual situations arise during decoding. If nil, be as conservative as possible. If t, ignore things that didn’t work, and overwrite existing files. Otherwise, ask each time.

gnus-uu-ignore-files-by-name

Files with name matching this regular expression won’t be viewed.

gnus-uu-ignore-files-by-type

Files with a mime type matching this variable won’t be viewed. Note that Gnus tries to guess what type the file is based on the name. gnus-uu is not a mime package (yet), so this is slightly kludgey.

gnus-uu-tmp-dir

Where gnus-uu does its work.

gnus-uu-do-not-unpack-archives

Non-nil means that gnus-uu won’t peek inside archives looking for files to display.

gnus-uu-view-and-save

Non-nil means that the user will always be asked to save a file after viewing it.

gnus-uu-ignore-default-view-rules

Non-nil means that gnus-uu will ignore the default viewing rules.

gnus-uu-ignore-default-archive-rules

Non-nil means that gnus-uu will ignore the default archive unpacking commands.

gnus-uu-kill-carriage-return

Non-nil means that gnus-uu will strip all carriage returns from articles.

gnus-uu-unmark-articles-not-decoded

Non-nil means that gnus-uu will mark unsuccessfully decoded articles as unread.

gnus-uu-correct-stripped-uucode

Non-nil means that gnus-uu will try to fix uuencoded files that have had trailing spaces deleted.

gnus-uu-pre-uudecode-hook

Hook run before sending a message to uudecode.

gnus-uu-view-with-metamail

Non-nil means that gnus-uu will ignore the viewing commands defined by the rule variables and just fudge a mime content type based on the file name. The result will be fed to metamail for viewing.

gnus-uu-save-in-digest

Non-nil means that gnus-uu, when asked to save without decoding, will save in digests. If this variable is nil, gnus-uu will just save everything in a file without any embellishments. The digesting almost conforms to RFC1153—no easy way to specify any meaningful volume and issue numbers were found, so I simply dropped them.

3.16.5.3 Uuencoding and Posting

gnus-uu-post-include-before-composing

Non-nil means that gnus-uu will ask for a file to encode before you compose the article. If this variable is t, you can either include an encoded file with C-c C-i or have one included for you when you post the article.

gnus-uu-post-length

Maximum length of an article. The encoded file will be split into how many articles it takes to post the entire file.

gnus-uu-post-threaded

Non-nil means that gnus-uu will post the encoded file in a thread. This may not be smart, as no other decoder I have seen is able to follow threads when collecting uuencoded articles. (Well, I have seen one package that does that— gnus-uu, but somehow, I don’t think that counts...) Default is nil.

gnus-uu-post-separate-description

Non-nil means that the description will be posted in a separate article. The first article will typically be numbered (0/x). If this variable is nil, the description the user enters will be included at the beginning of the first article, which will be numbered (1/x). Default is t.

3.16.6 Viewing Files

After decoding, if the file is some sort of archive, Gnus will attempt to unpack the archive and see if any of the files in the archive can be viewed. For instance, if you have a gzipped tar file ‘pics.tar.gz’ containing the files ‘pic1.jpg’ and ‘pic2.gif’, Gnus will uncompress and de-tar the main file, and then view the two pictures. This unpacking process is recursive, so if the archive contains archives of archives, it’ll all be unpacked.

Finally, Gnus will normally insert a pseudo-article for each extracted file into the summary buffer. If you go to these “articles”, you will be prompted for a command to run (usually Gnus will make a suggestion), and then the command will be run.

If gnus-view-pseudo-asynchronously is nil, Emacs will wait until the viewing is done before proceeding.

If gnus-view-pseudos is automatic, Gnus will not insert the pseudo-articles into the summary buffer, but view them immediately. If this variable is not-confirm, the user won’t even be asked for a confirmation before viewing is done.

If gnus-view-pseudos-separately is non-nil, one pseudo-article will be created for each file to be viewed. If nil, all files that use the same viewing command will be given as a list of parameters to that command.

If gnus-insert-pseudo-articles is non-nil, insert pseudo-articles when decoding. It is t by default.

So; there you are, reading your pseudo-articles in your virtual newsgroup from the virtual server; and you think: Why isn’t anything real anymore? How did we get here?

3.17 Article Treatment

Reading through this huge manual, you may have quite forgotten that the object of newsreaders is to actually, like, read what people have written. Reading articles. Unfortunately, people are quite bad at writing, so there are tons of functions and variables to make reading these articles easier.

3.17.1 Article Highlighting

Not only do you want your article buffer to look like fruit salad, but you want it to look like technicolor fruit salad.

W H a Do much highlighting of the current article (gnus-article-highlight). This function highlights header, cited text, the signature, and adds buttons to the body and the head.

Most users would prefer using gnus-article-maybe-highlight in

gnus-article-display-hook (see Section 4.3 [Customizing Articles], page 83) instead. This is a bit less agressive—it highlights only the headers, the signature and adds buttons.

W H h Highlight the headers (gnus-article-highlight-headers). The highlighting will be done according to the gnus-header-face-alist variable, which is a list where each element has the form (regexp name content). regexp is a regular expression for matching the header, name is the face used for highlighting the header name (see Section 8.6 [Faces and Fonts], page 160) and content is the face for highlighting the header value. The first match made will be used. Note that regexp shouldn’t have ‘^’ prepended—Gnus will add one.

W H c  Highlight cited text (gnus-article-highlight-citation).

Some variables to customize the citation highlights:

gnus-cite-parse-max-size

If the article size if bigger than this variable (which is 25000 by default), no citation highlighting will be performed.

gnus-cite-prefix-regexp

Regexp matching the longest possible citation prefix on a line.

gnus-cite-max-prefix

Maximum possible length for a citation prefix (default 20).

gnus-cite-face-list

List of faces used for highlighting citations (see Section 8.6 [Faces and Fonts], page 160). When there are citations from multiple articles in the same message, Gnus will try to give each citation from each article its own face. This should make it easier to see who wrote what.

gnus-supercite-regexp

Regexp matching normal Supercite attribution lines.

gnus-supercite-secondary-regexp

Regexp matching mangled Supercite attribution lines.

gnus-cite-minimum-match-count

Minimum number of identical prefixes we have to see before we believe that it’s a citation.

gnus-cite-attribution-prefix

Regexp matching the beginning of an attribution line.

gnus-cite-attribution-suffix

Regexp matching the end of an attribution line.

gnus-cite-attribution-face

Face used for attribution lines. It is merged with the face for the cited text belonging to the attribution.

W H s  Highlight the signature (gnus-article-highlight-signature). Everything

after gnus-signature-separator (see Section 3.17.7 [Article Signature], page 69) in an article will be considered a signature and will be highlighted

with gnus-signature-face, which is italic by default.

See Section 4.3 [Customizing Articles], page 83, for how to highlight articles automatically.

3.17.2 Article Fontisizing

People commonly add emphasis to words in news articles by writing things like ‘_this_’ or ‘*this*’. Gnus can make this look nicer by running the article through the W e (gnus-

article-emphasize) command.

How the emphasis is computed is controlled by the gnus-emphasis-alist variable. This is an alist where the first element is a regular expression to be matched. The second is a number that says what regular expression grouping is used to find the entire emphasized word. The third is a number that says what regexp grouping should be displayed and highlighted. (The text between these two groupings will be hidden.) The fourth is the face used for highlighting.

(setq gnus-article-emphasis
'(("_\\(\\w+\\)_" 0 1 gnus-emphasis-underline)
("\\*\\(\\w+\\)\\*" 0 1 gnus-emphasis-bold)))

By default, there are seven rules, and they use the following faces: gnus-emphasis-bold,

gnus-emphasis-italic, gnus-emphasis-underline, gnus-emphasis-bold-italic,
gnus-emphasis-underline-italic, gnus-emphasis-underline-bold, and gnus-
emphasis-underline-bold-italic.

If you want to change these faces, you can either use M-x customize, or you can use copy-face. For instance, if you want to make gnus-emphasis-italic use a red face instead, you could say something like:

     (copy-face 'red 'gnus-emphasis-italic)

See Section 4.3 [Customizing Articles], page 83, for how to fontize articles automatically.

3.17.3 Article Hiding

Or rather, hiding certain things in each article. There usually is much too much cruft in most articles.

W W a Do quite a lot of hiding on the article buffer (gnus-article-hide). In particular, this function will hide headers, PGP, cited text and the signature.

W W h Hide headers (gnus-article-hide-headers). See Section 4.1 [Hiding Headers], page 81.

W W b Hide headers that aren’t particularly interesting (gnus-article-hide-boring-headers). See Section 4.1 [Hiding Headers], page 81.

W W s Hide signature (gnus-article-hide-signature). See Section 3.17.7 [Article Signature], page 69.

W W p  Hide pgp signatures (gnus-article-hide-pgp). The gnus-article-hide-

pgp-hook hook will be run after a pgp signature has been hidden.

W W P Hide pem (privacy enhanced messages) cruft (gnus-article-hide-pem).

W W c Hide citation (gnus-article-hide-citation). Some variables for customizing the hiding:

          gnus-cited-opened-text-button-line-format
          gnus-cited-closed-text-button-line-format

Gnus adds buttons to show where the cited text has been hidden, and to allow toggle hiding the text. The format of the variable is specified by these format-like variable (see Section 8.4 [Formatting Variables], page 154). These specs are valid:

‘b’ Starting point of the hidden text.

‘e’ Ending point of the hidden text.

‘l’ Number of characters in the hidden region.

‘n’ Number of lines of hidden text.

          gnus-cited-lines-visible

The number of lines at the beginning of the cited text to leave shown.

W W C-c

Hide citation (gnus-article-hide-citation-maybe) depending on the following two variables:

gnus-cite-hide-percentage

If the cited text is of a bigger percentage than this variable (default 50), hide the cited text.

gnus-cite-hide-absolute

The cited text must have at least this length (default 10) before it is hidden.

W W C Hide cited text in articles that aren’t roots (gnus-article-hide-citation-in-followups). This isn’t very useful as an interactive command, but might be a handy function to stick in gnus-article-display-hook (see Section 4.3 [Customizing Articles], page 83).

All these “hiding” commands are toggles, but if you give a negative prefix to these commands, they will show what they have previously hidden. If you give a positive prefix, they will always hide.

Also see Section 3.17.1 [Article Highlighting], page 63 for further variables for citation customization.

See Section 4.3 [Customizing Articles], page 83, for how to hide article elements automatically.

3.17.4 Article Washing

We call this “article washing” for a really good reason. Namely, the A key was taken, so we had to use the W key instead.

Washing is defined by us as “changing something from something to something else”, but normally results in something looking better. Cleaner, perhaps.

W l Remove page breaks from the current article (gnus-summary-stop-page-breaking). See Section 4.5 [Misc Article], page 84, for page delimiters.

W r Do a Caesar rotate (rot13) on the article buffer (gnus-summary-caesar-message). Unreadable articles that tell you to read them with Caesar rotate or rot13. (Typically offensive jokes and such.) It’s commonly called “rot13” because each letter is rotated 13 positions in the alphabet, e. g. ‘B’ (letter #2) -> ‘O’ (letter #15). It is sometimes referred to as “Caesar rotate” because Caesar is rumored to have employed this form of, uh, somewhat weak encryption.

W t Toggle whether to display all headers in the article buffer (gnus-summary-

toggle-header).

W v Toggle whether to display all headers in the article buffer permanently (gnus-

summary-verbose-header).

W m Toggle whether to run the article through mime before displaying

(gnus-summary-toggle-mime).

W o  Treat overstrike (gnus-article-treat-overstrike).
W d  Treat M******** sm*rtq**t*s (gnus-article-treat-dumbquotes).
W w  Do word wrap (gnus-article-fill-cited-article). If you use this function

in gnus-article-display-hook, it should be run fairly late and certainly after any highlighting. You can give the command a numerical prefix to specify the width to use when filling.

W c Remove CR (i. e., ‘^M’s on the end of the lines) (gnus-article-remove-cr).

W q  Treat quoted-printable (gnus-article-de-quoted-unreadable). Quoted-

Printable is one common mime encoding employed when sending non-ASCII (i. e., 8-bit) articles. It typically makes strings like ‘dj vu’ look like ‘d=E9j=E0 vu’, which doesn’t look very readable to me.

W f Look for and display any X-Face headers (gnus-article-display-x-face). The command executed by this function is given by the gnus-article-x-face-command variable. If this variable is a string, this string will be executed in a sub-shell. If it is a function, this function will be called with the face as the argument. If the gnus-article-x-face-too-ugly (which is a regexp) matches the From header, the face will not be shown. The default action under Emacs is to fork off an xv to view the face; under XEmacs the default action is to display the face before the From header. (It’s nicer if XEmacs has been compiled with X-Face support—that will make display somewhat faster. If there’s no native X-Face support, Gnus will try to convert the X-Face header using external programs from the pbmplus package and friends.) If you want to have this function in the display hook, it should probably come last.

W b Add clickable buttons to the article (gnus-article-add-buttons). See Section 3.17.5 [Article Buttons], page 68.

W B Add clickable buttons to the article headers (gnus-article-add-buttons-to-

head).

W E l Remove all blank lines from the beginning of the article (gnus-article-strip-

leading-blank-lines).

W E m Replace all blank lines with empty lines and then all multiple empty lines with

a single empty line. (gnus-article-strip-multiple-blank-lines).

W E t Remove all blank lines at the end of the article (gnus-article-remove-

trailing-blank-lines).
W E a  Do all the three commands above (gnus-article-strip-blank-lines).
W E A  Remove all blank lines (gnus-article-strip-all-blank-lines).

W E s Remove all white space from the beginning of all lines of the article body (gnus-

article-strip-leading-space).

See Section 4.3 [Customizing Articles], page 83, for how to wash articles automatically.

3.17.5 Article Buttons

People often include references to other stuff in articles, and it would be nice if Gnus could just fetch whatever it is that people talk about with the minimum of fuzz when you hit RET or use the middle mouse button on these references.

Gnus adds buttons to certain standard references by default: Well-formed URLs, mail addresses and Message-IDs. This is controlled by two variables, one that handles article bodies and one that handles article heads:

gnus-button-alist

This is an alist where each entry has this form:

     (REGEXP BUTTON-PAR USE-P FUNCTION DATA-PAR)

regexp All text that match this regular expression will be considered an external reference. Here’s a typical regexp that matches embedded

          URLs: ‘<URL:\\([^\n\r>]*\\)>’.

button-par Gnus has to know which parts of the matches is to be highlighted. This is a number that says what sub-expression of the regexp is to be highlighted. If you want it all highlighted, you use 0 here.

use-p This form will be evaled, and if the result is non-nil, this is considered a match. This is useful if you want extra sifting to avoid false matches.

function This function will be called when you click on this button.

data-par As with button-par, this is a sub-expression number, but this one says which part of the match is to be sent as data to function.

So the full entry for buttonizing URLs is then

     ("<URL:\\([^\n\r>]*\\)>" 0 t gnus-button-url 1)
gnus-header-button-alist

This is just like the other alist, except that it is applied to the article head only, and that each entry has an additional element that is used to say what headers to apply the buttonize coding to:

     (HEADER REGEXP BUTTON-PAR USE-P FUNCTION DATA-PAR)

HEADER is a regular expression.

gnus-button-url-regexp

A regular expression that matches embedded URLs. It is used in the default values of the variables above.

gnus-article-button-face

Face used on buttons.

gnus-article-mouse-face

Face used when the mouse cursor is over a button.

See Section 4.3 [Customizing Articles], page 83, for how to buttonize articles automatically.

3.17.6 Article Date

The date is most likely generated in some obscure timezone you’ve never heard of, so it’s quite nice to be able to find out what the time was when the article was sent.

W T u Display the date in UT (aka. GMT, aka ZULU) (gnus-article-date-ut).

W T i Display the date in international format, aka. ISO 8601 (gnus-article-date-

          iso8601).

W T l Display the date in the local timezone (gnus-article-date-local).

W T s Display the date using a user-defined format (gnus-article-date-user). The format is specified by the gnus-article-time-format variable, and is a string that’s passed to format-time-string. See the documentation of that variable for a list of possible format specs.

W T e Say how much time has elapsed between the article was posted and now (gnusarticle-date-lapsed). If you want to have this line updated continually, you can put

               (gnus-start-date-timer)

in your ‘.gnus.el’ file, or you can run it off of some hook. If you want to stop the timer, you can use the gnus-stop-date-timer command.

W T o Display the original date (gnus-article-date-original). This can be useful if you normally use some other conversion function and are worried that it might be doing something totally wrong. Say, claiming that the article was posted in 1854. Although something like that is totally impossible. Don’t you trust me? *titter*

See Section 4.3 [Customizing Articles], page 83, for how to display the date in your preferred format automatically.

3.17.7 Article Signature

Each article is divided into two parts—the head and the body. The body can be divided into a signature part and a text part. The variable that says what is to be considered a signature is gnus-signature-separator. This is normally the standard ‘^-- $’ as mandated by son-of-RFC 1036. However, many people use non-standard signature separators, so this variable can also be a list of regular expressions to be tested, one by one. (Searches are done from the end of the body towards the beginning.) One likely value is:

     (setq gnus-signature-separator
           '("^-- $"         ; The standard
             "^-- *$"        ; A common mangling
             "^-------*$"    ; Many people just use a looong
                             ; line of dashes.  Shame!
             "^*--------*$" ; Double-shame!
             "^________*$"   ; Underscores are also popular
             "^========*$")) ; Pervert!

The more permissive you are, the more likely it is that you’ll get false positives.

gnus-signature-limit provides a limit to what is considered a signature.

1. If it is an integer, no signature may be longer (in characters) than that integer.

2. If it is a floating point number, no signature may be longer (in lines) than that number.

3. If it is a function, the function will be called without any parameters, and if it returns nil, there is no signature in the buffer.

4. If it is a string, it will be used as a regexp. If it matches, the text in question is not a signature.

This variable can also be a list where the elements may be of the types listed above. Here’s an example:

     (setq gnus-signature-limit
           '(200.0 "^---*Forwarded article"))

This means that if there are more than 200 lines after the signature separator, or the text after the signature separator is matched by the regular expression ‘^---*Forwarded article’, then it isn’t a signature after all.

3.18 Article Commands

A P Generate and print a PostScript image of the article buffer (gnus-summary-print-article). gnus-ps-print-hook will be run just before printing the buffer.

3.19 Summary Sorting

You can have the summary buffer sorted in various ways, even though I can’t really see why you’d want that.

C-c C-s C-n
          Sort by article number (gnus-summary-sort-by-number).
C-c C-s C-a
          Sort by author (gnus-summary-sort-by-author).
C-c C-s C-s
          Sort by subject (gnus-summary-sort-by-subject).
C-c C-s C-d
          Sort by date (gnus-summary-sort-by-date).
C-c C-s C-l
          Sort by lines (gnus-summary-sort-by-lines).
C-c C-s C-i
          Sort by score (gnus-summary-sort-by-score).

These functions will work both when you use threading and when you don’t use threading. In the latter case, all summary lines will be sorted, line by line. In the former case, sorting will be done on a root-by-root basis, which might not be what you were looking for. To toggle whether to use threading, type T T (see Section 3.9.2 [Thread Commands], page 51).

3.20 Finding the Parent

^If you’d like to read the parent of the current article, and it is not displayed in the summary buffer, you might still be able to. That is, if the current group is fetched by nntp, the parent hasn’t expired and the References in the current article are not mangled, you can just press ^or A r (gnus-summary-refer-parent-article). If everything goes well, you’ll get the parent. If the parent is already displayed in the summary buffer, point will just move to this article. If given a positive numerical prefix, fetch that many articles back into the ancestry. If given a negative numerical prefix, fetch just that ancestor. So if you say 3 ^, Gnus will fetch the parent, the grandparent and the grandgrandparent of the current article. If you say -3 ^, Gnus will only fetch the grandgrandparent of the current article.

A R (Summary)

Fetch all articles mentioned in the References header of the article (gnus-

summary-refer-references).
A T (Summary)

Display the full thread where the current article appears (gnus-summary-refer-thread). This command has to fetch all the headers in the current group to work, so it usually takes a while. If you do it often, you may consider setting gnus-fetch-old-headers to invisible (see Section 3.9.1.2 [Filling In Threads], page 49). This won’t have any visible effects normally, but it’ll make this command work a whole lot faster. Of course, it’ll make group entry somewhat slow. The gnus-refer-thread-limit variable says how many old (i. e., articles before the first displayed in the current group) headers to fetch when doing this command. The default is 200. If t, all the available headers will be fetched. This variable can be overridden by giving the A T command a numerical prefix.

M-^(Summary)

You can also ask the nntp server for an arbitrary article, no matter what group

it belongs to. M-^(gnus-summary-refer-article) will ask you for a Message-

ID, which is one of those long, hard-to-read thingies that look something like ‘<38o6up$6f2@hymir.ifi.uio.no>’. You have to get it all exactly right. No fuzzy searches, I’m afraid.

The current select method will be used when fetching by Message-ID from non-news select method, but you can override this by giving this command a prefix.

If the group you are reading is located on a backend that does not support fetching by

Message-ID very well (like nnspool), you can set gnus-refer-article-method to an nntp

method. It would, perhaps, be best if the nntp server you consult is the one updating the spool you are reading from, but that’s not really necessary.

Most of the mail backends support fetching by Message-ID, but do not do a particularly excellent job at it. That is, nnmbox and nnbabyl are able to locate articles from any groups, while nnml and nnfolder are only able to locate articles that have been posted to the current group. (Anything else would be too time consuming.) nnmh does not support this at all.

3.21 Alternative Approaches

Different people like to read news using different methods. This being Gnus, we offer a small selection of minor modes for the summary buffers.

3.21.1 Pick and Read

Some newsreaders (like nn and, uhm, Netnews on VM/CMS) use a two-phased reading interface. The user first marks in a summary buffer the articles she wants to read. Then she starts reading the articles with just an article buffer displayed.

Gnus provides a summary buffer minor mode that allows this—gnus-pick-mode. This basically means that a few process mark commands become one-keystroke commands to allow easy marking, and it provides one additional command for switching to the summary buffer.

Here are the available keystrokes when using pick mode:

. Pick the article or thread on the current line (gnus-pick-article-or-thread). If the variable gnus-thread-hide-subtree is true, then this key selects the entire thread when used at the first article of the thread. Otherwise, it selects just the article. If given a numerical prefix, go to that thread or article and pick it. (The line number is normally displayed at the beginning of the summary pick lines.)

SPACE Scroll the summary buffer up one page (gnus-pick-next-page). If at the end of the buffer, start reading the picked articles.

u  Unpick the thread or article (gnus-pick-unmark-article-or-thread). If the

variable gnus-thread-hide-subtree is true, then this key unpicks the thread if used at the first article of the thread. Otherwise it unpicks just the article. You can give this key a numerical prefix to unpick the thread or article at that line.

RET Start reading the picked articles (gnus-pick-start-reading). If given a prefix, mark all unpicked articles as read first. If gnus-pick-display-summary is nonnil, the summary buffer will still be visible when you are reading.

All the normal summary mode commands are still available in the pick-mode, with the exception of u. However ! is available which is mapped to the same function gnus-summary-

tick-article-forward.

If this sounds like a good idea to you, you could say:

     (add-hook 'gnus-summary-mode-hook 'gnus-pick-mode)

gnus-pick-mode-hook is run in pick minor mode buffers.

   If gnus-mark-unpicked-articles-as-read is non-nil, mark all unpicked articles as

read. The default is nil.

The summary line format in pick mode is slightly different from the standard format. At the beginning of each line the line number is displayed. The pick mode line format is controlled by the gnus-summary-pick-line-format variable (see Section 8.4 [Formatting Variables], page 154). It accepts the same format specs that gnus-summary-line-format does (see Section 3.1.1 [Summary Buffer Lines], page 33).

3.21.2 Binary Groups

If you spend much time in binary groups, you may grow tired of hitting X u, n, RET all the time. M-x gnus-binary-mode is a minor mode for summary buffers that makes all ordinary Gnus article selection functions uudecode series of articles and display the result instead of just displaying the articles the normal way.

The only way, in fact, to see the actual articles is the g command, when you have turned

on this mode (gnus-binary-show-article).

gnus-binary-mode-hook is called in binary minor mode buffers.

3.22 Tree Display

If you don’t like the normal Gnus summary display, you might try setting gnus-use-trees to t. This will create (by default) an additional tree buffer. You can execute all summary mode commands in the tree buffer.

There are a few variables to customize the tree display, of course:

gnus-tree-mode-hook

A hook called in all tree mode buffers.

gnus-tree-mode-line-format

A format string for the mode bar in the tree mode buffers (see Section 8.4.2 [Mode Line Formatting], page 155). The default is ‘Gnus: %%b %S %Z’. For a list of valid specs, see Section 3.1.2 [Summary Buffer Mode Line], page 35.

gnus-selected-tree-face

Face used for highlighting the selected article in the tree buffer. The default is

modeline.
gnus-tree-line-format

A format string for the tree nodes. The name is a bit of a misnomer, though—it doesn’t define a line, but just the node. The default value is ‘%(%[%3,3n%]%)’, which displays the first three characters of the name of the poster. It is vital that all nodes are of the same length, so you must use ‘%4,4n’-like specifiers. Valid specs are:

‘n’ The name of the poster.

‘f’ The From header.

‘N’ The number of the article.

‘[’ The opening bracket.

‘]’ The closing bracket.

‘s’ The subject.

See Section 8.4 [Formatting Variables], page 154. Variables related to the display are:

gnus-tree-brackets

This is used for differentiating between “real” articles and “sparse” articles. The format is ((real-open . real-close) (sparse-open . sparse-close) (dummy-open . dummy-close)), and the default is

          ((?[ . ?]) (?( . ?)) (?{ . ?}) (?< . ?>)).
gnus-tree-parent-child-edges

This is a list that contains the characters used for connecting parent nodes to their children. The default is (?- ?\\ ?|).

gnus-tree-minimize-window

If this variable is non-nil, Gnus will try to keep the tree buffer as small as possible to allow more room for the other Gnus windows. If this variable is a number, the tree buffer will never be higher than that number. The default is t. Note that if you have several windows displayed side-by-side in a frame and the tree buffer is one of these, minimizing the tree window will also resize all other windows displayed next to it.

gnus-generate-tree-function

The function that actually generates the thread tree. Two predefined func-

tions are available: gnus-generate-horizontal-tree and gnus-generate-

vertical-tree (which is the default).

Here’s an example from a horizontal tree buffer:

{***}-(***)-[odd]-[Gun]
|      \[Jan]
|      \[odd]-[Eri]
|      \(***)-[Eri]
|            \[odd]-[Paa]
\[Bjo]
\[Gun]
\[Gun]-[Jor]

Here’s the same thread displayed in a vertical tree buffer:

{***}
|--------------------------\-----\-----\
(***)                         [Bjo] [Gun] [Gun]
|--\-----\-----\                          |
[odd] [Jan] [odd] (***)                   [Jor]
|           |     |--\
[Gun]       [Eri] [Eri] [odd]
                     |
                   [Paa]

If you’re using horizontal trees, it might be nice to display the trees side-by-side with the summary buffer. You could add something like the following to your ‘.gnus.el’ file:

(setq gnus-use-trees t
 gnus-generate-tree-function 'gnus-generate-horizontal-tree
 gnus-tree-minimize-window nil)
(gnus-add-configuration
'(article

(vertical 1.0
        (horizontal 0.25
                    (summary 0.75 point)
                    (tree 1.0))
        (article 1.0))))

See Section 8.5 [Windows Configuration], page 157.

3.23 Mail Group Commands

Some commands only make sense in mail groups. If these commands are invalid in the current group, they will raise a hell and let you know.

All these commands (except the expiry and edit commands) use the process/prefix convention (see Section 8.1 [Process/Prefix], page 153).

B e Expire all expirable articles in the group (gnus-summary-expire-articles).

B M-C-e Delete all the expirable articles in the group (gnus-summary-expire-articles-now). This means that all articles eligible for expiry in the current group will disappear forever into that big ‘/dev/null’ in the sky.

B DEL Delete the mail article. This is “delete” as in “delete it from your disk forever and ever, never to return again.” Use with caution. (gnus-summary-delete-

article).

B m Move the article from one mail group to another (gnus-summary-move-

article).

B c Copy the article from one group (mail group or not) to a mail group (gnus-

summary-copy-article).

B B Crosspost the current article to some other group (gnus-summary-crosspost-article). This will create a new copy of the article in the other group, and the Xref headers of the article will be properly updated.

B i Import an arbitrary file into the current mail newsgroup (gnus-summary-import-article). You will be prompted for a file name, a From header and a

Subject header.
B r  Respool the mail article (gnus-summary-respool-article). gnus-summary-

respool-default-method will be used as the default select method when respooling. This variable is nil by default, which means that the current group select method will be used instead.

B w

e Edit the current article (gnus-summary-edit-article). To finish editing and

make the changes permanent, type C-c C-c (gnus-summary-edit-article-

done). If you give a prefix to the C-c C-c command, Gnus won’t re-highlight the article.

B q If you want to re-spool an article, you might be curious as to what group the article will end up in before you do the re-spooling. This command will tell you

(gnus-summary-respool-query).

B t Similarly, this command will display all fancy splitting patterns used when

repooling, if any (gnus-summary-respool-trace).

B p Some people have a tendency to send you "courtesy" copies when they follow up to articles you have posted. These usually have a Newsgroups header in them, but not always. This command (gnus-summary-article-posted-p) will try to fetch the current article from your news server (or rather, from gnus-

refer-article-method or gnus-select-method) and will report back whether

it found the article or not. Even if it says that it didn’t find the article, it may have been posted anyway—mail propagation is much faster than news propagation, and the news copy may just not have arrived yet.

If you move (or copy) articles regularly, you might wish to have Gnus suggest where to put the articles. gnus-move-split-methods is a variable that uses the same syntax as gnus-split-methods (see Section 3.15 [Saving Articles], page 56). You may customize that variable to create suggestions you find reasonable.

(setq gnus-move-split-methods
 '(("^From:.*Lars Magne" "nnml:junk")
   ("^Subject:.*gnus" "nnfolder:important")
   (".*" "nnml:misc")))

3.24 Various Summary Stuff

gnus-summary-mode-hook

This hook is called when creating a summary mode buffer.

gnus-summary-generate-hook

This is called as the last thing before doing the threading and the generation of the summary buffer. It’s quite convenient for customizing the threading variables based on what data the newsgroup has. This hook is called from the summary buffer after most summary buffer variables have been set.

gnus-summary-prepare-hook

It is called after the summary buffer has been generated. You might use it to, for instance, highlight lines or modify the look of the buffer in some other ungodly manner. I don’t care.

gnus-summary-ignore-duplicates

When Gnus discovers two articles that have the same Message-ID, it has to do something drastic. No articles are allowed to have the same Message-ID, but this may happen when reading mail from some sources. Gnus allows you to customize what happens with this variable. If it is nil (which is the default), Gnus will rename the Message-ID (for display purposes only) and display the article as any other article. If this variable is t, it won’t display the article—it’ll be as if it never existed.

3.24.1 Summary Group Information

H f Try to fetch the FAQ (list of frequently asked questions) for the current group

(gnus-summary-fetch-faq). Gnus will try to get the FAQ from gnus-group-

faq-directory, which is usually a directory on a remote machine. This variable can also be a list of directories. In that case, giving a prefix to this command will allow you to choose between the various sites. ange-ftp or efs will probably be used for fetching the file.

H d Give a brief description of the current group (gnus-summary-describe-group). If given a prefix, force rereading the description from the server.

H h Give an extremely brief description of the most important summary keystrokes

(gnus-summary-describe-briefly).

H i Go to the Gnus info node (gnus-info-find-node).

3.24.2 Searching for Articles

M-s Search through all subsequent articles for a regexp (gnus-summary-search-

article-forward).

M-r Search through all previous articles for a regexp (gnus-summary-search-

article-backward).

& This command will prompt you for a header field, a regular expression to match on this field, and a command to be executed if the match is made (gnussummary-execute-command). If given a prefix, search backward instead.

M-& Perform any operation on all articles that have been marked with the process

mark (gnus-summary-universal-argument).

3.24.3 Summary Generation Commands

Y g Regenerate the current summary buffer (gnus-summary-prepare).

Y c Pull all cached articles (for the current group) into the summary buffer (gnus-

summary-insert-cached-articles).

3.24.4 Really Various Summary Commands

C-d If the current article is a collection of other articles (for instance, a digest), you might use this command to enter a group based on the that article (gnussummary-enter-digest-group). Gnus will try to guess what article type is currently displayed unless you give a prefix to this command, which forces a “digest” interpretation. Basically, whenever you see a message that is a collection of other messages of some format, you C-d and read these messages in a more convenient fashion.

M-C-d This command is very similar to the one above, but lets you gather several documents into one biiig group (gnus-summary-read-document). It does this by opening several nndoc groups for each document, and then opening an nnvirtual group on top of these nndoc groups. This command understands the process/prefix convention (see Section 8.1 [Process/Prefix], page 153).

C-t  Toggle truncation of summary lines (gnus-summary-toggle-truncation).

This will probably confuse the line centering function in the summary buffer, so it’s not a good idea to have truncation switched off while reading articles.

= Expand the summary buffer window (gnus-summary-expand-window). If given a prefix, force an article window configuration.

M-C-e Edit the group parameters (see Section 2.10 [Group Parameters], page 19) of

          the current group (gnus-summary-edit-parameters).

3.25 Exiting the Summary Buffer

Exiting from the summary buffer will normally update all info on the group and return you to the group buffer.

Z Z

q Exit the current group and update all information on the group (gnus-

          summary-exit). gnus-summary-prepare-exit-hook is called before doing

much of the exiting, which calls gnus-summary-expire-articles by default. gnus-summary-exit-hook is called after finishing the exit process.

          gnus-group-no-more-groups-hook is run when returning to group mode

having no more (unread) groups.

Z E

Q Exit the current group without updating any information on the group (gnus-

          summary-exit-no-update).
Z c

c Mark all unticked articles in the group as read and then exit (gnus-summary-

          catchup-and-exit).

Z C Mark all articles, even the ticked ones, as read and then exit (gnus-summary-

          catchup-all-and-exit).

Z n Mark all articles as read and go to the next group (gnus-summary-catchup-

          and-goto-next-group).

Z R Exit this group, and then enter it again (gnus-summary-reselect-current-group). If given a prefix, select all articles, both read and unread.

Z G

M-g Exit the group, check for new articles in the group, and select the group (gnussummary-rescan-group). If given a prefix, select all articles, both read and unread.

Z N Exit the group and go to the next group (gnus-summary-next-group).

Z P Exit the group and go to the previous group (gnus-summary-prev-group).

Z s Save the current number of read/marked articles in the dribble buffer and then save the dribble buffer (gnus-summary-save-newsrc). If given a prefix, also save the ‘.newsrc’ file(s). Using this command will make exit without updating (the Q command) worthless. gnus-exit-group-hook is called when you exit the current group. If you’re in the habit of exiting groups, and then changing your mind about it, you might

set gnus-kill-summary-on-exit to nil. If you do that, Gnus won’t kill the summary buffer when you exit it. (Quelle surprise!) Instead it will change the name of the buffer to something like ‘*Dead Summary ... *’ and install a minor mode called gnus-dead-summary-mode. Now, if you switch back to this buffer, you’ll find that all keys are mapped to a function called gnus-summary-wake-up-the-dead. So tapping any keys in a dead summary buffer will result in a live, normal summary buffer.

There will never be more than one dead summary buffer at any one time.

The data on the current group will be updated (which articles you have read, which articles you have replied to, etc.) when you exit the summary buffer. If the gnus-use-cross-reference variable is t (which is the default), articles that are cross-referenced to this group and are marked as read, will also be marked as read in the other subscribed groups they were cross-posted to. If this variable is neither nil nor t, the article will be marked as read in both subscribed and unsubscribed groups (see Section 3.26 [Crosspost Handling], page 79).

3.26 Crosspost Handling

Marking cross-posted articles as read ensures that you’ll never have to read the same article more than once. Unless, of course, somebody has posted it to several groups separately. Posting the same article to several groups (not cross-posting) is called spamming, and you are by law required to send nasty-grams to anyone who perpetrates such a heinous crime. You may want to try NoCeM handling to filter out spam (see Section 8.12 [NoCeM], page 164).

Remember: Cross-posting is kinda ok, but posting the same article separately to several groups is not. Massive cross-posting (aka. velveeta) is to be avoided at all costs, and you

can even use the gnus-summary-mail-crosspost-complaint command to complain about

excessive crossposting (see Section 3.5.1 [Summary Mail Commands], page 39).

One thing that may cause Gnus to not do the cross-posting thing correctly is if you use an nntp server that supports xover (which is very nice, because it speeds things up considerably) which does not include the Xref header in its nov lines. This is Evil, but all too common, alas, alack. Gnus tries to Do The Right Thing even with xover by registering the Xref lines of all articles you actually read, but if you kill the articles, or just mark them as read without reading them, Gnus will not get a chance to snoop the Xref lines out of these articles, and will be unable to use the cross reference mechanism.

To check whether your nntp server includes the Xref header in its overview files,

try ‘telnet your.nntp.server nntp’, ‘MODE READER’ on inn servers, and then say ‘LIST

overview.fmt’. This may not work, but if it does, and the last line you get does not read ‘Xref:full’, then you should shout and whine at your news admin until she includes the Xref header in the overview files.

If you want Gnus to get the Xrefs right all the time, you have to set gnus-nov-is-evil to t, which slows things down considerably.

C’est la vie.

For an alternative approach, see Section 3.27 [Duplicate Suppression], page 80.

3.27 Duplicate Suppression

By default, Gnus tries to make sure that you don’t have to read the same article more than once by utilizing the crossposting mechanism (see Section 3.26 [Crosspost Handling], page 79). However, that simple and efficient approach may not work satisfactory for some users for various reasons.

1. The nntp server may fail to generate the Xref header. This is evil and not very common.

2. The nntp server may fail to include the Xref header in the ‘.overview’ data bases. This is evil and all too common, alas.

3. You may be reading the same group (or several related groups) from different nntp servers.

4. You may be getting mail that duplicates articles posted to groups.

I’m sure there are other situations where Xref handling fails as well, but these four are the most common situations.

If, and only if, Xref handling fails for you, then you may consider switching on duplicate suppression. If you do so, Gnus will remember the Message-IDs of all articles you have read or otherwise marked as read, and then, as if by magic, mark them as read all subsequent times you see them—in all groups. Using this mechanism is quite likely to be somewhat inefficient, but not overly so. It’s certainly preferable to reading the same articles more than once.

Duplicate suppression is not a very subtle instrument. It’s more like a sledge hammer than anything else. It works in a very simple fashion—if you have marked an article as read, it adds this Message-ID to a cache. The next time it sees this Message-ID, it will mark the article as read with the ‘M’ mark. It doesn’t care what group it saw the article in.

gnus-suppress-duplicates

If non-nil, suppress duplicates.

gnus-save-duplicate-list

If non-nil, save the list of duplicates to a file. This will make startup and shutdown take longer, so the default is nil. However, this means that only duplicate articles read in a single Gnus session are suppressed.

gnus-duplicate-list-length

This variable says how many Message-IDs to keep in the duplicate suppression list. The default is 10000.

gnus-duplicate-file

The name of the file to store the duplicate suppression list in. The default is

          ‘~/News/suppression’.

If you have a tendency to stop and start Gnus often, setting gnus-save-duplicate-list to t is probably a good idea. If you leave Gnus running for weeks on end, you may have it nil. On the other hand, saving the list makes startup and shutdown much slower, so that means that if you stop and start Gnus often, you should set gnus-save-duplicate-list to nil. Uhm. I’ll leave this up to you to figure out, I think.

4 The Article Buffer

The articles are displayed in the article buffer, of which there is only one. All the summary buffers share the same article buffer unless you tell Gnus otherwise.

4.1 Hiding Headers

The top section of each article is the head. (The rest is the body, but you may have guessed that already.)

There is a lot of useful information in the head: the name of the person who wrote the article, the date it was written and the subject of the article. That’s well and nice, but there’s also lots of information most people do not want to see—what systems the article has passed through before reaching you, the Message-ID, the References, etc. ad nauseum—and you’ll probably want to get rid of some of those lines. If you want to keep all those lines in the article buffer, you can set gnus-show-all-headers to t.

Gnus provides you with two variables for sifting headers:

gnus-visible-headers

If this variable is non-nil, it should be a regular expression that says what headers you wish to keep in the article buffer. All headers that do not match this variable will be hidden. For instance, if you only want to see the name of the person who wrote the article and the subject, you’d say:

               (setq gnus-visible-headers "^From:\\|^Subject:")

This variable can also be a list of regexps to match headers to remain visible.

gnus-ignored-headers

This variable is the reverse of gnus-visible-headers. If this variable is set (and gnus-visible-headers is nil), it should be a regular expression that matches all lines that you want to hide. All lines that do not match this variable will remain visible. For instance, if you just want to get rid of the References line and the Xref line, you might say:

               (setq gnus-ignored-headers "^References:\\|^Xref:")

This variable can also be a list of regexps to match headers to be removed. Note that if gnus-visible-headers is non-nil, this variable will have no effect.

Gnus can also sort the headers for you. (It does this by default.) You can control the sorting by setting the gnus-sorted-header-list variable. It is a list of regular expressions that says in what order the headers are to be displayed.

For instance, if you want the name of the author of the article first, and then the subject, you might say something like:

     (setq gnus-sorted-header-list '("^From:" "^Subject:"))

Any headers that are to remain visible, but are not listed in this variable, will be displayed in random order after all the headers listed in this variable.

You can hide further boring headers by entering gnus-article-hide-boring-headers

into gnus-article-display-hook. What this function does depends on the gnus-boring-

article-headers variable. It’s a list, but this list doesn’t actually contain header names. Instead is lists various boring conditions that Gnus can check and remove from sight.

These conditions are:

empty Remove all empty headers.

followup-to

Remove the Followup-To header if it is identical to the Newsgroups header.

reply-to Remove the Reply-To header if it lists the same address as the From header.

newsgroups

Remove the Newsgroups header if it only contains the current group name.

date Remove the Date header if the article is less than three days old.

long-to Remove the To header if it is very long.

many-to Remove all To headers if there are more than one.

To include the four three elements, you could say something like;

     (setq gnus-boring-article-headers
           '(empty followup-to reply-to))

This is also the default value for this variable.

4.2 Using mime

Mime is a standard for waving your hands through the air, aimlessly, while people stand around yawning.

mime, however, is a standard for encoding your articles, aimlessly, while all newsreaders die of fear.

mime may specify what character set the article uses, the encoding of the characters, and it also makes it possible to embed pictures and other naughty stuff in innocent-looking articles.

Gnus handles mime by pushing the articles through gnus-show-mime-method, which is metamail-buffer by default. This function calls the external metamail program to actually do the work. One common problem with this program is that is thinks that it can’t display 8-bit things in the Emacs buffer. To tell it the truth, put something like the following in your ‘.bash_profile’ file. (You do use bash, don’t you?)

     export MM_CHARSET="iso-8859-1"

For more information on metamail, see its manual page.

Set gnus-show-mime to t if you want to use mime all the time. However, if gnusstrict-mime is non-nil, the mime method will only be used if there are mime headers in the article. If you have gnus-show-mime set, then you’ll see some unfortunate display glitches in the article buffer. These can’t be avoided.

It might be best to just use the toggling functions from the summary buffer to avoid getting nasty surprises. (For instance, you enter the group ‘alt.sing-a-long’ and, before you know it, mime has decoded the sound file in the article and some horrible sing-a-long song comes screaming out your speakers, and you can’t find the volume button, because there isn’t one, and people are starting to look at you, and you try to stop the program, but you can’t, and you can’t find the program to control the volume, and everybody else in the room suddenly decides to look at you disdainfully, and you’ll feel rather stupid.)

Any similarity to real events and people is purely coincidental. Ahem.

4.3 Customizing Articles

The gnus-article-display-hook is called after the article has been inserted into the article buffer. It is meant to handle all treatment of the article before it is displayed.

   By default this hook just contains gnus-article-maybe-hide-headers,
gnus-hide-boring-headers, gnus-article-treat-overstrike, and gnus-article-
maybe-highlight (and under XEmacs, gnus-article-display-x-face), but there are

thousands, nay millions, of functions you can put in this hook. For an overview of functions see Section 3.17.1 [Article Highlighting], page 63, see Section 3.17.3 [Article Hiding], page 65, see Section 3.17.4 [Article Washing], page 66, see Section 3.17.5 [Article Buttons], page 68 and see Section 3.17.6 [Article Date], page 69. Note that the order of functions in this hook might affect things, so you may have to fiddle a bit to get the desired results.

You can, of course, write your own functions. The functions are called from the article buffer, and you can do anything you like, pretty much. There is no information that you have to keep in the buffer—you can change everything. However, you shouldn’t delete any headers. Instead make them invisible if you want to make them go away.

4.4 Article Keymap

Most of the keystrokes in the summary buffer can also be used in the article buffer. They should behave as if you typed them in the summary buffer, which means that you don’t actually have to have a summary buffer displayed while reading. You can do it all from the article buffer.

A few additional keystrokes are available:

SPACE  Scroll forwards one page (gnus-article-next-page).
DEL  Scroll backwards one page (gnus-article-prev-page).

C-c ^If point is in the neighborhood of a Message-ID and you press C-c ^, Gnus will try to get that article from the server (gnus-article-refer-article).

C-c C-m Send a reply to the address near point (gnus-article-mail). If given a prefix, include the mail.

s Reconfigure the buffers so that the summary buffer becomes visible (gnus-

          article-show-summary).

? Give a very brief description of the available keystrokes (gnus-article-

          describe-briefly).

TAB Go to the next button, if any (gnus-article-next-button). This only makes sense if you have buttonizing turned on.

M-TAB Go to the previous button, if any (gnus-article-prev-button).

4.5 Misc Article

gnus-single-article-buffer

If non-nil, use the same article buffer for all the groups. (This is the default.) If nil, each group will have its own article buffer.

gnus-article-prepare-hook

This hook is called right after the article has been inserted into the article buffer. It is mainly intended for functions that do something depending on the contents; it should probably not be used for changing the contents of the article buffer.

gnus-article-display-hook

This hook is called as the last thing when displaying an article, and is intended for modifying the contents of the buffer, doing highlights, hiding headers, and the like.

gnus-article-mode-hook

Hook called in article mode buffers.

gnus-article-mode-syntax-table

Syntax table used in article buffers. It is initialized from text-mode-syntax-

table.
gnus-article-mode-line-format

This variable is a format string along the same lines as gnus-summary-mode-line-format (see Section 8.4.2 [Mode Line Formatting], page 155). It accepts the same format specifications as that variable, with one extension:

‘w’ The wash status of the article. This is a short string with one character for each possible article wash operation that may have been performed.

gnus-break-pages

Controls whether page breaking is to take place. If this variable is non-nil, the articles will be divided into pages whenever a page delimiter appears in the article. If this variable is nil, paging will not be done.

gnus-page-delimiter

This is the delimiter mentioned above. By default, it is ‘^L’ (formfeed).

5 Composing Messages

All commands for posting and mailing will put you in a message buffer where you can edit the article all you like, before you send the article by pressing C-c C-c. See section “Top” in The Message Manual. If you are in a foreign news group, and you wish to post the article using the foreign server, you can give a prefix to C-c C-c to make Gnus try to post using the foreign server.

Also see see Section 3.6 [Canceling and Superseding], page 41 for information on how to remove articles you shouldn’t have posted.

5.1 Mail

Variables for customizing outgoing mail:

gnus-uu-digest-headers

List of regexps to match headers included in digested messages. The headers will be included in the sequence they are matched.

gnus-add-to-list

If non-nil, add a to-list group parameter to mail groups that have none when you do a a.

5.2 Post

Variables for composing news articles:

gnus-sent-message-ids-file

Gnus will keep a Message-ID history file of all the mails it has sent. If it discovers that it has already sent a mail, it will ask the user whether to re-send the mail. (This is primarily useful when dealing with soup packets and the like where one is apt to send the same packet multiple times.) This variable says what the name of this history file is. It is ‘~/News/Sent-Message-IDs’ by default. Set this variable to nil if you don’t want Gnus to keep a history file.

gnus-sent-message-ids-length

This variable says how many Message-IDs to keep in the history file. It is 1000 by default.

5.3 Posting Server

When you press those magical C-c C-c keys to ship off your latest (extremely intelligent, of course) article, where does it go?

Thank you for asking. I hate you.

It can be quite complicated. Normally, Gnus will use the same native server. However. If your native server doesn’t allow posting, just reading, you probably want to use some other server to post your (extremely intelligent and fabulously interesting) articles. You can then set the gnus-post-method to some other method:

     (setq gnus-post-method '(nnspool ""))

Now, if you’ve done this, and then this server rejects your article, or this server is down, what do you do then? To override this variable you can use a non-zero prefix to the C-c C-c command to force using the “current” server for posting.

If you give a zero prefix (i.e., C-u 0 C-c C-c) to that command, Gnus will prompt you for what method to use for posting.

You can also set gnus-post-method to a list of select methods. If that’s the case, Gnus will always prompt you for what method to use for posting.

Finally, if you want to always post using the same select method as you’re reading from (which might be convenient if you’re reading lots of groups from different private servers), you can set this variable to current.

5.4 Mail and Post

Here’s a list of variables relevant to both mailing and posting:

gnus-mailing-list-groups

If your news server offers groups that are really mailing lists gatewayed to the nntp server, you can read those groups without problems, but you can’t post/followup to them without some difficulty. One solution is to add a to-address to the group parameters (see Section 2.10 [Group Parameters], page 19). An easier thing to do is set the gnus-mailing-list-groups to a regexp that matches the groups that really are mailing lists. Then, at least, followups to the mailing lists will work most of the time. Posting to these groups (a) is still a pain, though.

You may want to do spell-checking on messages that you send out. Or, if you don’t want to spell-check by hand, you could add automatic spell-checking via the ispell package:

     (add-hook 'message-send-hook 'ispell-message)

5.5 Archived Messages

Gnus provides a few different methods for storing the mail and news you send. The default method is to use the archive virtual server to store the messages. If you want to disable this completely, the gnus-message-archive-group variable should be nil, which is the default.

gnus-message-archive-method says what virtual server Gnus is to use to store sent messages. The default is:

     (nnfolder "archive"
               (nnfolder-directory   "~/Mail/archive")
               (nnfolder-active-file "~/Mail/archive/active")
               (nnfolder-get-new-mail nil)
               (nnfolder-inhibit-expiry t))

You can, however, use any mail select method (nnml, nnmbox, etc.). nnfolder is a quite likeable select method for doing this sort of thing, though. If you don’t like the default directory chosen, you could say something like:

  (setq gnus-message-archive-method
        '(nnfolder "archive"
                   (nnfolder-inhibit-expiry t)
                   (nnfolder-active-file "~/News/sent-mail/active")
                   (nnfolder-directory "~/News/sent-mail/")))

Gnus will insert Gcc headers in all outgoing messages that point to one or more group(s) on that server. Which group to use is determined by the gnus-message-archive-group variable.

This variable can be used to do the following:

• a string Messages will be saved in that group.

• a list of strings Messages will be saved in all those groups.

• an alist of regexps, functions and forms When a key “matches”, the result is used.

• nil No message archiving will take place. This is the default.

Let’s illustrate:

Just saving to a single group called ‘MisK’:

  (setq gnus-message-archive-group "MisK")

Saving to two groups, ‘MisK’ and ‘safe’:

  (setq gnus-message-archive-group '("MisK" "safe"))

Save to different groups based on what group you are in:

  (setq gnus-message-archive-group
        '(("^alt" "sent-to-alt")
          ("mail" "sent-to-mail")
          (".*" "sent-to-misc")))

More complex stuff:

  (setq gnus-message-archive-group
        '((if (message-news-p)
              "misc-news"
            "misc-mail")))

How about storing all news messages in one file, but storing all mail messages in one file per month:

  (setq gnus-message-archive-group
        '((if (message-news-p)
              "misc-news"
            (concat "mail." (format-time-string
                             "%Y-%m" (current-time))))))

(XEmacs 19.13 doesn’t have format-time-string, so you’ll have to use a different value

for gnus-message-archive-group there.)

Now, when you send a message off, it will be stored in the appropriate group. (If you want to disable storing for just one particular message, you can just remove the Gcc header that has been inserted.) The archive group will appear in the group buffer the next time you start Gnus, or the next time you press F in the group buffer. You can enter it and read the articles in it just like you’d read any other group. If the group gets really big and annoying, you can simply rename if (using G r in the group buffer) to something nice— ‘misc-mail-september-1995’, or whatever. New messages will continue to be stored in the old (now empty) group.

That’s the default method of archiving sent messages. Gnus offers a different way for the people who don’t like the default method. In that case you should set gnus-message-archive-group to nil; this will disable archiving.

gnus-outgoing-message-group

All outgoing messages will be put in this group. If you want to store all your outgoing mail and articles in the group ‘nnml:archive’, you set this variable to that value. This variable can also be a list of group names. If you want to have greater control over what group to put each message in, you can set this variable to a function that checks the current newsgroup name and then returns a suitable group name (or list of names). This variable can be used instead of gnus-message-archive-group, but the latter is the preferred method.

5.6 Posting Styles

All them variables, they make my head swim.

So what if you want a different Organization and signature based on what groups you post to? And you post both from your home machine and your work machine, and you want different From lines, and so on?

One way to do stuff like that is to write clever hooks that change the variables you need to have changed. That’s a bit boring, so somebody came up with the bright idea of letting the user specify these things in a handy alist. Here’s an example of a gnus-posting-styles variable:

     ((".*"
       (signature "Peace and happiness")
       (organization "What me?"))
      ("^comp"
       (signature "Death to everybody"))
      ("comp.emacs.i-love-it"
       (organization "Emacs is it")))

As you might surmise from this example, this alist consists of several styles. Each style will be applicable if the first element “matches”, in some form or other. The entire alist will be iterated over, from the beginning towards the end, and each match will be applied, which means that attributes in later styles that match override the same attributes in earlier

matching styles. So ‘comp.programming.literate’ will have the ‘Death to everybody’

signature and the ‘What me?’ Organization header.

The first element in each style is called the match. If it’s a string, then Gnus will try to regexp match it against the group name. If it’s a function symbol, that function will be called with no arguments. If it’s a variable symbol, then the variable will be referenced. If it’s a list, then that list will be evaled. In any case, if this returns a non-nil value, then the style is said to match.

Each style may contain a arbitrary amount of attributes. Each attribute consists of a (name . value) pair. The attribute name can be one of signature, signature-file, organization, address, name or body. The attribute name can also be a string. In that case, this will be used as a header name, and the value will be inserted in the headers of the article.

The attribute value can be a string (used verbatim), a function (the return value will be used), a variable (its value will be used) or a list (it will be evaled and the return value will be used).

If you wish to check whether the message you are about to compose is meant to be a news article or a mail message, you can check the values of the two dynamically bound

variables message-this-is-news and message-this-is-mail.

So here’s a new example:

     (setq gnus-posting-styles
           '((".*"
              (signature-file "~/.signature")
              (name "User Name")
              ("X-Home-Page" (getenv "WWW_HOME"))
              (organization "People's Front Against MWM"))
             ("^rec.humor"
              (signature my-funny-signature-randomizer))
             ((equal (system-name) "gnarly")
              (signature my-quote-randomizer))
             (message-this-is-news
              (signature my-news-signature))
             (posting-from-work-p
              (signature-file "~/.work-signature")
              (address "user@bar.foo")
              (body "You are fired.\n\nSincerely, your boss.")
              (organization "Important Work, Inc"))
             ("^nn.+:"
              (signature-file "~/.mail-signature"))))

5.7 Drafts

If you are writing a message (mail or news) and suddenly remember that you have a steak in the oven (or some pesto in the food processor, you craaazy vegetarians), you’ll probably wish there was a method to save the message you are writing so that you can continue editing it some other day, and send it when you feel its finished.

Well, don’t worry about it. Whenever you start composing a message of some sort using the Gnus mail and post commands, the buffer you get will automatically associate to an article in a special draft group. If you save the buffer the normal way (C-x C-s, for instance), the article will be saved there. (Auto-save files also go to the draft group.)

The draft group is a special group (which is implemented as an nndraft group, if you absolutely have to know) called ‘nndraft:drafts’. The variable nndraft-directory says where nndraft is to store its files. What makes this group special is that you can’t tick any articles in it or mark any articles as read—all articles in the group are permanently unread.

If the group doesn’t exist, it will be created and you’ll be subscribed to it. The only way to make it disappear from the Group buffer is to unsubscribe it.

When you want to continue editing the article, you simply enter the draft group and push D e (gnus-draft-edit-message) to do that. You will be placed in a buffer where you left off.

Rejected articles will also be put in this draft group (see Section 5.8 [Rejected Articles], page 90).

If you have lots of rejected messages you want to post (or mail) without doing further editing, you can use the D s command (gnus-draft-send-message). This command understands the process/prefix convention (see Section 8.1 [Process/Prefix], page 153). The D S command (gnus-draft-send-all-messages) will ship off all messages in the buffer.

If you have some messages that you wish not to send, you can use the D t (gnusdraft-toggle-sending) command to mark the message as unsendable. This is a toggling command.

5.8 Rejected Articles

Sometimes a news server will reject an article. Perhaps the server doesn’t like your face. Perhaps it just feels miserable. Perhaps there be demons. Perhaps you have included too much cited text. Perhaps the disk is full. Perhaps the server is down.

These situations are, of course, totally beyond the control of Gnus. (Gnus, of course, loves the way you look, always feels great, has angels fluttering around inside of it, doesn’t care about how much cited text you include, never runs full and never goes down.) So Gnus saves these articles until some later time when the server feels better.

The rejected articles will automatically be put in a special draft group (see Section 5.7 [Drafts], page 89). When the server comes back up again, you’d then typically enter that group and send all the articles off.

6 Select Methods

A foreign group is a group not read by the usual (or default) means. It could be, for instance, a group from a different nntp server, it could be a virtual group, or it could be your own personal mail group.

A foreign group (or any group, really) is specified by a name and a select method. To take the latter first, a select method is a list where the first element says what backend to use (e.g. nntp, nnspool, nnml) and the second element is the server name. There may be additional elements in the select method, where the value may have special meaning for the backend in question.

One could say that a select method defines a virtual server—so we do just that (see Section 6.1 [The Server Buffer], page 91).

The name of the group is the name the backend will recognize the group as.

For instance, the group ‘soc.motss’ on the nntp server ‘some.where.edu’ will have the name ‘soc.motss’ and select method (nntp "some.where.edu"). Gnus will call this group ‘nntp+some.where.edu:soc.motss’, even though the nntp backend just knows this group

as ‘soc.motss’.

The different methods all have their peculiarities, of course.

6.1 The Server Buffer

Traditionally, a server is a machine or a piece of software that one connects to, and then requests information from. Gnus does not connect directly to any real servers, but does all transactions through one backend or other. But that’s just putting one layer more between the actual media and Gnus, so we might just as well say that each backend represents a virtual server.

For instance, the nntp backend may be used to connect to several different actual nntp servers, or, perhaps, to many different ports on the same actual nntp server. You tell Gnus which backend to use, and what parameters to set by specifying a select method.

These select method specifications can sometimes become quite complicated—say, for instance, that you want to read from the nntp server ‘news.funet.fi’ on port number 13, which hangs if queried for nov headers and has a buggy select. Ahem. Anyways, if you had to specify that for each group that used this server, that would be too much work, so Gnus offers a way of naming select methods, which is what you do in the server buffer.

To enter the server buffer, use the ^(gnus-group-enter-server-mode) command in the group buffer.

gnus-server-mode-hook is run when creating the server buffer.

6.1.1 Server Buffer Format

You can change the look of the server buffer lines by changing the gnus-server-line-format variable. This is a format-like variable, with some simple extensions:

‘h’ How the news is fetched—the backend name.

‘n’ The name of this server.

‘w’ Where the news is to be fetched from—the address. ‘s’ The opened/closed/denied status of the server.

The mode line can also be customized by using the gnus-server-mode-line-format variable (see Section 8.4.2 [Mode Line Formatting], page 155). The following specs are understood: ‘S’ Server name. ‘M’ Server method.

Also see Section 8.4 [Formatting Variables], page 154.

6.1.2 Server Commands

a  Add a new server (gnus-server-add-server).
e  Edit a server (gnus-server-edit-server).
SPACE  Browse the current server (gnus-server-read-server).

q Return to the group buffer (gnus-server-exit).

k  Kill the current server (gnus-server-kill-server).

y Yank the previously killed server (gnus-server-yank-server).

c  Copy the current server (gnus-server-copy-server).
l  List all servers (gnus-server-list-servers).

s Request that the server scan its sources for new articles (gnus-server-scan-

server). This is mainly sensible with mail servers.

g Request that the server regenerate all its data structures (gnus-server-regenerate-server). This can be useful if you have a mail backend that has gotten out of synch.

6.1.3 Example Methods

Most select methods are pretty simple and self-explanatory:

     (nntp "news.funet.fi")

Reading directly from the spool is even simpler:

     (nnspool "")

As you can see, the first element in a select method is the name of the backend, and the second is the address, or name, if you will.

After these two elements, there may be an arbitrary number of (variable form) pairs. To go back to the first example—imagine that you want to read from port 15 on that

machine. This is what the select method should look like then:

     (nntp "news.funet.fi" (nntp-port-number 15))

You should read the documentation to each backend to find out what variables are relevant, but here’s an nnmh example:

nnmh is a mail backend that reads a spool-like structure. Say you have two structures that you wish to access: One is your private mail spool, and the other is a public one. Here’s the possible spec for your private mail:

     (nnmh "private" (nnmh-directory "~/private/mail/"))

(This server is then called ‘private’, but you may have guessed that.)

Here’s the method for a public spool:

     (nnmh "public"
           (nnmh-directory "/usr/information/spool/")
           (nnmh-get-new-mail nil))

If you are behind a firewall and only have access to the nntp server from the firewall machine, you can instruct Gnus to rlogin on the firewall machine and telnet from there to the nntp server. Doing this can be rather fiddly, but your virtual server definition should probably look something like this:

     (nntp "firewall"
           (nntp-address "the.firewall.machine")
           (nntp-open-connection-function nntp-open-rlogin)
           (nntp-end-of-line "\n")
           (nntp-rlogin-parameters
            ("telnet" "the.real.nntp.host" "nntp")))

If you want to use the wonderful ssh program to provide a compressed connection over the modem line, you could create a virtual server that would look something like this:

     (nntp "news"
            (nntp-address "copper.uio.no")
            (nntp-rlogin-program "ssh")
            (nntp-open-connection-function nntp-open-rlogin)
            (nntp-end-of-line "\n")
            (nntp-rlogin-parameters
             ("telnet" "news.uio.no" "nntp")))

This means that you have to have set up ssh-agent correctly to provide automatic authorization, of course. And to get a compressed connection, you have to have the ‘Compression’ option in the ssh ‘config’ file.

6.1.4 Creating a Virtual Server

If you’re saving lots of articles in the cache by using persistent articles, you may want to create a virtual server to read the cache.

First you need to add a new server. The a command does that. It would probably be best to use nnspool to read the cache. You could also use nnml or nnmh, though.

   Type a nnspool RET cache RET.

You should now have a brand new nnspool virtual server called ‘cache’. You now need to edit it to have the right definitions. Type e to edit the server. You’ll be entered into a buffer that will contain the following:

     (nnspool "cache")

Change that to:

     (nnspool "cache"
              (nnspool-spool-directory "~/News/cache/")
              (nnspool-nov-directory "~/News/cache/")
              (nnspool-active-file "~/News/cache/active"))

Type C-c C-c to return to the server buffer. If you now press RET over this virtual server, you should be entered into a browse buffer, and you should be able to enter any of the groups displayed.

6.1.5 Server Variables

One sticky point when defining variables (both on backends and in Emacs in general) is that some variables are typically initialized from other variables when the definition of the variables is being loaded. If you change the "base" variable after the variables have been loaded, you won’t change the "derived" variables.

This typically affects directory and file variables. For instance, nnml-directory is ‘~/Mail/’ by default, and all nnml directory variables are initialized from that variable, so nnml-active-file will be ‘~/Mail/active’. If you define a new virtual nnml server, it will not suffice to set just nnml-directory—you have to explicitly set all the file variables to be what you want them to be. For a complete list of variables for each backend, see each backend’s section later in this manual, but here’s an example nnml definition:

     (nnml "public"
           (nnml-directory "~/my-mail/")
           (nnml-active-file "~/my-mail/active")
           (nnml-newsgroups-file "~/my-mail/newsgroups"))

6.1.6 Servers and Methods

Wherever you would normally use a select method (e.g. gnus-secondary-select-method, in the group select method, when browsing a foreign server) you can use a virtual server name instead. This could potentially save lots of typing. And it’s nice all over.

6.1.7 Unavailable Servers

If a server seems to be unreachable, Gnus will mark that server as denied. That means that any subsequent attempt to make contact with that server will just be ignored. “It can’t be opened,” Gnus will tell you, without making the least effort to see whether that is actually the case or not.

That might seem quite naughty, but it does make sense most of the time. Let’s say you have 10 groups subscribed to on server ‘nephelococcygia.com’. This server is located somewhere quite far away from you and the machine is quite slow, so it takes 1 minute just to find out that it refuses connection to you today. If Gnus were to attempt to do that 10 times, you’d be quite annoyed, so Gnus won’t attempt to do that. Once it has gotten a single “connection refused”, it will regard that server as “down”.

So, what happens if the machine was only feeling unwell temporarily? How do you test to see whether the machine has come up again?

You jump to the server buffer (see Section 6.1 [The Server Buffer], page 91) and poke it with the following commands:

O Try to establish connection to the server on the current line (gnus-server-

          open-server).

C Close the connection (if any) to the server (gnus-server-close-server).

D Mark the current server as unreachable (gnus-server-deny-server).

M-o Open the connections to all servers in the buffer (gnus-server-open-all-

          servers).

M-c Close the connections to all servers in the buffer (gnus-server-close-all-

          servers).

R Remove all marks to whether Gnus was denied connection from any servers

          (gnus-server-remove-denials).

6.2 Getting News

A newsreader is normally used for reading news. Gnus currently provides only two methods of getting news—it can read from an nntp server, or it can read from a local spool.

6.2.1 nntp

Subscribing to a foreign group from an nntp server is rather easy. You just specify nntp as method and the address of the nntp server as the, uhm, address.

If the nntp server is located at a non-standard port, setting the third element of the select method to this port number should allow you to connect to the right port. You’ll have to edit the group info for that (see Section 2.9 [Foreign Groups], page 18).

The name of the foreign group can be the same as a native group. In fact, you can subscribe to the same group from as many different servers you feel like. There will be no name collisions.

The following variables can be used to create a virtual nntp server:

nntp-server-opened-hook

is run after a connection has been made. It can be used to send commands to the nntp server after it has been contacted. By default it sends the command

          MODE READER to the server with the nntp-send-mode-reader function. This

function should always be present in this hook.

nntp-authinfo-function

This function will be used to send ‘AUTHINFO’ to the nntp server. The default function is nntp-send-authinfo, which looks through your ‘~/.authinfo’ (or whatever you’ve set the nntp-authinfo-file variable to) for applicable entries. If none are found, it will prompt you for a login name and a password. The format of the ‘~/.authinfo’ file is (almost) the same as the ftp ‘~/.netrc’ file, which is defined in the ftp manual page, but here are the salient facts:

1. The file contains one or more line, each of which define one server.

2. Each line may contain an arbitrary number of token/value pairs. The valid tokens include ‘machine’, ‘login’, ‘password’, ‘default’ and ‘force’. (The latter is not a valid ‘.netrc’/ftp token, which is the only way the ‘.authinfo’ file format deviates from the ‘.netrc’ file format.)

Here’s an example file:

     machine news.uio.no login larsi password geheimnis
     machine nntp.ifi.uio.no login larsi force yes

The token/value pairs may appear in any order; ‘machine’ doesn’t have to be first, for instance. In this example, both login name and password have been supplied for the former server, while the latter has only the login name listed, and the user will be prompted for the password. The latter also has the ‘force’ tag, which means that the authinfo will be sent to the nntp server upon connection; the default (i.e., when there is not ‘force’ tag) is to not send authinfo to the nntp server until the nntp server asks for it. You can also add ‘default’ lines that will apply to all servers that don’t have matching ‘machine’ lines.

     default force yes

This will force sending ‘AUTHINFO’ commands to all servers not previously mentioned. Remember to not leave the ‘~/.authinfo’ file world-readable.

nntp-server-action-alist

This is a list of regexps to match on server types and actions to be taken when matches are made. For instance, if you want Gnus to beep every time you connect to innd, you could say something like:

     (setq nntp-server-action-alist
           '(("innd" (ding))))

You probably don’t want to do that, though. The default value is

     '(("nntpd 1\\.5\\.11t"
        (remove-hook 'nntp-server-opened-hook 'nntp-send-mode-reader)))

This ensures that Gnus doesn’t send the MODE READER command to nntpd 1.5.11t, since that command chokes that server, I’ve been told.

nntp-maximum-request

If the nntp server doesn’t support nov headers, this backend will collect headers by sending a series of head commands. To speed things up, the backend sends lots of these commands without waiting for reply, and then reads all the replies. This is controlled by the nntp-maximum-request variable, and is 400 by default. If your network is buggy, you should set this to 1.

nntp-connection-timeout

If you have lots of foreign nntp groups that you connect to regularly, you’re sure to have problems with nntp servers not responding properly, or being too loaded to reply within reasonable time. This is can lead to awkward problems, which can be helped somewhat by setting nntp-connection-timeout. This is an integer that says how many seconds the nntp backend should wait for a connection before giving up. If it is nil, which is the default, no timeouts are done.

nntp-server-hook

This hook is run as the last step when connecting to an nntp server.

nntp-open-connection-function

This function is used to connect to the remote system. Four pre-made functions are supplied:

nntp-open-network-stream

This is the default, and simply connects to some port or other on the remote system.

nntp-open-rlogin

Does an ‘rlogin’ on the remote system, and then does a ‘telnet’ to the nntp server available there. nntp-open-rlogin-related variables:

nntp-rlogin-program

Program used to log in on remote machines. The default is ‘rsh’, but ‘ssh’ is a popular alternative.

nntp-rlogin-parameters

This list will be used as the parameter list given to rsh.

nntp-rlogin-user-name

User name on the remote system.

nntp-open-telnet

Does a ‘telnet’ to the remote system and then another ‘telnet’ to get to the nntp server. nntp-open-telnet-related variables:

nntp-telnet-command

Command used to start telnet.

nntp-telnet-switches

List of strings to be used as the switches to the telnet command.

nntp-telnet-user-name

User name for log in on the remote system.

nntp-telnet-passwd

Password to use when logging in.

nntp-telnet-parameters

A list of strings executed as a command after logging

          in via telnet.

          nntp-telnet-shell-prompt

Regexp matching the shell prompt on the remote ma-

                    chine. The default is ‘bash\\|\$ *\r?$\\|> *\r?’.
          nntp-open-telnet-envuser

If non-nil, the telnet session (client and server both) will support the ENVIRON option and not prompt for login name. This works for Solaris telnet, for instance.

nntp-open-ssl-stream

Opens a connection to a server over a secure channel. To use this you must have SSLay installed

          (‘ftp://ftp.psy.uq.oz.au/pub/Crypto/SSL’,  and  you

also need ‘ssl.el’ (from the W3 distributeion, for instance). You then define a server as follows:

               ;; Type `C-c C-c' after you've finished editing.
               ;;
               ;; "snews" is port 563 and is predefined in our /etc/services
               ;;
               (nntp "snews.bar.com"
                     (nntp-open-connection-function nntp-open-ssl-stream)
                     (nntp-port-number "snews")
                     (nntp-address "snews.bar.com"))
nntp-end-of-line

String to use as end-of-line marker when talking to the nntp server. This is ‘\r\n’ by default, but should be ‘\n’ when using rlogin to talk to the server.

nntp-rlogin-user-name

User name on the remote system when using the rlogin connect function.

nntp-address

The address of the remote system running the nntp server.

nntp-port-number

Port number to connect to when using the nntp-open-network-stream connect function.

nntp-buggy-select

Set this to non-nil if your select routine is buggy.

nntp-nov-is-evil

If the nntp server does not support nov, you could set this variable to t, but nntp usually checks automatically whether nov can be used.

nntp-xover-commands

List of strings used as commands to fetch nov lines from a server. The default value of this variable is ("XOVER" "XOVERVIEW").

nntp-nov-gap

nntp normally sends just one big request for nov lines to the server. The server responds with one huge list of lines. However, if you have read articles 2-5000 in the group, and only want to read article 1 and 5001, that means that nntp will fetch 4999 nov lines that you will not need. This variable says how big a gap between two consecutive articles is allowed to be before the XOVER request is split into several request. Note that if your network is fast, setting this variable to a really small number means that fetching will probably be slower. If this variable is nil, nntp will never split requests. The default is 5.

nntp-prepare-server-hook

A hook run before attempting to connect to an nntp server.

nntp-warn-about-losing-connection

If this variable is non-nil, some noise will be made when a server closes connection.

nntp-record-commands

If non-nil, nntp will log all commands it sends to the nntp server (along with a timestep) in the ‘*nntp-log*’ buffer. This is useful if you are debugging a Gnus/nntp connection that doesn’t seem to work.

6.2.2 News Spool

Subscribing to a foreign group from the local spool is extremely easy, and might be useful, for instance, to speed up reading groups that contain very big

articles—‘alt.binaries.pictures.furniture’, for instance.

Anyways, you just specify nnspool as the method and "" (or anything else) as the address.

If you have access to a local spool, you should probably use that as the native select method (see Section 1.1 [Finding the News], page 3). It is normally faster than using an nntp select method, but might not be. It depends. You just have to try to find out what’s best at your site.

nnspool-inews-program

Program used to post an article.

nnspool-inews-switches

Parameters given to the inews program when posting an article.

nnspool-spool-directory

Where nnspool looks for the articles. This is normally ‘/usr/spool/news/’.

nnspool-nov-directory

Where nnspool will look for nov files. This is normally

‘/usr/spool/news/over.view/’.
nnspool-lib-dir

Where the news lib dir is (‘/usr/lib/news/’ by default).

nnspool-active-file

The path to the active file.

nnspool-newsgroups-file

The path to the group descriptions file.

nnspool-history-file

The path to the news history file.

nnspool-active-times-file

The path to the active date file.

nnspool-nov-is-evil

If non-nil, nnspool won’t try to use any nov files that it finds.

nnspool-sift-nov-with-sed

If non-nil, which is the default, use sed to get the relevant portion from the overview file. If nil, nnspool will load the entire file into a buffer and process it there.

6.3 Getting Mail

Reading mail with a newsreader—isn’t that just plain WeIrD? But of course.

6.3.1 Getting Started Reading Mail

It’s quite easy to use Gnus to read your new mail. You just plonk the mail backend of your choice into gnus-secondary-select-methods, and things will happen automatically.

For instance, if you want to use nnml (which is a "one file per mail" backend), you could put the following in your ‘.gnus’ file:

     (setq gnus-secondary-select-methods
           '((nnml "private")))

Now, the next time you start Gnus, this backend will be queried for new articles, and it will move all the messages in your spool file to its directory, which is ~/Mail/ by default. The new group that will be created (‘mail.misc’) will be subscribed, and you can read it like any other group.

You will probably want to split the mail into several groups, though:

     (setq nnmail-split-methods
           '(("junk" "^From:.*Lars Ingebrigtsen")
             ("crazy" "^Subject:.*die\\|^Organization:.*flabby")
             ("other" "")))

This will result in three new nnml mail groups being created: ‘nnml:junk’, ‘nnml:crazy’, and ‘nnml:other’. All the mail that doesn’t fit into the first two groups will be placed in the last group.

This should be sufficient for reading mail with Gnus. You might want to give the other sections in this part of the manual a perusal, though. Especially see Section 6.3.11 [Choosing a Mail Backend], page 110 and see Section 6.3.7 [Expiring Mail], page 106.

6.3.2 Splitting Mail

The nnmail-split-methods variable says how the incoming mail is to be split into groups.

     (setq nnmail-split-methods
       '(("mail.junk" "^From:.*Lars Ingebrigtsen")
         ("mail.crazy" "^Subject:.*die\\|^Organization:.*flabby")
         ("mail.other" "")))

This variable is a list of lists, where the first element of each of these lists is the name of the mail group (they do not have to be called something beginning with ‘mail’, by the way), and the second element is a regular expression used on the header of each mail to determine if it belongs in this mail group. The first string may contain ‘\\1’ forms, like the ones used by replace-match to insert sub-expressions from the matched text. For instance:

     ("list.\\1" "From:.*\\(.*\\)-list@majordomo.com")

The second element can also be a function. In that case, it will be called narrowed to the headers with the first element of the rule as the argument. It should return a non-nil value if it thinks that the mail belongs in that group.

The last of these groups should always be a general one, and the regular expression should always be ‘’ so that it matches any mails that haven’t been matched by any of the other regexps. (These rules are processed from the beginning of the alist toward the end. The first rule to make a match will "win", unless you have crossposting enabled. In that case, all matching rules will "win".)

If you like to tinker with this yourself, you can set this variable to a function of your choice. This function will be called without any arguments in a buffer narrowed to the headers of an incoming mail message. The function should return a list of group names that it thinks should carry this mail message.

Note that the mail backends are free to maul the poor, innocent, incoming headers all they want to. They all add Lines headers; some add X-Gnus-Group headers; most rename the Unix mbox From<SPACE> line to something else.

The mail backends all support cross-posting. If several regexps match, the mail will be “cross-posted” to all those groups. nnmail-crosspost says whether to use this mechanism or not. Note that no articles are crossposted to the general (‘’) group.

nnmh and nnml makes crossposts by creating hard links to the crossposted articles. However, not all file systems support hard links. If that’s the case for you, set nnmail-

crosspost-link-function to copy-file. (This variable is add-name-to-file by default.)

If you wish to see where the previous mail split put the messages, you can use the M-x

nnmail-split-history command.

Gnus gives you all the opportunity you could possibly want for shooting yourself in the foot. Let’s say you create a group that will contain all the mail you get from your boss. And then you accidentally unsubscribe from the group. Gnus will still put all the mail from your boss in the unsubscribed group, and so, when your boss mails you “Have that report ready by Monday or you’re fired!”, you’ll never see it and, come Tuesday, you’ll still believe that you’re gainfully employed while you really should be out collecting empty bottles to save up for next month’s rent money.

6.3.3 Mail Backend Variables

These variables are (for the most part) pertinent to all the various mail backends.

nnmail-read-incoming-hook

The mail backends all call this hook after reading new mail. You can use this hook to notify any mail watch programs, if you want to.

nnmail-spool-file

The backends will look for new mail in this file. If this variable is nil, the mail backends will never attempt to fetch mail by themselves. If you are using a POP mail server and your name is ‘larsi’, you should set this variable to ‘po:larsi’. If your name is not ‘larsi’, you should probably modify that slightly, but you may have guessed that already, you smart & handsome devil! You can also set this variable to pop, and Gnus will try to figure out the POP mail string by itself. In any case, Gnus will call movemail which will contact the POP server named in the MAILHOST environment variable. If the POP server needs a password, you can either set nnmail-pop-password-required to t and be prompted for the password, or set nnmail-pop-password to the password itself.

nnmail-spool-file can also be a list of mailboxes.

Your Emacs has to have been configured with ‘--with-pop’ before compilation. This is the default, but some installations have it switched off.

When you use a mail backend, Gnus will slurp all your mail from your inbox and plonk it down in your home directory. Gnus doesn’t move any mail if you’re not using a mail backend—you have to do a lot of magic invocations first. At the time when you have finished drawing the pentagram, lightened the candles, and sacrificed the goat, you really shouldn’t be too surprised when Gnus moves your mail.

nnmail-use-procmail

If non-nil, the mail backends will look in nnmail-procmail-directory for incoming mail. All the files in that directory that have names ending in nnmailprocmail-suffix will be considered incoming mailboxes, and will be searched for new mail.

nnmail-crash-box

When a mail backend reads a spool file, mail is first moved to this file, which is ‘~/.gnus-crash-box’ by default. If this file already exists, it will always be read (and incorporated) before any other spool files.

nnmail-prepare-incoming-hook

This is run in a buffer that holds all the new incoming mail, and can be used for, well, anything, really.

nnmail-split-hook

Hook run in the buffer where the mail headers of each message is kept just before the splitting based on these headers is done. The hook is free to modify the buffer contents in any way it sees fit—the buffer is discarded after the splitting has been done, and no changes performed in the buffer will show up in any files. gnus-article-decode-rfc1522 is one likely function to add to this hook.

nnmail-pre-get-new-mail-hook
nnmail-post-get-new-mail-hook

These are two useful hooks executed when treating new incoming mail— nnmail-pre-get-new-mail-hook (is called just before starting to handle the new mail) and nnmail-post-get-new-mail-hook (is called when the mail handling is done). Here’s and example of using these two hooks to change the default file modes the new mail files get:

     (add-hook 'gnus-pre-get-new-mail-hook
               (lambda () (set-default-file-modes 511)))
     (add-hook 'gnus-post-get-new-mail-hook
               (lambda () (set-default-file-modes 551)))
nnmail-tmp-directory

This variable says where to move incoming mail to – while processing it. This is usually done in the same directory that the mail backend inhabits (e.g., ‘~/Mail/’), but if this variable is non-nil, it will be used instead.

nnmail-movemail-program

This program is executed to move mail from the user’s inbox to her home directory. The default is ‘movemail’. This can also be a function. In that case, the function will be called with two parameters – the name of the inbox, and the file to be moved to.

nnmail-delete-incoming

If non-nil, the mail backends will delete the temporary incoming file after splitting mail into the proper groups. This is t by default. (No Gnus release since (ding) Gnus 0.10 (or something like that) have lost mail, I think, but that’s not the point. (Except certain versions of Red Gnus.)) By not deleting the Incoming* files, one can be sure not to lose mail – if Gnus totally whacks out, one can always recover what was lost. You may delete the ‘Incoming*’ files at will.

nnmail-use-long-file-names

If non-nil, the mail backends will use long file and directory names. Groups like ‘mail.misc’ will end up in directories (assuming use of nnml backend) or files (assuming use of nnfolder backend) like ‘mail.misc’. If it is nil, the same group will end up in ‘mail/misc’.

nnmail-delete-file-function

Function called to delete files. It is delete-file by default.

nnmail-cache-accepted-message-ids

If non-nil, put the Message-IDs of articles imported into the backend (via Gcc, for instance) into the mail duplication discovery cache. The default is nil.

6.3.4 Fancy Mail Splitting

If the rather simple, standard method for specifying how to split mail doesn’t allow you

to do what you want, you can set nnmail-split-methods to nnmail-split-fancy. Then

you can play with the nnmail-split-fancy variable.

Let’s look at an example value of this variable first:

 ;; Messages from the mailer daemon are not crossposted to any of
 ;; the ordinary groups.  Warnings are put in a separate group
 ;; from real errors.
 (| ("from" mail (| ("subject" "warn.*" "mail.warning")
                    "mail.misc"))
    ;; Non-error messages are crossposted to all relevant
    ;; groups, but we don't crosspost between the group for the
    ;; (ding) list and the group for other (ding) related mail.
    (& (| (any "ding@ifi\\.uio\\.no" "ding.list")
          ("subject" "ding" "ding.misc"))
       ;; Other mailing lists...
       (any "procmail@informatik\\.rwth-aachen\\.de" "procmail.list")
       (any "SmartList@informatik\\.rwth-aachen\\.de" "SmartList.list")
       ;; People...
       (any "larsi@ifi\\.uio\\.no" "people.Lars_Magne_Ingebrigtsen"))
    ;; Unmatched mail goes to the catch all group.
    "misc.misc")

This variable has the format of a split. A split is a (possibly) recursive structure where each split may contain other splits. Here are the five possible split syntaxes:

1. ‘group’: If the split is a string, that will be taken as a group name. Normal regexp match expansion will be done. See below for examples.

2. (FIELD VALUE SPLIT): If the split is a list, the first element of which is a string, then store the message as specified by SPLIT, if header FIELD (a regexp) contains VALUE (also a regexp).

3. (| SPLIT...): If the split is a list, and the first element is | (vertical bar), then process each SPLIT until one of them matches. A SPLIT is said to match if it will cause the mail message to be stored in one or more groups.

4. (& SPLIT...): If the split is a list, and the first element is &, then process all SPLITs in the list.

5. junk: If the split is the symbol junk, then don’t save this message. Use with extreme caution.

6. (: function arg1 arg2 ...): If the split is a list, and the first element is :, then the second element will be called as a function with args given as arguments. The function should return a SPLIT.

7. nil: If the split is nil, it is ignored.

In these splits, FIELD must match a complete field name. VALUE must match a complete word according to the fundamental mode syntax table. You can use .* in the regexps to match partial field names or words. In other words, all VALUE’s are wrapped in ‘\<’ and ‘\>’ pairs.

FIELD and VALUE can also be lisp symbols, in that case they are expanded as specified by the variable nnmail-split-abbrev-alist. This is an alist of cons cells, where the car of a cell contains the key, and the cdr contains the associated value.

nnmail-split-fancy-syntax-table is the syntax table in effect when all this splitting is performed.

If you want to have Gnus create groups dynamically based on some information in the headers (i.e., do replace-match-like substitutions in the group names), you can say things like:

     (any "debian-\\b\\(\\w+\\)@lists.debian.org" "mail.debian.\\1")

If the string contains the element ‘\&’, then the previously matched string will be substituted. Similarly, the elements ‘\\1’ up to ‘\\9’ will be substituted with the text matched by the groupings 1 through 9.

6.3.5 Mail and Procmail

Many people use procmail (or some other mail filter program or external delivery agent—slocal, elm, etc) to split incoming mail into groups. If you do that, you should set nnmail-spool-file to procmail to ensure that the mail backends never ever try to fetch mail by themselves.

If you have a combined procmail/POP/mailbox setup, you can do something like the following:

     (setq nnmail-use-procmail t)
     (setq nnmail-spool-file
           '("/usr/spool/mail/my-name" "po:my-name"))

This also means that you probably don’t want to set nnmail-split-methods either, which has some, perhaps, unexpected side effects.

When a mail backend is queried for what groups it carries, it replies with the contents of that variable, along with any groups it has figured out that it carries by other means. None of the backends, except nnmh, actually go out to the disk and check what groups actually exist. (It’s not trivial to distinguish between what the user thinks is a basis for a newsgroup and what is just a plain old file or directory.)

This means that you have to tell Gnus (and the backends) by hand what groups exist.

Let’s take the nnmh backend as an example:

The folders are located in nnmh-directory, say, ‘~/Mail/’. There are three folders,

‘foo’, ‘bar’ and ‘mail.baz’.

Go to the group buffer and type G m. When prompted, answer ‘foo’ for the name and ‘nnmh’ for the method. Repeat twice for the two other groups, ‘bar’ and ‘mail.baz’. Be sure to include all your mail groups.

That’s it. You are now set to read your mail. An active file for this method will be created automatically.

If you use nnfolder or any other backend that store more than a single article in each file, you should never have procmail add mails to the file that Gnus sees. Instead, procmail should put all incoming mail in nnmail-procmail-directory. To arrive at the file name to put the incoming mail in, append nnmail-procmail-suffix to the group name. The mail backends will read the mail from these files.

When Gnus reads a file called ‘mail.misc.spool’, this mail will be put in the mail.misc, as one would expect. However, if you want Gnus to split the mail the normal way, you could

set nnmail-resplit-incoming to t.

If you use procmail to split things directly into an nnmh directory (which you shouldn’t do), you should set nnmail-keep-last-article to non-nil to prevent Gnus from ever expiring the final article (i.e., the article with the highest article number) in a mail newsgroup. This is quite, quite important.

Here’s an example setup: The incoming spools are located in ‘~/incoming/’ and have ‘""’ as suffixes (i.e., the incoming spool files have the same names as the equivalent groups). The nnfolder backend is to be used as the mail interface, and the nnfolder directory is

‘~/fMail/’.
     (setq nnfolder-directory "~/fMail/")
     (setq nnmail-spool-file 'procmail)
     (setq nnmail-procmail-directory "~/incoming/")
     (setq gnus-secondary-select-methods '((nnfolder "")))
     (setq nnmail-procmail-suffix "")

6.3.6 Incorporating Old Mail

Most people have lots of old mail stored in various file formats. If you have set up Gnus to read mail using one of the spiffy Gnus mail backends, you’ll probably wish to have that old mail incorporated into your mail groups.

Doing so can be quite easy.

To take an example: You’re reading mail using nnml (see Section 6.3.11.3 [Mail Spool], page 111), and have set nnmail-split-methods to a satisfactory value (see Section 6.3.2 [Splitting Mail], page 100). You have an old Unix mbox file filled with important, but old, mail. You want to move it into your nnml groups.

Here’s how:

1. Go to the group buffer.

2. Type ‘G f’ and give the path to the mbox file when prompted to create an nndoc group from the mbox file (see Section 2.9 [Foreign Groups], page 18).

3. Type ‘SPACE’ to enter the newly created group.

4. Type ‘M P b’ to process-mark all articles in this group’s buffer (see Section 3.7.5 [Setting Process Marks], page 44).

5. Type ‘B r’ to respool all the process-marked articles, and answer ‘nnml’ when prompted (see Section 3.23 [Mail Group Commands], page 75).

All the mail messages in the mbox file will now also be spread out over all your nnml groups. Try entering them and check whether things have gone without a glitch. If things look ok, you may consider deleting the mbox file, but I wouldn’t do that unless I was absolutely sure that all the mail has ended up where it should be.

Respooling is also a handy thing to do if you’re switching from one mail backend to another. Just respool all the mail in the old mail groups using the new mail backend.

6.3.7 Expiring Mail

Traditional mail readers have a tendency to remove mail articles when you mark them as read, in some way. Gnus takes a fundamentally different approach to mail reading.

Gnus basically considers mail just to be news that has been received in a rather peculiar manner. It does not think that it has the power to actually change the mail, or delete any mail messages. If you enter a mail group, and mark articles as “read”, or kill them in some other fashion, the mail articles will still exist on the system. I repeat: Gnus will not delete your old, read mail. Unless you ask it to, of course.

To make Gnus get rid of your unwanted mail, you have to mark the articles as expirable. This does not mean that the articles will disappear right away, however. In general, a mail article will be deleted from your system if, 1) it is marked as expirable, AND 2) it is more than one week old. If you do not mark an article as expirable, it will remain on your system until hell freezes over. This bears repeating one more time, with some spurious capitalizations: IF you do NOT mark articles as EXPIRABLE, Gnus will NEVER delete those ARTICLES.

You do not have to mark articles as expirable by hand. Groups that match the regular expression gnus-auto-expirable-newsgroups will have all articles that you read marked as expirable automatically. All articles marked as expirable have an ‘E’ in the first column in the summary buffer.

By default, if you have auto expiry switched on, Gnus will mark all the articles you read as expirable, no matter if they were read or unread before. To avoid having articles marked as read marked as expirable automatically, you can put something like the following in your ‘.gnus’ file:

     (remove-hook 'gnus-mark-article-hook
                  'gnus-summary-mark-read-and-unread-as-read)
     (add-hook 'gnus-mark-article-hook 'gnus-summary-mark-unread-as-read)

Note that making a group auto-expirable doesn’t mean that all read articles are expired— only the articles marked as expirable will be expired. Also note that using the d command won’t make groups expirable—only semi-automatic marking of articles as read will mark the articles as expirable in auto-expirable groups.

Let’s say you subscribe to a couple of mailing lists, and you want the articles you have read to disappear after a while:

     (setq gnus-auto-expirable-newsgroups
           "mail.nonsense-list\\|mail.nice-list")

Another way to have auto-expiry happen is to have the element auto-expire in the group parameters of the group.

If you use adaptive scoring (see Section 7.6 [Adaptive Scoring], page 140) and autoexpiring, you’ll have problems. Auto-expiring and adaptive scoring don’t really mix very well.

The nnmail-expiry-wait variable supplies the default time an expirable article has to live. Gnus starts counting days from when the message arrived, not from when it was sent. The default is seven days.

Gnus also supplies a function that lets you fine-tune how long articles are to live, based on what group they are in. Let’s say you want to have one month expiry period in the ‘mail.private’ group, a one day expiry period in the ‘mail.junk’ group, and a six day expiry period everywhere else:

     (setq nnmail-expiry-wait-function

           (lambda (group)
            (cond ((string= group "mail.private")
                    31)
                  ((string= group "mail.junk")
                    1)
                  ((string= group "important")
                    'never)
                  (t
                    6))))

The group names this function is fed are “unadorned” group names—no ‘nnml:’ prefixes and the like.

   The nnmail-expiry-wait variable and nnmail-expiry-wait-function function can

either be a number (not necessarily an integer) or one of the symbols immediate or never.

You can also use the expiry-wait group parameter to selectively change the expiry period (see Section 2.10 [Group Parameters], page 19).

If nnmail-keep-last-article is non-nil, Gnus will never expire the final article in a mail newsgroup. This is to make life easier for procmail users.

By the way: That line up there, about Gnus never expiring non-expirable articles, is a lie. If you put total-expire in the group parameters, articles will not be marked as expirable, but all read articles will be put through the expiry process. Use with extreme caution. Even more dangerous is the gnus-total-expirable-newsgroups variable. All groups that match this regexp will have all read articles put through the expiry process, which means that all old mail articles in the groups in question will be deleted after a while. Use with extreme caution, and don’t come crying to me when you discover that the regexp you used matched the wrong group and all your important mail has disappeared. Be a man! Or a woman! Whatever you feel more comfortable with! So there!

Most people make most of their mail groups total-expirable, though.

6.3.8 Washing Mail

Mailers and list servers are notorious for doing all sorts of really, really stupid things with mail. “Hey, RFC822 doesn’t explicitly prohibit us from adding the string wE aRe ElItE!!!!!1!! to the end of all lines passing through our server, so let’s do that!!!!1!” Yes, but RFC822 wasn’t designed to be read by morons. Things that were considered to be self-evident were not discussed. So. Here we are.

Case in point: The German version of Microsoft Exchange adds ‘AW: ’ to the subjects of replies instead of ‘Re: ’. I could pretend to be shocked and dismayed by this, but I haven’t got the energy. It is to laugh.

Gnus provides a plethora of functions for washing articles while displaying them, but it might be nicer to do the filtering before storing the mail to disc. For that purpose, we have three hooks and various functions that can be put in these hooks.

nnmail-prepare-incoming-hook

This hook is called before doing anything with the mail and is meant for grand, sweeping gestures. Functions to be used include:

          nnheader-ms-strip-cr

Remove trailing carriage returns from each line. This is default on Emacs running on MS machines.

nnmail-prepare-incoming-header-hook

This hook is called narrowed to each header. It can be used when cleaning up the headers. Functions that can be used include:

          nnmail-remove-leading-whitespace

Clear leading white space that “helpful” listservs have added to the headers to make them look nice. Aaah.

          nnmail-remove-list-identifiers

Some list servers add an identifier—for example, ‘(idm)’—to the beginning of all Subject headers. I’m sure that’s nice for people who use stone age mail readers. This function will remove strings that match the nnmail-list-identifiers regexp, which can also be a list of regexp. For instance, if you want to remove the ‘(idm)’ and the ‘nagnagnag’ identifiers:

                         (setq nnmail-list-identifiers
                               '("(idm)" "nagnagnag"))
          nnmail-remove-tabs

Translate all ‘TAB’ characters into ‘SPACE’ characters.

nnmail-prepare-incoming-message-hook

This hook is called narrowed to each message. Functions to be used include:

          article-de-quoted-unreadable

Decode Quoted Readable encoding.

6.3.9 Duplicates

If you are a member of a couple of mailing lists, you will sometimes receive two copies of the same mail. This can be quite annoying, so nnmail checks for and treats any duplicates it might find. To do this, it keeps a cache of old Message-IDs— nnmail-message-id-cache-file, which is ‘~/.nnmail-cache’ by default. The approximate maximum number of

Message-IDs stored there is controlled by the nnmail-message-id-cache-length variable,

which is 1000 by default. (So 1000 Message-IDs will be stored.) If all this sounds scary to you, you can set nnmail-treat-duplicates to warn (which is what it is by default), and nnmail won’t delete duplicate mails. Instead it will insert a warning into the head of the mail saying that it thinks that this is a duplicate of a different message.

This variable can also be a function. If that’s the case, the function will be called from a buffer narrowed to the message in question with the Message-ID as a parameter. The function must return either nil, warn, or delete.

You can turn this feature off completely by setting the variable to nil.

If you want all the duplicate mails to be put into a special duplicates group, you could do that using the normal mail split methods:

     (setq nnmail-split-fancy
           '(| ;; Messages duplicates go to a separate group.
               ("gnus-warning" "duplication of message" "duplicate")
               ;; Message from daemons, postmaster, and the like to another.
               (any mail "mail.misc")
               ;; Other rules.
               [ ... ] ))

Or something like:

     (setq nnmail-split-methods
           '(("duplicates" "^Gnus-Warning:")
             ;; Other rules.
             [...]))

Here’s a neat feature: If you know that the recipient reads her mail with Gnus, and that she has nnmail-treat-duplicates set to delete, you can send her as many insults as you like, just by using a Message-ID of a mail that you know that she’s already received. Think of all the fun! She’ll never see any of it! Whee!

6.3.10 Not Reading Mail

If you start using any of the mail backends, they have the annoying habit of assuming that you want to read mail with them. This might not be unreasonable, but it might not be what you want.

If you set nnmail-spool-file to nil, none of the backends will ever attempt to read incoming mail, which should help.

This might be too much, if, for instance, you are reading mail quite happily with nnml and just want to peek at some old rmail file you have stashed away with nnbabyl. All backends have variables called backend-get-new-mail. If you want to disable the nnbabyl mail reading, you edit the virtual server for the group to have a setting where nnbabyl-

get-new-mail to nil.

All the mail backends will call nn*-prepare-save-mail-hook narrowed to the article to be saved before saving it when reading incoming mail.

6.3.11 Choosing a Mail Backend

Gnus will read the mail spool when you activate a mail group. The mail file is first copied to your home directory. What happens after that depends on what format you want to store your mail in.

6.3.11.1 Unix Mail Box

The nnmbox backend will use the standard Un*x mbox file to store mail. nnmbox will add extra headers to each mail article to say which group it belongs in.

Virtual server settings:

nnmbox-mbox-file

The name of the mail box in the user’s home directory.

nnmbox-active-file

The name of the active file for the mail box.

nnmbox-get-new-mail

If non-nil, nnmbox will read incoming mail and split it into groups.

6.3.11.2 Rmail Babyl

The nnbabyl backend will use a babyl mail box (aka. rmail mbox) to store mail. nnbabyl will add extra headers to each mail article to say which group it belongs in.

Virtual server settings:

nnbabyl-mbox-file

The name of the rmail mbox file.

nnbabyl-active-file

The name of the active file for the rmail box.

nnbabyl-get-new-mail

If non-nil, nnbabyl will read incoming mail.

6.3.11.3 Mail Spool

The nnml spool mail format isn’t compatible with any other known format. It should be used with some caution.

If you use this backend, Gnus will split all incoming mail into files, one file for each mail, and put the articles into the corresponding directories under the directory specified by the nnml-directory variable. The default value is ‘~/Mail/’.

You do not have to create any directories beforehand; Gnus will take care of all that. If you have a strict limit as to how many files you are allowed to store in your account,

you should not use this backend. As each mail gets its own file, you might very well occupy thousands of inodes within a few weeks. If this is no problem for you, and it isn’t a problem for you having your friendly systems administrator walking around, madly, shouting “Who is eating all my inodes?! Who? Who!?!”, then you should know that this is probably the fastest format to use. You do not have to trudge through a big mbox file just to read your new mail.

nnml is probably the slowest backend when it comes to article splitting. It has to create lots of files, and it also generates nov databases for the incoming mails. This makes it the fastest backend when it comes to reading mail.

Virtual server settings:

nnml-directory

All nnml directories will be placed under this directory.

nnml-active-file

The active file for the nnml server.

nnml-newsgroups-file

The nnml group descriptions file. See Section 10.6.9.2 [Newsgroups File Format], page 227.

nnml-get-new-mail

If non-nil, nnml will read incoming mail.

nnml-nov-is-evil

If non-nil, this backend will ignore any nov files.

nnml-nov-file-name

The name of the nov files. The default is ‘.overview’.

nnml-prepare-save-mail-hook

Hook run narrowed to an article before saving.

If your nnml groups and nov files get totally out of whack, you can do a complete update by typing M-x nnml-generate-nov-databases. This command will trawl through the entire nnml hierarchy, looking at each and every article, so it might take a while to complete. A better interface to this functionality can be found in the server buffer (see Section 6.1.2 [Server Commands], page 92).

6.3.11.4 MH Spool

nnmh is just like nnml, except that is doesn’t generate nov databases and it doesn’t keep an active file. This makes nnmh a much slower backend than nnml, but it also makes it easier to write procmail scripts for.

Virtual server settings:

nnmh-directory

All nnmh directories will be located under this directory.

nnmh-get-new-mail

If non-nil, nnmh will read incoming mail.

nnmh-be-safe

If non-nil, nnmh will go to ridiculous lengths to make sure that the articles in the folder are actually what Gnus thinks they are. It will check date stamps and stat everything in sight, so setting this to t will mean a serious slow-down. If you never use anything but Gnus to read the nnmh articles, you do not have to set this variable to t.

6.3.11.5 Mail Folders

nnfolder is a backend for storing each mail group in a separate file. Each file is in the standard Un*x mbox format. nnfolder will add extra headers to keep track of article numbers and arrival dates.

Virtual server settings:

nnfolder-directory

All the nnfolder mail boxes will be stored under this directory.

nnfolder-active-file

The name of the active file.

nnfolder-newsgroups-file

The name of the group descriptions file. See Section 10.6.9.2 [Newsgroups File Format], page 227.

nnfolder-get-new-mail

If non-nil, nnfolder will read incoming mail.

nnfolder-save-buffer-hook

Hook run before saving the folders. Note that Emacs does the normal backup renaming of files even with the nnfolder buffers. If you wish to switch this off, you could say something like the following in your ‘.emacs’ file:

               (defun turn-off-backup ()
                 (set (make-local-variable 'backup-inhibited) t))
               (add-hook 'nnfolder-save-buffer-hook 'turn-off-backup)

If you have lots of nnfolder-like files you’d like to read with nnfolder, you can use the

M-x nnfolder-generate-active-file command to make nnfolder aware of all likely files
in nnfolder-directory.

6.4 Other Sources

Gnus can do more than just read news or mail. The methods described below allow Gnus to view directories and files as if they were newsgroups.

6.4.1 Directory Groups

If you have a directory that has lots of articles in separate files in it, you might treat it as a newsgroup. The files have to have numerical names, of course.

This might be an opportune moment to mention ange-ftp (and its successor efs), that most wonderful of all wonderful Emacs packages. When I wrote nndir, I didn’t think much about it—a backend to read directories. Big deal.

ange-ftp changes that picture dramatically. For instance, if you enter the ange-ftp

file name ‘/ftp.hpc.uh.edu:/pub/emacs/ding-list/’ as the directory name, ange-ftp or

efs will actually allow you to read this directory over at ‘sina’ as a newsgroup. Distributed news ahoy!

nndir will use nov files if they are present.

nndir is a “read-only” backend—you can’t delete or expire articles with this method. You can use nnmh or nnml for whatever you use nndir for, so you could switch to any of those methods if you feel the need to have a non-read-only nndir.

6.4.2 Anything Groups

From the nndir backend (which reads a single spool-like directory), it’s just a hop and a skip to nneething, which pretends that any arbitrary directory is a newsgroup. Strange, but true.

When nneething is presented with a directory, it will scan this directory and assign article numbers to each file. When you enter such a group, nneething must create “headers” that Gnus can use. After all, Gnus is a newsreader, in case you’re forgetting. nneething does this in a two-step process. First, it snoops each file in question. If the file looks like an article (i.e., the first few lines look like headers), it will use this as the head. If this is just some arbitrary file without a head (e.g. a C source file), nneething will cobble up a header out of thin air. It will use file ownership, name and date and do whatever it can with these elements.

All this should happen automatically for you, and you will be presented with something that looks very much like a newsgroup. Totally like a newsgroup, to be precise. If you select an article, it will be displayed in the article buffer, just as usual.

If you select a line that represents a directory, Gnus will pop you into a new summary buffer for this nneething group. And so on. You can traverse the entire disk this way, if you feel like, but remember that Gnus is not dired, really, and does not intend to be, either.

There are two overall modes to this action—ephemeral or solid. When doing the ephemeral thing (i.e., G D from the group buffer), Gnus will not store information on what files you have read, and what files are new, and so on. If you create a solid nneething group the normal way with G m, Gnus will store a mapping table between article numbers and file names, and you can treat this group like any other groups. When you activate a solid nneething group, you will be told how many unread articles it contains, etc., etc.

Some variables:

nneething-map-file-directory

All the mapping files for solid nneething groups will be stored in this directory, which defaults to ‘~/.nneething/’.

nneething-exclude-files

All files that match this regexp will be ignored. Nice to use to exclude auto-save files and the like, which is what it does by default.

nneething-map-file

Name of the map files.

6.4.3 Document Groups

nndoc is a cute little thing that will let you read a single file as a newsgroup. Several files types are supported:

babyl The babyl (rmail) mail box.

mbox The standard Unix mbox file.

mmdf The MMDF mail box format.

news Several news articles appended into a file.

rnews The rnews batch transport format.

forward Forwarded articles.

mime-parts

MIME multipart messages, besides digests.

mime-digest

MIME (RFC 1341) digest format.

standard-digest

The standard (RFC 1153) digest format.

slack-digest

Non-standard digest format—matches most things, but does it badly.

You can also use the special “file type” guess, which means that nndoc will try to guess what file type it is looking at. digest means that nndoc should guess what digest type the file is.

nndoc will not try to change the file or insert any extra headers into it—it will simply, like, let you use the file as the basis for a group. And that’s it.

If you have some old archived articles that you want to insert into your new & spiffy Gnus mail backend, nndoc can probably help you with that. Say you have an old ‘RMAIL’ file with mail that you now want to split into your new nnml groups. You look at that file using nndoc (using the G f command in the group buffer (see Section 2.9 [Foreign Groups], page 18)), set the process mark on all the articles in the buffer (M P b, for instance), and then re-spool (B r) using nnml. If all goes well, all the mail in the ‘RMAIL’ file is now also stored in lots of nnml directories, and you can delete that pesky ‘RMAIL’ file. If you have the guts!

Virtual server variables:

nndoc-article-type
          This should be one of mbox, babyl, digest, news, rnews, mmdf, forward,
          rfc934, rfc822-forward, mime-parts, mime-digest, standard-digest,
          slack-digest, clari-briefs or guess.
nndoc-post-type

This variable says whether Gnus is to consider the group a news group or a mail group. There are two valid values: mail (the default) and news.

6.4.3.1 Document Server Internals

Adding new document types to be recognized by nndoc isn’t difficult. You just have to whip up a definition of what the document looks like, write a predicate function to recognize that document type, and then hook into nndoc.

First, here’s an example document type definition:

     (mmdf
      (article-begin .  "^\^A\^A\^A\^A\n")
      (body-end .  "^\^A\^A\^A\^A\n"))

The definition is simply a unique name followed by a series of regexp pseudo-variable settings. Below are the possible variables—don’t be daunted by the number of variables; most document types can be defined with very few settings:

first-article

If present, nndoc will skip past all text until it finds something that match this regexp. All text before this will be totally ignored.

article-begin

This setting has to be present in all document type definitions. It says what the beginning of each article looks like.

head-begin-function

If present, this should be a function that moves point to the head of the article.

nndoc-head-begin

If present, this should be a regexp that matches the head of the article.

nndoc-head-end

This should match the end of the head of the article. It defaults to ‘^$’—the empty line.

body-begin-function

If present, this function should move point to the beginning of the body of the article.

body-begin

This should match the beginning of the body of the article. It defaults to ‘^\n’.

body-end-function

If present, this function should move point to the end of the body of the article.

body-end If present, this should match the end of the body of the article.

file-end If present, this should match the end of the file. All text after this regexp will be totally ignored.

So, using these variables nndoc is able to dissect a document file into a series of articles, each with a head and a body. However, a few more variables are needed since not all document types are all that news-like—variables needed to transform the head or the body into something that’s palatable for Gnus:

prepare-body-function

If present, this function will be called when requesting an article. It will be called with point at the start of the body, and is useful if the document has encoded some parts of its contents.

article-transform-function

If present, this function is called when requesting an article. It’s meant to be used for more wide-ranging transformation of both head and body of the article.

generate-head-function

If present, this function is called to generate a head that Gnus can understand. It is called with the article number as a parameter, and is expected to generate a nice head for the article in question. It is called when requesting the headers of all articles.

Let’s look at the most complicated example I can come up with—standard digests:

(standard-digest
(first-article . ,(concat "^" (make-string 70 ?-) "\n\n+"))
(article-begin . ,(concat "\n\n" (make-string 30 ?-) "\n\n+"))
(prepare-body-function . nndoc-unquote-dashes)

      (body-end-function . nndoc-digest-body-end)
      (head-end . "^?$")
      (body-begin . "^?\n")
      (file-end . "^End of .*digest.*[0-9].*\n\\*\\*\\|^End of.*Digest *$")
      (subtype digest guess))

We see that all text before a 70-width line of dashes is ignored; all text after a line that starts with that ‘^End of’ is also ignored; each article begins with a 30-width line of dashes; the line separating the head from the body may contain a single space; and that the body is run through nndoc-unquote-dashes before being delivered.

To hook your own document definition into nndoc, use the nndoc-add-type function. It takes two parameters—the first is the definition itself and the second (optional) parameter says where in the document type definition alist to put this definition. The alist is traversed sequentially, and nndoc-TYPE-type-p is called for a given type TYPE. So nndoc-mmdf-type-p is called to see whether a document is of mmdf type, and so on. These type predicates should return nil if the document is not of the correct type; t if it is of the correct type; and a number if the document might be of the correct type. A high number means high probability; a low number means low probability with ‘0’ being the lowest valid number.

6.4.4 SOUP

In the PC world people often talk about “offline” newsreaders. These are thingies that are combined reader/news transport monstrosities. With built-in modem programs. Yecchh!

Of course, us Unix Weenie types of human beans use things like uucp and, like, nntpd and set up proper news and mail transport things like Ghod intended. And then we just use normal newsreaders.

However, it can sometimes be convenient to do something a that’s a bit easier on the brain if you have a very slow modem, and you’re not really that interested in doing things properly.

A file format called soup has been developed for transporting news and mail from servers to home machines and back again. It can be a bit fiddly.

First some terminology:

server This is the machine that is connected to the outside world and where you get news and/or mail from.

home machine This is the machine that you want to do the actual reading and responding on. It is typically not connected to the rest of the world in any way.

packet Something that contains messages and/or commands. There are two kinds of packets:

message packets These are packets made at the server, and typically contain lots of messages for you to read. These are called ‘SoupoutX.tgz’ by default, where X is a number. response packets

These are packets made at the home machine, and typically contains replies that you’ve written. These are called ‘SoupinX.tgz’ by default, where X is a number.

1. You log in on the server and create a soup packet. You can either use a dedicated soup thingie (like the awk program), or you can use Gnus to create the packet with its soup commands (O s and/or G s b; and then G s p) (see Section 6.4.4.1 [SOUP Commands], page 118).

2. You transfer the packet home. Rail, boat, car or modem will do fine.

3. You put the packet in your home directory.

4. You fire up Gnus on your home machine using the nnsoup backend as the native or secondary server.

5. You read articles and mail and answer and followup to the things you want (see Section 6.4.4.3 [SOUP Replies], page 120).

6. You do the G s r command to pack these replies into a soup packet.

7. You transfer this packet to the server.

8. You use Gnus to mail this packet out with the G s s command.

9. You then repeat until you die.

So you basically have a bipartite system—you use nnsoup for reading and Gnus for packing/sending these soup packets.

6.4.4.1 SOUP Commands

These are commands for creating and manipulating soup packets.

G s b Pack all unread articles in the current group (gnus-group-brew-soup). This command understands the process/prefix convention.

G s w  Save all soup data files (gnus-soup-save-areas).

G s s Send all replies from the replies packet (gnus-soup-send-replies).

G s p Pack all files into a soup packet (gnus-soup-pack-packet).

G s r Pack all replies into a replies packet (nnsoup-pack-replies).

O s This summary-mode command adds the current article to a soup packet (gnussoup-add-article). It understands the process/prefix convention (see Section 8.1 [Process/Prefix], page 153).

There are a few variables to customize where Gnus will put all these thingies:

gnus-soup-directory

Directory where Gnus will save intermediate files while composing soup packets. The default is ‘~/SoupBrew/’.

gnus-soup-replies-directory

This is what Gnus will use as a temporary directory while sending our reply

          packets. ‘~/SoupBrew/SoupReplies/’ is the default.

gnus-soup-prefix-file

Name of the file where Gnus stores the last used prefix. The default is

‘gnus-prefix’.
gnus-soup-packer

A format string command for packing a soup packet. The default is ‘tar cf -

%s | gzip > $HOME/Soupout%d.tgz’.
gnus-soup-unpacker

Format string command for unpacking a soup packet. The default is ‘gunzip

-c %s | tar xvf -’.
gnus-soup-packet-directory

Where Gnus will look for reply packets. The default is ‘~/’.

gnus-soup-packet-regexp

Regular expression matching soup reply packets in gnus-soup-packet-

directory.

6.4.4.2 soup Groups

nnsoup is the backend for reading soup packets. It will read incoming packets, unpack them, and put them in a directory where you can read them at leisure.

These are the variables you can use to customize its behavior:

nnsoup-tmp-directory

When nnsoup unpacks a soup packet, it does it in this directory. (‘/tmp/’ by default.)

nnsoup-directory

nnsoup then moves each message and index file to this directory. The default

is ‘~/SOUP/’.
nnsoup-replies-directory

All replies will be stored in this directory before being packed into a reply packet. The default is ‘~/SOUP/replies/"’.

nnsoup-replies-format-type

The soup format of the replies packets. The default is ‘?n’ (rnews), and I don’t think you should touch that variable. I probably shouldn’t even have documented it. Drats! Too late!

nnsoup-replies-index-type

The index type of the replies packet. The default is ‘?n’, which means “none”. Don’t fiddle with this one either!

nnsoup-active-file

Where nnsoup stores lots of information. This is not an “active file” in the nntp sense; it’s an Emacs Lisp file. If you lose this file or mess it up in any way, you’re dead. The default is ‘~/SOUP/active’.

nnsoup-packer

Format string command for packing a reply soup packet. The default is ‘tar

cf - %s | gzip > $HOME/Soupin%d.tgz’.

nnsoup-unpacker

Format string command for unpacking incoming soup packets. The default is

          ‘gunzip -c %s | tar xvf -’.
nnsoup-packet-directory

Where nnsoup will look for incoming packets. The default is ‘~/’.

nnsoup-packet-regexp

Regular expression matching incoming soup packets. The default is ‘Soupout’.

nnsoup-always-save

If non-nil, save the replies buffer after each posted message.

6.4.4.3 SOUP Replies

Just using nnsoup won’t mean that your postings and mailings end up in soup reply packets automagically. You have to work a bit more for that to happen.

The nnsoup-set-variables command will set the appropriate variables to ensure that all your followups and replies end up in the soup system.

In specific, this is what it does:

     (setq message-send-news-function 'nnsoup-request-post)
     (setq message-send-mail-function 'nnsoup-request-mail)

And that’s it, really. If you only want news to go into the soup system you just use the first line. If you only want mail to be souped you use the second.

6.4.5 Web Searches

It’s, like, too neat to search the Usenet for articles that match a string, but it, like, totally sucks, like, totally, to use one of those, like, Web browsers, and you, like, have to, rilly, like, look at the commercials, so, like, with Gnus you can do rad, rilly, searches without having to use a browser.

The nnweb backend allows an easy interface to the mighty search engine. You create an nnweb group, enter a search pattern, and then enter the group and read the articles like you would any normal group. The G w command in the group buffer (see Section 2.9 [Foreign Groups], page 18) will do this in an easy-to-use fashion.

nnweb groups don’t really lend themselves to being solid groups—they have a very fleeting idea of article numbers. In fact, each time you enter an nnweb group (not even changing the search pattern), you are likely to get the articles ordered in a different manner. Not even using duplicate suppression (see Section 3.27 [Duplicate Suppression], page 80) will help, since nnweb doesn’t even know the Message-ID of the articles before reading them using some search engines (DejaNews, for instance). The only possible way to keep track of which articles you’ve read is by scoring on the Date header—mark all articles posted before the last date you read the group as read.

If the search engine changes its output substantially, nnweb won’t be able to parse it and will fail. One could hardly fault the Web providers if they were to do this—their raison d’tre is to make money off of advertisements, not to provide services to the community. Since nnweb washes the ads off all the articles, one might think that the providers might be somewhat miffed. We’ll see.

You must have the url and w3 package installed to be able to use nnweb. Virtual server variables:

nnweb-type

What search engine type is being used. The currently supported types are

dejanews, dejanewsold, altavista and reference.
nnweb-search

The search string to feed to the search engine.

nnweb-max-hits

Advisory maximum number of hits per search to display. The default is 100.

nnweb-type-definition

Type-to-definition alist. This alist says what nnweb should do with the various search engine types. The following elements must be present:

article Function to decode the article and provide something that Gnus understands.

map Function to create an article number to message header and URL alist.

search Function to send the search string to the search engine.

address The address the aforementioned function should send the search string to.

id Format string URL to fetch an article by Message-ID.

6.4.6 Mail-To-News Gateways

If your local nntp server doesn’t allow posting, for some reason or other, you can post using one of the numerous mail-to-news gateways. The nngateway backend provides the interface.

Note that you can’t read anything from this backend—it can only be used to post with. Server variables:

nngateway-address

This is the address of the mail-to-news gateway.

nngateway-header-transformation

News headers often have to be transformed in some odd way or other for the mail-to-news gateway to accept it. This variable says what transformation

should be called, and defaults to nngateway-simple-header-transformation.

The function is called narrowed to the headers to be transformed and with one parameter—the gateway address. This default function just inserts a new To header based on the Newsgroups header and the gateway address. For instance, an article with this Newsgroups header:

     Newsgroups: alt.religion.emacs

will get this From header inserted:

               To: alt-religion-emacs@GATEWAY

The following pre-defined functions exist:

          nngateway-simple-header-transformation

Creates a To header that looks like newsgroup@nngateway-

                    address.
          nngateway-mail2news-header-transformation

Creates a To header that looks like nngateway-address.

Here’s an example:

                         (setq gnus-post-method
                               '(nngateway "mail2news@replay.com"
                                           (nngateway-header-transformation
                                            nngateway-mail2news-header-transformation)))

So, to use this, simply say something like:

     (setq gnus-post-method '(nngateway "GATEWAY.ADDRESS"))

6.5 Combined Groups

Gnus allows combining a mixture of all the other group types into bigger groups.

6.5.1 Virtual Groups

An nnvirtual group is really nothing more than a collection of other groups.

For instance, if you are tired of reading many small groups, you can put them all in one big group, and then grow tired of reading one big, unwieldy group. The joys of computing!

You specify nnvirtual as the method. The address should be a regexp to match component groups.

All marks in the virtual group will stick to the articles in the component groups. So if you tick an article in a virtual group, the article will also be ticked in the component group from whence it came. (And vice versa—marks from the component groups will also be shown in the virtual group.)

Here’s an example nnvirtual method that collects all Andrea Dworkin newsgroups into one, big, happy newsgroup:

     (nnvirtual "^alt\\.fan\\.andrea-dworkin$\\|^rec\\.dworkin.*")

The component groups can be native or foreign; everything should work smoothly, but if your computer explodes, it was probably my fault.

Collecting the same group from several servers might actually be a good idea if users have set the Distribution header to limit distribution. If you would like to read ‘soc.motss’ both from a server in Japan and a server in Norway, you could use the following as the group regexp:

     "^nntp\\+server\\.jp:soc\\.motss$\\|^nntp\\+server\\.no:soc\\.motss$"

(Remember, though, that if you’re creating the group with G m, you shouldn’t double the backslashes, and you should leave off the quote characters at the beginning and the end of the string.)

This should work kinda smoothly—all articles from both groups should end up in this one, and there should be no duplicates. Threading (and the rest) will still work as usual, but there might be problems with the sequence of articles. Sorting on date might be an option here (see Section 2.3 [Selecting a Group], page 14).

One limitation, however—all groups included in a virtual group have to be alive (i.e., subscribed or unsubscribed). Killed or zombie groups can’t be component groups for nnvirtual groups.

   If the nnvirtual-always-rescan is non-nil, nnvirtual will always scan groups for

unread articles when entering a virtual group. If this variable is nil (which is the default) and you read articles in a component group after the virtual group has been activated, the read articles from the component group will show up when you enter the virtual group. You’ll also see this effect if you have two virtual groups that have a component group in common. If that’s the case, you should set this variable to t. Or you can just tap M-g on the virtual group every time before you enter it—it’ll have much the same effect.

nnvirtual can have both mail and news groups as component groups. When responding to articles in nnvirtual groups, nnvirtual has to ask the backend of the component group the article comes from whether it is a news or mail backend. However, when you do a ^, there is typically no sure way for the component backend to know this, and in that case nnvirtual tells Gnus that the article came from a not-news backend. (Just to be on the safe side.)

C-c C-t in the message buffer will insert the Newsgroups line from the article you respond to in these cases.

6.5.2 Kibozed Groups

Kibozing is defined by oed as “grepping through (parts of) the news feed”. nnkiboze is a backend that will do this for you. Oh joy! Now you can grind any nntp server down to a halt with useless requests! Oh happiness!

To create a kibozed group, use the G k command in the group buffer.

The address field of the nnkiboze method is, as with nnvirtual, a regexp to match groups to be “included” in the nnkiboze group. That’s where most similarities between

nnkiboze and nnvirtual end.

In addition to this regexp detailing component groups, an nnkiboze group must have a score file to say what articles are to be included in the group (see Chapter 7 [Scoring], page 131).

   You must run M-x nnkiboze-generate-groups after creating the nnkiboze groups you

want to have. This command will take time. Lots of time. Oodles and oodles of time. Gnus has to fetch the headers from all the articles in all the component groups and run them through the scoring process to determine if there are any articles in the groups that are to be part of the nnkiboze groups.

Please limit the number of component groups by using restrictive regexps. Otherwise your sysadmin may become annoyed with you, and the nntp site may throw you off and never let you back in again. Stranger things have happened.

nnkiboze component groups do not have to be alive—they can be dead, and they can be foreign. No restrictions.

The generation of an nnkiboze group means writing two files in nnkiboze-directory, which is ‘~/News/’ by default. One contains the nov header lines for all the articles in the group, and the other is an additional ‘.newsrc’ file to store information on what groups have been searched through to find component articles.

Articles marked as read in the nnkiboze group will have their nov lines removed from the nov file.

6.6 Gnus Unplugged

In olden times (ca. February ’88), people used to run their newsreaders on big machines with permanent connections to the net. News transport was dealt with by news servers, and all the newsreaders had to do was to read news. Believe it or not.

Nowadays most people read news and mail at home, and use some sort of modem to connect to the net. To avoid running up huge phone bills, it would be nice to have a way to slurp down all the news and mail, hang up the phone, read for several hours, and then upload any responses you have to make. And then you repeat the procedure.

Of course, you can use news servers for doing this as well. I’ve used inn together with slurp, pop and sendmail for some years, but doing that’s a bore. Moving the news server functionality up to the newsreader makes sense if you’re the only person reading news on a machine.

Using Gnus as an “offline” newsreader is quite simple.

• First, set up Gnus as you would do if you were running it on a machine that has full connection to the net. Go ahead. I’ll still be waiting here.

• Then, put the following magical incantation at the end of your ‘.gnus.el’ file:

         (gnus-agentize)

That’s it. Gnus is now an “offline” newsreader.

Of course, to use it as such, you have to learn a few new commands.

6.6.1 Agent Basics

First, let’s get some terminology out of the way.

The Gnus Agent is said to be unplugged when you have severed the connection to the net (and notified the Agent that this is the case). When the connection to the net is up again (and Gnus knows this), the Agent is plugged.

The local machine is the one you’re running on, and which isn’t connected to the net continuously.

Downloading means fetching things from the net to your local machine. Uploading is doing the opposite.

Let’s take a typical Gnus session using the Agent.

• You start Gnus with gnus-unplugged. This brings up the Gnus Agent in a disconnected state. You can read all the news that you have already fetched while in this mode.

• You then decide to see whether any new news has arrived. You connect your machine to the net (using PPP or whatever), and then hit J j to make Gnus become plugged.

• You can then read the new news immediately, or you can download the news onto your local machine. If you want to do the latter, you press J s to fetch all the eligible articles in all the groups. (To let Gnus know which articles you want to download, see Section 6.6.2 [Agent Categories], page 125.)

• After fetching the articles, you press J j to make Gnus become unplugged again, and you shut down the PPP thing (or whatever). And then you read the news offline.

• And then you go to step 2.

Here are some things you should do the first time (or so) that you use the Agent.

• Decide which servers should be covered by the Agent. If you have a mail backend, it would probably be nonsensical to have it covered by the Agent. Go to the server buffer (^in the group buffer) and press J a the server (or servers) that you wish to have covered by the Agent (see Section 6.6.3.3 [Server Agent Commands], page 128). This will typically be only the primary select method, which is listed on the bottom in the buffer.

• Decide on download policy. See Section 6.6.2 [Agent Categories], page 125.

• Uhm... that’s it.

6.6.2 Agent Categories

One of the main reasons to integrate the news transport layer into the newsreader is to allow greater control over what articles to download. There’s not much point in downloading huge amounts of articles, just to find out that you’re not interested in reading any of them. It’s better to be somewhat more conservative in choosing what to download, and then mark the articles for downloading manually if it should turn out that you’re interested in the articles anyway.

The main way to control what is to be downloaded is to create a category and then assign some (or all) groups to this category. Gnus has its own buffer for creating and managing categories.

6.6.2.1 Category Syntax

A category consists of two things.

1. A predicate which (generally) gives a rough outline of which articles are eligible for downloading; and

2. a score rule which (generally) gives you a finer granularity when deciding what articles to download. (Note that this download score is wholly unrelated to normal scores.)

A predicate consists of predicates with logical operators sprinkled in between. Perhaps some examples are in order. Here’s a simple predicate. (It’s the default predicate, in fact, used for all groups that

don’t belong to any other category.)

 short

Quite simple, eh? This predicate is true if and only if the article is short (for some value of “short”).

Here’s a more complex predicate:

     (or high
         (and
          (not low)
          (not long)))

This means that an article should be downloaded if it has a high score, or if the score is not low and the article is not long. You get the drift.

The available logical operators are or, and and not. (If you prefer, you can use the more “C”-ish operators ‘|’, & and ! instead.)

The following predicates are pre-defined, but if none of these fit what you want to do, you can write your own.

short True iff the article is shorter than gnus-agent-short-article lines; default 100.

long True iff the article is longer than gnus-agent-long-article lines; default 200.

low True iff the article has a download score less than gnus-agent-low-score; default 0.

high True iff the article has a download score greater than gnus-agent-high-score; default 0.

spam True iff the Gnus Agent guesses that the article is spam. The heuristics may change over time, but at present it just computes a checksum and sees whether articles match.

true Always true.

false Always false.

If you want to create your own predicate function, here’s what you have to know: The functions are called with no parameters, but the gnus-headers and gnus-score dynamic variables are bound to useful values.

Now, the syntax of the download score is the same as the syntax of normal score files, except that all elements that require actually seeing the article itself are verboten. This means that only the following headers can be scored on: From, Subject, Date, Xref, Lines,

Chars, Message-ID, and References.

6.6.2.2 The Category Buffer

You’d normally do all category maintenance from the category buffer. When you enter it for the first time (with the J c command from the group buffer), you’ll only see the default category.

The following commands are available in this buffer:

q Return to the group buffer (gnus-category-exit).

k Kill the current category (gnus-category-kill).

c Copy the current category (gnus-category-copy).

a  Add a new category (gnus-category-add).

p Edit the predicate of the current category (gnus-category-edit-predicate).

g Edit the list of groups belonging to the current category (gnus-category-edit-

          groups).

s Edit the download score rule of the current category (gnus-category-edit-

          score).

l List all the categories (gnus-category-list).

6.6.2.3 Category Variables

gnus-category-mode-hook

Hook run in category buffers.

gnus-category-line-format

Format of the lines in the category buffer (see Section 8.4 [Formatting Variables], page 154). Valid elements are:

‘c’ The name of the category.

‘g’ The number of groups in the category.

gnus-category-mode-line-format

Format of the category mode line (see Section 8.4.2 [Mode Line Formatting], page 155).

gnus-agent-short-article

Articles that have fewer lines than this are short. Default 100.

gnus-agent-long-article

Articles that have more lines than this are long. Default 200.

gnus-agent-low-score

Articles that have a score lower than this have a low score. Default 0.

gnus-agent-high-score

Articles that have a score higher than this have a high score. Default 0.

6.6.3 Agent Commands

All the Gnus Agent commands are on the J submap. The J j (gnus-agent-toggle-plugged command works in all modes, and toggles the plugged/unplugged state of the Gnus Agent.

You can run a complete batch fetch from the command line with the following incantation:

     $ emacs -batch -l ~/.gnus.el -f gnus-agent-batch-fetch

6.6.3.1 Group Agent Commands

J u Fetch all eligible articles in the current group (gnus-agent-fetch-groups).

J c Enter the Agent category buffer (gnus-enter-category-buffer).

J s Fetch all eligible articles in all groups (gnus-agent-fetch-session).

J S Send all sendable messages in the draft group (gnus-agent-fetch-session). See Section 5.7 [Drafts], page 89.

J a Add the current group to an Agent category (gnus-agent-add-group).

6.6.3.2 Summary Agent Commands

J # Mark the article for downloading (gnus-agent-mark-article).

J M-# Remove the downloading mark from the article (gnus-agent-unmark-

          article).

@ Toggle whether to download the article (gnus-agent-toggle-mark).

J c Mark all undownloaded articles as read (gnus-agent-catchup).

6.6.3.3 Server Agent Commands

J a Add the current server to the list of servers covered by the Gnus Agent (gnus-

          agent-add-server).

J r Remove the current server from the list of servers covered by the Gnus Agent

          (gnus-agent-remove-server).

6.6.4 Agent Expiry

nnagent doesn’t handle expiry. Instead, there’s a special gnus-agent-expire command that will expire all read articles that are older than gnus-agent-expire-days days. It can be run whenever you feel that you’re running out of space. It’s not particularly fast or efficient, and it’s not a particularly good idea to interrupt it (with C-g or anything else) once you’ve started it.

if gnus-agent-expire-all is non-nil, this command will expire all articles—unread, read, ticked and dormant. If nil (which is the default), only read articles are eligible for expiry, and unread, ticked and dormant articles will be kept indefinitely.

6.6.5 Outgoing Messages

When Gnus is unplugged, all outgoing messages (both mail and news) are stored in the draft groups (see Section 5.7 [Drafts], page 89). You can view them there after posting, and edit them at will.

When Gnus is plugged again, you can send the messages either from the draft group with the special commands available there, or you can use the J S command in the group buffer to send all the sendable messages in the draft group.

6.6.6 Agent Variables

gnus-agent-directory

Where the Gnus Agent will store its files. The default is ‘~/News/agent/’.

gnus-agent-handle-level

Groups on levels (see Section 2.6 [Group Levels], page 16) higher than this variable will be ignored by the Agent. The default is gnus-level-subscribed, which means that only subscribed group will be considered by the Agent by default.

gnus-agent-plugged-hook

Hook run when connecting to the network.

gnus-agent-unplugged-hook

Hook run when disconnecting from the network.

6.6.7 Example Setup

If you don’t want to read this manual, and you have a fairly standard setup, you may be able to use something like the following as your ‘.gnus.el’ file to get started.

     ;;; Define how Gnus is to fetch news.  We do this over NNTP
     ;;; from your ISP's server.
     (setq gnus-select-method '(nntp "nntp.your-isp.com"))
     ;;; Define how Gnus is to read your mail.  We read mail from
     ;;; your ISP's POP server.
     (setenv "MAILHOST" "pop.your-isp.com")
     (setq nnmail-spool-file "po:username")
     ;;; Say how Gnus is to store the mail.  We use nnml groups.
     (setq gnus-secondary-select-methods '((nnml "")))
     ;;; Make Gnus into an offline newsreader.
     (gnus-agentize)

That should be it, basically. Put that in your ‘~/.gnus.el’ file, edit to suit your needs, start up PPP (or whatever), and type M-x gnus.

If this is the first time you’ve run Gnus, you will be subscribed automatically to a few default newsgroups. You’ll probably want to subscribe to more groups, and to do that, you have to query the nntp server for a complete list of groups with the A A command. This usually takes quite a while, but you only have to do it once.

After reading and parsing a while, you’ll be presented with a list of groups. Subscribe to the ones you want to read with the u command. l to make all the killed groups disappear after you’ve subscribe to all the groups you want to read. (A k will bring back all the killed groups.)

You can now read the groups at once, or you can download the articles with the J s command. And then read the rest of this manual to find out which of the other gazillion things you want to customize.

6.6.8 Batching Agents

Having the Gnus Agent fetch articles (and post whatever messages you’ve written) is quite easy once you’ve gotten things set up properly. The following shell script will do everything that is necessary:

     #!/bin/sh
     emacs -batch -l ~/.emacs -f gnus-agent-batch >/dev/null

7 Scoring

Other people use kill files, but we here at Gnus Towers like scoring better than killing, so we’d rather switch than fight. They do something completely different as well, so sit up straight and pay attention!

All articles have a default score (gnus-summary-default-score), which is 0 by default. This score may be raised or lowered either interactively or by score files. Articles that have a score lower than gnus-summary-mark-below are marked as read.

Gnus will read any score files that apply to the current group before generating the summary buffer.

There are several commands in the summary buffer that insert score entries based on the current article. You can, for instance, ask Gnus to lower or increase the score of all articles with a certain subject.

There are two sorts of scoring entries: Permanent and temporary. Temporary score entries are self-expiring entries. Any entries that are temporary and have not been used for, say, a week, will be removed silently to help keep the sizes of the score files down.

7.1 Summary Score Commands

The score commands that alter score entries do not actually modify real score files. That would be too inefficient. Gnus maintains a cache of previously loaded score files, one of which is considered the current score file alist. The score commands simply insert entries into this list, and upon group exit, this list is saved.

The current score file is by default the group’s local score file, even if no such score file actually exists. To insert score commands into some other score file (e.g. ‘all.SCORE’), you must first make this score file the current one.

General score commands that don’t actually change the score file:

V s Set the score of the current article (gnus-summary-set-score).

V S Display the score of the current article (gnus-summary-current-score).

V t Display all score rules that have been used on the current article (gnus-score-

          find-trace).

V R Run the current summary through the scoring process (gnus-summary-rescore). This might be useful if you’re playing around with your score files behind Gnus’ back and want to see the effect you’re having.

V c Make a different score file the current (gnus-score-change-score-file).

V e Edit the current score file (gnus-score-edit-current-scores). You will be popped into a gnus-score-mode buffer (see Section 7.5 [Score File Editing], page 140).

V f Edit a score file and make this score file the current one (gnus-score-edit-

          file).

V F Flush the score cache (gnus-score-flush-cache). This is useful after editing score files.

V C Customize a score file in a visually pleasing manner (gnus-score-customize).

The rest of these commands modify the local score file. V m Prompt for a score, and mark all articles with a score below this as read (gnus-

      score-set-mark-below).

V x Prompt for a score, and add a score rule to the current score file to expunge all

      articles below this score (gnus-score-set-expunge-below).

The keystrokes for actually making score entries follow a very regular pattern, so there’s no need to list all the commands. (Hundreds of them.)

1. The first key is either I (upper case i) for increasing the score or L for lowering the score.

2. The second key says what header you want to score on. The following keys are available: a Score on the author name. s Score on the subject line. x Score on the Xref line—i.e., the cross-posting line. r Score on the References line. d Score on the date. l Score on the number of lines. i Score on the Message-ID. f Score on followups. b Score on the body. h Score on the head. t Score on thead.

3. The third key is the match type. Which match types are valid depends on what headers you are scoring on.

strings

e Exact matching. s Substring matching. f Fuzzy matching (see Section 8.16 [Fuzzy Matching], page 171). r Regexp matching

date

b Before date. a After date. n This date.

number

< Less than number.

= Equal to number.

> Greater than number.

4. The fourth and final key says whether this is a temporary (i.e., expiring) score entry, or a permanent (i.e., non-expiring) score entry, or whether it is to be done immediately, without adding to the score file.

t Temporary score entry.

p Permanent score entry.

i Immediately scoring.

So, let’s say you want to increase the score on the current author with exact matching permanently: I a e p. If you want to lower the score based on the subject line, using substring matching, and make a temporary score entry: L s s t. Pretty easy.

To make things a bit more complicated, there are shortcuts. If you use a capital letter on either the second or third keys, Gnus will use defaults for the remaining one or two keystrokes. The defaults are “substring” and “temporary”. So I A is the same as I a s t, and I a R is the same as I a r t.

These functions take both the numerical prefix and the symbolic prefix (see Section 8.3 [Symbolic Prefixes], page 154). A numerical prefix says how much to lower (or increase) the score of the article. A symbolic prefix of a says to use the ‘all.SCORE’ file for the command instead of the current score file.

The gnus-score-mimic-keymap says whether these commands will pretend they are keymaps or not.

7.2 Group Score Commands

There aren’t many of these as yet, I’m afraid.

W f Gnus maintains a cache of score alists to avoid having to reload them all the time. This command will flush the cache (gnus-score-flush-cache).

You can do scoring from the command line by saying something like:

     $ emacs -batch -l ~/.emacs -l gnus -f gnus-batch-score

7.3 Score Variables

gnus-use-scoring

If nil, Gnus will not check for score files, and will not, in general, do any score-related work. This is t by default.

gnus-kill-killed

If this variable is nil, Gnus will never apply score files to articles that have already been through the kill process. While this may save you lots of time, it also means that if you apply a kill file to a group, and then change the kill file and want to run it over you group again to kill more articles, it won’t work. You have to set this variable to t to do that. (It is t by default.)

gnus-kill-files-directory

All kill and score files will be stored in this directory, which is initialized from the SAVEDIR environment variable by default. This is ‘~/News/’ by default.

gnus-score-file-suffix

Suffix to add to the group name to arrive at the score file name (‘SCORE’ by default.)

gnus-score-uncacheable-files

All score files are normally cached to avoid excessive re-loading of score files. However, if this might make you Emacs grow big and bloated, so this regexp can be used to weed out score files unlikely to be needed again. It would be a bad idea to deny caching of ‘all.SCORE’, while it might be a good idea to not

cache ‘comp.infosystems.www.authoring.misc.ADAPT’. In fact, this variable

is ‘ADAPT$’ by default, so no adaptive score files will be cached.

gnus-save-score

If you have really complicated score files, and do lots of batch scoring, then you might set this variable to t. This will make Gnus save the scores into the

‘.newsrc.eld’ file.
gnus-score-interactive-default-score

Score used by all the interactive raise/lower commands to raise/lower score with. Default is 1000, which may seem excessive, but this is to ensure that the adaptive scoring scheme gets enough room to play with. We don’t want the small changes from the adaptive scoring to overwrite manually entered data.

gnus-summary-default-score

Default score of an article, which is 0 by default.

gnus-summary-expunge-below

Don’t display the summary lines of articles that have scores lower than this variable. This is nil by default, which means that no articles will be hidden. This variable is local to the summary buffers, and has to be set from gnus-

summary-mode-hook.
gnus-score-over-mark

Mark (in the third column) used for articles with a score over the default. Default is ‘+’.

gnus-score-below-mark

Mark (in the third column) used for articles with a score below the default. Default is ‘-’.

gnus-score-find-score-files-function

Function used to find score files for the current group. This function is called with the name of the group as the argument. Predefined functions available are:

gnus-score-find-single

Only apply the group’s own score file.

gnus-score-find-bnews

Apply all score files that match, using bnews syntax. This is the default. If the current group is ‘gnu.emacs.gnus’, for

          instance, ‘all.emacs.all.SCORE’, ‘not.alt.all.SCORE’ and

‘gnu.all.SCORE’ would all apply. In short, the instances of ‘all’ in the score file names are translated into ‘.*’, and then a regexp match is done.

This means that if you have some score entries that you want to apply to all groups, then you put those entries in the ‘all.SCORE’ file.

The score files are applied in a semi-random order, although Gnus will try to apply the more general score files before the more specific score files. It does this by looking at the number of elements in the score file names—discarding the ‘all’ elements.

gnus-score-find-hierarchical

Apply all score files from all the parent groups. This means that you can’t have score files like ‘all.SCORE’, but you can have ‘SCORE’,

          ‘comp.SCORE’ and ‘comp.emacs.SCORE’.

This variable can also be a list of functions. In that case, all these functions will be called, and all the returned lists of score files will be applied. These functions can also return lists of score alists directly. In that case, the functions that return these non-file score alists should probably be placed before the “real” score file functions, to ensure that the last score file returned is the local score file. Phu.

gnus-score-expiry-days

This variable says how many days should pass before an unused score file entry is expired. If this variable is nil, no score file entries are expired. It’s 7 by default.

gnus-update-score-entry-dates

If this variable is non-nil, matching score entries will have their dates updated. (This is how Gnus controls expiry—all non-matching entries will become too old while matching entries will stay fresh and young.) However, if you set this variable to nil, even matching entries will grow old and will have to face that oh-so grim reaper.

gnus-score-after-write-file-function

Function called with the name of the score file just written.

gnus-score-thread-simplify

If this variable is non-nil, article subjects will be simplified for subject scoring purposes in the same manner as with threading—according to the current value of gnus-simplify-subject-functions. If the scoring entry uses substring or exact matching, the match will also be simplified in this manner.

7.4 Score File Format

A score file is an emacs-lisp file that normally contains just a single form. Casual users are not expected to edit these files; everything can be changed from the summary buffer.

Anyway, if you’d like to dig into it yourself, here’s an example:

(("from"
 ("Lars Ingebrigtsen" -10000)
 ("Per Abrahamsen")
 ("larsi\\|lmi" -50000 nil R))
("subject"
 ("Ding is Badd" nil 728373))
("xref"
 ("alt.politics" -1000 728372 s))
("lines"
 (2 -100 nil <))
(mark 0)
(expunge -1000)
(mark-and-expunge -10)
(read-only nil)
(orphan -10)
(adapt t)
(files "/hom/larsi/News/gnu.SCORE")
(exclude-files "all.SCORE")
(local (gnus-newsgroup-auto-expire t)
       (gnus-summary-make-false-root empty))
(eval (ding)))

This example demonstrates most score file elements. For a different approach, see see Section 7.15 [Advanced Scoring], page 149.

Even though this looks much like lisp code, nothing here is actually evaled. The lisp reader is used to read this form, though, so it has to be valid syntactically, if not semantically.

Six keys are supported by this alist:

STRING If the key is a string, it is the name of the header to perform the match on. Scoring can only be performed on these eight headers: From, Subject, References, Message-ID, Xref, Lines, Chars and Date. In addition to these headers, there are three strings to tell Gnus to fetch the entire article and do the match on larger parts of the article: Body will perform the match on the body of the article, Head will perform the match on the head of the article, and All will perform the match on the entire article. Note that using any of these last three keys will slow down group entry considerably. The final “header” you can score on is Followup. These score entries will result in new score entries being added for all follow-ups to articles that matches these score entries.

Following this key is a arbitrary number of score entries, where each score entry has one to four elements.

1. The first element is the match element. On most headers this will be a string, but on the Lines and Chars headers, this must be an integer.

2. If the second element is present, it should be a number—the score element. This number should be an integer in the neginf to posinf interval. This number is added to the score of the article if the match is successful. If

this element is not present, the gnus-score-interactive-default-score

number will be used instead. This is 1000 by default.

3. If the third element is present, it should be a number—the date element. This date says when the last time this score entry matched, which provides a mechanism for expiring the score entries. It this element is not present, the score entry is permanent. The date is represented by the number of days since December 31, 1 BCE.

4. If the fourth element is present, it should be a symbol—the type element. This element specifies what function should be used to see whether this score entry matches the article. What match types that can be used depends on what header you wish to perform the match on.

From, Subject, References, Xref, Message-ID For most header types, there are the r and R (regexp), as well as s and S (substring) types, and e and E (exact match), and w (word match) types. If this element is not present, Gnus will assume that substring matching should be used. R, S, and E differ from the others in that the matches will be done in a case-sensitive manner. All these one-letter types are really just abbreviations for the regexp, string, exact, and word types, which you can use instead, if you feel like.

Lines, Chars These two headers use different match types: <, >, =, >= and

<=.

These predicates are true if

     (PREDICATE HEADER MATCH)

evaluates to non-nil. For instance, the advanced match ("lines" 4 <) (see Section 7.15 [Advanced Scoring], page 149) will result in the following form:

     (< header-value 4)

Or to put it another way: When using < on Lines with 4 as the match, we get the score added if the article has less than 4 lines. (It’s easy to get confused and think it’s the other way around. But it’s not. I think.) When matching on Lines, be careful because some backends (like nndir) do not generate Lines header, so every article ends up being marked as having 0 lines. This can lead to strange results if you happen to lower score of the articles with few lines.

Date For the Date header we have three kinda silly match types: before, at and after. I can’t really imagine this ever being useful, but, like, it would feel kinda silly not to provide this function. Just in case. You never know. Better safe than sorry. Once burnt, twice shy. Don’t judge a book by its cover. Never not have sex on a first date. (I have been told that at least one person, and I quote, “found this function indispensable”, however.)

A more useful match type is regexp. With it, you can match the date string using a regular expression. The date is normalized to ISO8601 compact format first—YYYYMMDDTHHMMSS. If you want to match all articles that have been posted on April 1st in every year, you could use ‘....0401.........’ as a match string, for instance. (Note that the date is kept in its original time zone, so this will match articles that were posted when it was April 1st where the article was posted from. Time zones are such wholesome fun for the whole family, eh?)

Head, Body, All These three match keys use the same match types as the From (etc) header uses.

Followup This match key is somewhat special, in that it will match the From header, and affect the score of not only the matching articles, but also all followups to the matching articles. This allows you e.g. increase the score of followups to your own articles, or decrease the score of followups to the articles of some known trouble-maker. Uses the same match types as the From header uses. (Using this match key will lead to creation of ‘ADAPT’ files.)

Thread This match key works along the same lines as the Followup match key. If you say that you want to score on a (sub- )thread started by an article with a Message-ID X, then you add a ‘thread’ match. This will add a new ‘thread’ match for each article that has X in its References header. (These new ‘thread’ matches will use the Message-IDs of these matching articles.) This will ensure that you can raise/lower the score of an entire thread, even though some articles in the thread may not have complete References headers. Note that using this may lead to undeterministic scores of the articles in the thread. (Using this match key will lead to creation of ‘ADAPT’ files.)

mark The value of this entry should be a number. Any articles with a score lower than this number will be marked as read.

expunge The value of this entry should be a number. Any articles with a score lower than this number will be removed from the summary buffer.

mark-and-expunge

The value of this entry should be a number. Any articles with a score lower than this number will be marked as read and removed from the summary buffer.

thread-mark-and-expunge

The value of this entry should be a number. All articles that belong to a thread that has a total score below this number will be marked as read and removed from the summary buffer. gnus-thread-score-function says how to compute the total score for a thread.

files The value of this entry should be any number of file names. These files are assumed to be score files as well, and will be loaded the same way this one was.

exclude-files

The clue of this entry should be any number of files. These files will not be loaded, even though they would normally be so, for some reason or other.

eval The value of this entry will be evalel. This element will be ignored when handling global score files.

read-only

Read-only score files will not be updated or saved. Global score files should feature this atom (see Section 7.11 [Global Score Files], page 145). (Note: Global here really means global; not your personal apply-to-all-groups score files.)

orphan The value of this entry should be a number. Articles that do not have parents will get this number added to their scores. Imagine you follow some high-volume newsgroup, like ‘comp.lang.c’. Most likely you will only follow a few of the threads, also want to see any new threads. You can do this with the following two score file entries:

             (orphan -500)
             (mark-and-expunge -100)

When you enter the group the first time, you will only see the new threads. You then raise the score of the threads that you find interesting (with I T or I S), and ignore (C y) the rest. Next time you enter the group, you will see new articles in the interesting threads, plus any new threads. I.e.—the orphan score atom is for high-volume groups where there exist a few interesting threads which can’t be found automatically by ordinary scoring rules.

adapt This entry controls the adaptive scoring. If it is t, the default adaptive scoring rules will be used. If it is ignore, no adaptive scoring will be performed on this group. If it is a list, this list will be used as the adaptive scoring rules. If it isn’t present, or is something other than t or ignore, the default adaptive scoring rules will be used. If you want to use adaptive scoring on most groups,

you’d set gnus-use-adaptive-scoring to t, and insert an (adapt ignore) in

the groups where you do not want adaptive scoring. If you only want adaptive scoring in a few groups, you’d set gnus-use-adaptive-scoring to nil, and insert (adapt t) in the score files of the groups where you want it.

adapt-file

All adaptive score entries will go to the file named by this entry. It will also be applied when entering the group. This atom might be handy if you want to adapt on several groups at once, using the same adaptive file for a number of groups.

local The value of this entry should be a list of (VAR VALUE) pairs. Each var will be made buffer-local to the current summary buffer, and set to the value specified. This is a convenient, if somewhat strange, way of setting variables in some groups if you don’t like hooks much. Note that the value won’t be evaluated.

7.5 Score File Editing

You normally enter all scoring commands from the summary buffer, but you might feel the urge to edit them by hand as well, so we’ve supplied you with a mode for that.

It’s simply a slightly customized emacs-lisp mode, with these additional commands:

C-c C-c Save the changes you have made and return to the summary buffer (gnus-

          score-edit-done).

C-c C-d Insert the current date in numerical format (gnus-score-edit-insert-date). This is really the day number, if you were wondering.

C-c C-p The adaptive score files are saved in an unformatted fashion. If you intend to read one of these files, you want to pretty print it first. This command

          (gnus-score-pretty-print) does that for you.
   Type M-x gnus-score-mode to use this mode.

gnus-score-menu-hook is run in score mode buffers.

In the summary buffer you can use commands like V f and V e to begin editing score files.

7.6 Adaptive Scoring

If all this scoring is getting you down, Gnus has a way of making it all happen automatically—as if by magic. Or rather, as if by artificial stupidity, to be precise.

When you read an article, or mark an article as read, or kill an article, you leave marks behind. On exit from the group, Gnus can sniff these marks and add score elements depending on what marks it finds. You turn on this ability by setting gnus-use-adaptive-scoring to t or (line). If you want score adaptively on separate words appearing in the subjects, you should set this variable to (word). If you want to use both adaptive methods, set this variable to (word line).

To give you complete control over the scoring process, you can customize the gnusdefault-adaptive-score-alist variable. For instance, it might look something like this:

     (defvar gnus-default-adaptive-score-alist
       '((gnus-unread-mark)
         (gnus-ticked-mark (from 4))
         (gnus-dormant-mark (from 5))

         (gnus-del-mark (from -4) (subject -1))
         (gnus-read-mark (from 4) (subject 2))
         (gnus-expirable-mark (from -1) (subject -1))
         (gnus-killed-mark (from -1) (subject -3))
         (gnus-kill-file-mark)
         (gnus-ancient-mark)
         (gnus-low-score-mark)
         (gnus-catchup-mark (from -1) (subject -1))))

As you see, each element in this alist has a mark as a key (either a variable name or a “real” mark—a character). Following this key is a arbitrary number of header/score pairs. If there are no header/score pairs following the key, no adaptive scoring will be done on articles that have that key as the article mark. For instance, articles with gnus-unread-mark in the example above will not get adaptive score entries.

Each article can have only one mark, so just a single of these rules will be applied to each article.

To take gnus-del-mark as an example—this alist says that all articles that have that mark (i.e., are marked with ‘D’) will have a score entry added to lower based on the From header by -4, and lowered by Subject by -1. Change this to fit your prejudices.

If you have marked 10 articles with the same subject with gnus-del-mark, the rule for that mark will be applied ten times. That means that that subject will get a score of ten times -1, which should be, unless I’m much mistaken, -10.

If you have auto-expirable (mail) groups (see Section 6.3.7 [Expiring Mail], page 106), all the read articles will be marked with the ‘E’ mark. This’ll probably make adaptive scoring slightly impossible, so auto-expiring and adaptive scoring doesn’t really mix very well.

   The headers you can score on are from, subject, message-id, references, xref, lines,

chars and date. In addition, you can score on followup, which will create an adaptive score entry that matches on the References header using the Message-ID of the current article, thereby matching the following thread.

You can also score on thread, which will try to score all articles that appear in a thread. thread matches uses a Message-ID to match on the References header of the article. If the match is made, the Message-ID of the article is added to the thread rule. (Think about it. I’d recommend two aspirins afterwards.)

If you use this scheme, you should set the score file atom mark to something small—like -300, perhaps, to avoid having small random changes result in articles getting marked as read.

After using adaptive scoring for a week or so, Gnus should start to become properly trained and enhance the authors you like best, and kill the authors you like least, without you having to say so explicitly.

You can control what groups the adaptive scoring is to be performed on by using the score files (see Section 7.4 [Score File Format], page 136). This will also let you use different rules in different groups.

The adaptive score entries will be put into a file where the name is the group name with

gnus-adaptive-file-suffix appended. The default is ‘ADAPT’.

When doing adaptive scoring, substring or fuzzy matching would probably give you the best results in most cases. However, if the header one matches is short, the possibility for false positives is great, so if the length of the match is less than gnus-score-exact-adapt-limit, exact matching will be used. If this variable is nil, exact matching will always be used to avoid this problem.

As mentioned above, you can adapt either on individual words or entire headers. If you

adapt on words, the gnus-default-adaptive-word-score-alist variable says what score

each instance of a word should add given a mark.

     (setq gnus-default-adaptive-word-score-alist
           `((,gnus-read-mark . 30)
             (,gnus-catchup-mark . -10)
             (,gnus-killed-mark . -20)
             (,gnus-del-mark . -15)))

This is the default value. If you have adaption on words enabled, every word that appears in subjects of articles marked with gnus-read-mark will result in a score rule that increase the score with 30 points.

Words that appear in the gnus-default-ignored-adaptive-words list will be ignored. If you wish to add more words to be ignored, use the gnus-ignored-adaptive-words list instead.

When the scoring is done, gnus-adaptive-word-syntax-table is the syntax table in effect. It is similar to the standard syntax table, but it considers numbers to be non-word-constituent characters.

If gnus-adaptive-word-minimum is set to a number, the adaptive word scoring process will never bring down the score of an article to below this number. The default is nil.

After using this scheme for a while, it might be nice to write a gnus-psychoanalyze-user command to go through the rules and see what words you like and what words you don’t like. Or perhaps not.

Note that the adaptive word scoring thing is highly experimental and is likely to change in the future. Initial impressions seem to indicate that it’s totally useless as it stands. Some more work (involving more rigorous statistical methods) will have to be done to make this useful.

7.7 Home Score File

The score file where new score file entries will go is called the home score file. This is normally (and by default) the score file for the group itself. For instance, the home score

file for ‘gnu.emacs.gnus’ is ‘gnu.emacs.gnus.SCORE’.

However, this may not be what you want. It is often convenient to share a common home score file among many groups—all ‘emacs’ groups could perhaps use the same home score file.

The variable that controls this is gnus-home-score-file. It can be:

1. A string. Then this file will be used as the home score file for all groups.

2. A function. The result of this function will be used as the home score file. The function will be called with the name of the group as the parameter.

3. A list. The elements in this list can be:

1. (regexp file-name). If the regexp matches the group name, the file-name will will be used as the home score file.

2. A function. If the function returns non-nil, the result will be used as the home score file.

3. A string. Use the string as the home score file.

The list will be traversed from the beginning towards the end looking for matches.

So, if you want to use just a single score file, you could say:

     (setq gnus-home-score-file
           "my-total-score-file.SCORE")

If you want to use ‘gnu.SCORE’ for all ‘gnu’ groups and ‘rec.SCORE’ for all ‘rec’ groups (and so on), you can say:

     (setq gnus-home-score-file
           'gnus-hierarchial-home-score-file)

This is a ready-made function provided for your convenience. Other functions include

gnus-current-home-score-file

Return the “current” regular score file. This will make scoring commands add entry to the “innermost” matching score file.

If you want to have one score file for the ‘emacs’ groups and another for the ‘comp’ groups, while letting all other groups use their own home score files:

     (setq gnus-home-score-file
           ;; All groups that match the regexp "\\.emacs"
           '(("\\.emacs" "emacs.SCORE")
             ;; All the comp groups in one score file
             ("^comp" "comp.SCORE")))
   gnus-home-adapt-file works exactly the same way as gnus-home-score-file, but

says what the home adaptive score file is instead. All new adaptive file entries will go into the file specified by this variable, and the same syntax is allowed.

   In addition to using gnus-home-score-file and gnus-home-adapt-file, you can also

use group parameters (see Section 2.10 [Group Parameters], page 19) and topic parameters (see Section 2.16.5 [Topic Parameters], page 29) to achieve much the same. Group and topic parameters take precedence over this variable.

7.8 Followups To Yourself

Gnus offers two commands for picking out the Message-ID header in the current buffer. Gnus will then add a score rule that scores using this Message-ID on the References header of other articles. This will, in effect, increase the score of all articles that respond to the article in the current buffer. Quite useful if you want to easily note when people answer what you’ve said.

gnus-score-followup-article

This will add a score to articles that directly follow up your own article.

gnus-score-followup-thread

This will add a score to all articles that appear in a thread “below” your own article.

These two functions are both primarily meant to be used in hooks like message-sent-

hook.

If you look closely at your own Message-ID, you’ll notice that the first two or three characters are always the same. Here’s two of mine:

<x6u3u47icf.fsf@eyesore.no>
<x6sp9o7ibw.fsf@eyesore.no>

So “my” ident on this machine is ‘x6’. This can be exploited—the following rule will raise the score on all followups to myself:

("references"
("<x6[0-9a-z]+\\.fsf\\(_-_\\)?@.*eyesore.no>"
1000 nil r))

Whether it’s the first two or first three characters that are “yours” is system-dependent.

7.9 Scoring Tips

Crossposts If you want to lower the score of crossposts, the line to match on is the Xref header.

     ("xref" (" talk.politics.misc:" -1000))

Multiple crossposts If you want to lower the score of articles that have been crossposted to more than, say, 3 groups:

     ("xref" ("[^:\n]+:[0-9]+ +[^:\n]+:[0-9]+ +[^:\n]+:[0-9]+" -1000 nil r))

Matching on the body This is generally not a very good idea—it takes a very long time. Gnus actually has to fetch each individual article from the server. But you might want to anyway, I guess. Even though there are three match keys (Head, Body and All), you should choose one and stick with it in each score file. If you use any two, each article will be fetched twice. If you want to match a bit on the Head and a bit on the Body, just use All for all the matches.

Marking as read You will probably want to mark articles that has a score below a certain number as read. This is most easily achieved by putting the following in your

‘all.SCORE’ file:
     ((mark -100))

You may also consider doing something similar with expunge.

Negated character classes If you say stuff like [^abcd]*, you may get unexpected results. That will match newlines, which might lead to, well, The Unknown. Say [^abcd\n]* instead.

7.10 Reverse Scoring

If you want to keep just articles that have ‘Sex with Emacs’ in the subject header, and expunge all other articles, you could put something like this in your score file:

     (("subject"
       ("Sex with Emacs" 2))
      (mark 1)
      (expunge 1))

So, you raise all articles that match ‘Sex with Emacs’ and mark the rest as read, and expunge them to boot.

7.11 Global Score Files

Sure, other newsreaders have “global kill files”. These are usually nothing more than a single kill file that applies to all groups, stored in the user’s home directory. Bah! Puny, weak newsreaders!

What I’m talking about here are Global Score Files. Score files from all over the world, from users everywhere, uniting all nations in one big, happy score file union! Ange-score! New and untested!

All you have to do to use other people’s score files is to set the gnus-global-score-files variable. One entry for each score file, or each score file directory. Gnus will decide by itself what score files are applicable to which group.

   Say you want to use the score file ‘/ftp@ftp.gnus.org:/pub/larsi/ding/score/soc.motss.SCORE’

and all score files in the ‘/ftp@ftp.some-where:/pub/score’ directory:

     (setq gnus-global-score-files
           '("/ftp@ftp.gnus.org:/pub/larsi/ding/score/soc.motss.SCORE"
             "/ftp@ftp.some-where:/pub/score/"))

Simple, eh? Directory names must end with a ‘/’. These directories are typically scanned only once during each Gnus session. If you feel the need to manually re-scan the remote

directories, you can use the gnus-score-search-global-directories command.

Note that, at present, using this option will slow down group entry somewhat. (That is—a lot.)

If you want to start maintaining score files for other people to use, just put your score file up for anonymous ftp and announce it to the world. Become a retro-moderator! Participate in the retro-moderator wars sure to ensue, where retro-moderators battle it out for the sympathy of the people, luring them to use their score files on false premises! Yay! The net is saved!

Here are some tips for the would-be retro-moderator, off the top of my head:

• Articles heavily crossposted are probably junk.

• To lower a single inappropriate article, lower by Message-ID.

• Particularly brilliant authors can be raised on a permanent basis.

• Authors that repeatedly post off-charter for the group can safely be lowered out of existence.

• Set the mark and expunge atoms to obliterate the nastiest articles completely.

• Use expiring score entries to keep the size of the file down. You should probably have a long expiry period, though, as some sites keep old articles for a long time.

... I wonder whether other newsreaders will support global score files in the future. Snicker. Yup, any day now, newsreaders like Blue Wave, xrn and 1stReader are bound to implement scoring. Should we start holding our breath yet?

7.12 Kill Files

Gnus still supports those pesky old kill files. In fact, the kill file entries can now be expiring, which is something I wrote before Daniel Quinlan thought of doing score files, so I’ve left the code in there.

In short, kill processing is a lot slower (and I do mean a lot) than score processing, so it might be a good idea to rewrite your kill files into score files.

Anyway, a kill file is a normal emacs-lisp file. You can put any forms into this file, which means that you can use kill files as some sort of primitive hook function to be run on group entry, even though that isn’t a very good idea.

Normal kill files look like this:

     (gnus-kill "From" "Lars Ingebrigtsen")
     (gnus-kill "Subject" "ding")
     (gnus-expunge "X")

This will mark every article written by me as read, and remove the marked articles from the summary buffer. Very useful, you’ll agree.

Other programs use a totally different kill file syntax. If Gnus encounters what looks like a rn kill file, it will take a stab at interpreting it.

Two summary functions for editing a GNUS kill file:

M-k  Edit this group’s kill file (gnus-summary-edit-local-kill).
M-K  Edit the general kill file (gnus-summary-edit-global-kill).

Two group mode functions for editing the kill files:

M-k  Edit this group’s kill file (gnus-group-edit-local-kill).
M-K  Edit the general kill file (gnus-group-edit-global-kill).

Kill file variables:

gnus-kill-file-name

A kill file for the group ‘soc.motss’ is normally called ‘soc.motss.KILL’. The suffix appended to the group name to get this file name is detailed by the gnuskill-file-name variable. The “global” kill file (not in the score file sense of “global”, of course) is just called ‘KILL’.

gnus-kill-save-kill-file

If this variable is non-nil, Gnus will save the kill file after processing, which is necessary if you use expiring kills.

gnus-apply-kill-hook

A hook called to apply kill files to a group. It is (gnus-apply-kill-file) by default. If you want to ignore the kill file if you have a score file for the same

          group, you can set this hook to (gnus-apply-kill-file-unless-scored). If

you don’t want kill files to be processed, you should set this variable to nil.

gnus-kill-file-mode-hook

A hook called in kill-file mode buffers.

7.13 Converting Kill Files

If you have loads of old kill files, you may want to convert them into score files. If they are “regular”, you can use the ‘gnus-kill-to-score.el’ package; if not, you’ll have to do it by hand.

The kill to score conversion package isn’t included in Gnus by default. You can fetch it

from ‘http://www.stud.ifi.uio.no/~larsi/ding-other/gnus-kill-to-score’.

If your old kill files are very complex—if they contain more non-gnus-kill forms than not, you’ll have to convert them by hand. Or just let them be as they are. Gnus will still use them as before.

7.14 GroupLens

GroupLens is a collaborative filtering system that helps you work together with other people to find the quality news articles out of the huge volume of news articles generated every day.

To accomplish this the GroupLens system combines your opinions about articles you have already read with the opinions of others who have done likewise and gives you a personalized prediction for each unread news article. Think of GroupLens as a matchmaker. GroupLens watches how you rate articles, and finds other people that rate articles the same way. Once it has found some people you agree with it tells you, in the form of a prediction, what they thought of the article. You can use this prediction to help you decide whether or not you want to read the article.

7.14.1 Using GroupLens

To use GroupLens you must register a pseudonym with your local Better Bit Bureau

(BBB). ‘http://www.cs.umn.edu/Research/GroupLens/bbb.html’ is the only better bit

in town at the moment.

Once you have registered you’ll need to set a couple of variables.

gnus-use-grouplens

Setting this variable to a non-nil value will make Gnus hook into all the relevant GroupLens functions.

grouplens-pseudonym

This variable should be set to the pseudonym you got when registering with the Better Bit Bureau.

grouplens-newsgroups

A list of groups that you want to get GroupLens predictions for.

That’s the minimum of what you need to get up and running with GroupLens. Once you’ve registered, GroupLens will start giving you scores for articles based on the average of what other people think. But, to get the real benefit of GroupLens you need to start rating articles yourself. Then the scores GroupLens gives you will be personalized for you, based on how the people you usually agree with have already rated.

7.14.2 Rating Articles

In GroupLens, an article is rated on a scale from 1 to 5, inclusive. Where 1 means something like this article is a waste of bandwidth and 5 means that the article was really good. The basic question to ask yourself is, "on a scale from 1 to 5 would I like to see more articles like this one?"

There are four ways to enter a rating for an article in GroupLens.

r This function will prompt you for a rating on a scale of one to five.

k This function will prompt you for a rating, and rate all the articles in the thread. This is really useful for some of those long running giant threads in rec.humor.

The next two commands, n and , take a numerical prefix to be the score of the article you’re reading.

1-5 n Rate the article and go to the next unread article.

1-5 , Rate the article and go to the next unread article with the highest score.

If you want to give the current article a score of 4 and then go to the next article, just type 4 n.

7.14.3 Displaying Predictions

GroupLens makes a prediction for you about how much you will like a news article. The predictions from GroupLens are on a scale from 1 to 5, where 1 is the worst and 5 is the best. You can use the predictions from GroupLens in one of three ways controlled by the

variable gnus-grouplens-override-scoring.

There are three ways to display predictions in grouplens. You may choose to have the GroupLens scores contribute to, or override the regular gnus scoring mechanism. override is the default; however, some people prefer to see the Gnus scores plus the grouplens scores. To get the separate scoring behavior you need to set gnus-grouplens-override-scoring to 'separate. To have the GroupLens predictions combined with the grouplens scores

set it to 'override and to combine the scores set gnus-grouplens-override-scoring

to 'combine. When you use the combine option you will also want to set the values for

grouplens-prediction-offset and grouplens-score-scale-factor.

In either case, GroupLens gives you a few choices for how you would like to see your predictions displayed. The display of predictions is controlled by the grouplens-prediction-display variable.

The following are valid values for that variable.

prediction-spot

The higher the prediction, the further to the right an ‘*’ is displayed.

confidence-interval

A numeric confidence interval.

prediction-bar

The higher the prediction, the longer the bar.

confidence-bar

Numerical confidence.

confidence-spot

The spot gets bigger with more confidence.

prediction-num

Plain-old numeric value.

confidence-plus-minus

Prediction +/- confidence.

7.14.4 GroupLens Variables

gnus-summary-grouplens-line-format

The summary line format used in GroupLens-enhanced summary buffers. It accepts the same specs as the normal summary line format (see Section 3.1.1 [Summary Buffer Lines], page 33). The default is ‘%U%R%z%l%I%(%[%4L:

%-20,20n%]%) %s\n’.
grouplens-bbb-host

Host running the bbbd server. ‘grouplens.cs.umn.edu’ is the default.

grouplens-bbb-port

Port of the host running the bbbd server. The default is 9000.

grouplens-score-offset

Offset the prediction by this value. In other words, subtract the prediction value by this number to arrive at the effective score. The default is 0.

grouplens-score-scale-factor

This variable allows the user to magnify the effect of GroupLens scores. The scale factor is applied after the offset. The default is 1.

7.15 Advanced Scoring

Scoring on Subjects and From headers is nice enough, but what if you’re really interested in what a person has to say only when she’s talking about a particular subject? Or what if you really don’t want to read what person A has to say when she’s following up to person B, but want to read what she says when she’s following up to person C?

By using advanced scoring rules you may create arbitrarily complex scoring patterns.

7.15.1 Advanced Scoring Syntax

Ordinary scoring rules have a string as the first element in the rule. Advanced scoring rules have a list as the first element. The second element is the score to be applied if the first element evaluated to a non-nil value.

These lists may consist of three logical operators, one redirection operator, and various match operators.

Logical operators:

&

and This logical operator will evaluate each of its arguments until it finds one that evaluates to false, and then it’ll stop. If all arguments evaluate to true values, then this operator will return true.

|

or This logical operator will evaluate each of its arguments until it finds one that evaluates to true. If no arguments are true, then this operator will return

          false.
!
not

This logical operator only takes a single argument. It returns the logical negation of the value of its argument.

There is an indirection operator that will make its arguments apply to the ancestors of the current article being scored. For instance, 1- will make score rules apply to the parent of the current article. 2- will make score rules apply to the grandparent of the current article. Alternatively, you can write ^^, where the number of ^s (carets) says how far back into the ancestry you want to go.

Finally, we have the match operators. These are the ones that do the real work. Match operators are header name strings followed by a match and a match type. A typical match operator looks like ‘("from" "Lars Ingebrigtsen" s)’. The header names are the same as when using simple scoring, and the match types are also the same.

7.15.2 Advanced Scoring Examples

Let’s say you want to increase the score of articles written by Lars when he’s talking about Gnus:

     ((&
       ("from" "Lars Ingebrigtsen")
       ("subject" "Gnus"))
      1000)

Quite simple, huh?

When he writes long articles, he sometimes has something nice to say:

     ((&
       ("from" "Lars Ingebrigtsen")
       (|
        ("subject" "Gnus")

        ("lines" 100 >)))
      1000)

However, when he responds to things written by Reig Eigil Logge, you really don’t want to read what he’s written:

     ((&
       ("from" "Lars Ingebrigtsen")
       (1- ("from" "Reig Eigir Logge")))
      -100000)

Everybody that follows up Redmondo when he writes about disappearing socks should have their scores raised, but only when they talk about white socks. However, when Lars talks about socks, it’s usually not very interesting:

     ((&
       (1-
        (&
         ("from" "redmondo@.*no" r)
         ("body" "disappearing.*socks" t)))
       (! ("from" "Lars Ingebrigtsen"))
       ("body" "white.*socks"))
      1000)

The possibilities are endless.

7.15.3 Advanced Scoring Tips

The & and | logical operators do short-circuit logic. That is, they stop processing their arguments when it’s clear what the result of the operation will be. For instance, if one of the arguments of an & evaluates to false, there’s no point in evaluating the rest of the arguments. This means that you should put slow matches (‘body’, ‘header’) last and quick matches (‘from’, ‘subject’) first.

The indirection arguments (1- and so on) will make their arguments work on previous generations of the thread. If you say something like:

     ...
     (1-
      (1-
       ("from" "lars")))
     ...

Then that means "score on the from header of the grandparent of the current article". An indirection is quite fast, but it’s better to say:

     (1-
      (&
       ("from" "Lars")
       ("subject" "Gnus")))

than it is to say:

     (&
      (1- ("from" "Lars"))
      (1- ("subject" "Gnus")))

7.16 Score Decays

You may find that your scores have a tendency to grow without bounds, especially if you’re using adaptive scoring. If scores get too big, they lose all meaning—they simply max out and it’s difficult to use them in any sensible way.

Gnus provides a mechanism for decaying scores to help with this problem. When score files are loaded and gnus-decay-scores is non-nil, Gnus will run the score files through the decaying mechanism thereby lowering the scores of all non-permanent score rules. The decay itself if performed by the gnus-decay-score-function function, which is gnusdecay-score by default. Here’s the definition of that function:

     (defun gnus-decay-score (score)
       "Decay SCORE.
     This is done according to `gnus-score-decay-constant'
     and `gnus-score-decay-scale'."
       (floor
        (- score
           (* (if (< score 0) 1 -1)
              (min (abs score)
                   (max gnus-score-decay-constant
                        (* (abs score)
                           gnus-score-decay-scale)))))))
   gnus-score-decay-constant is 3 by default and gnus-score-decay-scale is 0.05.

This should cause the following:

1. Scores between -3 and 3 will be set to 0 when this function is called.

2. Scores with magnitudes between 3 and 60 will be shrunk by 3.

3. Scores with magnitudes greater than 60 will be shrunk by 5% of the score.

If you don’t like this decay function, write your own. It is called with the score to be decayed as its only parameter, and it should return the new score, which should be an integer.

Gnus will try to decay scores once a day. If you haven’t run Gnus for four days, Gnus will decay the scores four times, for instance.

8 Various

8.1 Process/Prefix

Many functions, among them functions for moving, decoding and saving articles, use what is known as the Process/Prefix convention.

This is a method for figuring out what articles the user wants the command to be performed on.

It goes like this:

If the numeric prefix is N, perform the operation on the next N articles, starting with the current one. If the numeric prefix is negative, perform the operation on the previous N articles, starting with the current one.

If transient-mark-mode in non-nil and the region is active, all articles in the region will be worked upon.

If there is no numeric prefix, but some articles are marked with the process mark, perform the operation on the articles marked with the process mark.

If there is neither a numeric prefix nor any articles marked with the process mark, just perform the operation on the current article.

Quite simple, really, but it needs to be made clear so that surprises are avoided.

Commands that react to the process mark will push the current list of process marked articles onto a stack and will then clear all process marked articles. You can restore the previous configuration with the M P y command (see Section 3.7.5 [Setting Process Marks], page 44).

One thing that seems to shock & horrify lots of people is that, for instance, 3 d does exactly the same as d d d. Since each d (which marks the current article as read) by default goes to the next unread article after marking, this means that 3 d will mark the next three unread articles as read, no matter what the summary buffer looks like. Set gnus-summary-goto-unread to nil for a more straightforward action.

8.2 Interactive

gnus-novice-user

If this variable is non-nil, you are either a newcomer to the World of Usenet, or you are very cautious, which is a nice thing to be, really. You will be given questions of the type “Are you sure you want to do this?” before doing anything dangerous. This is t by default.

gnus-expert-user

If this variable is non-nil, you will seldom be asked any questions by Gnus. It will simply assume you know what you’re doing, no matter how strange.

gnus-interactive-catchup

Require confirmation before catching up a group if non-nil. It is t by default.

gnus-interactive-exit

Require confirmation before exiting Gnus. This variable is t by default.

8.3 Symbolic Prefixes

Quite a lot of Emacs commands react to the (numeric) prefix. For instance, C-u 4 Cf moves point four characters forward, and C-u 9 0 0 I s s p adds a permanent Subject substring score rule of 900 to the current article.

This is all nice and well, but what if you want to give a command some additional information? Well, what most commands do is interpret the “raw” prefix in some special way. C-u 0 C-x C-s means that one doesn’t want a backup file to be created when saving the current buffer, for instance. But what if you want to save without making a backup file, and you want Emacs to flash lights and play a nice tune at the same time? You can’t, and you’re probably perfectly happy that way.

I’m not, so I’ve added a second prefix—the symbolic prefix. The prefix key is M-i (gnussymbolic-argument), and the next character typed in is the value. You can stack as many M-i prefixes as you want. M-i a M-C-u means “feed the M-C-u command the symbolic prefix a”. M-i a M-i b M-C-u means “feed the M-C-u command the symbolic prefixes a and b”. You get the drift.

Typing in symbolic prefixes to commands that don’t accept them doesn’t hurt, but it doesn’t do any good either. Currently not many Gnus functions make use of the symbolic prefix.

If you’re interested in how Gnus implements this, see Section 10.6.7 [Extended Interactive], page 226.

8.4 Formatting Variables

Throughout this manual you’ve probably noticed lots of variables called things like gnus-

group-line-format and gnus-summary-mode-line-format. These control how Gnus is to

output lines in the various buffers. There’s quite a lot of them. Fortunately, they all use the same syntax, so there’s not that much to be annoyed by.

Here’s an example format spec (from the group buffer): ‘%M%S%5y: %(%g%)\n’. We see that it is indeed extremely ugly, and that there are lots of percentages everywhere.

Currently Gnus uses the following formatting variables: gnus-group-line-format,

gnus-summary-line-format, gnus-server-line-format, gnus-topic-line-format,
gnus-group-mode-line-format, gnus-summary-mode-line-format, gnus-article-
mode-line-format, gnus-server-mode-line-format, and gnus-summary-pick-line-
format.

All these format variables can also be arbitrary elisp forms. In that case, they will be evaled to insert the required lines.

Gnus includes a command to help you while creating your own format specs. M-x gnusupdate-format will eval the current form, update the spec in question and pop you to a buffer where you can examine the resulting lisp code to be run to generate the line.

8.4.1 Formatting Basics

Each ‘%’ element will be replaced by some string or other when the buffer in question is generated. ‘%5y’ means “insert the ‘y’ spec, and pad with spaces to get a 5-character field”.

As with normal C and Emacs Lisp formatting strings, the numerical modifier between the ‘%’ and the formatting type character will pad the output so that it is always at least that long. ‘%5y’ will make the field always (at least) five characters wide by padding with spaces to the left. If you say ‘%-5y’, it will pad to the right instead.

You may also wish to limit the length of the field to protect against particularly wide values. For that you can say ‘%4,6y’, which means that the field will never be more than 6 characters wide and never less than 4 characters wide.

8.4.2 Mode Line Formatting

Mode line formatting variables (e.g., gnus-summary-mode-line-format) follow the same rules as other, buffer line oriented formatting variables (see Section 8.4.1 [Formatting Basics], page 154) with the following two differences:

1. There must be no newline (‘\n’) at the end.

2. The special ‘%%b’ spec can be used to display the buffer name. Well, it’s no spec at all, really—‘%%’ is just a way to quote ‘%’ to allow it to pass through the formatting machinery unmangled, so that Emacs receives ‘%b’, which is something the Emacs mode line display interprets to mean “show the buffer name”. For a full list of mode line specs Emacs understands, see the documentation of the mode-line-format variable.

8.4.3 Advanced Formatting

It is frequently useful to post-process the fields in some way. Padding, limiting, cutting off parts and suppressing certain values can be achieved by using tilde modifiers. A typical tilde spec might look like ‘%~(cut 3)~(ignore "0")y’.

These are the valid modifiers:

pad

pad-left Pad the field to the left with spaces until it reaches the required length.

pad-right

Pad the field to the right with spaces until it reaches the required length.

max

max-left Cut off characters from the left until it reaches the specified length.

max-right

Cut off characters from the right until it reaches the specified length.

cut

cut-left Cut off the specified number of characters from the left.

cut-right

Cut off the specified number of characters from the right.

ignore Return an empty string if the field is equal to the specified value.

form Use the specified form as the field value when the ‘@’ spec is used.

Let’s take an example. The ‘%o’ spec in the summary mode lines will return a date in compact ISO8601 format—‘19960809T230410’. This is quite a mouthful, so we want to shave off the century number and the time, leaving us with a six-character date. That would be ‘%~(cut-left 2)~(max-right 6)~(pad 6)o’. (Cutting is done before maxing, and we need the padding to ensure that the date is never less than 6 characters to make it look nice in columns.)

Ignoring is done first; then cutting; then maxing; and then as the very last operation, padding.

If you use lots of these advanced thingies, you’ll find that Gnus gets quite slow. This can be helped enormously by running M-x gnus-compile when you are satisfied with the look of your lines. See Section 8.7 [Compilation], page 160.

8.4.4 User-Defined Specs

All the specs allow for inserting user defined specifiers—‘u’. The next character in the format string should be a letter. Gnus will call the function gnus-user-format-function- ‘X’, where ‘X’ is the letter following ‘%u’. The function will be passed a single parameter— what the parameter means depends on what buffer it’s being called from. The function should return a string, which will be inserted into the buffer just like information from any other specifier. This function may also be called with dummy values, so it should protect against that.

You can also use tilde modifiers (see Section 8.4.3 [Advanced Formatting], page 155 to achieve much the same without defining new functions. Here’s an example: ‘%~(form (count-lines (point-min) (point)))@’. The form given here will be evaluated to yield the current line number, and then inserted.

8.4.5 Formatting Fonts

There are specs for highlighting, and these are shared by all the format variables. Text inside the ‘%(’ and ‘%)’ specifiers will get the special mouse-face property set, which means that it will be highlighted (with gnus-mouse-face) when you put the mouse pointer over it.

Text inside the ‘%{’ and ‘%}’ specifiers will have their normal faces set using gnus-face-0, which is bold by default. If you say ‘%1{’, you’ll get gnus-face-1 instead, and so on. Create as many faces as you wish. The same goes for the mouse-face specs—you can say

‘%3(hello%)’ to have ‘hello’ mouse-highlighted with gnus-mouse-face-3.

Here’s an alternative recipe for the group buffer:

     ;; Create three face types.
     (setq gnus-face-1 'bold)
     (setq gnus-face-3 'italic)
     ;; We want the article count to be in
     ;; a bold and green face.  So we create
     ;; a new face called `my-green-bold'.
     (copy-face 'bold 'my-green-bold)
     ;; Set the color.
     (set-face-foreground 'my-green-bold "ForestGreen")
     (setq gnus-face-2 'my-green-bold)

     ;; Set the new & fancy format.
     (setq gnus-group-line-format
           "%M%S%3{%5y%}%2[:%] %(%1{%g%}%)\n")

I’m sure you’ll be able to use this scheme to create totally unreadable and extremely vulgar displays. Have fun!

Note that the ‘%(’ specs (and friends) do not make any sense on the mode-line variables.

8.5 Windows Configuration

No, there’s nothing here about X, so be quiet.

If gnus-use-full-window non-nil, Gnus will delete all other windows and occupy the entire Emacs screen by itself. It is t by default.

gnus-buffer-configuration describes how much space each Gnus buffer should be given. Here’s an excerpt of this variable:

     ((group (vertical 1.0 (group 1.0 point)
                           (if gnus-carpal (group-carpal 4))))
      (article (vertical 1.0 (summary 0.25 point)
                             (article 1.0))))

This is an alist. The key is a symbol that names some action or other. For instance, when displaying the group buffer, the window configuration function will use group as the key. A full list of possible names is listed below.

The value (i.e., the split) says how much space each buffer should occupy. To take the article split as an example -

     (article (vertical 1.0 (summary 0.25 point)
                            (article 1.0)))

This split says that the summary buffer should occupy 25% of upper half of the screen, and that it is placed over the article buffer. As you may have noticed, 100% + 25% is actually 125% (yup, I saw y’all reaching for that calculator there). However, the special number 1.0 is used to signal that this buffer should soak up all the rest of the space available after the rest of the buffers have taken whatever they need. There should be only one buffer with the 1.0 size spec per split.

Point will be put in the buffer that has the optional third element point. In a frame split, the last subsplit having a leaf split where the tag frame-focus is a member (i.e. is the third or fourth element in the list, depending on whether the point tag is present) gets focus.

Here’s a more complicated example:

     (article (vertical 1.0 (group 4)
                            (summary 0.25 point)
                            (if gnus-carpal (summary-carpal 4))
                            (article 1.0)))

If the size spec is an integer instead of a floating point number, then that number will be used to say how many lines a buffer should occupy, not a percentage.

If the split looks like something that can be evaled (to be precise—if the car of the split is a function or a subr), this split will be evaled. If the result is non-nil, it will be used as a split. This means that there will be three buffers if gnus-carpal is nil, and four buffers if gnus-carpal is non-nil.

Not complicated enough for you? Well, try this on for size:

     (article (horizontal 1.0
                  (vertical 0.5
                      (group 1.0)
                      (gnus-carpal 4))
                  (vertical 1.0
                      (summary 0.25 point)
                      (summary-carpal 4)
                      (article 1.0))))

Whoops. Two buffers with the mystery 100% tag. And what’s that horizontal thingie?

If the first element in one of the split is horizontal, Gnus will split the window horizontally, giving you two windows side-by-side. Inside each of these strips you may carry on all you like in the normal fashion. The number following horizontal says what percentage of the screen is to be given to this strip.

For each split, there must be one element that has the 100% tag. The splitting is never accurate, and this buffer will eat any leftover lines from the splits.

To be slightly more formal, here’s a definition of what a valid split may look like:

     split       = frame | horizontal | vertical | buffer | form
     frame       = "(frame " size *split ")"
     horizontal  = "(horizontal " size *split ")"
     vertical    = "(vertical " size *split ")"
     buffer      = "(" buffer-name " " size *[ "point" ] *[ "frame-focus"] ")"
     size        = number | frame-params
     buffer-name = group | article | summary ...

The limitations are that the frame split can only appear as the top-level split. form should be an Emacs Lisp form that should return a valid split. We see that each split is fully recursive, and may contain any number of vertical and horizontal splits.

Finding the right sizes can be a bit complicated. No window may be less than gnuswindow-min-height (default 1) characters high, and all windows must be at least gnuswindow-min-width (default 1) characters wide. Gnus will try to enforce this before applying the splits. If you want to use the normal Emacs window width/height limit, you can just set these two variables to nil.

If you’re not familiar with Emacs terminology, horizontal and vertical splits may work the opposite way of what you’d expect. Windows inside a horizontal split are shown side-by-side, and windows within a vertical split are shown above each other.

If you want to experiment with window placement, a good tip is to call gnus-configure-frame directly with a split. This is the function that does all the real work when splitting buffers. Below is a pretty nonsensical configuration with 5 windows; two for the group buffer and three for the article buffer. (I said it was nonsensical.) If you eval the statement below, you can get an idea of how that would look straight away, without going through the normal Gnus channels. Play with it until you’re satisfied, and then use gnus-add-configuration to add your new creation to the buffer configuration list.

     (gnus-configure-frame
      '(horizontal 1.0
         (vertical 10
           (group 1.0)
           (article 0.3 point))
         (vertical 1.0
           (article 1.0)
           (horizontal 4
             (group 1.0)
             (article 10)))))

You might want to have several frames as well. No prob—just use the frame split:

     (gnus-configure-frame
      '(frame 1.0
              (vertical 1.0
                        (summary 0.25 point frame-focus)
                        (article 1.0))
              (vertical ((height . 5) (width . 15)
                         (user-position . t)
                         (left . -1) (top . 1))
                        (picon 1.0))))

This split will result in the familiar summary/article window configuration in the first (or “main”) frame, while a small additional frame will be created where picons will be shown. As you can see, instead of the normal 1.0 top-level spec, each additional split should have a frame parameter alist as the size spec. See section “Frame Parameters” in The GNU Emacs Lisp Reference Manual. Under XEmacs, a frame property list will be accepted, too—for

instance, (height 5 width 15 left -1 top 1) is such a plist.

Here’s a list of all possible keys for gnus-buffer-configuration:

   group, summary, article, server, browse, message, pick, info, summary-faq, edit-
group, edit-server, edit-score, post, reply, forward, reply-yank, mail-bounce,
draft, pipe, bug, compose-bounce, and score-trace.
   Note that the message key is used for both gnus-group-mail and gnus-summary-mail-

other-window. If it is desirable to distinguish between the two, something like this might be used:

     (message (horizontal 1.0
                          (vertical 1.0 (message 1.0 point))
                          (vertical 0.24
                                    (if (buffer-live-p gnus-summary-buffer)
                                        '(summary 0.5))
                                    (group 1.0)))))

Since the gnus-buffer-configuration variable is so long and complicated, there’s a function you can use to ease changing the config of a single setting: gnus-add-configuration. If, for instance, you want to change the article setting, you could say:

     (gnus-add-configuration
      '(article (vertical 1.0

                    (group 4)
                    (summary .25 point)
                    (article 1.0))))

You’d typically stick these gnus-add-configuration calls in your ‘.gnus.el’ file or in some startup hook—they should be run after Gnus has been loaded.

If all windows mentioned in the configuration are already visible, Gnus won’t change the window configuration. If you always want to force the “right” window configuration, you

can set gnus-always-force-window-configuration to non-nil.

8.6 Faces and Fonts

Fiddling with fonts and faces used to be very difficult, but these days it is very simple. You simply say M-x customize-face, pick out the face you want to alter, and alter it via the standard Customize interface.

8.7 Compilation

Remember all those line format specification variables? gnus-summary-line-format, gnus-group-line-format, and so on. Now, Gnus will of course heed whatever these variables are, but, unfortunately, changing them will mean a quite significant slow-down. (The default values of these variables have byte-compiled functions associated with them, while the user-generated versions do not, of course.)

To help with this, you can run M-x gnus-compile after you’ve fiddled around with the variables and feel that you’re (kind of) satisfied. This will result in the new specs being byte-compiled, and you’ll get top speed again. Gnus will save these compiled specs in the ‘.newsrc.eld’ file. (User-defined functions aren’t compiled by this function, though—you should compile them yourself by sticking them into the .gnus.el file and byte-compiling that file.)

8.8 Mode Lines

gnus-updated-mode-lines says what buffers should keep their mode lines updated. It is a list of symbols. Supported symbols include group, article, summary, server, browse, and tree. If the corresponding symbol is present, Gnus will keep that mode line updated with information that may be pertinent. If this variable is nil, screen refresh may be quicker.

By default, Gnus displays information on the current article in the mode lines of the summary and article buffers. The information Gnus wishes to display (e.g. the subject of the article) is often longer than the mode lines, and therefore have to be cut off at some point. The gnus-mode-non-string-length variable says how long the other elements on the line is (i.e., the non-info part). If you put additional elements on the mode line (e.g. a clock), you should modify this variable:

     (add-hook 'display-time-hook
               (lambda () (setq gnus-mode-non-string-length
                                (+ 21
                                   (if line-number-mode 5 0)

                                   (if column-number-mode 4 0)
                                   (length display-time-string)))))

If this variable is nil (which is the default), the mode line strings won’t be chopped off, and they won’t be padded either. Note that the default is unlikely to be desirable, as even the percentage complete in the buffer may be crowded off the mode line; the user should configure this variable appropriately for her configuration.

8.9 Highlighting and Menus

The gnus-visual variable controls most of the Gnus-prettifying aspects. If nil, Gnus won’t attempt to create menus or use fancy colors or fonts. This will also inhibit loading

the ‘gnus-vis.el’ file.

This variable can be a list of visual properties that are enabled. The following elements are valid, and are all included by default:

group-highlight

Do highlights in the group buffer.

summary-highlight

Do highlights in the summary buffer.

article-highlight

Do highlights in the article buffer.

highlight

Turn on highlighting in all buffers.

group-menu

Create menus in the group buffer.

summary-menu

Create menus in the summary buffers.

article-menu

Create menus in the article buffer.

browse-menu

Create menus in the browse buffer.

server-menu

Create menus in the server buffer.

score-menu

Create menus in the score buffers.

menu Create menus in all buffers.

So if you only want highlighting in the article buffer and menus in all buffers, you could say something like:

     (setq gnus-visual '(article-highlight menu))

If you want highlighting only and no menus whatsoever, you’d say:

     (setq gnus-visual '(highlight))

If gnus-visual is t, highlighting and menus will be used in all Gnus buffers.

Other general variables that influence the look of all buffers include:

gnus-mouse-face

This is the face (i.e., font) used for mouse highlighting in Gnus. No mouse highlights will be done if gnus-visual is nil.

There are hooks associated with the creation of all the different menus:

gnus-article-menu-hook

Hook called after creating the article mode menu.

gnus-group-menu-hook

Hook called after creating the group mode menu.

gnus-summary-menu-hook

Hook called after creating the summary mode menu.

gnus-server-menu-hook

Hook called after creating the server mode menu.

gnus-browse-menu-hook

Hook called after creating the browse mode menu.

gnus-score-menu-hook

Hook called after creating the score mode menu.

8.10 Buttons

Those new-fangled mouse contraptions is very popular with the young, hep kids who don’t want to learn the proper way to do things these days. Why, I remember way back in the summer of ’89, when I was using Emacs on a Tops 20 system. Three hundred users on one single machine, and every user was running Simula compilers. Bah!

Right.

Well, you can make Gnus display bufferfuls of buttons you can click to do anything by setting gnus-carpal to t. Pretty simple, really. Tell the chiropractor I sent you.

gnus-carpal-mode-hook

Hook run in all carpal mode buffers.

gnus-carpal-button-face

Face used on buttons.

gnus-carpal-header-face

Face used on carpal buffer headers.

gnus-carpal-group-buffer-buttons

Buttons in the group buffer.

gnus-carpal-summary-buffer-buttons

Buttons in the summary buffer.

gnus-carpal-server-buffer-buttons

Buttons in the server buffer.

gnus-carpal-browse-buffer-buttons

Buttons in the browse buffer.

All the buttons variables are lists. The elements in these list are either cons cells where the car contains a text to be displayed and the cdr contains a function symbol, or a simple string.

8.11 Daemons

Gnus, being larger than any program ever written (allegedly), does lots of strange stuff that you may wish to have done while you’re not present. For instance, you may want it to check for new mail once in a while. Or you may want it to close down all connections to all servers when you leave Emacs idle. And stuff like that.

Gnus will let you do stuff like that by defining various handlers. Each handler consists of three elements: A function, a time, and an idle parameter.

Here’s an example of a handler that closes connections when Emacs has been idle for thirty minutes:

     (gnus-demon-close-connections nil 30)

Here’s a handler that scans for PGP headers every hour when Emacs is idle:

     (gnus-demon-scan-pgp 60 t)

This time parameter and than idle parameter work together in a strange, but wonderful fashion. Basically, if idle is nil, then the function will be called every time minutes.

If idle is t, then the function will be called after time minutes only if Emacs is idle. So if Emacs is never idle, the function will never be called. But once Emacs goes idle, the function will be called every time minutes.

If idle is a number and time is a number, the function will be called every time minutes only when Emacs has been idle for idle minutes.

If idle is a number and time is nil, the function will be called once every time Emacs has been idle for idle minutes.

And if time is a string, it should look like ‘07:31’, and the function will then be called once every day somewhere near that time. Modified by the idle parameter, of course.

(When I say “minute” here, I really mean gnus-demon-timestep seconds. This is 60 by default. If you change that variable, all the timings in the handlers will be affected.)

To set the whole thing in motion, though, you have to set gnus-use-demon to t.

So, if you want to add a handler, you could put something like this in your ‘.gnus’ file:

     (gnus-demon-add-handler 'gnus-demon-close-connections 30 t)

Some ready-made functions to do this have been created: gnus-demon-add-nocem,

gnus-demon-add-disconnection, gnus-demon-add-nntp-close-connection, gnus-
demon-add-scan-timestamps, gnus-demon-add-rescan, and gnus-demon-add-scanmail.

Just put those functions in your ‘.gnus’ if you want those abilities.

If you add handlers to gnus-demon-handlers directly, you should run gnus-demon-init to make the changes take hold. To cancel all daemons, you can use the gnus-demon-cancel function.

Note that adding daemons can be pretty naughty if you over do it. Adding functions that scan all news and mail from all servers every two seconds is a sure-fire way of getting booted off any respectable system. So behave.

8.12 NoCeM

Spamming is posting the same article lots and lots of times. Spamming is bad. Spamming is evil.

Spamming is usually canceled within a day or so by various anti-spamming agencies. These agencies usually also send out NoCeM messages. NoCeM is pronounced “no see-’em”, and means what the name implies—these are messages that make the offending articles, like, go away.

What use are these NoCeM messages if the articles are canceled anyway? Some sites do not honor cancel messages and some sites just honor cancels from a select few people. Then you may wish to make use of the NoCeM messages, which are distributed in the

‘alt.nocem.misc’ newsgroup.

Gnus can read and parse the messages in this group automatically, and this will make spam disappear.

There are some variables to customize, of course:

gnus-use-nocem

Set this variable to t to set the ball rolling. It is nil by default.

gnus-nocem-groups

Gnus will look for NoCeM messages in the groups in this list. The

          default is ("news.lists.filters" "news.admin.net-abuse.bulletins"
          "alt.nocem.misc" "news.admin.net-abuse.announce").
gnus-nocem-issuers

There are many people issuing NoCeM messages. This list says what people you want to listen to. The default is ("Automoose-1"

          "rbraver@ohww.norman.ok.us" "clewis@ferret.ocunix.on.ca"
          "jem@xpat.com" "snowhare@xmission.com" "red@redpoll.mrfs.oh.us

(Richard E. Depew)"); fine, upstanding citizens all of them. Known despammers that you can put in this list include:

          ‘clewis@ferret.ocunix.on.ca;’

Chris Lewis—Major Canadian despammer who has probably canceled more usenet abuse than anybody else.

          ‘Automoose-1’

The CancelMoose[tm] on autopilot. The CancelMoose[tm] is reputed to be Norwegian, and was the person(s) who invented No-CeM.

‘jem@xpat.com;’

John Milburn—despammer located in Korea who is getting very busy these days.

‘red@redpoll.mrfs.oh.us (Richard E. Depew)’

Richard E. Depew—lone American despammer. He mostly cancels binary postings to non-binary groups and removes spews (regurgitated articles).

You do not have to heed NoCeM messages from all these people—just the ones you want to listen to. You also don’t have to accept all NoCeM messages from the people you like. Each NoCeM message has a type header that gives the message a (more or less, usually less) rigorous definition. Common types are ‘spam’, ‘spew’, ‘mmf’, ‘binary’, and ‘troll’. To specify this, you have to use (issuer conditions ...) elements in the list. Each condition is either a string (which is a regexp that matches types you want to use) or a list on the form (not STRING), where string is a regexp that matches types you don’t want to use.

For instance, if you want all NoCeM messages from Chris Lewis except his ‘troll’ messages, you’d say:

     ("clewis@ferret.ocunix.on.ca" ".*" (not "troll"))

On the other hand, if you just want nothing but his ‘spam’ and ‘spew’ messages, you’d say:

     ("clewis@ferret.ocunix.on.ca" (not ".*") "spew" "spam")

The specs are applied left-to-right.

gnus-nocem-verifyer

This should be a function for verifying that the NoCeM issuer is who she says she is. The default is mc-verify, which is a Mailcrypt function. If this is too slow and you don’t care for verification (which may be dangerous), you can set this variable to nil.

If you want signed NoCeM messages to be verified and unsigned messages not to be verified (but used anyway), you could do something like:

     (setq gnus-nocem-verifyer 'my-gnus-mc-verify)
     (defun my-gnus-mc-verify ()
       (not (eq 'forged
                (ignore-errors
                  (if (mc-verify)
                      t
                    'forged)))))

This might be dangerous, though.

gnus-nocem-directory

This is where Gnus will store its NoCeM cache files. The default is

‘~/News/NoCeM/’.

gnus-nocem-expiry-wait

The number of days before removing old NoCeM entries from the cache. The default is 15. If you make it shorter Gnus will be faster, but you might then see old spam.

Using NoCeM could potentially be a memory hog. If you have many living (i. e., subscribed or unsubscribed groups), your Emacs process will grow big. If this is a problem, you should kill off all (or most) of your unsubscribed groups (see Section 2.4 [Subscription Commands], page 15).

8.13 Undo

It is very useful to be able to undo actions one has done. In normal Emacs buffers, it’s easy enough—you just push the undo button. In Gnus buffers, however, it isn’t that simple.

The things Gnus displays in its buffer is of no value whatsoever to Gnus—it’s all just data designed to look nice to the user. Killing a group in the group buffer with C-k makes the line disappear, but that’s just a side-effect of the real action—the removal of the group in question from the internal Gnus structures. Undoing something like that can’t be done by the normal Emacs undo function.

Gnus tries to remedy this somewhat by keeping track of what the user does and coming up with actions that would reverse the actions the user takes. When the user then presses the undo key, Gnus will run the code to reverse the previous action, or the previous actions. However, not all actions are easily reversible, so Gnus currently offers a few key functions to be undoable. These include killing groups, yanking groups, and changing the list of read articles of groups. That’s it, really. More functions may be added in the future, but each added function means an increase in data to be stored, so Gnus will never be totally undoable.

The undoability is provided by the gnus-undo-mode minor mode. It is used if gnus-use-undo is non-nil, which is the default. The M-C-_key performs the gnus-undo command command, which should feel kinda like the normal Emacs undo command.

8.14 Moderation

If you are a moderator, you can use the ‘gnus-mdrtn.el’ package. It is not included in the standard Gnus package. Write a mail to ‘larsi@gnus.org’ and state what group you moderate, and you’ll get a copy.

The moderation package is implemented as a minor mode for summary buffers. Put

     (add-hook 'gnus-summary-mode-hook 'gnus-moderate)

in your ‘.gnus.el’ file.

If you are the moderator of ‘rec.zoofle’, this is how it’s supposed to work:

1. You split your incoming mail by matching on ‘Newsgroups:.*rec.zoofle’, which will put all the to-be-posted articles in some mail group—for instance, ‘nnml:rec.zoofle’.

2. You enter that group once in a while and post articles using the e (edit-and-post) or s (just send unedited) commands.

3. If, while reading the ‘rec.zoofle’ newsgroup, you happen upon some articles that weren’t approved by you, you can cancel them with the c command.

To use moderation mode in these two groups, say:

     (setq gnus-moderated-list
           "^nnml:rec.zoofle$\\|^rec.zoofle$")

8.15 XEmacs Enhancements

XEmacs is able to display pictures and stuff, so Gnus has taken advantage of that.

8.15.1 Picons

So... You want to slow down your news reader even more! This is a good way to do so. Its also a great way to impress people staring over your shoulder as you read news.

8.15.1.1 Picon Basics

What are Picons? To quote directly from the Picons Web site:

Picons is short for “personal icons”. They’re small, constrained images used to represent users and domains on the net, organized into databases so that the appropriate image for a given e-mail address can be found. Besides users and domains, there are picon databases for Usenet newsgroups and weather forecasts. The picons are in either monochrome XBM format or color XPM and GIF formats.

If you have a permanent connection to the Internet you can use Steve Kinzler’s Picons Search engine by setting gnus-picons-piconsearch-url to the string

‘http://www.cs.indiana.edu/picons/search.html’.

Otherwise you need a local copy of his database. For instructions on obtaining and installing the picons databases, point your Web browser at

‘http://www.cs.indiana.edu/picons/ftp/index.html’. Gnus expects picons to be in-

stalled into a location pointed to by gnus-picons-database.

8.15.1.2 Picon Requirements

To have Gnus display Picons for you, you must be running XEmacs 19.13 or greater since all other versions of Emacs aren’t yet able to display images.

Additionally, you must have x support compiled into XEmacs. To display color picons which are much nicer than the black & white one, you also need one of xpm or gif compiled into XEmacs.

If you want to display faces from X-Face headers, you should have the xface support compiled into XEmacs. Otherwise you must have the netpbm utilities installed, or munge the gnus-picons-convert-x-face variable to use something else.

8.15.1.3 Easy Picons

To enable displaying picons, simply put the following line in your ‘~/.gnus’ file and start Gnus.

(setq gnus-use-picons t)
(add-hook 'gnus-article-display-hook
     'gnus-article-display-picons t)
(add-hook 'gnus-article-display-hook
     'gnus-picons-article-display-x-face)

and make sure gnus-picons-database points to the directory containing the Picons databases.

Alternatively if you want to use the web piconsearch engine add this:

(setq gnus-picons-piconsearch-url
 "http://www.cs.indiana.edu:800/piconsearch")

8.15.1.4 Hard Picons

Gnus can display picons for you as you enter and leave groups and articles. It knows how to interact with three sections of the picons database. Namely, it can display the picons newsgroup pictures, author’s face picture(s), and the authors domain. To enable this feature, you need to select where to get the picons from, and where to display them.

gnus-picons-database

The location of the picons database. Should point to a directory containing the ‘news’, ‘domains’, ‘users’ (and so on) subdirectories. This is only useful if

gnus-picons-piconsearch-url is nil. Defaults to ‘/usr/local/faces/’.
gnus-picons-piconsearch-url

The URL for the web picons search engine. The only currently known engine

is ‘http://www.cs.indiana.edu:800/piconsearch’. To workaround network

delays, icons will be fetched in the background. If this is nil ’the default), then picons are fetched from local database indicated by gnus-picons-database.

gnus-picons-display-where

Where the picon images should be displayed. It is picons by default (which by default maps to the buffer ‘*Picons*’). Other valid places could be article, summary, or ‘*scratch*’ for all I care. Just make sure that you’ve made the buffer visible using the standard Gnus window configuration routines—see Section 8.5 [Windows Configuration], page 157.

gnus-picons-group-excluded-groups

Groups that are matched by this regexp won’t have their group icons displayed.

Note: If you set gnus-use-picons to t, it will set up your window configuration for you to include the picons buffer.

Now that you’ve made those decision, you need to add the following functions to the appropriate hooks so these pictures will get displayed at the right time.

gnus-article-display-picons

Looks up and displays the picons for the author and the author’s domain in the

gnus-picons-display-where buffer. Should be added to the gnus-article-
display-hook.
gnus-picons-article-display-x-face

Decodes and displays the X-Face header if present. This function should be

added to gnus-article-display-hook.

Note: You must append them to the hook, so make sure to specify ’t’ for the append

flag of add-hook:
(add-hook 'gnus-article-display-hook 'gnus-article-display-picons t)

8.15.1.5 Picon Useless Configuration

The following variables offer further control over how things are done, where things are located, and other useless stuff you really don’t need to worry about.

gnus-picons-news-directories

List of subdirectories to search in gnus-picons-database for newsgroups faces. ("news") is the default.

gnus-picons-user-directories

List of subdirectories to search in gnus-picons-database for user faces.

("local" "users" "usenix" "misc") is the default.
gnus-picons-domain-directories

List of subdirectories to search in gnus-picons-database for domain name faces. Defaults to ("domains"). Some people may want to add ‘"unknown"’ to this list.

gnus-picons-convert-x-face

If you don’t have xface support builtin XEmacs, this is the command to use to convert the X-Face header to an X bitmap (xbm). Defaults to (format

"{ echo '/* Width=48, Height=48 */'; uncompface; } | icontopbm |
pbmtoxbm > %s" gnus-picons-x-face-file-name)
gnus-picons-x-face-file-name

Names a temporary file to store the X-Face bitmap in. Defaults to (format

"/tmp/picon-xface.%s.xbm" (user-login-name)).
gnus-picons-has-modeline-p
If you have set gnus-picons-display-where to picons, your XEmacs frame

will become really cluttered. To alleviate this a bit you can set gnus-picons-has-modeline-p to nil; this will remove the mode line from the Picons buffer.

This is only useful if gnus-picons-display-where is picons.
gnus-picons-refresh-before-display

If non-nil, display the article buffer before computing the picons. Defaults to

nil.
gnus-picons-display-as-address

If t display textual email addresses along with pictures. Defaults to t.

gnus-picons-file-suffixes

Ordered list of suffixes on picon file names to try. Defaults to ("xpm" "gif" "xbm") minus those not builtin your XEmacs.

gnus-picons-display-article-move-p

Whether to move point to first empty line when displaying picons. This has only an effect if ‘gnus-picons-display-where’ has value ‘article’.

gnus-picons-clear-cache-on-shutdown

Whether to clear the picons cache when exiting gnus. Gnus caches every picons it finds while it is running. This saves some time in the search process but eats some memory. If this variable is set to nil, Gnus will never clear the cache itself; you will have to manually call gnus-picons-clear-cache to clear it. Otherwise the cache will be cleared every time you exit Gnus. Defaults to t.

8.15.2 Smileys

Smiley is a package separate from Gnus, but since Gnus is currently the only package that uses Smiley, it is documented here.

In short—to use Smiley in Gnus, put the following in your ‘.gnus.el’ file:

     (add-hook 'gnus-article-display-hook 'gnus-smiley-display t)

Smiley maps text smiley faces—‘:-)’, ‘:-=’, ‘:-(’ and the like—to pictures and displays those instead of the text smiley faces. The conversion is controlled by a list of regexps that matches text and maps that to file names.

Smiley supplies two example conversion alists by default: smiley-deformed-regexp-alist (which matches ‘:)’, ‘:(’ and so on), and smiley-nosey-regexp-alist (which matches ‘:-)’, ‘:-(’ and so on).

The alist used is specified by the smiley-regexp-alist variable, which defaults to the

value of smiley-deformed-regexp-alist.

The first item in each element is the regexp to be matched; the second element is the regexp match group that is to be replaced by the picture; and the third element is the name of the file to be displayed.

The following variables customize where Smiley will look for these files, as well as the color to be used and stuff:

smiley-data-directory

Where Smiley will look for smiley faces files.

smiley-flesh-color

Skin color. The default is ‘yellow’, which is really racist.

smiley-features-color

Color of the features of the face. The default is ‘black’.

smiley-tongue-color

Color of the tongue. The default is ‘red’.

smiley-circle-color

Color of the circle around the face. The default is ‘black’.

smiley-mouse-face

Face used for mouse highlighting over the smiley face.

8.15.3 Toolbar

gnus-use-toolbar

If nil, don’t display toolbars. If non-nil, it should be one of default-toolbar,

          top-toolbar, bottom-toolbar, right-toolbar, or left-toolbar.
gnus-group-toolbar

The toolbar in the group buffer.

gnus-summary-toolbar

The toolbar in the summary buffer.

gnus-summary-mail-toolbar

The toolbar in the summary buffer of mail groups.

8.15.4 Various XEmacs Variables

gnus-xmas-glyph-directory

This is where Gnus will look for pictures. Gnus will normally auto-detect this directory, but you may set it manually if you have an unusual directory structure.

gnus-xmas-logo-color-alist

This is an alist where the key is a type symbol and the values are the foreground and background color of the splash page glyph.

gnus-xmas-logo-color-style

This is the key used to look up the color in the alist described above. Valid

          values include flame, pine, moss, irish, sky, tin, velvet, grape, labia,
          berry, neutral, and september.
gnus-xmas-modeline-glyph

A glyph displayed in all Gnus mode lines. It is a tiny gnu head by default.

8.16 Fuzzy Matching

Gnus provides fuzzy matching of Subject lines when doing things like scoring, thread gathering and thread comparison.

As opposed to regular expression matching, fuzzy matching is very fuzzy. It’s so fuzzy that there’s not even a definition of what fuzziness means, and the implementation has changed over time.

Basically, it tries to remove all noise from lines before comparing. ‘Re: ’, parenthetical remarks, white space, and so on, are filtered out of the strings before comparing the results. This often leads to adequate results—even when faced with strings generated by text manglers masquerading as newsreaders.

8.17 Thwarting Email Spam

In these last days of the Usenet, commercial vultures are hanging about and grepping through news like crazy to find email addresses they can foist off their scams and products to. As a reaction to this, many people have started putting nonsense addresses into their From lines. I think this is counterproductive—it makes it difficult for people to send you legitimate mail in response to things you write, as well as making it difficult to see who wrote what. This rewriting may perhaps be a bigger menace than the unsolicited commercial email itself in the end.

The biggest problem I have with email spam is that it comes in under false pretenses. I press g and Gnus merrily informs me that I have 10 new emails. I say “Golly gee! Happy is me!” and select the mail group, only to find two pyramid schemes, seven advertisements (“New! Miracle tonic for growing full, lustrous hair on your toes!”) and one mail asking me to repent and find some god.

This is annoying.

The way to deal with this is having Gnus split out all spam into a ‘spam’ mail group (see Section 6.3.2 [Splitting Mail], page 100).

First, pick one (1) valid mail address that you can be reached at, and put it in your From header of all your news articles. (I’ve chosen ‘larsi@trym.ifi.uio.no’, but for many addresses on the form ‘larsi+usenet@ifi.uio.no’ will be a better choice. Ask your sysadm whether your sendmail installation accepts keywords in the local part of the mail address.)

     (setq message-default-news-headers
           "From: Lars Magne Ingebrigtsen <larsi@trym.ifi.uio.no>\n")

Then put the following split rule in nnmail-split-fancy (see Section 6.3.4 [Fancy Mail Splitting], page 103):

     (
      ...
      (to "larsi@trym.ifi.uio.no"
           (| ("subject" "re:.*" "misc")
              ("references" ".*@.*" "misc")
              "spam"))
      ...
     )

This says that all mail to this address is suspect, but if it has a Subject that starts with a ‘Re:’ or has a References header, it’s probably ok. All the rest goes to the ‘spam’ group. (This idea probably comes from Tim Pierce.)

In addition, many mail spammers talk directly to your smtp server and do not include your email address explicitly in the To header. Why they do this is unknown—perhaps it’s to thwart this thwarting scheme? In any case, this is trivial to deal with—you just put anything not addressed to you in the ‘spam’ group by ending your fancy split rule in this way:

     (
      ...
      (to "larsi" "misc")
      "spam")

In my experience, this will sort virtually everything into the right group. You still have to check the ‘spam’ group from time to time to check for legitimate mail, though. If you feel like being a good net citizen, you can even send off complaints to the proper authorities on each unsolicited commercial email—at your leisure.

If you are also a lazy net citizen, you will probably prefer complaining automatically with the ‘gnus-junk.el’ package, available FOR FREE at

‘<URL:http://stud2.tuwien.ac.at/~e9426626/gnus-junk.html>’. Since most e-mail

spam is sent automatically, this may reconcile the cosmic balance somewhat.

This works for me. It allows people an easy way to contact me (they can just press r in the usual way), and I’m not bothered at all with spam. It’s a win-win situation. Forging From headers to point to non-existent domains is yucky, in my opinion.

8.18 Various Various

gnus-home-directory

All Gnus path variables will be initialized from this variable, which defaults to ‘~/’.

gnus-directory

Most Gnus storage path variables will be initialized from this variable, which defaults to the ‘SAVEDIR’ environment variable, or ‘~/News/’ if that variable isn’t set. Note that Gnus is mostly loaded when the ‘.gnus.el’ file is read. This means that other directory variables that are initialized from this variable won’t be set properly if you set this variable in ‘.gnus.el’. Set this variable in ‘.emacs’ instead.

gnus-default-directory

Not related to the above variable at all—this variable says what the default directory of all Gnus buffers should be. If you issue commands like C-x C-f, the prompt you’ll get starts in the current buffer’s default directory. If this variable is nil (which is the default), the default directory will be the default directory of the buffer you were in when you started Gnus.

gnus-verbose

This variable is an integer between zero and ten. The higher the value, the more messages will be displayed. If this variable is zero, Gnus will never flash any messages, if it is seven (which is the default), most important messages will be shown, and if it is ten, Gnus won’t ever shut up, but will flash so many messages it will make your head swim.

gnus-verbose-backends

This variable works the same way as gnus-verbose, but it applies to the Gnus backends instead of Gnus proper.

nnheader-max-head-length

When the backends read straight heads of articles, they all try to read as little as possible. This variable (default 4096) specifies the absolute max length the backends will try to read before giving up on finding a separator line between the head and the body. If this variable is nil, there is no upper read bound. If it is t, the backends won’t try to read the articles piece by piece, but read the entire articles. This makes sense with some versions of ange-ftp or efs.

nnheader-head-chop-length

This variable (default 2048) says how big a piece of each article to read when doing the operation described above.

nnheader-file-name-translation-alist

This is an alist that says how to translate characters in file names. For instance, if ‘:’ is invalid as a file character in file names on your system (you OS/2 user you), you could say something like:

     (setq nnheader-file-name-translation-alist
           '((?: . ?_)))

In fact, this is the default value for this variable on OS/2 and MS Windows (phooey) systems.

gnus-hidden-properties

This is a list of properties to use to hide “invisible” text. It is (invisible t intangible t) by default on most systems, which makes invisible text invisible and intangible.

gnus-parse-headers-hook

A hook called before parsing headers. It can be used, for instance, to gather statistics on the headers fetched, or perhaps you’d like to prune some headers. I don’t see why you’d want that, though.

gnus-shell-command-separator

String used to separate two shell commands. The default is ‘;’.

9 The End

Well, that’s the manual—you can get on with your life now. Keep in touch. Say hello to your cats from me.

My ghod—I just can’t stand goodbyes. Sniffle. Ol’ Charles Reznikoff said it pretty well, so I leave the floor to him:

Te Deum

Not because of victories I sing, having none, but for the common sunshine, the breeze, the largess of the spring.

Not for victory but for the day’s work done as well as I was able; not for a seat upon the dais but at the common table.

176 Gnus Manual

10 Appendices

10.1 History

gnus was written by Masanobu Umeda. When autumn crept up in ’94, Lars Magne Ingebrigtsen grew bored and decided to rewrite Gnus.

If you want to investigate the person responsible for this outrage, you can point your (feh!) web browser to ‘http://www.stud.ifi.uio.no/~larsi/’. This is also the primary distribution point for the new and spiffy versions of Gnus, and is known as The Site That Destroys Newsrcs And Drives People Mad.

During the first extended alpha period of development, the new Gnus was called “(ding) Gnus”. (ding) is, of course, short for ding is not Gnus, which is a total and utter lie, but who cares? (Besides, the “Gnus” in this abbreviation should probably be pronounced “news” as Umeda intended, which makes it a more appropriate name, don’t you think?)

In any case, after spending all that energy on coming up with a new and spunky name, we decided that the name was too spunky, so we renamed it back again to “Gnus”. But in mixed case. “Gnus” vs. “gnus”. New vs. old.

The first “proper” release of Gnus 5 was done in November 1995 when it was included in the Emacs 19.30 distribution (132 (ding) Gnus releases plus 15 Gnus 5.0 releases).

In May 1996 the next Gnus generation (aka. “September Gnus” (after 99 releases)) was released under the name “Gnus 5.2” (40 releases).

On July 28th 1996 work on Red Gnus was begun, and it was released on January 25th 1997 (after 84 releases) as “Gnus 5.4” (67 releases).

On September 13th 1997, Quassia Gnus was started and lasted 37 releases. If was released as “Gnus 5.6 on March 8th 1998.

If you happen upon a version of Gnus that has a prefixed name – “(ding) Gnus”, “September Gnus”, “Red Gnus”, “Quassia Gnus” – don’t panic. Don’t let it know that you’re frightened. Back away. Slowly. Whatever you do, don’t run. Walk away, calmly, until you’re out of its reach. Find a proper released version of Gnus and snuggle up to that instead.

10.1.1 Why?

What’s the point of Gnus?

I want to provide a “rad”, “happening”, “way cool” and “hep” newsreader, that lets you do anything you can think of. That was my original motivation, but while working on Gnus, it has become clear to me that this generation of newsreaders really belong in the stone age. Newsreaders haven’t developed much since the infancy of the net. If the volume continues to rise with the current rate of increase, all current newsreaders will be pretty much useless. How do you deal with newsgroups that have thousands of new articles each day? How do you keep track of millions of people who post?

Gnus offers no real solutions to these questions, but I would very much like to see Gnus being used as a testing ground for new methods of reading and fetching news. Expanding on Umeda-san’s wise decision to separate the newsreader from the backends, Gnus now offers a simple interface for anybody who wants to write new backends for fetching mail and news from different sources. I have added hooks for customizations everywhere I could imagine it being useful. By doing so, I’m inviting every one of you to explore and invent.

   May Gnus never be complete. C-u 100 M-x all-hail-emacs and C-u 100 M-x all-
hail-xemacs.

10.1.2 Compatibility

Gnus was designed to be fully compatible with gnus. Almost all key bindings have been kept. More key bindings have been added, of course, but only in one or two obscure cases have old bindings been changed.

Our motto is:

In a cloud bones of steel.

All commands have kept their names. Some internal functions have changed their names. The gnus-uu package has changed drastically. See Section 3.16 [Decoding Articles],

page 59.

One major compatibility question is the presence of several summary buffers. All variables relevant while reading a group are buffer-local to the summary buffer they belong in. Although many important variables have their values copied into their global counterparts whenever a command is executed in the summary buffer, this change might lead to incorrect values being used unless you are careful.

All code that relies on knowledge of gnus internals will probably fail. To take two examples: Sorting gnus-newsrc-alist (or changing it in any way, as a matter of fact) is strictly verboten. Gnus maintains a hash table that points to the entries in this alist (which speeds up many functions), and changing the alist directly will lead to peculiar results.

Old hilit19 code does not work at all. In fact, you should probably remove all hilit

code from all Gnus hooks (gnus-group-prepare-hook and gnus-summary-prepare-hook).

Gnus provides various integrated functions for highlighting. These are faster and more accurate. To make life easier for everybody, Gnus will by default remove all hilit calls from all hilit hooks. Uncleanliness! Away!

Packages like expire-kill will no longer work. As a matter of fact, you should probably remove all old gnus packages (and other code) when you start using Gnus. More likely than not, Gnus already does what you have written code to make gnus do. (Snicker.)

Even though old methods of doing things are still supported, only the new methods are documented in this manual. If you detect a new method of doing something while reading this manual, that does not mean you have to stop doing it the old way.

Gnus understands all gnus startup files.

Overall, a casual user who hasn’t written much code that depends on gnus internals should suffer no problems. If problems occur, please let me know by issuing that magic

command M-x gnus-bug.

If you are in the habit of sending bug reports very often, you may find the helpful help buffer annoying after a while. If so, set gnus-bug-create-help-buffer to nil to avoid having it pop up at you.

10.1.3 Conformity

No rebels without a clue here, ma’am. We conform to all standards known to (wo)man. Except for those standards and/or conventions we disagree with, of course.

RFC 822 There are no known breaches of this standard.

RFC 1036 There are no known breaches of this standard, either.

Son-of-RFC 1036 We do have some breaches to this one.

MIME Gnus does no MIME handling, and this standard-to-be seems to think that MIME is the bees’ knees, so we have major breakage here.

X-Newsreader This is considered to be a “vanity header”, while I consider it to be consumer information. After seeing so many badly formatted articles coming from tin and Netscape I know not to use either of those for posting articles. I would not have known that if it wasn’t for the X-Newsreader header.

If you ever notice Gnus acting non-compliant with regards to the texts mentioned above, don’t hesitate to drop a note to Gnus Towers and let us know.

10.1.4 Emacsen

Gnus should work on :

• Emacs 19.32 and up.

• XEmacs 19.14 and up.

• Mule versions based on Emacs 19.32 and up.

Gnus will absolutely not work on any Emacsen older than that. Not reliably, at least. There are some vague differences between Gnus on the various platforms—XEmacs fea-

tures more graphics (a logo and a toolbar)—but other than that, things should look pretty much the same under all Emacsen.

10.1.5 Contributors

The new Gnus version couldn’t have been done without the help of all the people on the (ding) mailing list. Every day for over a year I have gotten billions of nice bug reports from them, filling me with joy, every single one of them. Smooches. The people on the list have been tried beyond endurance, what with my “oh, that’s a neat idea <type type>, yup, I’ll release it right away <ship off> no wait, that doesn’t work at all <type type>, yup, I’ll ship that one off right away <ship off> no, wait, that absolutely does not work” policy for releases. Micro$oft—bah. Amateurs. I’m much worse. (Or is that “worser”? “much worser”? “worsest”?)

I would like to take this opportunity to thank the Academy for... oops, wrong show.

• Masanobu Umeda—the writer of the original gnus.

• Per Abrahamsen—custom, scoring, highlighting and soup code (as well as numerous other things).

• Luis Fernandes—design and graphics.

• Erik Naggum—help, ideas, support, code and stuff.

• Wes Hardaker—‘gnus-picon.el’ and the manual section on picons (see Section 8.15.1 [Picons], page 167).

• Kim-Minh Kaplan—further work on the picon code.

• Brad Miller—‘gnus-gl.el’ and the GroupLens manual section (see Section 7.14 [GroupLens], page 147).

• Sudish Joseph—innumerable bug fixes.

• Ilja Weis—‘gnus-topic.el’.

• Steven L. Baur—lots and lots and lots of bugs detections and fixes.

• Vladimir Alexiev—the refcard and reference booklets.

• Felix Lee & Jamie Zawinski—I stole some pieces from the XGnus distribution by Felix Lee and JWZ.

• Scott Byer—‘nnfolder.el’ enhancements & rewrite.

• Peter Mutsaers—orphan article scoring code.

• Ken Raeburn—POP mail support.

• Hallvard B Furuseth—various bits and pieces, especially dealing with .newsrc files.

• Brian Edmonds—‘gnus-bbdb.el’.

• David Moore—rewrite of ‘nnvirtual.el’ and many other things.

• Kevin Davidson—came up with the name ding, so blame him.

• Franois Pinard—many, many interesting and thorough bug reports, as well as autoconf support. This manual was proof-read by Adrian Aichner, with Ricardo Nassif, Mark Borges, and

Jost Krieger proof-reading parts of the manual.

The following people have contributed many patches and suggestions: Christopher Davis, Andrew Eskilsson, Kai Grossjohann, David Kgedal, Richard Pieri,

Fabrice Popineau, Daniel Quinlan, Jason L. Tibbitts, III, and Jack Vinson.

Also thanks to the following for patches and stuff: Jari Aalto, Adrian Aichner, Vladimir Alexiev, Russ Allbery, Peter Arius, Matt Arm-

strong, Marc Auslander, Frank Bennett, Robert Bihlmeyer, Chris Bone, Mark Borges, Mark Boyns, Lance A. Brown, Kees de Bruin, Martin Buchholz, Joe Buehler, Kevin Buhr, Alastair Burt, Joao Cachopo, Zlatko Calusic, Massimo Campostrini, Castor, David Charlap, Dan Christensen, Kevin Christian, Michael R. Cook, Glenn Coombs, Frank D. Cringle, Geoffrey T. Dairiki, Andre Deparade, Ulrik Dickow, Dave Disser, Rui-Tao Dong, Joev Dubach, Michael Welsh Duggan, Dave Edmondson, Paul Eggert, Enami Tsugutomo, Michael Ernst, Luc Van Eycken, Sam Falkner, Nelson Jose dos Santos Ferreira, Sigbjorn Finne, Decklin Foster, Gary D. Foster, Paul Franklin, Guy Geens, Arne Georg Gleditsch, David S. Goldberg, Michelangelo Grigni, D. Hall, Magnus Hammerin, Kenichi Handa, Raja R. Harinath, Yoshiki Hayashi, P. E. Jareth Hein, Hisashige Kenji, Marc Horowitz, Gunnar Horrigmo, Richard Hoskins, Brad Howes, Franois Felix Ingrand, Ishikawa Ichiro, Lee Iverson, Iwamuro Motonori, Rajappa Iyer, Andreas Jaeger, Randell Jesup, Fred Johansen, Gareth Jones, Simon Josefsson, Greg Klanderman, Karl Kleinpaste, Peter Skov Knudsen, Shuhei Kobayashi, Koseki Yoshinori, Thor Kristoffersen, Jens Lautenbacher, Martin Larose, Seokchan Lee, Carsten Leonhardt, James LewisMoss, Christian Limpach, Markus Linnala, Dave Love, Mike McEwan, Tonny Madsen, Shlomo Mahlab, Nat Makarevitch, Istvan Marko, David Martin, Jason R. Mastaler, Gordon Matzigkeit, Timo Metzemakers, Richard Mlynarik, Lantz Moore, Morioka Tomohiko, Erik Toubro Nielsen, Hrvoje Niksic, Andy Norman, Fred Oberhauser, C. R. Oldham, Alexandre Oliva, Ken Olstad, Masaharu Onishi, Hideki Ono, William Perry, Stephen Peters, Jens-Ulrik Holger Petersen, Ulrich Pfeifer, Matt Pharr, John McClary Prevost, Bill Pringlemeir, Mike Pullen, Jim Radford, Colin Rafferty, Lasse Rasinen, Lars Balker Rasmussen, Joe Reiss, Renaud Rioboo, Roland B. Roberts, Bart Robinson, Christian von Roques, Jason Rumney, Wolfgang Rupprecht, Jay Sachs, Dewey M. Sasser, Loren Schall, Dan Schmidt, Ralph Schleicher, Philippe Schnoebelen, Andreas Schwab, Randal L. Schwartz, Justin Sheehy, Danny Siu, Matt Simmons, Paul D. Smith, Jeff Sparkes, Toby Speight, Michael Sperber, Darren Stalder, Richard Stallman, Greg Stark, Sam Steingold, Paul Stodghill, Kurt Swanson, Samuel Tardieu, Teddy, Chuck Thompson, Philippe Troin, James Troup, Trung Tran-Duc, Aaron M. Ucko, Aki Vehtari, Didier Verna, Jan Vroonhof, Stefan Waldherr, Pete Ware, Barry A. Warsaw, Christoph Wedler, Joe Wells, Katsumi Yamaoka, and Shenghuo Zhu.

For a full overview of what each person has done, the ChangeLogs included in the Gnus alpha distributions should give ample reading (550kB and counting).

Apologies to everybody that I’ve forgotten, of which there are many, I’m sure.

Gee, that’s quite a list of people. I guess that must mean that there actually are people who are using Gnus. Who’d’a thunk it!

10.1.6 New Features

These lists are, of course, just short overviews of the most important new features. No, really. There are tons more. Yes, we have feeping creaturism in full effect.

10.1.6.1 (ding) Gnus

New features in Gnus 5.0/5.1:

• The look of all buffers can be changed by setting format-like variables (see Section 2.1 [Group Buffer Format], page 11 and see Section 3.1 [Summary Buffer Format], page 33).

• Local spool and several nntp servers can be used at once (see Chapter 6 [Select Methods], page 91).

• You can combine groups into virtual groups (see Section 6.5.1 [Virtual Groups], page 122).

• You can read a number of different mail formats (see Section 6.3 [Getting Mail], page 100). All the mail backends implement a convenient mail expiry scheme (see Section 6.3.7 [Expiring Mail], page 106).

• Gnus can use various strategies for gathering threads that have lost their roots (thereby gathering loose sub-threads into one thread) or it can go back and retrieve enough headers to build a complete thread (see Section 3.9.1 [Customizing Threading], page 47).

• Killed groups can be displayed in the group buffer, and you can read them as well (see Section 2.11 [Listing Groups], page 22).

• Gnus can do partial group updates—you do not have to retrieve the entire active file just to check for new articles in a few groups (see Section 1.10 [The Active File], page 9).

• Gnus implements a sliding scale of subscribedness to groups (see Section 2.6 [Group Levels], page 16).

• You can score articles according to any number of criteria (see Chapter 7 [Scoring], page 131). You can even get Gnus to find out how to score articles for you (see Section 7.6 [Adaptive Scoring], page 140).

• Gnus maintains a dribble buffer that is auto-saved the normal Emacs manner, so it should be difficult to lose much data on what you have read if your machine should go down (see Section 1.9 [Auto Save], page 9).

• Gnus now has its own startup file (‘.gnus’) to avoid cluttering up the ‘.emacs’ file.

• You can set the process mark on both groups and articles and perform operations on all the marked items (see Section 8.1 [Process/Prefix], page 153).

• You can grep through a subset of groups and create a group from the results (see Section 6.5.2 [Kibozed Groups], page 123).

• You can list subsets of groups according to, well, anything (see Section 2.11 [Listing Groups], page 22).

• You can browse foreign servers and subscribe to groups from those servers (see Section 2.14 [Browse Foreign Server], page 24).

• Gnus can fetch articles, asynchronously, on a second connection to the server (see Section 3.11 [Asynchronous Fetching], page 53).

• You can cache articles locally (see Section 3.12 [Article Caching], page 54).

• The uudecode functions have been expanded and generalized (see Section 3.16 [Decoding Articles], page 59).

• You can still post uuencoded articles, which was a little-known feature of gnus’ past (see Section 3.16.5.3 [Uuencoding and Posting], page 62).

• Fetching parents (and other articles) now actually works without glitches (see Section 3.20 [Finding the Parent], page 71).

• Gnus can fetch FAQs and group descriptions (see Section 2.17.2 [Group Information], page 30).

• Digests (and other files) can be used as the basis for groups (see Section 6.4.3 [Document Groups], page 114).

• Articles can be highlighted and customized (see Section 4.3 [Customizing Articles], page 83).

• URLs and other external references can be buttonized (see Section 3.17.5 [Article Buttons], page 68).

• You can do lots of strange stuff with the Gnus window & frame configuration (see Section 8.5 [Windows Configuration], page 157).

• You can click on buttons instead of using the keyboard (see Section 8.10 [Buttons], page 162).

10.1.6.2 September Gnus

New features in Gnus 5.2/5.3:

• A new message composition mode is used. All old customization variables for mail-

  mode, rnews-reply-mode and gnus-msg are now obsolete.

• Gnus is now able to generate sparse threads—threads where missing articles are rep-

resented by empty nodes (see Section 3.9.1 [Customizing Threading], page 47).

       (setq gnus-build-sparse-threads 'some)

• Outgoing articles are stored on a special archive server (see Section 5.5 [Archived Messages], page 86).

• Partial thread regeneration now happens when articles are referred.

• Gnus can make use of GroupLens predictions (see Section 7.14 [GroupLens], page 147).

• Picons (personal icons) can be displayed under XEmacs (see Section 8.15.1 [Picons], page 167).

• A trn-like tree buffer can be displayed (see Section 3.22 [Tree Display], page 73).

       (setq gnus-use-trees t)

• An nn-like pick-and-read minor mode is available for the summary buffers (see Sec-

tion 3.21.1 [Pick and Read], page 72).

       (add-hook 'gnus-summary-mode-hook 'gnus-pick-mode)

• In binary groups you can use a special binary minor mode (see Section 3.21.2 [Binary Groups], page 73).

• Groups can be grouped in a folding topic hierarchy (see Section 2.16 [Group Topics],

page 25).

       (add-hook 'gnus-group-mode-hook 'gnus-topic-mode)

• Gnus can re-send and bounce mail (see Section 3.5.1 [Summary Mail Commands], page 39).

• Groups can now have a score, and bubbling based on entry frequency is possible (see

Section 2.7 [Group Score], page 17).

       (add-hook 'gnus-summary-exit-hook 'gnus-summary-bubble-group)

• Groups can be process-marked, and commands can be performed on groups of groups (see Section 2.8 [Marking Groups], page 17).

• Caching is possible in virtual groups.

• nndoc now understands all kinds of digests, mail boxes, rnews news batches, ClariNet briefs collections, and just about everything else (see Section 6.4.3 [Document Groups], page 114).

• Gnus has a new backend (nnsoup) to create/read SOUP packets (see Section 6.4.4 [SOUP], page 117).

• The Gnus cache is much faster.

• Groups can be sorted according to many criteria (see Section 2.12 [Sorting Groups], page 23).

• New group parameters have been introduced to set list-addresses and expiry times (see Section 2.10 [Group Parameters], page 19).

• All formatting specs allow specifying faces to be used (see Section 8.4.5 [Formatting Fonts], page 156).

• There are several more commands for setting/removing/acting on process marked articles on the M P submap (see Section 3.7.5 [Setting Process Marks], page 44).

• The summary buffer can be limited to show parts of the available articles based on a wide range of criteria. These commands have been bound to keys on the / submap (see Section 3.8 [Limiting], page 45).

• Articles can be made persistent with the * command (see Section 3.13 [Persistent Articles], page 55).

• All functions for hiding article elements are now toggles.

• Article headers can be buttonized (see Section 3.17.4 [Article Washing], page 66).

       (add-hook 'gnus-article-display-hook
                 'gnus-article-add-buttons-to-head)

• All mail backends support fetching articles by Message-ID.

• Duplicate mail can now be treated properly (see Section 6.3.9 [Duplicates], page 109).

• All summary mode commands are available directly from the article buffer (see Section 4.4 [Article Keymap], page 83).

• Frames can be part of gnus-buffer-configuration (see Section 8.5 [Windows Configuration], page 157).

• Mail can be re-scanned by a daemonic process (see Section 8.11 [Daemons], page 163).

• Gnus can make use of NoCeM files to weed out spam (see Section 8.12 [NoCeM],

page 164).

       (setq gnus-use-nocem t)

• Groups can be made permanently visible (see Section 2.11 [Listing Groups], page 22).

       (setq gnus-permanently-visible-groups "^nnml:")

• Many new hooks have been introduced to make customizing easier.

• Gnus respects the Mail-Copies-To header.

• Threads can be gathered by looking at the References header (see Section 3.9.1 [Cus-

tomizing Threading], page 47).

       (setq gnus-summary-thread-gathering-function
             'gnus-gather-threads-by-references)

• Read articles can be stored in a special backlog buffer to avoid refetching (see Sec-

tion 3.14 [Article Backlog], page 55).

       (setq gnus-keep-backlog 50)

• A clean copy of the current article is always stored in a separate buffer to allow easier treatment.

• Gnus can suggest where to save articles (see Section 3.15 [Saving Articles], page 56).

• Gnus doesn’t have to do as much prompting when saving (see Section 3.15 [Saving

Articles], page 56).

       (setq gnus-prompt-before-saving t)

• gnus-uu can view decoded files asynchronously while fetching articles (see Section 3.16.5.2 [Other Decode Variables], page 61).

       (setq gnus-uu-grabbed-file-functions 'gnus-uu-grab-view)

• Filling in the article buffer now works properly on cited text (see Section 3.17.4 [Article Washing], page 66).

• Hiding cited text adds buttons to toggle hiding, and how much cited text to hide is

now customizable (see Section 3.17.3 [Article Hiding], page 65).

       (setq gnus-cited-lines-visible 2)

• Boring headers can be hidden (see Section 3.17.3 [Article Hiding], page 65).

       (add-hook 'gnus-article-display-hook
                 'gnus-article-hide-boring-headers t)

• Default scoring values can now be set from the menu bar.

• Further syntax checking of outgoing articles have been added.

10.1.6.3 Red Gnus

New features in Gnus 5.4/5.5:

• ‘nntp.el’ has been totally rewritten in an asynchronous fashion.

• Article prefetching functionality has been moved up into Gnus (see Section 3.11 [Asynchronous Fetching], page 53).

• Scoring can now be performed with logical operators like and, or, not, and parent redirection (see Section 7.15 [Advanced Scoring], page 149).

• Article washing status can be displayed in the article mode line (see Section 4.5 [Misc Article], page 84).

• ‘gnus.el’ has been split into many smaller files.

• Suppression of duplicate articles based on Message-ID can be done (see Section 3.27

[Duplicate Suppression], page 80).

       (setq gnus-suppress-duplicates t)

• New variables for specifying what score and adapt files are to be considered home score and adapt files (see Section 7.7 [Home Score File], page 142) have been added.

• nndoc was rewritten to be easily extendable (see Section 6.4.3.1 [Document Server Internals], page 115).

• Groups can inherit group parameters from parent topics (see Section 2.16.5 [Topic Parameters], page 29).

• Article editing has been revamped and is now actually usable.

• Signatures can be recognized in more intelligent fashions (see Section 3.17.7 [Article Signature], page 69).

• Summary pick mode has been made to look more nn-like. Line numbers are displayed and the . command can be used to pick articles (Pick and Read).

• Commands for moving the ‘.newsrc.eld’ from one server to another have been added (see Section 1.7 [Changing Servers], page 7).

• There’s a way now to specify that “uninteresting” fields be suppressed when generating lines in buffers (see Section 8.4.3 [Advanced Formatting], page 155).

• Several commands in the group buffer can be undone with M-C-_(see Section 8.13 [Undo], page 166).

• Scoring can be done on words using the new score type w (see Section 7.4 [Score File Format], page 136).

• Adaptive scoring can be done on a Subject word-by-word basis (see Section 7.6 [Adap-

tive Scoring], page 140).

       (setq gnus-use-adaptive-scoring '(word))

• Scores can be decayed (see Section 7.16 [Score Decays], page 152).

       (setq gnus-decay-scores t)

• Scoring can be performed using a regexp on the Date header. The Date is normalized to compact ISO 8601 format first (see Section 7.4 [Score File Format], page 136).

• A new command has been added to remove all data on articles from the native server (see Section 1.7 [Changing Servers], page 7).

• A new command for reading collections of documents (nndoc with nnvirtual on top) has been added—M-C-d (see Section 3.24.4 [Really Various Summary Commands], page 77).

• Process mark sets can be pushed and popped (see Section 3.7.5 [Setting Process Marks], page 44).

• A new mail-to-news backend makes it possible to post even when the NNTP server doesn’t allow posting (see Section 6.4.6 [Mail-To-News Gateways], page 121).

• A new backend for reading searches from Web search engines (DejaNews, Alta Vista, InReference) has been added (see Section 6.4.5 [Web Searches], page 120).

• Groups inside topics can now be sorted using the standard sorting functions, and each topic can be sorted independently (see Section 2.16.3 [Topic Sorting], page 28).

• Subsets of the groups can be sorted independently (Sorting Groups).

• Cached articles can be pulled into the groups (see Section 3.24.3 [Summary Generation Commands], page 77).

• Score files are now applied in a more reliable order (see Section 7.3 [Score Variables], page 133).

• Reports on where mail messages end up can be generated (see Section 6.3.2 [Splitting Mail], page 100).

• More hooks and functions have been added to remove junk from incoming mail before saving the mail (see Section 6.3.8 [Washing Mail], page 108).

• Emphasized text can be properly fontisized:

       (add-hook 'gnus-article-display-hook
                 'gnus-article-emphasize)

10.1.6.4 Quassia Gnus

New features in Gnus 5.6:

• New functionality for using Gnus as an offline newsreader has been added. A plethora of new commands and modes have been added. See see Section 6.6 [Gnus Unplugged], page 124 for the full story.

• The nndraft backend has returned, but works differently than before. All Message buffers are now also articles in the nndraft group, which is created automatically.

• gnus-alter-header-function can now be used to alter header values.

• gnus-summary-goto-article now accept Message-ID’s.

• A new Message command for deleting text in the body of a message outside the region:

  C-c C-v.

• You can now post to component group in nnvirtual groups with C-u C-c C-c.

• nntp-rlogin-program—new variable to ease customization.

• C-u C-c C-c in gnus-article-edit-mode will now inhibit re-highlighting of the article buffer.

• New element in gnus-boring-article-headers—long-to.

• M-i symbolic prefix command. See the section "Symbolic Prefixes" in the Gnus manual for details.

• L and I in the summary buffer now take the symbolic prefix a to add the score rule to the "all.SCORE" file.

• gnus-simplify-subject-functions variable to allow greater control over simplification.

• A T—new command for fetching the current thread.

• / T—new command for including the current thread in the limit.

• M-RET is a new Message command for breaking cited text.

• ‘\\1’-expressions are now valid in nnmail-split-methods.

• The custom-face-lookup function has been removed. If you used this function in your initialization files, you must rewrite them to use face-spec-set instead.

• Canceling now uses the current select method. Symbolic prefix a forces normal posting method.

• New command to translate M******** sm*rtq**t*s into proper text—W d.

• For easier debugging of nntp, you can set nntp-record-commands to a non-nil value.

• nntp now uses ‘~/.authinfo’, a ‘.netrc’-like file, for controlling where and how to send authinfo to nntp servers.

• A command for editing group parameters from the summary buffer has been added.

• A history of where mails have been split is available.

• A new article date command has been added—article-date-iso8601.

• Subjects can be simplified when threading by setting gnus-score-thread-simplify.

• A new function for citing in Message has been added—message-cite-original-

  without-signature.
• article-strip-all-blank-lines—new article command.

• A new Message command to kill to the end of the article has been added.

• A minimum adaptive score can be specified by using the gnus-adaptive-word-minimum variable.

• The "lapsed date" article header can be kept continually updated by the gnus-start-

  date-timer command.

• Web listserv archives can be read with the nnlistserv backend.

• Old dejanews archives can now be read by nnweb.

10.1.7 Newest Features

Also known as the todo list. Sure to be implemented before the next millennium. Be afraid. Be very afraid. (That a feature appears in this list doesn’t necessarily mean that I’ve decided to actually

implement it. It just means that I think it sounds interesting.)

(Yes, this is the actual, up-to-the-second todo list.)

• Native mime support is something that should be done.

• Really do unbinhexing.

• I would like the zombie-page to contain an URL to the source of the latest version of gnus or some explanation on where to find it.

• A way to continue editing the latest Message composition.

• http://www.sonicnet.com/feature/ari3/

• facep is not declared.

• Include a section in the manual on why the number of articles isn’t the same in the group buffer and on the SPC prompt.

• Interacting with rmail fcc isn’t easy. •

        Hypermail:
       <URL:http://www.falch.no/people/pepper/DSSSL-Lite/archives/>
       <URL:http://www.eit.com/software/hypermail/hypermail.html>
       <URL:http://homer.ncm.com/>
       <URL:http://www.yahoo.com/Computers_and_Internet/Internet/World_Wide_Web/HTML_Converters/>
       http://www.uwsg.indiana.edu/hypermail/linux/kernel/9610/index.html
       <URL:http://union.ncsa.uiuc.edu/HyperNews/get/www/html/converters.html>
       http://www.miranova.com/gnus-list/

• ‘^-- ’ is made into - in LaTeX.

• gnus-kill is much slower than it was in GNUS 4.1.3.

• when expunging articles on low score, the sparse nodes keep hanging on?

• starting the first time seems to hang Gnus on some systems. Does NEWGROUPS answer too fast?

• nndir doesn’t read gzipped files.

• FAQ doesn’t have an up node?

• when moving mail from a procmail spool to the crash-box, the crash-box is only appropriate to one specific group.

• ‘t’ ‘t’ makes X-Faces disappear.

• nnmh-be-safe means that crossposted articles will be marked as unread.

• Orphan score entries don’t show on "V t" score trace

• when clearing out data, the cache data should also be reset.

• rewrite gnus-summary-limit-children to be non-recursive to avoid exceeding lisp nesting on huge groups.

• expunged articles are counted when computing scores.

• implement gnus-batch-brew-soup

• ticked articles aren’t easy to read in pick mode – ‘n’ and stuff just skips past them. Read articles are the same.

• topics that contain just groups with ticked articles aren’t displayed.

• nndoc should always allocate unique Message-IDs.

• If there are mail groups the first time you use Gnus, Gnus’ll make the mail groups killed.

• no "no news is good news" when using topics.

• when doing crosspost marking, the cache has to be consulted and articles have to be removed.

• nnweb should fetch complete articles when they are split into several parts.

• scoring on head immediate doesn’t work.

• finding short score file names takes forever.

• canceling articles in foreign groups.

• nntp-open-rlogin no longer works.

• C-u C-x C-s (Summary) switches to the group buffer.

• move nnmail-split-history out to the backends.

• nnweb doesn’t work properly.

• using a virtual server name as ‘gnus-select-method’ doesn’t work?

• when killing/yanking a group from one topic to another in a slave, the master will yank it first to one topic and then add it to another. Perhaps.

• warn user about ‘=’ redirection of a group in the active file?

• really unbinhex binhex files.

• take over the XEmacs menubar and offer a toggle between the XEmacs bar and the Gnus bar.

•

        push active file and NOV file parsing down into C code.
       `(canonize-message-id id)'
       `(mail-parent-message-id references n)'
       `(parse-news-nov-line &optional dependency-hashtb)'
       `(parse-news-nov-region beg end &optional dependency-hashtb fullp)'
       `(parse-news-active-region beg end hashtb)'

• nnml .overview directory with splits.

• asynchronous cache

• postponed commands.

• the selected article show have its Subject displayed in its summary line.

• when entering groups, get the real number of unread articles from the server?

• sort after gathering threads – make false roots have the headers of the oldest orphan with a 0 article number?

• nndoc groups should inherit the score files of their parents? Also inherit copy prompts and save files.

• command to start up Gnus (if not running) and enter a mail mode buffer.

• allow editing the group description from the group buffer for backends that support that.

• gnus-hide,show-all-topics

• groups and sub-topics should be allowed to mingle inside each topic, and not just list all subtopics at the end.

• a command to remove all read articles that are not needed to connect threads – ‘gnussummary-limit-to-sparse-unread’?

• a variable to turn off limiting/cutting of threads in the tree buffer.

• a variable to limit how many files are uudecoded.

• add zombie groups to a special "New Groups" topic.

• server mode command: close/open all connections

• put a file date in gnus-score-alist and check whether the file has been changed before using it.

• on exit from a digest group, go to the next article in the parent group.

• hide (sub)threads with low score.

• when expiring, remove all marks from expired articles.

• gnus-summary-limit-to-body

• a regexp alist that says what level groups are to be subscribed on. Eg. – ‘(("nnml:" . 1))’.

• easier interface to nnkiboze to create ephemeral groups that contain groups that match a regexp.

• allow newlines in <URL:> urls, but remove them before using the URL.

• If there is no From line, the mail backends should fudge one from the "From " line.

• fuzzy simplifying should strip all non-alpha-numerical info from subject lines.

• gnus-soup-brew-soup-with-high-scores.

• nntp-ping-before-connect

• command to check whether NOV is evil. "list overview.fmt".

• when entering a group, Gnus should look through the score files very early for ‘local’ atoms and set those local variables.

• message annotations.

• topics are always yanked before groups, and that’s not good.

• (set-extent-property extent ’help-echo "String to display in minibuf") to display help in the minibuffer on buttons under XEmacs.

• allow group line format spec to say how many articles there are in the cache.

• AUTHINFO GENERIC

• support qmail maildir spools

• ‘run-with-idle-timer’ in gnus-demon.

• stop using invisible text properties and start using overlays instead

• C-c C-f C-e to add an Expires header.

• go from one group to the next; everything is expunged; go to the next group instead of going to the group buffer.

• gnus-renumber-cache – to renumber the cache using "low" numbers.

• record topic changes in the dribble buffer.

• ‘nnfolder-generate-active-file’ should look at the folders it finds and generate proper active ranges.

• nneething-look-in-files-for-article-heads variable to control whether nneething should sniff all files in the directories.

• gnus-fetch-article – start Gnus, enter group, display article

• gnus-dont-move-articles-to-same-group variable when respooling.

• when messages are crossposted between several auto-expirable groups, articles aren’t properly marked as expirable.

• nneething should allow deletion/moving.

• TAB on the last button should go to the first button.

• if the car of an element in ‘mail-split-methods’ is a function, and the function returns non-nil, use that as the name of the group(s) to save mail in.

• command for listing all score files that have been applied.

• a command in the article buffer to return to ‘summary’ config.

• ‘gnus-always-post-using-current-server’ – variable to override ‘C-c C-c’ when posting.

• nnmail-group-spool-alist – says where each group should use as a spool file.

• when an article is crossposted to an auto-expirable group, the article should be marker as expirable.

• article mode command/menu for "send region as URL to browser".

• on errors, jump to info nodes that explain the error. For instance, on invalid From headers, or on error messages from the nntp server.

• when gathering threads, make the article that has no "Re: " the parent. Also consult Date headers.

• a token in splits to call shrink-window-if-larger-than-buffer

• ‘1 0 A M’ to do matches on the active hashtb.

• duplicates – command to remove Gnus-Warning header, use the read Message-ID, delete the "original".

• when replying to several messages at once, put the "other" message-ids into a See-Also header.

• support setext: URL:http://www.bsdi.com/setext/

• support ProleText: <URL:http://proletext.clari.net/prole/proletext.html>

• when browsing a foreign server, the groups that are already subscribed should be listed as such and not as "K".

• generate font names dynamically.

• score file mode auto-alist.

• allow nndoc to change/add/delete things from documents. Implement methods for each format for adding an article to the document.

• ‘gnus-fetch-old-headers’ ‘all’ value to incorporate absolutely all headers there is.

• function like ‘|’, but concatenate all marked articles and pipe them to the process.

• cache the list of killed (or active) groups in a separate file. Update the file whenever we read the active file or the list of killed groups in the .eld file reaches a certain length.

• function for starting to edit a file to put into the current mail group.

• score-find-trace should display the total score of the article.

• "ghettozie" – score on Xref header and nix it out after using it to avoid marking as read in other groups it has been crossposted to.

• look at procmail splitting. The backends should create the groups automatically if a spool file exists for that group.

• function for backends to register themselves with Gnus.

• when replying to several process-marked articles, have all the From end up in Cc headers? Variable to toggle.

• command to delete a crossposted mail article from all groups it has been mailed to.

• ‘B c’ and ‘B m’ should be crosspost aware.

• hide-pgp should also hide PGP public key blocks.

• Command in the group buffer to respool process-marked groups.

• ‘gnus-summary-find-matching’ should accept pseudo-"headers" like "body", "head" and "all"

• When buttifying <URL: > things, all white space (including newlines) should be ignored.

• Process-marking all groups in a topic should process-mark groups in subtopics as well.

• Add non-native groups to the list of killed groups when killing them.

• nntp-suggest-kewl-config to probe the nntp server and suggest variable settings.

• add edit and forward secondary marks.

• nnml shouldn’t visit its .overview files.

• allow customizing sorting within gathered threads.

• ‘B q’ shouldn’t select the current article.

• nnmbox should support a newsgroups file for descriptions.

• allow fetching mail from several pop servers.

• Be able to specify whether the saving commands save the original or the formatted article.

• a command to reparent with the child process-marked (cf. ‘T ^’.).

• I think the possibility to send a password with nntp-open-rlogin should be a feature in Red Gnus.

• The ‘Z n’ command should be possible to execute from a mouse click.

• more limiting functions – date, etc.

• be able to limit on a random header; on body; using reverse matches.

• a group parameter (‘absofucking-total-expiry’) that will make Gnus expire even unread articles.

• a command to print the article buffer as postscript.

• variable to disable password fetching when opening by nntp-open-telnet.

• manual: more example servers – nntp with rlogin, telnet

• checking for bogus groups should clean topic alists as well.

• canceling articles in foreign groups.

• article number in folded topics isn’t properly updated by Xref handling.

• Movement in the group buffer to the next unread group should go to the next closed topic with unread messages if no group can be found.

• Extensive info pages generated on the fly with help everywhere – in the "*Gnus edit*" buffers, for instance.

• Topic movement commands – like thread movement. Up, down, forward, next.

• a way to tick/mark as read Gcc’d articles.

• a way to say that all groups within a specific topic comes from a particular server? Hm.

• ‘gnus-article-fill-if-long-lines’ – a function to fill the article buffer if there are any looong lines there.

• ‘T h’ should jump to the parent topic and fold it.

• a command to create an ephemeral nndoc group out of a file, and then splitting it/moving it to some other group/backend.

• a group parameter for nnkiboze groups that says that all kibozed articles should be entered into the cache.

• It should also probably be possible to delimit what ‘gnus-jog-cache’ does – for instance, work on just some groups, or on some levels, and entering just articles that have a score higher than a certain number.

• nnfolder should append to the folder instead of re-writing the entire folder to disk when accepting new messages.

• allow all backends to do the proper thing with .gz files.

• a backend for reading collections of babyl files nnbabylfolder?

• a command for making the native groups into foreign groups.

• server mode command for clearing read marks from all groups from a server.

• when following up multiple articles, include all To, Cc, etc headers from all articles.

• a command for deciding what the total score of the current thread is. Also a way to highlight based on this.

• command to show and edit group scores

• a gnus-tree-minimize-horizontal to minimize tree buffers horizontally.

• command to generate nnml overview file for one group.

• ‘C-u C-u a’ – prompt for many crossposted groups.

• keep track of which mail groups have received new articles (in this session). Be able to generate a report and perhaps do some marking in the group buffer.

• gnus-build-sparse-threads to a number – build only sparse threads that are of that length.

• have nnmh respect mh’s unseen sequence in .mh profile.

• cache the newsgroups descriptions locally.

• asynchronous posting under nntp.

• be able to control word adaptive scoring from the score files.

• a variable to make ‘C-c C-c’ post using the "current" select method.

• ‘limit-exclude-low-scored-articles’.

• if ‘gnus-summary-show-thread’ is a number, hide threads that have a score lower than this number.

• split newsgroup subscription variable up into "order" and "method".

• buttonize ange-ftp file names.

• a command to make a duplicate copy of the current article so that each copy can be edited separately.

• nnweb should allow fetching from the local nntp server.

• record the sorting done in the summary buffer so that it can be repeated when limiting/regenerating the buffer.

• nnml-generate-nov-databses should generate for all nnml servers.

• when the user does commands in the group buffer, check the modification time of the .newsrc.eld file and use ask-user-about-supersession-threat. Also warn when trying to save .newsrc.eld and it has changed.

• M-g on a topic will display all groups with 0 articles in the topic.

• command to remove all topic stuff.

• allow exploding incoming digests when reading incoming mail and splitting the resulting digests.

• nnsoup shouldn’t set the ‘message-’ variables.

• command to nix out all nnoo state information.

• nnmail-process-alist that calls functions if group names matches an alist – before saving.

• use buffer-invisibility-spec everywhere for hiding text.

• variable to activate each group before entering them to get the (new) number of articles. ‘gnus-activate-before-entering’.

• command to fetch a Message-ID from any buffer, even starting Gnus first if necessary.

• when posting and checking whether a group exists or not, just ask the nntp server instead of relying on the active hashtb.

• buttonize the output of ‘C-c C-a’ in an apropos-like way.

• ‘G p’ should understand process/prefix, and allow editing of several groups at once.

• command to create an ephemeral nnvirtual group that matches some regexp(s).

• nndoc should understand "Content-Type: message/rfc822" forwarded messages.

• it should be possible to score "thread" on the From header.

• hitting RET on a "gnus-uu-archive" pseudo article should unpack it.

• ‘B i’ should display the article at once in the summary buffer.

• remove the "*" mark at once when unticking an article.

• ‘M-s’ should highlight the matching text.

• when checking for duplicated mails, use Resent-Message-ID if present.

• killing and yanking groups in topics should be better. If killing one copy of a group that exists in multiple topics, only that copy should be removed. Yanking should insert the copy, and yanking topics should be possible to be interspersed with the other yankings.

• command for enter a group just to read the cached articles. A way to say "ignore the nntp connection; just read from the cache."

• ‘X u’ should decode base64 articles.

• a way to hide all "inner" cited text, leaving just the most recently cited text.

• nnvirtual should be asynchronous.

• after editing an article, gnus-original-article-buffer should be invalidated.

• there should probably be a way to make Gnus not connect to the server and just read the articles in the server

• allow a ‘set-default’ (or something) to change the default value of nnoo variables.

• a command to import group infos from a .newsrc.eld file.

• groups from secondary servers have the entire select method listed in each group info.

• a command for just switching from the summary buffer to the group buffer.

• a way to specify that some incoming mail washing functions should only be applied to some groups.

• Message ‘C-f C-t’ should ask the user whether to heed mail-copies-to: never.

• new group parameter – ‘post-to-server’ that says to post using the current server. Also a variable to do the same.

• the slave dribble files should autosave to the slave file names.

• a group parameter that says what articles to display on group entry, based on article marks.

• a way to visually distinguish slave Gnusae from masters. (Whip instead of normal logo?)

• Use DJ Bernstein "From " quoting/dequoting, where applicable.

• Why is hide-citation-maybe and hide-citation different? Also clear up info.

• group user-defined meta-parameters.

From: John Griffith <griffith@sfs.nphil.uni-tuebingen.de>

• I like the option for trying to retrieve the FAQ for a group and I was thinking it would be great if for those newsgroups that had archives you could also try to read the archive for that group. Part of the problem is that archives are spread all over the net, unlike FAQs. What would be best I suppose is to find the one closest to your site. In any case, there is a list of general news group archives at ftp://ftp.neosoft.com/pub/users/claird/news.lists/newsgroup archives.html

•

       From: Jason L Tibbitts III <tibbs@hpc.uh.edu>
       (add-hook 'gnus-select-group-hook
         (lambda ()

         (gnus-group-add-parameter group
            (cons 'gnus-group-date-last-entered (list (current-time-string))))))
     (defun gnus-user-format-function-d (headers)
       "Return the date the group was last read."
       (cond ((car (gnus-group-get-parameter gnus-tmp-group 'gnus-group-date-last-entered)))
           (t "")))

• tanken var at nr du bruker ‘gnus-startup-file’ som prefix (FOO) til lete opp en fil FOO-SERVER, FOO-SERVER.el, FOO-SERVER.eld, kan du la den vre en liste hvor du bruker hvert element i listen som FOO, istedet. da kunne man hatt forskjellige serveres startup-filer forskjellige steder.

• LMI> Well, nnbabyl could alter the group info to heed labels like LMI> answered and read, I guess. It could also keep them updated (the same for the Status: header of unix mbox files). They could be used like this:

     `M l <name> RET' add label <name> to current message.
     `M u <name> RET' remove label <name> from current message.
     `/ l <expr> RET' limit summary buffer according to <expr>.
     <expr> would be a boolean expression on the labels, e.g.
     `/ l bug & !fixed RET'

would show all the messages which are labeled ‘bug’ but not labeled ‘fixed’. One could also imagine the labels being used for highlighting, or affect the summary line format.

• Sender: abraham@dina.kvl.dk I’d like a gnus-find-file which work like find file, except that it would recognize things that looks like messages or folders:

- If it is a directory containing numbered files, create an nndir summary buffer.

- For other directories, create a nneething summary buffer.

- For files matching "\\‘From ", create a nndoc/mbox summary.

- For files matching "\\‘BABYL OPTIONS:", create a nndoc/baby summary.

- For files matching "\\‘[^\t\n]+:", create an *Article* buffer.

- For other files, just find them normally. I’d like ‘nneething’ to use this function, so it would work on a directory potentially containing mboxes or babyl files.

• Please send a mail to bwarsaw@cnri.reston.va.us (Barry A. Warsaw) and tell him what you are doing.

• Currently, I get prompted:

decend into sci? - type y decend into sci.something ? - type n decend into ucd? The problem above is that since there is really only one subsection of science, shouldn’t it prompt you for only descending sci.something? If there was a sci.somethingelse group or section, then it should prompt for sci? first the sci.something? then sci.somethingelse?...

• Ja, det burde vre en mte si slikt. Kanskje en ny variabel? ‘gnus-use-few-score-files’? S kunne score-regler legges til den "mest" lokale score-fila. F. eks. ville no-gruppene betjenes av "no.all.SCORE", osv.

• What i want is for Gnus to treat any sequence or combination of the following as a single spoiler warning and hide it all, replacing it with a "Next Page" button: ^L’s more than n blank lines more than m identical lines (which should be replaced with button to show them) any whitespace surrounding any of the above

• Well, we could allow a new value to ‘gnus-thread-ignore-subject’ – ‘spaces’, or something. (We could even default to that.) And then subjects that differ in white space only could be considered the "same" subject for threading purposes.

• Modes to preprocess the contents (e.g. jka-compr) use the second form "(REGEXP FUNCTION NON-NIL)" while ordinary modes (e.g. tex) use the first form "(REGEXP . FUNCTION)", so you could use it to distinguish between those two types of modes. (auto-modes-alist, insert-file-contents-literally.)

• Under XEmacs – do funny article marks: tick - thumb tack killed - skull soup - bowl of soup score below - dim light bulb score over - bright light bulb

• Yes. I think the algorithm is as follows:

     Group-mode
        show-list-of-articles-in-group
             if (key-pressed == SPACE)
                     if (no-more-articles-in-group-to-select)
                             if (articles-selected)
                                     start-reading-selected-articles;
                             junk-unread-articles;
                             next-group;
                      else
                             show-next-page;
              else if (key-pressed = '.')
                     if (consolidated-menus)         # same as hide-thread in Gnus
                             select-thread-under-cursor;
                     else
                             select-article-under-cursor;

     Article-mode
             if (key-pressed == SPACE)
                     if (more-pages-in-article)
                             next-page;
                     else if (more-selected-articles-to-read)
                             next-article;
                     else
                             next-group;

• My precise need here would have been to limit files to Incoming*. One could think of some ‘nneething-only-files’ variable, but I guess it would have been unacceptable if one was using many unrelated such nneething groups. A more useful approach would be to, in response to the ‘G D’ prompt, be allowed to say something like: ‘~/.mail/Incoming*’, somewhat limiting the top-level directory only (in case directories would be matched by the wildcard expression).

• It would be nice if it also handled <URL:news://sunsite.auc.dk/> which should correspond to ‘B nntp RET sunsite.auc.dk’ in *Group*.

• Take a look at w3-menu.el in the Emacs-W3 distribution - this works out really well. Each menu is ’named’ by a symbol that would be on a gnus-*-menus (where * would be whatever, but at least group, summary, and article versions) variable. So for gnus-summary-menus, I would set to ’(sort mark dispose ...) A value of ’1’ would just put all the menus in a single ’GNUS’ menu in the main menubar. This approach works really well for Emacs-W3 and VM.

• nndoc should take care to create unique Message-IDs for all its articles.

• gnus-score-followup-article only works when you have a summary buffer active. Make it work when posting from the group buffer as well. (message-sent-hook).

• rewrite gnus-demon to use run-with-idle-timers.

• * Enhancements to Gnus:

Add two commands:

* gnus-servers (gnus-start-server-buffer?)–enters Gnus and goes straight to the server buffer, without opening any connections to servers first.

* gnus-server-read-server-newsrc–produces a buffer very similar to the group buffer, but with only groups from that server listed; quitting this buffer returns to the server buffer.

• add a command to check the integrity of an nnfolder folder – go through the article numbers and see that there are no duplicates, and stuff.

• ‘unsmileyfy-buffer’ to undo smileification.

• a command to give all relevant info on an article, including all secondary marks.

• when doing ‘-request-accept-article’, the backends should do the nnmail duplicate checking.

• allow ‘message-signature-file’ to be a function to return the value of the signature file.

• In addition, I would love it if I could configure message-tab so that it could call ‘bbdbcomplete-name’ in other headers. So, some sort of interface like (setq message-tab-alist ’((message-header-regexp message-expand-group) ("^\\(To\\|[cC]c\\|[bB]cc\\)" bbdb-complete-name))) then you could run the relevant function to complete the information in the header

• cache the newsgroups file locally to avoid reloading it all the time.

• a command to import a buffer into a group.

• nnweb should allow fetching by Message-ID from servers.

• point in the article buffer doesn’t always go to the beginning of the buffer when selecting new articles.

• a command to process mark all unread articles.

• ‘gnus-gather-threads-by-references-and-subject’ – first do gathering by references, and then go through the dummy roots and do more gathering by subject.

• gnus-uu-mark-in-numerical-order – process mark articles in article numerical order.

• (gnus-thread-total-score (gnus-id-to-thread (mail-header-id (gnus-summary-article-header)))) bind to a key.

• sorting by score is wrong when using sparse threads.

• a command to fetch an arbitrary article – without having to be in the summary buffer.

• a new nncvs backend. Each group would show an article, using version branches as threading, checkin date as the date, etc.

• http://www.dejanews.com/forms/dnsetfilter exp.html ? This filter allows one to construct advance queries on the Dejanews database such as specifying start and end dates, subject, author, and/or newsgroup name.

• new Date header scoring type – older, newer

• use the summary toolbar in the article buffer.

• a command to fetch all articles that are less than X days old.

• in pick mode, ‘q’ should save the list of selected articles in the group info. The next time the group is selected, these articles will automatically get the process mark.

• Isn’t it possible to (also?) allow M-^to automatically try the default server if it fails on the current server? (controlled by a user variable, (nil, t, ’ask)).

• make it possible to cancel articles using the select method for the current group.

• ‘gnus-summary-select-article-on-entry’ or something. It’ll default to t and will select whatever article decided by ‘gnus-auto-select-first’.

• a new variable to control which selection commands should be unselecting. ‘first’, ‘best’, ‘next’, ‘prev’, ‘next-unread’, ‘prev-unread’ are candidates.

• be able to select groups that have no articles in them to be able to post in them (using the current select method).

• be able to post via DejaNews.

• ‘x’ should retain any sortings that have been performed.

• allow the user to specify the precedence of the secondary marks. Also allow them to be displayed separately.

• gnus-summary-save-in-pipe should concatenate the results from the processes when doing a process marked pipe.

• a new match type, like Followup, but which adds Thread matches on all articles that match a certain From header.

• a function that can be read from kill-emacs-query-functions to offer saving living summary buffers.

• a function for selecting a particular group which will contain the articles listed in a list of article numbers/id’s.

• a battery of character translation functions to translate common Mac, MS (etc) char-

acters into ISO 8859-1.

(defun article-fix-m$word ()
  "Fix M$Word smartquotes in an article."
  (interactive)
  (save-excursion
    (let ((buffer-read-only nil))
      (goto-char (point-min))
      (while (search-forward "\221" nil t)
        (replace-match "`" t t))
      (goto-char (point-min))
      (while (search-forward "\222" nil t)
        (replace-match "'" t t))
      (goto-char (point-min))
      (while (search-forward "\223" nil t)
        (replace-match "\"" t t))
      (goto-char (point-min))
      (while (search-forward "\224" nil t)
        (replace-match "\"" t t)))))

•

 (add-hook 'gnus-exit-query-functions
'(lambda ()
   (if (and (file-exists-p nnmail-spool-file)
            (> (nnheader-file-size nnmail-spool-file) 0))
       (yes-or-no-p "New mail has arrived.  Quit Gnus anyways? ")
       (y-or-n-p "Are you sure you want to quit Gnus? "))))

• allow message-default-headers to be a function.

• new Date score match types – < > = (etc) that take floating point numbers and match on the age of the article.

•

>  > > If so, I've got one gripe: It seems that when I fire up gnus 5.2.25
>  > > under xemacs-19.14, it's creating a new frame, but is erasing the
>  > > buffer in the frame that it was called from =:-O
>
>  > Hm.  How do you start up Gnus?  From the toolbar or with
>  > `M-x gnus-other-frame'?
>
>    I normally start it up from the toolbar; at
> least that's the way I've caught it doing the
> deed before.

• all commands that react to the process mark should push the current process mark set onto the stack.

• gnus-article-hide-pgp Selv ville jeg nok ha valgt slette den dersom teksten matcher

"\\(This\s+\\)?[^]+ has been automatically signed by"

og det er maks hundre tegn mellom match-end og —-linja. Men -det- er min type heuristikk og langt fra alles.

• ‘gnus-subscribe-sorted’ – insert new groups where they would have been sorted to if ‘gnus-group-sort-function’ were run.

• gnus-(group,summary)-highlight should respect any ‘face’ text props set on the lines.

• use run-with-idle-timer for gnus-demon instead of the home-brewed stuff for better reliability.

• add a way to select which NoCeM type to apply – spam, troll, etc.

• nndraft-request-group should tally autosave files.

• implement nntp-retry-on-break and nntp-command-timeout.

• gnus-article-highlight-limit that says when not to highlight (long) articles.

• (nnoo-set SERVER VARIABLE VALUE)

• nn*-spool-methods

• interrupitng agent fetching of articles should save articles.

• command to open a digest group, and copy all the articles there to the current group.

• a variable to disable article body highlights if there’s more than X characters in the body.

• handle 480/381 authinfo requests separately.

• include the texi/dir file in the distribution.

• format spec to "tab" to a position.

• Move all prompting to the new ‘M-n’ default style.

• command to display all dormant articles.

• gnus-auto-select-next makeover – list of things it should do.

• a score match type that adds scores matching on From if From has replied to something someone else has said.

• Read Netscape discussion groups: snews://secnews.netscape.com/netscape.communicator.unix • One command to edit the original version if an article, and one to edit the displayed

version.

• T v – make all process-marked articles the children of the current article.

• Switch from initial text to the new default text mechanism.

• How about making it possible to expire local articles? Will it be possible to make various constraints on when an article can be expired, e.g. (read), (age > 14 days), or the more interesting (read & age > 14 days)?

• New limit command—limit to articles that have a certain string in the head or body.

• Allow breaking lengthy NNTP commands.

• gnus-article-highlight-limit, to disable highlighting in big articles.

• Editing an article should put the article to be edited in a special, unique buffer.

• A command to send a mail to the admin-address group param.

• A Date scoring type that will match if the article is less than a certain number of days old.

• New spec: %~(tab 56) to put point on column 56

• Allow Gnus Agent scoring to use normal score files.

• Rething the Agent active file thing. ‘M-g’ doesn’t update the active file, for instance.

• With dummy roots, ‘^’ and then selecing the first article in any other dummy thread will make Gnus highlight the dummy root instead of the first article.

• Propagate all group properties (marks, article numbers, etc) up to the topics for displaying.

• ‘n’ in the group buffer with topics should go to the next group with unread articles, even if that group is hidden in a topic.

• gnus-posting-styles doesn’t work in drafts.

• gnus-summary-limit-include-cached is slow when there are many articles in the cache, since it regenerates big parts of the summary buffer for each article.

• Implement gnus-batch-brew-soup.

• Group parameters and summary commands for un/subscribing to mailing lists.

• Introduce nnmail-home-directory.

• gnus-fetch-group and friends should exit Gnus when the user exits the group.

• Solve the halting problem.

10.2 The Manual

This manual was generated from a TeXinfo file and then run through either texi2dvi to get what you hold in your hands now.

The following conventions have been used:

1. This is a ‘string’

2. This is a keystroke

3. This is a ‘file’

4. This is a symbol

So if I were to say “set flargnoze to ‘yes’”, that would mean:

  (setq flargnoze "yes")

If I say “set flumphel to yes”, that would mean:

  (setq flumphel 'yes)

‘yes’ and yes are two very different things—don’t ever get them confused.

10.3 Terminology

news This is what you are supposed to use this thing for—reading news. News is generally fetched from a nearby nntp server, and is generally publicly available to everybody. If you post news, the entire world is likely to read just what you have written, and they’ll all snigger mischievously. Behind your back.

mail Everything that’s delivered to you personally is mail. Some news/mail readers (like Gnus) blur the distinction between mail and news, but there is a difference. Mail is private. News is public. Mailing is not posting, and replying is not following up.

reply Send a mail to the person who has written what you are reading.

follow up Post an article to the current newsgroup responding to the article you are reading.

backend Gnus gets fed articles from a number of backends, both news and mail backends. Gnus does not handle the underlying media, so to speak—this is all done by the backends.

native Gnus will always use one method (and backend) as the native, or default, way of getting news.

foreign You can also have any number of foreign groups active at the same time. These are groups that use non-native non-secondary backends for getting news.

secondary Secondary backends are somewhere half-way between being native and being foreign, but they mostly act like they are native.

article A message that has been posted as news.

mail message A message that has been mailed.

message A mail message or news article

head The top part of a message, where administrative information (etc.) is put.

body The rest of an article. Everything not in the head is in the body.

header A line from the head of an article.

headers A collection of such lines, or a collection of heads. Or even a collection of nov lines.

nov When Gnus enters a group, it asks the backend for the headers of all unread articles in the group. Most servers support the News OverView format, which is more compact and much faster to read and parse than the normal head format.

level Each group is subscribed at some level or other (1-9). The ones that have a lower level are “more” subscribed than the groups with a higher level. In fact, groups on levels 1-5 are considered subscribed; 6-7 are unsubscribed; 8 are zombies; and 9 are killed. Commands for listing groups and scanning for new articles will all use the numeric prefix as working level. killed groups

No information on killed groups is stored or updated, which makes killed groups much easier to handle than subscribed groups.

zombie groups Just like killed groups, only slightly less dead.

active file The news server has to keep track of what articles it carries, and what groups exist. All this information in stored in the active file, which is rather large, as you might surmise.

bogus groups A group that exists in the ‘.newsrc’ file, but isn’t known to the server (i.e., it isn’t in the active file), is a bogus group. This means that the group probably doesn’t exist (any more).

activating The act of asking the server for info on a group and computing the number of unread articles is called activating the group. Un-activated groups are listed with ‘*’ in the group buffer.

server A machine one can connect to and get news (or mail) from.

select method A structure that specifies the backend, the server and the virtual server settings.

virtual server A named select method. Since a select method defines all there is to know about connecting to a (physical) server, taking the thing as a whole is a virtual server.

washing Taking a buffer and running it through a filter of some sort. The result will (more often than not) be cleaner and more pleasing than the original.

ephemeral groups Most groups store data on what articles you have read. Ephemeral groups are groups that will have no data stored—when you exit the group, it’ll disappear into the aether.

solid groups This is the opposite of ephemeral groups. All groups listed in the group buffer are solid groups.

sparse articles These are article placeholders shown in the summary buffer when gnus-build-sparse-threads has been switched on.

threading To put responses to articles directly after the articles they respond to—in a hierarchical fashion.

root The first article in a thread is the root. It is the ancestor of all articles in the thread.

parent An article that has responses.

child An article that responds to a different article—its parent.

206 Gnus Manual digest A collection of messages in one file. The most common digest format is specified

by RFC1153.

10.4 Customization

All variables are properly documented elsewhere in this manual. This section is designed to give general pointers on how to customize Gnus for some quite common situations.

10.4.1 Slow/Expensive nntp Connection

If you run Emacs on a machine locally, and get your news from a machine over some very thin strings, you want to cut down on the amount of data Gnus has to get from the nntp server.

gnus-read-active-file

Set this to nil, which will inhibit Gnus from requesting the entire active file from the server. This file is often v. large. You also have to set gnus-check-

new-newsgroups and gnus-check-bogus-newsgroups to nil to make sure that

Gnus doesn’t suddenly decide to fetch the active file anyway.

gnus-nov-is-evil

This one has to be nil. If not, grabbing article headers from the nntp server will not be very fast. Not all nntp servers support xover; Gnus will detect this by itself.

10.4.2 Slow Terminal Connection

Let’s say you use your home computer for dialing up the system that runs Emacs and Gnus. If your modem is slow, you want to reduce (as much as possible) the amount of data sent over the wires.

gnus-auto-center-summary

Set this to nil to inhibit Gnus from re-centering the summary buffer all the time. If it is vertical, do only vertical re-centering. If it is neither nil nor vertical, do both horizontal and vertical recentering.

gnus-visible-headers

Cut down on the headers included in the articles to the minimum. You can, in fact, make do without them altogether—most of the useful data is in the summary buffer, anyway. Set this variable to ‘^NEVVVVER’ or ‘From:’, or whatever you feel you need.

gnus-article-display-hook

Set this hook to all the available hiding commands:

     (setq gnus-article-display-hook
           '(gnus-article-hide-headers
             gnus-article-hide-signature
             gnus-article-hide-citation))
gnus-use-full-window

By setting this to nil, you can make all the windows smaller. While this doesn’t really cut down much generally, it means that you have to see smaller portions of articles before deciding that you didn’t want to read them anyway.

gnus-thread-hide-subtree

If this is non-nil, all threads in the summary buffer will be hidden initially.

gnus-updated-mode-lines

If this is nil, Gnus will not put information in the buffer mode lines, which might save some time.

10.4.3 Little Disk Space

The startup files can get rather large, so you may want to cut their sizes a bit if you are running out of space.

gnus-save-newsrc-file

If this is nil, Gnus will never save ‘.newsrc’—it will only save ‘.newsrc.eld’. This means that you will not be able to use any other newsreaders than Gnus. This variable is t by default.

gnus-save-killed-list

If this is nil, Gnus will not save the list of dead groups. You should also

set gnus-check-new-newsgroups to ask-server and gnus-check-bogus-

newsgroups to nil if you set this variable to nil. This variable is t by default.

10.4.4 Slow Machine

If you have a slow machine, or are just really impatient, there are a few things you can do to make Gnus run faster.

Set gnus-check-new-newsgroups and gnus-check-bogus-newsgroups to nil to make

startup faster.

Set gnus-show-threads, gnus-use-cross-reference and gnus-nov-is-evil to nil

to make entering and exiting the summary buffer faster.

Set gnus-article-display-hook to nil to make article processing a bit faster.

10.5 Troubleshooting

Gnus works so well straight out of the box—I can’t imagine any problems, really.

Ahem.

1. Make sure your computer is switched on.

2. Make sure that you really load the current Gnus version. If you have been running gnus, you need to exit Emacs and start it up again before Gnus will work.

3. Try doing an M-x gnus-version. If you get something that looks like ‘Gnus v5.46; nntp 4.0’ you have the right files loaded. If, on the other hand, you get something like ‘NNTP 3.x’ or ‘nntp flee’, you have some old ‘.el’ files lying around. Delete these.

4. Read the help group (G h in the group buffer) for a FAQ and a how-to.

5. Gnus works on many recursive structures, and in some extreme (and very rare) cases Gnus may recurse down “too deeply” and Emacs will beep at you. If this happens to you, set max-lisp-eval-depth to 500 or something like that.

If all else fails, report the problem as a bug.

If you find a bug in Gnus, you can report it with the M-x gnus-bug command. M-x

set-variable RET debug-on-error RET t RET, and send me the backtrace. I will fix bugs,

but I can only fix them if you send me a precise description as to how to reproduce the bug.

You really can never be too detailed in a bug report. Always use the M-x gnus-bug command when you make bug reports, even if it creates a 10Kb mail each time you use it, and even if you have sent me your environment 500 times before. I don’t care. I want the full info each time.

It is also important to remember that I have no memory whatsoever. If you send a bug report, and I send you a reply, and then you just send back “No, it’s not! Moron!”, I will have no idea what you are insulting me about. Always over-explain everything. It’s much easier for all of us—if I don’t have all the information I need, I will just mail you and ask for more info, and everything takes more time.

If the problem you’re seeing is very visual, and you can’t quite explain it, copy the Emacs window to a file (with xwd, for instance), put it somewhere it can be reached, and include the URL of the picture in the bug report.

If you just need help, you are better off asking on ‘gnu.emacs.gnus’. I’m not very helpful.

You can also ask on the ding mailing list—‘ding@gnus.org’. Write to

‘ding-request@gnus.org’ to subscribe.

10.6 A Programmerś Guide to Gnus

It is my hope that other people will figure out smart stuff that Gnus can do, and that other people will write those smart things as well. To facilitate that I thought it would be a good idea to describe the inner workings of Gnus. And some of the not-so-inner workings, while I’m at it.

You can never expect the internals of a program not to change, but I will be defining (in some details) the interface between Gnus and its backends (this is written in stone), the format of the score files (ditto), data structures (some are less likely to change than others) and general methods of operation.

10.6.1 Gnus Utility Functions

When writing small functions to be run from hooks (and stuff), it’s vital to have access to the Gnus internal functions and variables. Below is a list of the most common ones.

gnus-newsgroup-name

This variable holds the name of the current newsgroup.

gnus-find-method-for-group

A function that returns the select method for group.

gnus-group-real-name

Takes a full (prefixed) Gnus group name, and returns the unprefixed name.

gnus-group-prefixed-name

Takes an unprefixed group name and a select method, and returns the full (prefixed) Gnus group name.

gnus-get-info

Returns the group info list for group.

gnus-group-unread

The number of unread articles in group, or t if that is unknown.

gnus-active

The active entry for group.

gnus-set-active

Set the active entry for group.

gnus-add-current-to-buffer-list

Adds the current buffer to the list of buffers to be killed on Gnus exit.

gnus-continuum-version

Takes a Gnus version string as a parameter and returns a floating point number. Earlier versions will always get a lower number than later versions.

gnus-group-read-only-p

Says whether group is read-only or not.

gnus-news-group-p

Says whether group came from a news backend.

gnus-ephemeral-group-p

Says whether group is ephemeral or not.

gnus-server-to-method

Returns the select method corresponding to server.

gnus-server-equal

Says whether two virtual servers are equal.

gnus-group-native-p

Says whether group is native or not.

gnus-group-secondary-p

Says whether group is secondary or not.

gnus-group-foreign-p

Says whether group is foreign or not.

group-group-find-parameter

Returns the parameter list of group. If given a second parameter, returns the value of that parameter for group.

gnus-group-set-parameter

Takes three parameters; group, parameter and value.

gnus-narrow-to-body

Narrows the current buffer to the body of the article.

gnus-check-backend-function

Takes two parameters, function and group. If the backend group comes from supports function, return non-nil.

               (gnus-check-backend-function "request-scan" "nnml:misc")
               => t
gnus-read-method

Prompts the user for a select method.

10.6.2 Backend Interface

Gnus doesn’t know anything about nntp, spools, mail or virtual groups. It only knows how to talk to virtual servers. A virtual server is a backend and some backend variables. As examples of the first, we have nntp, nnspool and nnmbox. As examples of the latter we

have nntp-port-number and nnmbox-directory.

When Gnus asks for information from a backend—say nntp—on something, it will normally include a virtual server name in the function parameters. (If not, the backend should use the “current” virtual server.) For instance, nntp-request-list takes a virtual server as its only (optional) parameter. If this virtual server hasn’t been opened, the function should fail.

Note that a virtual server name has no relation to some physical server name. Take this example:

     (nntp "odd-one"
           (nntp-address "ifi.uio.no")

 (nntp-port-number 4324))

Here the virtual server name is ‘odd-one’ while the name of the physical server is

‘ifi.uio.no’.

The backends should be able to switch between several virtual servers. The standard backends implement this by keeping an alist of virtual server environments that they pull down/push up when needed.

There are two groups of interface functions: required functions, which must be present, and optional functions, which Gnus will always check for presence before attempting to call ’em.

All these functions are expected to return data in the buffer nntp-server-buffer (‘ *nntpd*’), which is somewhat unfortunately named, but we’ll have to live with it. When I talk about resulting data, I always refer to the data in that buffer. When I talk about return value, I talk about the function value returned by the function call. Functions that fail should return nil as the return value.

Some backends could be said to be server-forming backends, and some might be said not to be. The latter are backends that generally only operate on one group at a time, and have no concept of “server” – they have a group, and they deliver info on that group and nothing more.

In the examples and definitions I will refer to the imaginary backend nnchoke.

10.6.2.1 Required Backend Functions

(nnchoke-retrieve-headers ARTICLES &optional GROUP SERVER FETCH-OLD)

articles is either a range of article numbers or a list of Message-IDs. Current backends do not fully support either—only sequences (lists) of article numbers, and most backends do not support retrieval of Message-IDs. But they should try for both. The result data should either be HEADs or NOV lines, and the result value should either be headers or nov to reflect this. This might later be expanded to various, which will be a mixture of HEADs and NOV lines, but this is currently not supported by Gnus. If fetch-old is non-nil it says to try fetching "extra headers", in some meaning of the word. This is generally done by fetching (at most) fetch-old extra headers less than the smallest article number in articles, and filling the gaps as well. The presence of this parameter can be ignored if the backend finds it cumbersome to follow the request. If this is non-nil and not a number, do maximum fetches. Here’s an example HEAD:

     221 1056 Article retrieved.
     Path: ifi.uio.no!sturles
     From: sturles@ifi.uio.no (Sturle Sunde)
     Newsgroups: ifi.discussion
     Subject: Re: Something very droll
     Date: 27 Oct 1994 14:02:57 +0100
     Organization: Dept. of Informatics, University of Oslo, Norway

     Lines: 26
     Message-ID: <38o8e1$a0o@holmenkollen.ifi.uio.no>
     References: <38jdmq$4qu@visbur.ifi.uio.no>
     NNTP-Posting-Host: holmenkollen.ifi.uio.no
     .

So a headers return value would imply that there’s a number of these in the data buffer. Here’s a BNF definition of such a buffer:

     headers        = *head
     head           = error / valid-head
     error-message  = [ "4" / "5" ] 2number " " <error message> eol
     valid-head     = valid-message *header "." eol
     valid-message  = "221 " <number> " Article retrieved." eol
     header         = <text> eol

If the return value is nov, the data buffer should contain network overview database lines. These are basically fields separated by tabs.

     nov-buffer = *nov-line
     nov-line   = 8*9 [ field <TAB> ] eol
     field      = <text except TAB>

For a closer look at what should be in those fields, see Section 10.6.4 [Headers], page 223.

(nnchoke-open-server SERVER &optional DEFINITIONS)

server is here the virtual server name. definitions is a list of (VARIABLE VALUE) pairs that define this virtual server. If the server can’t be opened, no error should be signaled. The backend may then choose to refuse further attempts at connecting to this server. In fact, it should do so. If the server is opened already, this function should return a non-nil value. There should be no data returned.

(nnchoke-close-server &optional SERVER)

Close connection to server and free all resources connected to it. Return nil if the server couldn’t be closed for some reason. There should be no data returned.

(nnchoke-request-close)

Close connection to all servers and free all resources that the backend have reserved. All buffers that have been created by that backend should be killed. (Not the nntp-server-buffer, though.) This function is generally only called when Gnus is shutting down. There should be no data returned.

(nnchoke-server-opened &optional SERVER)

If server is the current virtual server, and the connection to the physical server is alive, then this function should return a non-nil vlue. This function should under no circumstances attempt to reconnect to a server we have lost connection to.

There should be no data returned.

(nnchoke-status-message &optional SERVER)

This function should return the last error message from server. There should be no data returned.

(nnchoke-request-article ARTICLE &optional GROUP SERVER TO-BUFFER)

The result data from this function should be the article specified by article. This might either be a Message-ID or a number. It is optional whether to implement retrieval by Message-ID, but it would be nice if that were possible. If to-buffer is non-nil, the result data should be returned in this buffer instead of the normal data buffer. This is to make it possible to avoid copying large amounts of data from one buffer to another, while Gnus mainly requests articles to be inserted directly into its article buffer. If it is at all possible, this function should return a cons cell where the car is the group name the article was fetched from, and the cdr is the article number. This will enable Gnus to find out what the real group and article numbers are when fetching articles by Message-ID. If this isn’t possible, t should be returned on successful article retrieval.

(nnchoke-request-group GROUP &optional SERVER FAST)

Get data on group. This function also has the side effect of making group the current group. If FAST, don’t bother to return useful data, just make group the current group. Here’s an example of some result data and a definition of the same:

     211 56 1000 1059 ifi.discussion

The first number is the status, which should be 211. Next is the total number of articles in the group, the lowest article number, the highest article number, and finally the group name. Note that the total number of articles may be less than one might think while just considering the highest and lowest article numbers, but some articles may have been canceled. Gnus just discards the total-number, so whether one should take the bother to generate it properly (if that is a problem) is left as an exercise to the reader.

     group-status = [ error / info ] eol
     error        = [ "4" / "5" ] 2<number> " " <Error message>
     info         = "211 " 3* [ <number> " " ] <string>
(nnchoke-close-group GROUP &optional SERVER)

Close group and free any resources connected to it. This will be a no-op on most backends. There should be no data returned.

(nnchoke-request-list &optional SERVER)

Return a list of all groups available on server. And that means all. Here’s an example from a server that only carries two groups:

     ifi.test 0000002200 0000002000 y
     ifi.discussion 3324 3300 n

On each line we have a group name, then the highest article number in that group, the lowest article number, and finally a flag.

     active-file = *active-line
     active-line = name " " <number> " " <number> " " flags eol
     name        = <string>
     flags       = "n" / "y" / "m" / "x" / "j" / "=" name

The flag says whether the group is read-only (‘n’), is moderated (‘m’), is dead (‘x’), is aliased to some other group (‘=other-group’) or none of the above (‘y’).

(nnchoke-request-post &optional SERVER)

This function should post the current buffer. It might return whether the posting was successful or not, but that’s not required. If, for instance, the posting is done asynchronously, it has generally not been completed by the time this function concludes. In that case, this function should set up some kind of sentinel to beep the user loud and clear if the posting could not be completed. There should be no result data from this function.

10.6.2.2 Optional Backend Functions

(nnchoke-retrieve-groups GROUPS &optional SERVER)

groups is a list of groups, and this function should request data on all those groups. How it does it is of no concern to Gnus, but it should attempt to do this in a speedy fashion. The return value of this function can be either active or group, which says what the format of the result data is. The former is in the same format as the data from nnchoke-request-list, while the latter is a buffer full of lines in

the same format as nnchoke-request-group gives.
     group-buffer = *active-line / *group-status
(nnchoke-request-update-info GROUP INFO &optional SERVER)

A Gnus group info (see Section 10.6.6 [Group Info], page 224) is handed to the backend for alterations. This comes in handy if the backend really carries all the information (as is the case with virtual and imap groups). This function should destructively alter the info to suit its needs, and should return the (altered) group info. There should be no result data from this function.

(nnchoke-request-type GROUP &optional ARTICLE)

When the user issues commands for “sending news” (F in the summary buffer, for instance), Gnus has to know whether the article the user is following up on is news or mail. This function should return news if article in group is news, mail if it is mail and unknown if the type can’t be decided. (The article parameter is necessary in nnvirtual groups which might very well combine mail groups and news groups.) Both group and article may be nil. There should be no result data from this function.

(nnchoke-request-update-mark GROUP ARTICLE MARK)

If the user tries to set a mark that the backend doesn’t like, this function may change the mark. Gnus will use whatever this function returns as the mark for article instead of the original mark. If the backend doesn’t care, it must return the original mark, and not nil or any other type of garbage. The only use for this I can see is what nnvirtual does with it—if a component group is auto-expirable, marking an article as read in the virtual group should result in the article being marked as expirable. There should be no result data from this function.

(nnchoke-request-scan &optional GROUP SERVER)

This function may be called at any time (by Gnus or anything else) to request that the backend check for incoming articles, in one way or another. A mail backend will typically read the spool file or query the POP server when this function is invoked. The group doesn’t have to be heeded—if the backend decides that it is too much work just scanning for a single group, it may do a total scan of all groups. It would be nice, however, to keep things local if that’s practical. There should be no result data from this function.

(nnchoke-request-group-description GROUP &optional SERVER)

The result data from this function should be a description of group.

     description-line = name <TAB> description eol
     name             = <string>
     description      = <text>
(nnchoke-request-list-newsgroups &optional SERVER)

The result data from this function should be the description of all groups available on the server.

     description-buffer = *description-line
(nnchoke-request-newgroups DATE &optional SERVER)

The result data from this function should be all groups that were created after ‘date’, which is in normal human-readable date format. The data should be in the active buffer format.

(nnchoke-request-create-group GROUP &optional SERVER)

This function should create an empty group with name group. There should be no return data.

(nnchoke-request-expire-articles ARTICLES &optional GROUP SERVER FORCE)

This function should run the expiry process on all articles in the articles range (which is currently a simple list of article numbers.) It is left up to the backend to decide how old articles should be before they are removed by this function. If force is non-nil, all articles should be deleted, no matter how new they are. This function should return a list of articles that it did not/was not able to delete. There should be no result data returned.

(nnchoke-request-move-article ARTICLE GROUP SERVER ACCEPT-FORM

&optional LAST) This function should move article (which is a number) from group by calling accept-form. This function should ready the article in question for moving by removing any header lines it has added to the article, and generally should “tidy up” the article. Then it should eval accept-form in the buffer where the “tidy” article is. This will do the actual copying. If this eval returns a non-nil value, the article should be removed. If last is nil, that means that there is a high likelihood that there will be more requests issued shortly, so that allows some optimizations. The function should return a cons where the car is the group name and the cdr is the article number that the article was entered as. There should be no data returned.

(nnchoke-request-accept-article GROUP &optional SERVER LAST)

This function takes the current buffer and inserts it into group. If last in nil, that means that there will be more calls to this function in short order. The function should return a cons where the car is the group name and the cdr is the article number that the article was entered as. There should be no data returned.

(nnchoke-request-replace-article ARTICLE GROUP BUFFER)

This function should remove article (which is a number) from group and insert buffer there instead. There should be no data returned.

(nnchoke-request-delete-group GROUP FORCE &optional SERVER)

This function should delete group. If force, it should really delete all the articles in the group, and then delete the group itself. (If there is such a thing as “the group itself”.) There should be no data returned.

(nnchoke-request-rename-group GROUP NEW-NAME &optional SERVER)

This function should rename group into new-name. All articles in group should move to new-name. There should be no data returned.

10.6.2.3 Error Messaging

The backends should use the function nnheader-report to report error conditions— they should not raise errors when they aren’t able to perform a request. The first argument to this function is the backend symbol, and the rest are interpreted as arguments to format if there are multiple of them, or just a string if there is one of them. This function must always returns nil.

(nnheader-report 'nnchoke "You did something totally bogus")

     (nnheader-report 'nnchoke "Could not request group %s" group)

Gnus, in turn, will call nnheader-get-report when it gets a nil back from a server, and this function returns the most recently reported message for the backend in question. This function takes one argument—the server symbol.

Internally, these functions access backend-status-string, so the nnchoke backend will have its error message stored in nnchoke-status-string.

10.6.2.4 Writing New Backends

Many backends are quite similar. nnml is just like nnspool, but it allows you to edit the articles on the server. nnmh is just like nnml, but it doesn’t use an active file, and it doesn’t maintain overview databases. nndir is just like nnml, but it has no concept of “groups”, and it doesn’t allow editing articles.

It would make sense if it were possible to “inherit” functions from backends when writing new backends. And, indeed, you can do that if you want to. (You don’t have to if you don’t want to, of course.)

All the backends declare their public variables and functions by using a package called

nnoo.

To inherit functions from other backends (and allow other backends to inherit functions from the current backend), you should use the following macros:

nnoo-declare

This macro declares the first parameter to be a child of the subsequent parameters. For instance:

               (nnoo-declare nndir
                 nnml nnmh)

nndir has declared here that it intends to inherit functions from both nnml and

          nnmh.

defvoo This macro is equivalent to defvar, but registers the variable as a public server variable. Most state-oriented variables should be declared with defvoo instead

          of defvar.

In addition to the normal defvar parameters, it takes a list of variables in the parent backends to map the variable to when executing a function in those backends.

               (defvoo nndir-directory nil
                 "Where nndir will look for groups."
                 nnml-current-directory nnmh-current-directory)
          This means that nnml-current-directory will be set to nndir-directory

when an nnml function is called on behalf of nndir. (The same with nnmh.)

nnoo-define-basics

This macro defines some common functions that almost all backends should have.

               (nnoo-define-basics nndir)

deffoo This macro is just like defun and takes the same parameters. In addition to doing the normal defun things, it registers the function as being public so that other backends can inherit it.

nnoo-map-functions

This macro allows mapping of functions from the current backend to functions from the parent backends.

     (nnoo-map-functions nndir
       (nnml-retrieve-headers 0 nndir-current-group 0 0)
       (nnmh-request-article 0 nndir-current-group 0 0))

This means that when nndir-retrieve-headers is called, the first, third, and fourth parameters will be passed on to nnml-retrieve-headers, while the second parameter is set to the value of nndir-current-group.

nnoo-import

This macro allows importing functions from backends. It should be the last thing in the source file, since it will only define functions that haven’t already been defined.

     (nnoo-import nndir
       (nnmh
        nnmh-request-list
        nnmh-request-newgroups)
       (nnml))

This means that calls to nndir-request-list should just be passed on to nnmh-request-list, while all public functions from nnml that haven’t been defined in nndir yet should be defined now.

Below is a slightly shortened version of the nndir backend.

;;; nndir.el --- single directory newsgroup access for Gnus
;; Copyright (C) 1995,96 Free Software Foundation, Inc.
;;; Code:
(require 'nnheader)
(require 'nnmh)
(require 'nnml)
(require 'nnoo)
(eval-when-compile (require 'cl))
(nnoo-declare nndir
nnml nnmh)
(defvoo nndir-directory nil
"Where nndir will look for groups."
nnml-current-directory nnmh-current-directory)
(defvoo nndir-nov-is-evil nil
"*Non-nil means that nndir will never retrieve NOV headers."
nnml-nov-is-evil)

(defvoo nndir-current-group "" nil nnml-current-group nnmh-current-group)
(defvoo nndir-top-directory nil nil nnml-directory nnmh-directory)
(defvoo nndir-get-new-mail nil nil nnml-get-new-mail nnmh-get-new-mail)
(defvoo nndir-status-string "" nil nnmh-status-string)
(defconst nndir-version "nndir 1.0")
;;; Interface functions.
(nnoo-define-basics nndir)
(deffoo nndir-open-server (server &optional defs)
  (setq nndir-directory
        (or (cadr (assq 'nndir-directory defs))
            server))
  (unless (assq 'nndir-directory defs)
    (push `(nndir-directory ,server) defs))
  (push `(nndir-current-group
          ,(file-name-nondirectory (directory-file-name nndir-directory)))
        defs)
  (push `(nndir-top-directory
          ,(file-name-directory (directory-file-name nndir-directory)))
        defs)
  (nnoo-change-server 'nndir server defs))
(nnoo-map-functions nndir
  (nnml-retrieve-headers 0 nndir-current-group 0 0)
  (nnmh-request-article 0 nndir-current-group 0 0)
  (nnmh-request-group nndir-current-group 0 0)
  (nnmh-close-group nndir-current-group 0))
(nnoo-import nndir
  (nnmh
   nnmh-status-message
   nnmh-request-list
   nnmh-request-newgroups))
(provide 'nndir)

10.6.2.5 Hooking New Backends Into Gnus

Having Gnus start using your new backend is rather easy—you just declare it with the gnus-declare-backend functions. This will enter the backend into the gnus-valid-

select-methods variable.

gnus-declare-backend takes two parameters—the backend name and an arbitrary number of abilities.

Here’s an example:

(gnus-declare-backend "nnchoke" 'mail 'respool 'address)

The abilities can be:

mail This is a mailish backend—followups should (probably) go via mail.

post This is a newsish backend—followups should (probably) go via news.

post-mail

This backend supports both mail and news.

none This is neither a post nor mail backend—it’s something completely different.

respool It supports respooling—or rather, it is able to modify its source articles and groups.

address The name of the server should be in the virtual server name. This is true for almost all backends.

prompt-address

The user should be prompted for an address when doing commands like B in the group buffer. This is true for backends like nntp, but not nnmbox, for instance.

10.6.2.6 Mail-like Backends

One of the things that separate the mail backends from the rest of the backends is the heavy dependence by the mail backends on common functions in ‘nnmail.el’. For instance, here’s the definition of nnml-request-scan:

     (deffoo nnml-request-scan (&optional group server)
       (setq nnml-article-file-alist nil)
       (nnmail-get-new-mail 'nnml 'nnml-save-nov nnml-directory group))

It simply calls nnmail-get-new-mail with a few parameters, and nnmail takes care of all the moving and splitting of the mail.

This function takes four parameters.

method This should be a symbol to designate which backend is responsible for the call.

exit-function This function should be called after the splitting has been performed.

temp-directory Where the temporary files should be stored.

group This optional argument should be a group name if the splitting is to be performed for one group only.

nnmail-get-new-mail will call backend-save-mail to save each article. backendactive-number will be called to find the article number assigned to this article.

The function also uses the following variables: backend-get-new-mail (to see whether to get new mail for this backend); and backend-group-alist and backend-active-file to generate the new active file. backend-group-alist should be a group-active alist, like this:

     (("a-group" (1 . 10))
      ("some-group" (34 . 39)))

10.6.3 Score File Syntax

Score files are meant to be easily parseable, but yet extremely mallable. It was decided that something that had the same read syntax as an Emacs Lisp list would fit that spec.

Here’s a typical score file:

(("summary"
  ("win95" -10000 nil s)
  ("Gnus"))
 ("from"
  ("Lars" -1000))
 (mark -100))

BNF definition of a score file:

score-file       = "" / "(" *element ")"
element          = rule / atom
rule             = string-rule / number-rule / date-rule
string-rule      = "(" quote string-header quote space *string-match ")"
number-rule      = "(" quote number-header quote space *number-match ")"
date-rule        = "(" quote date-header quote space *date-match ")"
quote            = <ascii 34>
string-header    = "subject" / "from" / "references" / "message-id" /
                   "xref" / "body" / "head" / "all" / "followup"
number-header    = "lines" / "chars"
date-header      = "date"
string-match     = "(" quote <string> quote [ "" / [ space score [ "" /
                   space date [ "" / [ space string-match-t ] ] ] ] ] ")"
score            = "nil" / <integer>
date             = "nil" / <natural number>
string-match-t   = "nil" / "s" / "substring" / "S" / "Substring" /
                   "r" / "regex" / "R" / "Regex" /
                   "e" / "exact" / "E" / "Exact" /
                   "f" / "fuzzy" / "F" / "Fuzzy"
number-match     = "(" <integer> [ "" / [ space score [ "" /
                   space date [ "" / [ space number-match-t ] ] ] ] ] ")"
number-match-t   = "nil" / "=" / "<" / ">" / ">=" / "<="
date-match       = "(" quote <string> quote [ "" / [ space score [ "" /
                   space date [ "" / [ space date-match-t ] ] ] ] ")"
date-match-t     = "nil" / "at" / "before" / "after"
atom             = "(" [ required-atom / optional-atom ] ")"
required-atom    = mark / expunge / mark-and-expunge / files /
                   exclude-files / read-only / touched
optional-atom    = adapt / local / eval
mark             = "mark" space nil-or-number
nil-or-number    = "nil" / <integer>
expunge          = "expunge" space nil-or-number
mark-and-expunge = "mark-and-expunge" space nil-or-number
files            = "files" *[ space <string> ]
exclude-files    = "exclude-files" *[ space <string> ]
read-only        = "read-only" [ space "nil" / space "t" ]

     adapt            = "adapt" [ space "ignore" / space "t" / space adapt-rule ]
     adapt-rule       = "(" *[ <string> *[ "(" <string> <integer> ")" ] ")"
     local            = "local" *[ space "(" <string> space <form> ")" ]
     eval             = "eval" space <form>
     space            = *[ " " / <TAB> / <NEWLINE> ]

Any unrecognized elements in a score file should be ignored, but not discarded.

As you can see, white space is needed, but the type and amount of white space is irrelevant. This means that formatting of the score file is left up to the programmer—if it’s simpler to just spew it all out on one looong line, then that’s ok.

The meaning of the various atoms are explained elsewhere in this manual (see Section 7.4 [Score File Format], page 136).

10.6.4 Headers

Internally Gnus uses a format for storing article headers that corresponds to the nov format in a mysterious fashion. One could almost suspect that the author looked at the nov specification and just shamelessly stole the entire thing, and one would be right.

Header is a severely overloaded term. “Header” is used in RFC1036 to talk about lines in the head of an article (e.g., From). It is used by many people as a synonym for “head”— “the header and the body”. (That should be avoided, in my opinion.) And Gnus uses a format internally that it calls “header”, which is what I’m talking about here. This is a 9-element vector, basically, with each header (ouch) having one slot.

   These slots are, in order: number, subject, from, date, id, references, chars, lines,

xref. There are macros for accessing and setting these slots—they all have predictable names beginning with mail-header- and mail-header-set-, respectively.

The xref slot is really a misc slot. Any extra info will be put in there.

10.6.5 Ranges

gnus introduced a concept that I found so useful that I’ve started using it a lot and have elaborated on it greatly.

The question is simple: If you have a large amount of objects that are identified by numbers (say, articles, to take a wild example) that you want to qualify as being “included”, a normal sequence isn’t very useful. (A 200,000 length sequence is a bit long-winded.)

The solution is as simple as the question: You just collapse the sequence.

     (1 2 3 4 5 6 10 11 12)

is transformed into

     ((1 . 6) (10 . 12))

To avoid having those nasty ‘(13 . 13)’ elements to denote a lonesome object, a ‘13’ is a valid element:

     ((1 . 6) 7 (10 . 12))

This means that comparing two ranges to find out whether they are equal is slightly tricky:

     ((1 . 5) 7 8 (10 . 12))

and

     ((1 . 5) (7 . 8) (10 . 12))

are equal. In fact, any non-descending list is a range:

     (1 2 3 4 5)

is a perfectly valid range, although a pretty long-winded one. This is also valid:

     (1 . 5)

and is equal to the previous range.

Here’s a BNF definition of ranges. Of course, one must remember the semantic requirement that the numbers are non-descending. (Any number of repetition of the same number is allowed, but apt to disappear in range handling.)

     range           = simple-range / normal-range
     simple-range    = "(" number " . " number ")"
     normal-range    = "(" start-contents ")"
     contents        = "" / simple-range *[ " " contents ] /
                       number *[ " " contents ]

Gnus currently uses ranges to keep track of read articles and article marks. I plan on implementing a number of range operators in C if The Powers That Be are willing to let me. (I haven’t asked yet, because I need to do some more thinking on what operators I need to make life totally range-based without ever having to convert back to normal sequences.)

10.6.6 Group Info

Gnus stores all permanent info on groups in a group info list. This list is from three to six elements (or more) long and exhaustively describes the group.

Here are two example group infos; one is a very simple group while the second is a more complex one:

     ("no.group" 5 (1 . 54324))
     ("nnml:my.mail" 3 ((1 . 5) 9 (20 . 55))
                     ((tick (15 . 19)) (replied 3 6 (19 . 3)))
                     (nnml "")
                     ((auto-expire . t) (to-address . "ding@gnus.org")))

The first element is the group name—as Gnus knows the group, anyway. The second element is the subscription level, which normally is a small integer. (It can also be the rank, which is a cons cell where the car is the level and the cdr is the score.) The third element is a list of ranges of read articles. The fourth element is a list of lists of article marks of various kinds. The fifth element is the select method (or virtual server, if you like). The sixth element is a list of group parameters, which is what this section is about.

Any of the last three elements may be missing if they are not required. In fact, the vast majority of groups will normally only have the first three elements, which saves quite a lot of cons cells.

Here’s a BNF definition of the group info format:

     info          = "(" group space ralevel space read
                     [ "" / [ space marks-list [ "" / [ space method [ "" /
                     space parameters ] ] ] ] ] ")"
     group         = quote <string> quote

     ralevel       = rank / level
     level         = <integer in the range of 1 to inf>
     rank          = "(" level "." score ")"
     score         = <integer in the range of 1 to inf>
     read          = range
     marks-lists   = nil / "(" *marks ")"
     marks         = "(" <string> range ")"
     method        = "(" <string> *elisp-forms ")"
     parameters    = "(" *elisp-forms ")"

Actually that ‘marks’ rule is a fib. A ‘marks’ is a ‘<string>’ consed on to a ‘range’, but that’s a bitch to say in pseudo-BNF.

If you have a Gnus info and want to access the elements, Gnus offers a series of macros for getting/setting these elements.

gnus-info-group
gnus-info-set-group

Get/set the group name.

gnus-info-rank
gnus-info-set-rank

Get/set the group rank (see Section 2.7 [Group Score], page 17).

gnus-info-level
gnus-info-set-level

Get/set the group level.

gnus-info-score
gnus-info-set-score

Get/set the group score (see Section 2.7 [Group Score], page 17).

gnus-info-read
gnus-info-set-read

Get/set the ranges of read articles.

gnus-info-marks
gnus-info-set-marks

Get/set the lists of ranges of marked articles.

gnus-info-method
gnus-info-set-method

Get/set the group select method.

gnus-info-params
gnus-info-set-params

Get/set the group parameters.

All the getter functions take one parameter—the info list. The setter functions take two parameters—the info list and the new value.

The last three elements in the group info aren’t mandatory, so it may be necessary to extend the group info before setting the element. If this is necessary, you can just pass on a non-nil third parameter to the three final setter functions to have this happen automatically.

10.6.7 Extended Interactive

Gnus extends the standard Emacs interactive specification slightly to allow easy use of the symbolic prefix (see Section 8.3 [Symbolic Prefixes], page 154). Here’s an example of how this is used:

     (defun gnus-summary-increase-score (&optional score symp)
       (interactive (gnus-interactive "P\ny"))
       ...
       )

The best thing to do would have been to implement gnus-interactive as a macro which would have returned an interactive form, but this isn’t possible since Emacs checks whether a function is interactive or not by simply doing an assq on the lambda form. So, instead we have gnus-interactive function that takes a string and returns values that are

usable to interactive.

This function accepts (almost) all normal interactive specs, but adds a few more.

‘y’ The current symbolic prefix—the gnus-current-prefix-symbol variable.

‘Y’ A list of the current symbolic prefixes—the gnus-current-prefix-symbol variable.

‘A’ The current article number—the gnus-summary-article-number function.

‘H’ The current article header—the gnus-summary-article-header function.

‘g’ The current group name—the gnus-group-group-name function.

10.6.8 Emacs/XEmacs Code

While Gnus runs under Emacs, XEmacs and Mule, I decided that one of the platforms must be the primary one. I chose Emacs. Not because I don’t like XEmacs or Mule, but because it comes first alphabetically.

This means that Gnus will byte-compile under Emacs with nary a warning, while XEmacs will pump out gigabytes of warnings while byte-compiling. As I use byte-compilation warnings to help me root out trivial errors in Gnus, that’s very useful.

I’ve also consistently used Emacs function interfaces, but have used Gnusey aliases for the functions. To take an example: Emacs defines a run-at-time function while XEmacs defines a start-itimer function. I then define a function called gnus-run-at-time that takes the same parameters as the Emacs run-at-time. When running Gnus under Emacs, the former function is just an alias for the latter. However, when running under XEmacs, the former is an alias for the following function:

     (defun gnus-xmas-run-at-time (time repeat function &rest args)
       (start-itimer
        "gnus-run-at-time"
        `(lambda ()
           (,function ,@args))
        time repeat))

This sort of thing has been done for bunches of functions. Gnus does not redefine any native Emacs functions while running under XEmacs—it does this defalias thing with Gnus equivalents instead. Cleaner all over.

In the cases where the XEmacs function interface was obviously cleaner, I used it instead. For example gnus-region-active-p is an alias for region-active-p in XEmacs, whereas in Emacs it is a function.

Of course, I could have chosen XEmacs as my native platform and done mapping functions the other way around. But I didn’t. The performance hit these indirections impose on Gnus under XEmacs should be slight.

10.6.9 Various File Formats

10.6.9.1 Active File Format

The active file lists all groups available on the server in question. It also lists the highest and lowest current article numbers in each group.

Here’s an excerpt from a typical active file:

soc.motss 296030 293865 y
alt.binaries.pictures.fractals 3922 3913 n
comp.sources.unix 1605 1593 m
comp.binaries.ibm.pc 5097 5089 y
no.general 1000 900 y

Here’s a pseudo-BNF definition of this file:

active      = *group-line
group-line  = group space high-number space low-number space flag <NEWLINE>
group       = <non-white-space string>
space       = " "
high-number = <non-negative integer>
low-number  = <positive integer>
flag        = "y" / "n" / "m" / "j" / "x" / "=" group

For a full description of this file, see the manual pages for ‘innd’, in particular

‘active(5)’.

10.6.9.2 Newsgroups File Format

The newsgroups file lists groups along with their descriptions. Not all groups on the server have to be listed, and not all groups in the file have to exist on the server. The file is meant purely as information to the user.

The format is quite simple; a group name, a tab, and the description. Here’s the definition:

newsgroups    = *line
line          = group tab description <NEWLINE>
group         = <non-white-space string>
tab           = <TAB>
description   = <string>

10.7 Emacs for Heathens

Believe it or not, but some people who use Gnus haven’t really used Emacs much before they embarked on their journey on the Gnus Love Boat. If you are one of those unfortunates whom “M-C-a”, “kill the region”, and “set gnus-flargblossen to an alist where the key is a regexp that is used for matching on the group name” are magical phrases with little or no meaning, then this appendix is for you. If you are already familiar with Emacs, just ignore this and go fondle your cat instead.

10.7.1 Keystrokes

• Q: What is an experienced Emacs user?

• A: A person who wishes that the terminal had pedals.

Yes, when you use Emacs, you are apt to use the control key, the shift key and the meta key a lot. This is very annoying to some people (notably vile users), and the rest of us just love the hell out of it. Just give up and submit. Emacs really does stand for “Escape-Meta-Alt-Control-Shift”, and not “Editing Macros”, as you may have heard from other disreputable sources (like the Emacs author).

The shift keys are normally located near your pinky fingers, and are normally used to get capital letters and stuff. You probably use it all the time. The control key is normally marked “CTRL” or something like that. The meta key is, funnily enough, never marked as such on any keyboard. The one I’m currently at has a key that’s marked “Alt”, which is the meta key on this keyboard. It’s usually located somewhere to the left hand side of the keyboard, usually on the bottom row.

Now, us Emacs people don’t say “press the meta-control-m key”, because that’s just too inconvenient. We say “press the M-C-m key”. M- is the prefix that means “meta” and “C-” is the prefix that means “control”. So “press C-k” means “press down the control key, and hold it down while you press k”. “Press M-C-k” means “press down and hold down the meta key and the control key and then press k”. Simple, ay?

This is somewhat complicated by the fact that not all keyboards have a meta key. In that case you can use the “escape” key. Then M-k means “press escape, release escape, press k”. That’s much more work than if you have a meta key, so if that’s the case, I respectfully suggest you get a real keyboard with a meta key. You can’t live without it.

10.7.2 Emacs Lisp

Emacs is the King of Editors because it’s really a Lisp interpreter. Each and every key you tap runs some Emacs Lisp code snippet, and since Emacs Lisp is an interpreted language, that means that you can configure any key to run any arbitrary code. You just, like, do it.

Gnus is written in Emacs Lisp, and is run as a bunch of interpreted functions. (These are byte-compiled for speed, but it’s still interpreted.) If you decide that you don’t like the way Gnus does certain things, it’s trivial to have it do something a different way. (Well, at least if you know how to write Lisp code.) However, that’s beyond the scope of this manual, so we are simply going to talk about some common constructs that you normally use in your ‘.emacs’ file to customize Gnus.

If you want to set the variable gnus-florgbnize to four (4), you write the following:

     (setq gnus-florgbnize 4)

This function (really “special form”) setq is the one that can set a variable to some value. This is really all you need to know. Now you can go and fill your .emacs file with lots of these to change how Gnus works.

If you have put that thing in your .emacs file, it will be read and evaled (which is lisp-ese for “run”) the next time you start Emacs. If you want to change the variable right away, simply say C-x C-e after the closing parenthesis. That will eval the previous “form”, which is a simple setq statement here.

Go ahead—just try it, if you’re located at your Emacs. After you C-x C-e, you will see ‘4’ appear in the echo area, which is the return value of the form you evaled.

Some pitfalls:

If the manual says “set gnus-read-active-file to some”, that means:

     (setq gnus-read-active-file 'some)

On the other hand, if the manual says “set gnus-nntp-server to ‘nntp.ifi.uio.no’”, that means:

     (setq gnus-nntp-server "nntp.ifi.uio.no")

So be careful not to mix up strings (the latter) with symbols (the former). The manual is unambiguous, but it can be confusing.

\input texinfo

10.8 Frequently Asked Questions

This is the Gnus Frequently Asked Questions list. If you have a Web browser, the

official hypertext version is at ‘http://www.miranova.com/~steve/gnus-faq.html>’, and

has probably been updated since you got this manual.

10.8.1 Installation

• Q1.1 What is the latest version of Gnus?

The latest (and greatest) version is 5.0.10. You might also run across something called September Gnus. September Gnus is the alpha version of the next major release of Gnus. It is currently not stable enough to run unless you are prepared to debug lisp.

• Q1.2 Where do I get Gnus?

Any of the following locations:

 − ‘ftp://ftp.ifi.uio.no/pub/emacs/gnus/gnus.tar.gz’
 − ‘ftp://ftp.pilgrim.umass.edu/pub/misc/ding/’
 − ‘gopher://gopher.pilgrim.umass.edu/11/pub/misc/ding/’
 − ‘ftp://aphrodite.nectar.cs.cmu.edu/pub/ding-gnus/’
 − ‘ftp://ftp.solace.mh.se:/pub/gnu/elisp/’

• Q1.3 Which version of Emacs do I need?

At least GNU Emacs 19.28, or XEmacs 19.12 is recommended. GNU Emacs 19.25 has been reported to work under certain circumstances, but it doesn’t officially work on it. 19.27 has also been reported to work. Gnus has been reported to work under OS/2 as well as Unix.

• Q1.4 Where is timezone.el?

Upgrade to XEmacs 19.13. In earlier versions of XEmacs this file was placed with Gnus 4.1.3, but that has been corrected.

• Q1.5 When I run Gnus on XEmacs 19.13 I get weird error messages. You’re running an old version of Gnus. Upgrade to at least version 5.0.4.

• Q1.6 How do I unsubscribe from the Mailing List?

Send an e-mail message to ‘ding-request@ifi.uio.no’ with the magic word unsubscribe somewhere in it, and you will be removed. If you are reading the digest version of the list, send an e-mail message to

‘ding-rn-digests-d-request@moe.shore.net’ with unsubscribe as the subject and

you will be removed.

• Q1.7 How do I run Gnus on both Emacs and XEmacs?

The basic answer is to byte-compile under XEmacs, and then you can run under either Emacsen. There is, however, a potential version problem with easymenu.el with Gnu Emacs prior to 19.29. Per Abrahamsen <abraham@dina.kvl.dk> writes : The internal easymenu.el interface changed between 19.28 and 19.29 in order to make it possible to create byte compiled files that can be shared between Gnu Emacs and XEmacs. The change is upward compatible, but not downward compatible. This gives the following compatibility table:

     Compiled with:  | Can be used with:
     ----------------+--------------------------------------
     19.28           | 19.28         19.29
     19.29           |               19.29           XEmacs
     XEmacs          |               19.29           XEmacs

If you have Gnu Emacs 19.28 or earlier, or XEmacs 19.12 or earlier, get a recent version

of auc-menu.el from ‘ftp://ftp.iesd.auc.dk/pub/emacs-lisp/auc-menu.el’, and

install it under the name easymenu.el somewhere early in your load path.

• Q1.8 What resources are available?

There is the newsgroup Gnu.emacs.gnus. Discussion of Gnus 5.x is now taking place there. There is also a mailing list, send mail to ‘ding-request@ifi.uio.no’ with the magic word subscribe somewhere in it. NOTE: the traffic on this list is heavy so you may not want to be on it (unless you use Gnus as your mailer reader, that is). The mailing list is mainly for developers and testers. Gnus has a home World Wide Web page at

‘http://www.ifi.uio.no/~larsi/ding.html’.

Gnus has a write up in the X Windows Applications FAQ at

‘http://www.ee.ryerson.ca:8080/~elf/xapps/Q-III.html’.

The Gnus manual is also available on the World Wide Web. The canonical source is in Norway at

‘http://www.ifi.uio.no/~larsi/ding-manual/gnus_toc.html’.

There are three mirrors in the United States:

 1. ‘http://www.miranova.com/gnus-man/’
 2. ‘http://www.pilgrim.umass.edu/pub/misc/ding/manual/gnus_toc.html’
 3. ‘http://www.rtd.com/~woo/gnus/’

PostScript copies of the Gnus Reference card are available from

‘ftp://ftp.cs.ualberta.ca/pub/oolog/gnus/’. They are mirrored at
‘ftp://ftp.pilgrim.umass.edu/pub/misc/ding/refcard/’ in the United States.

And

‘ftp://marvin.fkphy.uni-duesseldorf.de/pub/gnus/’ in Germany.

An online version of the Gnus FAQ is available at

‘http://www.miranova.com/~steve/gnus-faq.html’. Off-line formats are also avail-

able:

ASCII: ‘ftp://ftp.miranova.com/pub/gnus/gnus-faq’
PostScript: ‘ftp://ftp.miranova.com/pub/gnus/gnus-faq.ps’.

• Q1.9 Gnus hangs on connecting to NNTP server I am running XEmacs on SunOS and Gnus prints a message about Connecting to NNTP server and then just hangs. Ben Wing <wing@netcom.com> writes : I wonder if you’re hitting the infamous libresolv problem. The basic problem is that under SunOS you can compile either with DNS or NIS name lookup libraries but not both. Try substituting the IP address and see if that works; if so, you need to download the sources and recompile.

• Q1.10 Mailcrypt 3.4 doesn’t work This problem is verified to still exist in Gnus 5.0.9 and MailCrypt 3.4. The answer comes from Peter Arius <arius@immd2.informatik.uni-erlangen.de>. I found out that mailcrypt uses gnus-eval-in-buffer-window, which is a macro. It seems as if you have compiled mailcrypt with plain old GNUS in load path, and the XEmacs byte compiler has inserted that macro definition into ‘mc-toplev.elc’. The solution is to recompile ‘mc-toplev.el’ with Gnus 5 in load-path, and it works fine. Steve Baur <steve@miranova.com> adds : The problem also manifests itself if neither GNUS 4 nor Gnus 5 is in the load-path.

• Q1.11 What other packages work with Gnus?

− Mailcrypt. Mailcrypt is an Emacs interface to PGP. It works, it installs without hassle, and integrates very easily. Mailcrypt can be obtained from

    ‘ftp://cag.lcs.mit.edu/pub/patl/mailcrypt-3.4.tar.gz’.

− Tiny Mime. Tiny Mime is an Emacs MUA interface to MIME. Installation is a two-step process unlike most other packages, so you should be prepared to move the byte-compiled code somewhere. There are currently two versions of this package available. It can be obtained from

    ‘ftp://ftp.jaist.ac.jp/pub/GNU/elisp/’. Be sure to apply the supplied patch.

It works with Gnus through version 5.0.9. In order for all dependencies to work correctly the load sequence is as follows:

           (load "tm-setup")
           (load "gnus")
           (load "mime-compose")

NOTE: Loading the package disables citation highlighting by default. To get the old behavior back, use the M-t command.

10.8.2 Customization

• Q2.1 Custom Edit does not work under XEmacs The custom package has not been ported to XEmacs.

• Q2.2 How do I quote messages?

I see lots of messages with quoted material in them. I am wondering how to have Gnus do it for me. This is Gnus, so there are a number of ways of doing this. You can use the built-in commands to do this. There are the F and R keys from the summary buffer which automatically include the article being responded to. These commands are also selectable as Followup and Yank and Reply and Yank in the Post menu. C-c C-y grabs the previous message and prefixes each line with ail-indentation-spaces spaces or mail-yank-prefix if that is non-nil, unless you have set your own mail-citation-hook, which will be called to to do the job.

You might also consider the Supercite package, which allows for pretty arbitrarily complex quoting styles. Some people love it, some people hate it.

• Q2.3 How can I keep my nnvirtual:* groups sorted? How can I most efficiently arrange matters so as to keep my nnvirtual:* (etc) groups at the top of my group selection buffer, whilst keeping everything sorted in alphabetical order. If you don’t subscribe often to new groups then the easiest way is to first sort the groups and then manually kill and yank the virtuals wherever you want them.

• Q2.4 Any good suggestions on stuff for an all.SCORE file? Here is a collection of suggestions from the Gnus mailing list.

1. From “Dave Disser” <disser@sdd.hp.com> I like blasting anything without lowercase letters. Weeds out most of the make $$ fast, as well as the lame titles like “IBM” and “HP-UX” with no further description.

(("Subject"
("^\\(Re: \\)?[^a-z]*$" -200 nil R)))

2. From “Peter Arius” <arius@immd2.informatik.uni-erlangen.de>

The most vital entries in my (still young) all.SCORE:

(("xref"
("alt.fan.oj-simpson" -1000 nil s))
("subject"
("\\<\\(make\\|fast\\|big\\)\\s-*\\(money\\|cash\\|bucks?\\)\\>" -1000 nil r)
("$$$$" -1000 nil s)))

3. From “Per Abrahamsen” <abraham@dina.kvl.dk>

(("subject"
;; CAPS OF THE WORLD, UNITE
("^..[^a-z]+$" -1 nil R)
;; $$$ Make Money $$$ (Try work)
("$" -1 nil s)
;; I'm important! And I have exclamation marks to prove it!
("!" -1 nil s)))

4. From “heddy boubaker” <boubaker@cenatls.cena.dgac.fr>

I would like to contribute with mine.

(
(read-only t)
("subject"
;; ALL CAPS SUBJECTS
("^\\([Rr][Ee]: +\\)?[^a-z]+$" -1 nil R)
;; $$$ Make Money $$$
("$$" -10 nil s)
;; Empty subjects are worthless!
("^*\\([(<]none[>)]\\|(no subject\\( given\\)?)\\)? *$" -10 nil r)
;; Sometimes interesting announces occur!
("ANN?OU?NC\\(E\\|ING\\)" +10 nil r)
;; Some people think they're on mailing lists

("\\(un\\)?sub?scribe" -100 nil r)
;; Stop Micro$oft NOW!!
("\\(m\\(icro\\)?[s$]\\(oft\\|lot\\)?-?\\)?wind?\\(ows\\|aube\\|oze\\)?[- ]*\\('?95\\|NT\\|3[.]1\\|32\\)" -1001 nil r)
;; I've nothing to buy
("\\(for\\|4\\)[- ]*sale" -100 nil r)
;; SELF-DISCIPLINED people
("\\[[^a-z0-9 \t\n][^a-z0-9 \t\n]\\]" +100 nil r)
)
("from"
;; To keep track of posters from my site
(".dgac.fr" +1000 nil s))
("followup"
;; Keep track of answers to my posts
("boubaker" +1000 nil s))
("lines"
;; Some people have really nothing to say!!
(1 -10 nil <=))
(mark -100)
(expunge -1000)
)

5. From “Christopher Jones” <cjones@au.oracle.com>

The sample ‘all.SCORE’ files from Per and boubaker could be augmented with:

(("subject"
;; No junk mail please!
("please ignore" -500 nil s)
("test" -500 nil e))
)

6. From “Brian Edmonds” <edmonds@cs.ubc.ca> Augment any of the above with a fast method of scoring down excessively cross posted articles.

("xref"
;; the more cross posting, the exponentially worse the article
("^xref: \\S-+ \\S-+ \\S-+ \\S-+" -1 nil r)
("^xref: \\S-+ \\S-+ \\S-+ \\S-+ \\S-+" -2 nil r)
("^xref: \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+" -4 nil r)
("^xref: \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+" -8 nil r)
("^xref: \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+" -16 nil r)
("^xref: \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+" -32 nil r)
("^xref: \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+" -64 nil r)
("^xref: \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+" -128 nil r)
("^xref: \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+" -256 nil r)
("^xref: \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+ \\S-+" -512 nil r))

• Q2.5 What do I use to yank-through when replying? You should probably reply and followup with R and F, instead of r and f, which solves your problem. But you could try something like:

(defconst mail-yank-ignored-headers
"^.*:"
"Delete these headers from old message when it's inserted in a reply.")

• Q2.6 I don’t like the default WWW browser Now when choosing an URL Gnus starts up a W3 buffer, I would like it to always use Netscape (I don’t browse in text-mode ;-).

1. Activate ‘Customize...’ from the ‘Help’ menu.

2. Scroll down to the ‘WWW Browser’ field.

3. Click ‘mouse-2’ on ‘WWW Browser’.

4. Select ‘Netscape’ from the pop up menu.

5. Press ‘C-c C-c’ If you are using XEmacs then to specify Netscape do

       (setq gnus-button-url 'gnus-netscape-open-url)

• Q2.7 What, if any, relation is between “ask-server” and “(setq gnus-read-active-file ’some)”? In order for Gnus to show you the complete list of newsgroups, it will either have to either store the list locally, or ask the server to transmit the list. You enable the first with

             (setq gnus-save-killed-list t)

and the second with

             (setq gnus-read-active-file t)

If both are disabled, Gnus will not know what newsgroups exists. There is no option to get the list by casting a spell.

• Q2.8 Moving between groups is slow.

Per Abrahamsen <abraham@dina.kvl.dk> writes:

Do you call define-key or something like that in one of the summary mode hooks? This would force Emacs to recalculate the keyboard shortcuts. Removing the call should speed up M-x gnus-summary-mode RET by a couple of orders of magnitude. You can use

     (define-key gnus-summary-mode-map KEY COMMAND)

in your ‘.gnus’ instead.

10.8.3 Reading News

• Q3.1 How do I convert my kill files to score files?

A kill-to-score translator was written by Ethan Bradford <ethanb@ptolemy.astro.washington.edu>. It is available from

‘http://baugi.ifi.uio.no/~larsi/ding-various/gnus-kill-to-score.el’.

• Q3.2 My news server has a lot of groups, and killing groups is painfully slow. Don’t do that then. The best way to get rid of groups that should be dead is to edit your newsrc directly. This problem will be addressed in the near future.

• Q3.3 How do I use an NNTP server with authentication?

Put the following into your .gnus:

      (add-hook 'nntp-server-opened-hook 'nntp-send-authinfo)

• Q3.4 Not reading the first article.

How do I avoid reading the first article when a group is selected?

1. Use RET to select the group instead of SPC.

 2. (setq gnus-auto-select first nil)

3. Luis Fernandes <elf@mailhost.ee.ryerson.ca>writes:

This is what I use...customize as necessary...

         ;;; Don't auto-select first article if reading sources, or archives or
         ;;; jobs postings, etc. and just display the summary buffer
         (add-hook 'gnus-select-group-hook
           (function
            (lambda ()
              (cond ((string-match "sources" gnus-newsgroup-name)
             (setq gnus-auto-select-first nil))
            ((string-match "jobs" gnus-newsgroup-name)
         (setq gnus-auto-select-first nil))
            ((string-match "comp\\.archives" gnus-newsgroup-name)
             (setq gnus-auto-select-first nil))
            ((string-match "reviews" gnus-newsgroup-name)
             (setq gnus-auto-select-first nil))
            ((string-match "announce" gnus-newsgroup-name)
             (setq gnus-auto-select-first nil))
            ((string-match "binaries" gnus-newsgroup-name)
             (setq gnus-auto-select-first nil))
            (t
             (setq gnus-auto-select-first t))))))

4. Per Abrahamsen <abraham@dina.kvl.dk> writes:

Another possibility is to create an ‘all.binaries.all.SCORE’ file like this:

         ((local
           (gnus-auto-select-first nil)))

and insert

         (setq gnus-auto-select-first t)

in your ‘.gnus’.

• Q3.5 Why aren’t BBDB known posters marked in the summary buffer? Brian Edmonds <edmonds@cs.ubc.ca> writes: Due to changes in Gnus 5.0, ‘bbdb-gnus.el’ no longer marks known posters in the summary buffer. An updated version, ‘gnus-bbdb.el’ is available at the locations listed below. This package also supports autofiling of incoming mail to folders specified in the BBDB. Extensive instructions are included as comments in the file. Send mail to ‘majordomo@edmonds.home.cs.ubc.ca’ with the following line in the body of the message: get misc gnus-bbdb.el. Or get it from the World Wide Web:

‘http://www.cs.ubc.ca/spider/edmonds/gnus-bbdb.el’.

10.8.4 Reading Mail

• Q4.1 What does the message “Buffer has changed on disk” mean in a mail group? Your filter program should not deliver mail directly to your folders, instead it should put the mail into spool files. Gnus will then move the mail safely from the spool files into the folders. This will eliminate the problem. Look it up in the manual, in the section entitled “Mail & Procmail”.

• Q4.2 How do you make articles un-expirable?

I am using nnml to read news and have used gnus-auto-expirable-newsgroups to automagically expire articles in some groups (Gnus being one of them). Sometimes there are interesting articles in these groups that I want to keep. Is there any way of explicitly marking an article as un-expirable - that is mark it as read but not expirable? Use u, !, d or M-u in the summary buffer. You just remove the E mark by setting some other mark. It’s not necessary to tick the articles.

• Q4.3 How do I delete bogus nnml: groups?

My problem is that I have various mail (nnml) groups generated while experimenting with Gnus. How do I remove them now? Setting the level to 9 does not help. Also

gnus-group-check-bogus-groups does not recognize them.

Removing mail groups is tricky at the moment. (It’s on the to-do list, though.) You basically have to kill the groups in Gnus, shut down Gnus, edit the active file to exclude these groups, and probably remove the nnml directories that contained these groups as well. Then start Gnus back up again.

• Q4.4 What happened to my new mail groups?

I got new mail, but I have never seen the groups they should have been placed in. They are probably there, but as zombies. Press A z to list zombie groups, and then subscribe to the groups you want with u. This is all documented quite nicely in the user’s manual.

• Q4.5 Not scoring mail groups How do you totally turn off scoring in mail groups? Use an nnbabyl:all.SCORE (or nnmh, or nnml, or whatever) file containing:

     ((adapt ignore)
      (local (gnus-use-scoring nil))
      (exclude-files "all.SCORE"))

238 Gnus Manual

Chapter 11: Index 239

11 Index

%

% . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

(

(ding) archive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

*

* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

.

.newsrc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 .newsrc.el . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 .newsrc.eld . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

>

> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

<

< . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

1

1153 digest . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115

A

activating groups . . . . . . . . . . . . . . . . . . . . . . . . 30, 205 active file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9, 205 adapt file group parameter . . . . . . . . . . . . . . . . . . . 21 adaptive scoring . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 adopting articles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 Agent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 Agent expire . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 Alta Vista . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 120 ange-ftp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 archive group. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 archived messages. . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 article . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 article backlog . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 article buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 article caching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 article customization . . . . . . . . . . . . . . . . . . . . . . . . . 83 article emphasis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 article expiry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 article hiding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 article history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 article marking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 article pre-fetch. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 article scrolling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 article series . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 article signature . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 article threading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 article ticking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 article washing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 article-de-quoted-unreadable . . . . . . . . . . . . . . . . . 109 article-decode-rfc1522 . . . . . . . . . . . . . . . . . . . . . . . 102 asynchronous article fetching . . . . . . . . . . . . . . . . . 53 authentification. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 authinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 auto-expire . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 auto-save. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

B

babyl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 backend. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 backlog . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 backup files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 batch scoring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 bbb-summary-rate-article. . . . . . . . . . . . . . . . . . . . 148 binary groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 body . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 bogus groups. . . . . . . . . . . . . . . . . . . . . . . . . . . . 24, 205 bookmarks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 bouncing mail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 broken-reply-to . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 browsing servers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178, 209 buttons. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68, 162 byte-compilation. . . . . . . . . . . . . . . . . . . . . . . . . . . . 160

240 Gnus Manual

C

caching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 canceling articles . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 CancelMoose[tm] . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 changing servers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 characters in file names . . . . . . . . . . . . . . . . . . . . . 174 child . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 Chris Lewis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 ClariNet Briefs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 click . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 colophon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 colors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 compatibility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 compilation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 composing mail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 composing messages. . . . . . . . . . . . . . . . . . . . . . . . . . 85 composing news . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 contributors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 converting kill files . . . . . . . . . . . . . . . . . . . . . . . . . . 147 copy mail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 cross-posting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 crosspost . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 crosspost mail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 crossposting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 crossposts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 customizing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 customizing threading . . . . . . . . . . . . . . . . . . . . . . . . 47

D

daemons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 decays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 decoding articles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 DejaNews. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 120 delete-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 deleting headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 deleting incoming files. . . . . . . . . . . . . . . . . . . . . . . 103 demons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 describing groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 digest . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115, 206 digests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 ding mailing list . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 directory groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 disk space . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 display-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 documentation group . . . . . . . . . . . . . . . . . . . . . . . 114 drafts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 dribble file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 duplicate mails . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109

E

elm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 Emacs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 Emacsen. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179, 226 email spam. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 emphasis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 ephemeral groups . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 excessive crossposting . . . . . . . . . . . . . . . . . . . . . . . . 40 exiting Gnus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 exiting groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 expirable mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 expiry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 expiry-wait . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

F

faces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 fancy mail splitting . . . . . . . . . . . . . . . . . . . . . . . . . 103 FAQ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 fetching a group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 fetching by Message-ID. . . . . . . . . . . . . . . . . . . . . . . 71 file commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 file names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 finding news . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 first time usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 follow up. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 followup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 fonts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 foreign . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 foreign groups. . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 91 foreign servers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 format-time-string . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 formatting variables. . . . . . . . . . . . . . . . . . . . . . . . . 154 forwarded messages . . . . . . . . . . . . . . . . . . . . . . . . . 114 functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 fuzzy article gathering . . . . . . . . . . . . . . . . . . . . . . . 48 fuzzy matching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171

G

gateways . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 Gcc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 gcc-self. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 general customization . . . . . . . . . . . . . . . . . . . . . . . 207 global score files . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 gnu.emacs.gnus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 gnus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Gnus Agent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 Gnus Agent expire . . . . . . . . . . . . . . . . . . . . . . . . . . 128 Gnus Unplugged . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 Gnus utility functions . . . . . . . . . . . . . . . . . . . . . . . 210 gnus-activate-all-groups . . . . . . . . . . . . . . . . . . . . . . 30 gnus-activate-foreign-newsgroups . . . . . . . . . . . . . . 19 gnus-activate-level . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 gnus-active . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 gnus-adaptive-file-suffix . . . . . . . . . . . . . . . . . . . . . 141 gnus-adaptive-word-minimum. . . . . . . . . . . . . . . . 142 gnus-adaptive-word-syntax-table . . . . . . . . . . . . . 142 gnus-add-configuration . . . . . . . . . . . . . . . . . . . . . . 159 gnus-add-current-to-buffer-list . . . . . . . . . . . . . . . 210 gnus-add-to-list . . . . . . . . . . . . . . . . . . . . . . . . . . 20, 85 gnus-after-getting-new-news-hook . . . . . . . . . . . . . 30 gnus-agent-add-group . . . . . . . . . . . . . . . . . . . . . . . 128 gnus-agent-add-server . . . . . . . . . . . . . . . . . . . . . . . 128 gnus-agent-batch-fetch . . . . . . . . . . . . . . . . . . . . . . 127 gnus-agent-catchup . . . . . . . . . . . . . . . . . . . . . . . . . 128 gnus-agent-directory . . . . . . . . . . . . . . . . . . . . . . . . 129 gnus-agent-expire . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 gnus-agent-expire-all . . . . . . . . . . . . . . . . . . . . . . . . 128 gnus-agent-expire-days . . . . . . . . . . . . . . . . . . . . . . 128 gnus-agent-fetch-groups . . . . . . . . . . . . . . . . . . . . . 127 gnus-agent-fetch-session . . . . . . . . . . . . . . . . . . . . . 128 gnus-agent-handle-level. . . . . . . . . . . . . . . . . . . . . . 129 gnus-agent-high-score . . . . . . . . . . . . . . . . . . . . . . . 127 gnus-agent-long-article . . . . . . . . . . . . . . . . . . . . . . 127 gnus-agent-low-score . . . . . . . . . . . . . . . . . . . . . . . . 127 gnus-agent-mark-article . . . . . . . . . . . . . . . . . . . . . 128 gnus-agent-plugged-hook . . . . . . . . . . . . . . . . . . . . 129 gnus-agent-remove-server . . . . . . . . . . . . . . . . . . . . 128 gnus-agent-short-article . . . . . . . . . . . . . . . . . . . . . 127 gnus-agent-toggle-mark. . . . . . . . . . . . . . . . . . . . . . 128 gnus-agent-unmark-article . . . . . . . . . . . . . . . . . . . 128 gnus-agent-unplugged-hook . . . . . . . . . . . . . . . . . . 129 gnus-alter-header-function . . . . . . . . . . . . . . . . . . . . 50 gnus-always-force-window-configuration. . . . . . . 160 gnus-always-read-dribble-file . . . . . . . . . . . . . . . . . . . 9 gnus-ancient-mark . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 gnus-apply-kill-file . . . . . . . . . . . . . . . . . . . . . . . . . . 147 gnus-apply-kill-file-unless-scored . . . . . . . . . . . . . 147 gnus-apply-kill-hook . . . . . . . . . . . . . . . . . . . . . . . . 147 gnus-article-add-buttons. . . . . . . . . . . . . . . . . . . . . . 67 gnus-article-add-buttons-to-head . . . . . . . . . . . . . . 67 gnus-article-button-face . . . . . . . . . . . . . . . . . . . . . . 68 gnus-article-date-iso8601 . . . . . . . . . . . . . . . . . . . . . 69 gnus-article-date-lapsed . . . . . . . . . . . . . . . . . . . . . . 69 gnus-article-date-local . . . . . . . . . . . . . . . . . . . . . . . . 69 gnus-article-date-original . . . . . . . . . . . . . . . . . . . . . 69 gnus-article-date-user . . . . . . . . . . . . . . . . . . . . . . . . 69 gnus-article-date-ut . . . . . . . . . . . . . . . . . . . . . . . . . . 69 gnus-article-de-quoted-unreadable . . . . . . . . . . . . . 67 gnus-article-describe-briefly . . . . . . . . . . . . . . . . . . . 83 gnus-article-display-hook . . . . . . . . . . 81, 83, 84, 168 gnus-article-display-picons . . . . . . . . . . . . . . . . . . . 169 gnus-article-display-x-face . . . . . . . . . . . . . . . . . . . . 67 gnus-article-emphasize . . . . . . . . . . . . . . . . . . . . . . . 64 gnus-article-fill-cited-article . . . . . . . . . . . . . . . . . . . 67 gnus-article-hide . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 gnus-article-hide-boring-headers . . . . . . . . . . . 65, 81 gnus-article-hide-citation . . . . . . . . . . . . . . . . . . . . . 65 gnus-article-hide-citation-in-followups . . . . . . . . . 66 gnus-article-hide-citation-maybe . . . . . . . . . . . . . . 66 gnus-article-hide-headers . . . . . . . . . . . . . . . . . . . . . 65 gnus-article-hide-pem . . . . . . . . . . . . . . . . . . . . . . . . 65 gnus-article-hide-pgp . . . . . . . . . . . . . . . . . . . . . . . . . 65 gnus-article-hide-pgp-hook. . . . . . . . . . . . . . . . . . . . 65 gnus-article-hide-signature . . . . . . . . . . . . . . . . . . . . 65 gnus-article-highlight. . . . . . . . . . . . . . . . . . . . . . . . . 63 gnus-article-highlight-citation . . . . . . . . . . . . . . . . . 63 gnus-article-highlight-headers . . . . . . . . . . . . . . . . . 63 gnus-article-highlight-signature . . . . . . . . . . . . . . . 64 gnus-article-mail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 gnus-article-maybe-hide-headers. . . . . . . . . . . . . . . 83 gnus-article-maybe-highlight. . . . . . . . . . . . . . . 63, 83 gnus-article-menu-hook. . . . . . . . . . . . . . . . . . . . . . 162 gnus-article-mode-hook. . . . . . . . . . . . . . . . . . . . . . . 84 gnus-article-mode-line-format . . . . . . . . . . . . . . . . . 84 gnus-article-mode-syntax-table . . . . . . . . . . . . . . . . 84 gnus-article-mouse-face . . . . . . . . . . . . . . . . . . . . . . . 68 gnus-article-next-button . . . . . . . . . . . . . . . . . . . . . . 83 gnus-article-next-page . . . . . . . . . . . . . . . . . . . . . . . . 83 gnus-article-prepare-hook . . . . . . . . . . . . . . . . . . . . . 84 gnus-article-prev-button . . . . . . . . . . . . . . . . . . . . . . 83 gnus-article-prev-page . . . . . . . . . . . . . . . . . . . . . . . . 83 gnus-article-refer-article . . . . . . . . . . . . . . . . . . . . . . 83 gnus-article-remove-cr . . . . . . . . . . . . . . . . . . . . . . . . 67 gnus-article-remove-trailing-blank-lines . . . . . . . . 67 gnus-article-save-directory . . . . . . . . . . . . . . . . . . . . 57 gnus-article-show-summary . . . . . . . . . . . . . . . . . . . 83 gnus-article-sort-by-author. . . . . . . . . . . . . . . . . . . . 52 gnus-article-sort-by-date. . . . . . . . . . . . . . . . . . . . . . 52 gnus-article-sort-by-number. . . . . . . . . . . . . . . . . . . 52 gnus-article-sort-by-score . . . . . . . . . . . . . . . . . . . . . 52 gnus-article-sort-by-subject . . . . . . . . . . . . . . . . . . . 52 gnus-article-sort-functions . . . . . . . . . . . . . . . . . . . . 52 gnus-article-strip-all-blank-lines . . . . . . . . . . . . . . . 67 gnus-article-strip-blank-lines . . . . . . . . . . . . . . . . . . 67 gnus-article-strip-leading-blank-lines. . . . . . . . . . . 67 gnus-article-strip-leading-space. . . . . . . . . . . . . . . . 67 gnus-article-strip-multiple-blank-lines. . . . . . . . . . 67 gnus-article-time-format . . . . . . . . . . . . . . . . . . . . . . 69 gnus-article-treat-dumbquotes . . . . . . . . . . . . . . . . 67 gnus-article-treat-overstrike . . . . . . . . . . . . . . . . . . . 67 gnus-article-x-face-command . . . . . . . . . . . . . . . . . . 67 gnus-article-x-face-too-ugly . . . . . . . . . . . . . . . . . . . 67 gnus-async-prefetch-article-p . . . . . . . . . . . . . . . . . . 53 gnus-async-read-p. . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 gnus-asynchronous . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 gnus-auto-center-summary. . . . . . . . . . . . . . . . . . . . 37 gnus-auto-expirable-newsgroups . . . . . . . . . . . . . . 107 gnus-auto-extend-newsgroup . . . . . . . . . . . . . . . . . . 38 gnus-auto-select-first . . . . . . . . . . . . . . . . . . . . . . . . . 15 gnus-auto-select-next. . . . . . . . . . . . . . . . . . . . . . . . . 36 gnus-auto-select-same . . . . . . . . . . . . . . . . . . . . . . . . 37 gnus-auto-subscribed-groups . . . . . . . . . . . . . . . . . . . 7 gnus-batch-score . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 gnus-before-startup-hook . . . . . . . . . . . . . . . . . . . . . 10 gnus-binary-mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 gnus-binary-mode-hook . . . . . . . . . . . . . . . . . . . . . . 73 gnus-binary-show-article. . . . . . . . . . . . . . . . . . . . . . 73 gnus-boring-article-headers . . . . . . . . . . . . . . . . . . . 81 gnus-break-pages . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 gnus-browse-describe-briefly . . . . . . . . . . . . . . . . . . 25 gnus-browse-exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 gnus-browse-menu-hook . . . . . . . . . . . . . . . . . . . . . 162 gnus-browse-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 gnus-browse-read-group . . . . . . . . . . . . . . . . . . . . . . 25 gnus-browse-select-group . . . . . . . . . . . . . . . . . . . . . 25 gnus-browse-unsubscribe-current-group . . . . . . . . 25 gnus-buffer-configuration . . . . . . . . . . . . . . . . . . . . 157 gnus-bug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178, 209 gnus-bug-create-help-buffer . . . . . . . . . . . . . . . . . . 178 gnus-build-sparse-threads . . . . . . . . . . . . . . . . . . . . . 49 gnus-button-alist . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 gnus-button-url-regexp . . . . . . . . . . . . . . . . . . . . . . . 68 gnus-cache-active-file . . . . . . . . . . . . . . . . . . . . . . . . . 55 gnus-cache-directory . . . . . . . . . . . . . . . . . . . . . . . . . 54 gnus-cache-enter-article . . . . . . . . . . . . . . . . . . . . . . 55 gnus-cache-enter-articles. . . . . . . . . . . . . . . . . . . . . . 54 gnus-cache-generate-active . . . . . . . . . . . . . . . . . . . . 55 gnus-cache-generate-nov-databases . . . . . . . . . . . . 55 gnus-cache-remove-article. . . . . . . . . . . . . . . . . . . . . 55 gnus-cache-remove-articles . . . . . . . . . . . . . . . . . . . . 54 gnus-cacheable-groups . . . . . . . . . . . . . . . . . . . . . . . . 54 gnus-cached-mark. . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 gnus-canceled-mark . . . . . . . . . . . . . . . . . . . . . . . . . . 42 gnus-carpal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 gnus-carpal-browse-buffer-buttons . . . . . . . . . . . . 163 gnus-carpal-button-face . . . . . . . . . . . . . . . . . . . . . 162 gnus-carpal-group-buffer-buttons . . . . . . . . . . . . . 162 gnus-carpal-header-face . . . . . . . . . . . . . . . . . . . . . 162 gnus-carpal-mode-hook . . . . . . . . . . . . . . . . . . . . . . 162 gnus-carpal-server-buffer-buttons. . . . . . . . . . . . . 163 gnus-carpal-summary-buffer-buttons. . . . . . . . . . 162 gnus-catchup-mark. . . . . . . . . . . . . . . . . . . . . . . . . . . 42 gnus-category-add . . . . . . . . . . . . . . . . . . . . . . . . . . 126 gnus-category-copy . . . . . . . . . . . . . . . . . . . . . . . . . 126 gnus-category-edit-groups . . . . . . . . . . . . . . . . . . . 127 gnus-category-edit-predicate . . . . . . . . . . . . . . . . . 127 gnus-category-edit-score . . . . . . . . . . . . . . . . . . . . . 127 gnus-category-exit . . . . . . . . . . . . . . . . . . . . . . . . . . 126 gnus-category-kill . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 gnus-category-line-format. . . . . . . . . . . . . . . . . . . . 127 gnus-category-list . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 gnus-category-mode-hook. . . . . . . . . . . . . . . . . . . . 127 gnus-category-mode-line-format . . . . . . . . . . . . . . 127 gnus-change-server . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 gnus-check-backend-function . . . . . . . . . . . . . . . . . 211 gnus-check-bogus-newsgroups . . . . . . . . . . . . . . . . . 10 gnus-check-new-newsgroups . . . . . . . . . . . . . . . . . . . . 5 gnus-cite-attribution-face . . . . . . . . . . . . . . . . . . . . . 64 gnus-cite-attribution-prefix . . . . . . . . . . . . . . . . . . . 64 gnus-cite-attribution-suffix. . . . . . . . . . . . . . . . . . . . 64 gnus-cite-face-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 gnus-cite-hide-absolute . . . . . . . . . . . . . . . . . . . . . . . 66 gnus-cite-hide-percentage . . . . . . . . . . . . . . . . . . . . . 66 gnus-cite-max-prefix . . . . . . . . . . . . . . . . . . . . . . . . . 64 gnus-cite-minimum-match-count . . . . . . . . . . . . . . 64 gnus-cite-parse-max-size . . . . . . . . . . . . . . . . . . . . . . 63 gnus-cite-prefix-regexp . . . . . . . . . . . . . . . . . . . . . . . 63 gnus-cited-closed-text-button-line-format . . . . . . 65 gnus-cited-lines-visible . . . . . . . . . . . . . . . . . . . . . . . 65 gnus-cited-opened-text-button-line-format . . . . . 65 gnus-compile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 gnus-configure-frame . . . . . . . . . . . . . . . . . . . . . . . . 158 gnus-continuum-version . . . . . . . . . . . . . . . . . . . . . 210 gnus-crosspost-complaint . . . . . . . . . . . . . . . . . . . . . 40 gnus-current-home-score-file . . . . . . . . . . . . . . . . . 143 gnus-current-prefix-symbol . . . . . . . . . . . . . . . . . . 226 gnus-current-prefix-symbols. . . . . . . . . . . . . . . . . . 226 gnus-dead-summary-mode . . . . . . . . . . . . . . . . . . . . 79 gnus-decay-score . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 gnus-decay-score-function . . . . . . . . . . . . . . . . . . . 152 gnus-decay-scores . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 gnus-default-adaptive-score-alist . . . . . . . . . . . . . 140 gnus-default-adaptive-word-score-alist . . . . . . . . 142 gnus-default-article-saver . . . . . . . . . . . . . . . . . . . . . 56 gnus-default-directory . . . . . . . . . . . . . . . . . . . . . . . 173 gnus-default-ignored-adaptive-words. . . . . . . . . . 142 gnus-default-subscribed-newsgroups . . . . . . . . . . . . 4 gnus-del-mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 gnus-demon-add-disconnection . . . . . . . . . . . . . . . 163 gnus-demon-add-handler . . . . . . . . . . . . . . . . . . . . 163 gnus-demon-add-nocem . . . . . . . . . . . . . . . . . . . . . 163 gnus-demon-add-rescan. . . . . . . . . . . . . . . . . . . . . . 163 gnus-demon-add-scan-timestamps . . . . . . . . . . . . 163 gnus-demon-add-scanmail . . . . . . . . . . . . . . . . . . . 163 gnus-demon-cancel . . . . . . . . . . . . . . . . . . . . . . . . . . 163 gnus-demon-handlers. . . . . . . . . . . . . . . . . . . . . . . . 163 gnus-demon-init . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 gnus-demon-timestep. . . . . . . . . . . . . . . . . . . . . . . . 163 gnus-directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 gnus-dormant-mark . . . . . . . . . . . . . . . . . . . . . . . . . . 42 gnus-draft-edit-message . . . . . . . . . . . . . . . . . . . . . . 90 gnus-draft-send-all-messages . . . . . . . . . . . . . . . . . . 90 gnus-draft-send-message . . . . . . . . . . . . . . . . . . . . . . 90 gnus-dribble-directory . . . . . . . . . . . . . . . . . . . . . . . . . 9 gnus-duplicate-file . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 gnus-duplicate-list-length . . . . . . . . . . . . . . . . . . . . . 80 gnus-duplicate-mark . . . . . . . . . . . . . . . . . . . . . . . . . 42 gnus-emphasis-alist . . . . . . . . . . . . . . . . . . . . . . . . . . 64 gnus-emphasis-bold . . . . . . . . . . . . . . . . . . . . . . . . . . 64 gnus-emphasis-bold-italic . . . . . . . . . . . . . . . . . . . . . 64 gnus-emphasis-italic. . . . . . . . . . . . . . . . . . . . . . . . . . 64 gnus-emphasis-underline. . . . . . . . . . . . . . . . . . . . . . 64 gnus-emphasis-underline-bold . . . . . . . . . . . . . . . . . 64 gnus-emphasis-underline-bold-italic. . . . . . . . . . . . 64 gnus-emphasis-underline-italic . . . . . . . . . . . . . . . . 64 gnus-empty-thread-mark . . . . . . . . . . . . . . . . . . . . . 43 gnus-enter-category-buffer . . . . . . . . . . . . . . . . . . . 127 gnus-ephemeral-group-p . . . . . . . . . . . . . . . . . . . . . 211 gnus-exit-gnus-hook . . . . . . . . . . . . . . . . . . . . . . . . . . 25 gnus-exit-group-hook. . . . . . . . . . . . . . . . . . . . . . . . . 79 gnus-expert-user . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 gnus-expirable-mark . . . . . . . . . . . . . . . . . . . . . . . . . 42 gnus-extract-address-components. . . . . . . . . . . . . . 33 gnus-fetch-group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 gnus-fetch-old-headers. . . . . . . . . . . . . . . . . . . . . . . . 49 gnus-file-save-name . . . . . . . . . . . . . . . . . . . . . . . . . . 57 gnus-find-method-for-group . . . . . . . . . . . . . . . . . . 210 gnus-folder-save-name . . . . . . . . . . . . . . . . . . . . . . . . 57 gnus-Folder-save-name . . . . . . . . . . . . . . . . . . . . . . . 57 gnus-gather-threads-by-references . . . . . . . . . . . . . 49 gnus-gather-threads-by-subject. . . . . . . . . . . . . . . . 49 gnus-generate-horizontal-tree . . . . . . . . . . . . . . . . . 74 gnus-generate-tree-function . . . . . . . . . . . . . . . . . . . 74 gnus-generate-vertical-tree . . . . . . . . . . . . . . . . . . . . 74 gnus-get-info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 gnus-get-new-news-hook . . . . . . . . . . . . . . . . . . . . . . 30 gnus-global-score-files . . . . . . . . . . . . . . . . . . . . . . . 145 gnus-goto-next-group-when-activating . . . . . . . . . 30 gnus-group-add-to-virtual . . . . . . . . . . . . . . . . . . . . 19 gnus-group-apropos . . . . . . . . . . . . . . . . . . . . . . . . . . 22 gnus-group-archive-directory . . . . . . . . . . . . . . . . . . 18 gnus-group-best-unread-group . . . . . . . . . . . . . . . . 14 gnus-group-brew-soup . . . . . . . . . . . . . . . . . . . . . . . 118 gnus-group-browse-foreign-server. . . . . . . . . . . . 3, 24 gnus-group-catchup-current . . . . . . . . . . . . . . . . . . . 16 gnus-group-catchup-current-all . . . . . . . . . . . . . . . . 16 gnus-group-catchup-group-hook . . . . . . . . . . . . . . . 16 gnus-group-check-bogus-groups . . . . . . . . . . . . . . . 24 gnus-group-clear-data . . . . . . . . . . . . . . . . . . . . . . . . 16 gnus-group-clear-data-on-native-groups. . . . . . 7, 16 gnus-group-customize . . . . . . . . . . . . . . . . . . . . . . . . 18 gnus-group-default-list-level. . . . . . . . . . . . . . . . . . . 17 gnus-group-delete-group . . . . . . . . . . . . . . . . . . . . . . 19 gnus-group-describe-all-groups . . . . . . . . . . . . . . . . 31 gnus-group-describe-briefly . . . . . . . . . . . . . . . . . . . 31 gnus-group-describe-group . . . . . . . . . . . . . . . . . . . . 30 gnus-group-description-apropos . . . . . . . . . . . . . . . 23 gnus-group-edit-global-kill . . . . . . . . . . . . . . . . . . . 146 gnus-group-edit-group . . . . . . . . . . . . . . . . . . . . . . . . 18 gnus-group-edit-group-method . . . . . . . . . . . . . . . . 18 gnus-group-edit-group-parameters . . . . . . . . . . . . . 18 gnus-group-edit-local-kill . . . . . . . . . . . . . . . . . . . . 146 gnus-group-enter-directory. . . . . . . . . . . . . . . . . . . . 19 gnus-group-enter-server-mode . . . . . . . . . . . . . . . . . 29 gnus-group-exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 gnus-group-expire-all-groups . . . . . . . . . . . . . . . . . . 24 gnus-group-expire-articles . . . . . . . . . . . . . . . . . . . . 24 gnus-group-faq-directory . . . . . . . . . . . . . . . . . . 30, 77 gnus-group-fetch-faq . . . . . . . . . . . . . . . . . . . . . . . . . 30 gnus-group-find-new-groups. . . . . . . . . . . . . . . . . . . 24 gnus-group-first-unread-group. . . . . . . . . . . . . . . . . 14 gnus-group-foreign-p . . . . . . . . . . . . . . . . . . . . . . . . 211 gnus-group-get-new-news . . . . . . . . . . . . . . . . . . . . . 30 gnus-group-get-new-news-this-group . . . . . . . . . . . 30 gnus-group-goto-unread . . . . . . . . . . . . . . . . . . . . . . 14 gnus-group-highlight . . . . . . . . . . . . . . . . . . . . . . . . . 12 gnus-group-highlight-line . . . . . . . . . . . . . . . . . . . . . 13 gnus-group-jump-to-group . . . . . . . . . . . . . . . . . . . . 14 gnus-group-kill-all-zombies. . . . . . . . . . . . . . . . . . . . 15 gnus-group-kill-group . . . . . . . . . . . . . . . . . . . . . . . . 15 gnus-group-kill-level. . . . . . . . . . . . . . . . . . . . . . . . . . 16 gnus-group-kill-region . . . . . . . . . . . . . . . . . . . . . . . . 15 gnus-group-line-format . . . . . . . . . . . . . . . . . . . . . . . 11 gnus-group-list-active . . . . . . . . . . . . . . . . . . . . . . . . 22 gnus-group-list-all-groups . . . . . . . . . . . . . . . . . . . . . 22 gnus-group-list-all-matching . . . . . . . . . . . . . . . . . . 22 gnus-group-list-groups. . . . . . . . . . . . . . . . . . . . . . . . 22 gnus-group-list-inactive-groups . . . . . . . . . . . . . . . . 17 gnus-group-list-killed . . . . . . . . . . . . . . . . . . . . . . . . . 22 gnus-group-list-level. . . . . . . . . . . . . . . . . . . . . . . . . . 22 gnus-group-list-matching . . . . . . . . . . . . . . . . . . . . . 22 gnus-group-list-zombies. . . . . . . . . . . . . . . . . . . . . . . 22 gnus-group-mail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 gnus-group-make-archive-group . . . . . . . . . . . . . . . 18 gnus-group-make-directory-group . . . . . . . . . . . . . 18 gnus-group-make-doc-group. . . . . . . . . . . . . . . . . . . 19 gnus-group-make-empty-virtual . . . . . . . . . . . . . . . 19 gnus-group-make-group . . . . . . . . . . . . . . . . . . . . . . 18 gnus-group-make-help-group . . . . . . . . . . . . . . . . . . 18 gnus-group-make-kiboze-group . . . . . . . . . . . . . . . . 19 gnus-group-make-useful-group . . . . . . . . . . . . . . . . 19 gnus-group-make-web-group . . . . . . . . . . . . . . . . . . 19 gnus-group-mark-buffer . . . . . . . . . . . . . . . . . . . . . . 18 gnus-group-mark-group. . . . . . . . . . . . . . . . . . . . . . . 18 gnus-group-mark-regexp . . . . . . . . . . . . . . . . . . . . . . 18 gnus-group-mark-region . . . . . . . . . . . . . . . . . . . . . . 18 gnus-group-menu-hook . . . . . . . . . . . . . . . . . . . . . . 162 gnus-group-mode-hook . . . . . . . . . . . . . . . . . . . . . . . 30 gnus-group-mode-line-format . . . . . . . . . . . . . . . . . 12 gnus-group-move-group-to-server . . . . . . . . . . . . . . . 7 gnus-group-native-p . . . . . . . . . . . . . . . . . . . . . . . . . 211 gnus-group-next-group . . . . . . . . . . . . . . . . . . . . 14, 25 gnus-group-next-unread-group . . . . . . . . . . . . . . . . 13 gnus-group-next-unread-group-same-level . . . . . . 14 gnus-group-post-news . . . . . . . . . . . . . . . . . . . . . . . . 29 gnus-group-prefixed-name . . . . . . . . . . . . . . . . . . . 210 gnus-group-prepare-hook . . . . . . . . . . . . . . . . . . . . . 30 gnus-group-prev-group . . . . . . . . . . . . . . . . . . . . 14, 25 gnus-group-prev-unread-group . . . . . . . . . . . . . . . . 14 gnus-group-prev-unread-group-same-level . . . . . . 14 gnus-group-quick-select-group. . . . . . . . . . . . . . . . . 14 gnus-group-quit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 gnus-group-read-group . . . . . . . . . . . . . . . . . . . . . . . 14 gnus-group-read-init-file . . . . . . . . . . . . . . . . . . . . . . 31 gnus-group-read-only-p . . . . . . . . . . . . . . . . . . . . . . 210 gnus-group-real-name . . . . . . . . . . . . . . . . . . . . . . . 210 gnus-group-recent-archive-directory . . . . . . . . . . . 18 gnus-group-rename-group. . . . . . . . . . . . . . . . . . . . . 18 gnus-group-restart . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 gnus-group-save-newsrc. . . . . . . . . . . . . . . . . . . . . . . 31 gnus-group-secondary-p . . . . . . . . . . . . . . . . . . . . . 211 gnus-group-select-group . . . . . . . . . . . . . . . . . . . . . . 14 gnus-group-select-group-ephemerally . . . . . . . . . . 15 gnus-group-send-drafts . . . . . . . . . . . . . . . . . . . . . . 128 gnus-group-set-current-level. . . . . . . . . . . . . . . . . . . 16 gnus-group-set-parameter. . . . . . . . . . . . . . . . . . . . 211 gnus-group-sort-by-alphabet . . . . . . . . . . . . . . . . . . 23 gnus-group-sort-by-level . . . . . . . . . . . . . . . . . . . . . . 23 gnus-group-sort-by-method . . . . . . . . . . . . . . . . . . . 23 gnus-group-sort-by-rank . . . . . . . . . . . . . . . . . . . . . . 23 gnus-group-sort-by-real-name . . . . . . . . . . . . . . . . . 23 gnus-group-sort-by-score. . . . . . . . . . . . . . . . . . . . . . 23 gnus-group-sort-by-unread . . . . . . . . . . . . . . . . . . . . 23 gnus-group-sort-function . . . . . . . . . . . . . . . . . . . . . 23 gnus-group-sort-groups . . . . . . . . . . . . . . . . . . . . . . . 23 gnus-group-sort-groups-by-alphabet . . . . . . . . . . . 23 gnus-group-sort-groups-by-level . . . . . . . . . . . . . . . 23 gnus-group-sort-groups-by-method . . . . . . . . . . . . 24 gnus-group-sort-groups-by-rank . . . . . . . . . . . . . . . 24 gnus-group-sort-groups-by-score . . . . . . . . . . . . . . . 23 gnus-group-sort-groups-by-unread . . . . . . . . . . . . . 23 gnus-group-sort-selected-groups-by-alphabet … 24 gnus-group-sort-selected-groups-by-level . . . . . . . 24 gnus-group-sort-selected-groups-by-method …. 24 gnus-group-sort-selected-groups-by-rank . . . . . . . 24 gnus-group-sort-selected-groups-by-score . . . . . . . 24 gnus-group-sort-selected-groups-by-unread . . . . . 24 gnus-group-suspend . . . . . . . . . . . . . . . . . . . . . . . . . . 25 gnus-group-toolbar. . . . . . . . . . . . . . . . . . . . . . . . . . 171 gnus-group-transpose-groups . . . . . . . . . . . . . . . . . . 15 gnus-group-uncollapsed-levels . . . . . . . . . . . . . . . . . 12 gnus-group-universal-argument. . . . . . . . . . . . . . . . 18 gnus-group-unmark-all-groups . . . . . . . . . . . . . . . . 18 gnus-group-unmark-group . . . . . . . . . . . . . . . . . . . . 18 gnus-group-unread . . . . . . . . . . . . . . . . . . . . . . . . . . 210 gnus-group-unsubscribe-current-group . . . . . . . . . 15 gnus-group-unsubscribe-group . . . . . . . . . . . . . . . . 15 gnus-group-update-hook . . . . . . . . . . . . . . . . . . . . . . 13 gnus-group-use-permanent-levels . . . . . . . . . . . . . . 17 gnus-group-visible-select-group . . . . . . . . . . . . . . . . 15 gnus-group-yank-group . . . . . . . . . . . . . . . . . . . . . . . 15 gnus-grouplens-override-scoring . . . . . . . . . . . . . . 148 gnus-header-button-alist. . . . . . . . . . . . . . . . . . . . . . 68 gnus-header-face-alist . . . . . . . . . . . . . . . . . . . . . . . . 63 gnus-hidden-properties . . . . . . . . . . . . . . . . . . . . . . 174 gnus-hierarchial-home-score-file . . . . . . . . . . . . . . 143 gnus-home-adapt-file . . . . . . . . . . . . . . . . . . . . . . . . 143 gnus-home-score-file. . . . . . . . . . . . . . . . . . . . . . . . . 142 gnus-ignored-adaptive-words . . . . . . . . . . . . . . . . . 142 gnus-ignored-headers . . . . . . . . . . . . . . . . . . . . . . . . . 81 gnus-ignored-newsgroups . . . . . . . . . . . . . . . . . . . . . . 9 gnus-info-find-node . . . . . . . . . . . . . . . . . . . . . . . 31, 77 gnus-info-group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 gnus-info-level . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 gnus-info-marks . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 gnus-info-method . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 gnus-info-params . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 gnus-info-rank . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 gnus-info-read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 gnus-info-score . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 gnus-info-set-group . . . . . . . . . . . . . . . . . . . . . . . . . 225 gnus-info-set-level. . . . . . . . . . . . . . . . . . . . . . . . . . . 225 gnus-info-set-marks . . . . . . . . . . . . . . . . . . . . . . . . . 225 gnus-info-set-method . . . . . . . . . . . . . . . . . . . . . . . . 225 gnus-info-set-params . . . . . . . . . . . . . . . . . . . . . . . . 225 gnus-info-set-rank. . . . . . . . . . . . . . . . . . . . . . . . . . . 225 gnus-info-set-read . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 gnus-info-set-score . . . . . . . . . . . . . . . . . . . . . . . . . . 225 gnus-inhibit-startup-message. . . . . . . . . . . . . . . . . . 10 gnus-init-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8, 31 gnus-insert-pseudo-articles . . . . . . . . . . . . . . . . . . . . 63 gnus-interactive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 gnus-interactive-catchup. . . . . . . . . . . . . . . . . . . . . 153 gnus-interactive-exit . . . . . . . . . . . . . . . . . . . . . . . . 153 gnus-jog-cache . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 gnus-keep-backlog. . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 gnus-keep-same-level . . . . . . . . . . . . . . . . . . . . . . . . . 17 gnus-kill-file-mark. . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 gnus-kill-file-mode-hook . . . . . . . . . . . . . . . . . . . . . 147 gnus-kill-file-name . . . . . . . . . . . . . . . . . . . . . . . . . . 146 gnus-kill-files-directory . . . . . . . . . . . . . . . . . . . . . . 134 gnus-kill-killed. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 gnus-kill-save-kill-file . . . . . . . . . . . . . . . . . . . . . . . . 146 gnus-kill-summary-on-exit . . . . . . . . . . . . . . . . . . . . 79 gnus-killed-mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 gnus-large-newsgroup . . . . . . . . . . . . . . . . . . . . . . . . 15 gnus-level-default-subscribed. . . . . . . . . . . . . . . . . . 16 gnus-level-default-unsubscribed . . . . . . . . . . . . . . . 16 gnus-level-killed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 gnus-level-subscribed . . . . . . . . . . . . . . . . . . . . . . . . . 16 gnus-level-unsubscribed . . . . . . . . . . . . . . . . . . . . . . 16 gnus-level-zombie . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 gnus-list-groups-with-ticked-articles . . . . . . . . . . . 23 gnus-load-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 gnus-low-score-mark . . . . . . . . . . . . . . . . . . . . . . . . . 42 gnus-mail-save-name . . . . . . . . . . . . . . . . . . . . . . . . . 57 gnus-mailing-list-groups . . . . . . . . . . . . . . . . . . . . . . 86 gnus-mark-article-hook . . . . . . . . . . . . . . . . . . 38, 107 gnus-mark-unpicked-articles-as-read . . . . . . . . . . . 72 gnus-message-archive-group. . . . . . . . . . . . . . . . . . . 87 gnus-message-archive-method . . . . . . . . . . . . . . . . . 86 gnus-mode-non-string-length. . . . . . . . . . . . . . . . . 160 gnus-mouse-face . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 gnus-move-split-methods . . . . . . . . . . . . . . . . . . . . . 76 gnus-narrow-to-body . . . . . . . . . . . . . . . . . . . . . . . . 211 gnus-new-mail-mark . . . . . . . . . . . . . . . . . . . . . . . . . 12 gnus-news-group-p . . . . . . . . . . . . . . . . . . . . . . . . . . 210 gnus-newsgroup-name . . . . . . . . . . . . . . . . . . . . . . . 210 gnus-nntp-server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 gnus-nntpserver-file . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 gnus-no-groups-message . . . . . . . . . . . . . . . . . . . . . . 10 gnus-no-server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 gnus-nocem-directory . . . . . . . . . . . . . . . . . . . . . . . 165 gnus-nocem-expiry-wait . . . . . . . . . . . . . . . . . . . . . 166 gnus-nocem-groups . . . . . . . . . . . . . . . . . . . . . . . . . 164 gnus-nocem-issuers. . . . . . . . . . . . . . . . . . . . . . . . . . 164 gnus-nocem-verifyer . . . . . . . . . . . . . . . . . . . . . . . . . 165 gnus-not-empty-thread-mark. . . . . . . . . . . . . . . . . . 43 gnus-nov-is-evil . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 gnus-novice-user . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 gnus-numeric-save-name . . . . . . . . . . . . . . . . . . . . . . 57 gnus-Numeric-save-name . . . . . . . . . . . . . . . . . . . . . 57 gnus-options-not-subscribe. . . . . . . . . . . . . . . . . . . . . 7 gnus-options-subscribe . . . . . . . . . . . . . . . . . . . . . . . . 7 gnus-other-frame. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 gnus-outgoing-message-group . . . . . . . . . . . . . . . . . 88 gnus-page-delimiter . . . . . . . . . . . . . . . . . . . . . . . . . . 84 gnus-parse-headers-hook . . . . . . . . . . . . . . . . . 50, 174 gnus-permanently-visible-groups . . . . . . . . . . . 23, 30 gnus-pick-article-or-thread . . . . . . . . . . . . . . . . . . . . 72 gnus-pick-display-summary . . . . . . . . . . . . . . . . . . . 72 gnus-pick-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 gnus-pick-mode-hook. . . . . . . . . . . . . . . . . . . . . . . . . 72 gnus-pick-next-page . . . . . . . . . . . . . . . . . . . . . . . . . . 72 gnus-pick-start-reading . . . . . . . . . . . . . . . . . . . . . . . 72 gnus-pick-unmark-article-or-thread. . . . . . . . . . . . 72 gnus-picons-clear-cache-on-shutdown . . . . . . . . . 170 gnus-picons-convert-x-face . . . . . . . . . . . . . . 167, 169 gnus-picons-database . . . . . . . . . . . . . . . . . . . 167, 168 gnus-picons-display-article-move-p . . . . . . . . . . . 170 gnus-picons-display-as-address . . . . . . . . . . . . . . . 169 gnus-picons-display-where . . . . . . . . . . . . . . . . . . . 168 gnus-picons-domain-directories . . . . . . . . . . . . . . . 169 gnus-picons-file-suffixes . . . . . . . . . . . . . . . . . . . . . . 170 gnus-picons-group-excluded-groups . . . . . . . . . . . 168 gnus-picons-has-modeline-p . . . . . . . . . . . . . . . . . . 169 gnus-picons-news-directories . . . . . . . . . . . . . . . . . 169 gnus-picons-piconsearch-url . . . . . . . . . . . . . 167, 168 gnus-picons-refresh-before-display . . . . . . . . . . . . 169 gnus-picons-user-directories . . . . . . . . . . . . . . . . . . 169 gnus-picons-x-face-file-name . . . . . . . . . . . . . . . . . 169 gnus-plain-save-name. . . . . . . . . . . . . . . . . . . . . . . . . 57 gnus-Plain-save-name . . . . . . . . . . . . . . . . . . . . . . . . 57 gnus-play-startup-jingle . . . . . . . . . . . . . . . . . . . . . . 10 gnus-post-method. . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 gnus-posting-styles . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 gnus-prefetched-article-deletion-strategy . . . . . . . 54 gnus-process-mark . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 gnus-prompt-before-saving. . . . . . . . . . . . . . . . . . . . 56 gnus-ps-print-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 gnus-read-active-file . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 gnus-read-mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 gnus-read-method. . . . . . . . . . . . . . . . . . . . . . . . . . . 211 gnus-refer-article-method . . . . . . . . . . . . . . . . . . . . . 71 gnus-refer-thread-limit . . . . . . . . . . . . . . . . . . . . . . . 71 gnus-replied-mark. . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 gnus-rmail-save-name . . . . . . . . . . . . . . . . . . . . . . . . 57 gnus-save-all-headers . . . . . . . . . . . . . . . . . . . . . . . . . 56 gnus-save-duplicate-list . . . . . . . . . . . . . . . . . . . . . . . 80 gnus-save-killed-list . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 gnus-save-newsrc-file . . . . . . . . . . . . . . . . . . . . . . . . . . 8 gnus-save-newsrc-hook . . . . . . . . . . . . . . . . . . . . . . . . 8 gnus-save-quick-newsrc-hook . . . . . . . . . . . . . . . . . . . 8 gnus-save-score . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 gnus-save-standard-newsrc-hook. . . . . . . . . . . . . . . . 8 gnus-saved-headers. . . . . . . . . . . . . . . . . . . . . . . . . . . 56 gnus-saved-mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 gnus-score-after-write-file-function. . . . . . . . . . . . 135 gnus-score-below-mark . . . . . . . . . . . . . . . . . . . . . . 134 gnus-score-change-score-file . . . . . . . . . . . . . . . . . . 131 gnus-score-customize . . . . . . . . . . . . . . . . . . . . . . . . 132 gnus-score-decay-constant . . . . . . . . . . . . . . . . . . . 152 gnus-score-decay-scale . . . . . . . . . . . . . . . . . . . . . . . 152 gnus-score-edit-current-scores . . . . . . . . . . . . . . . . 131 gnus-score-edit-done . . . . . . . . . . . . . . . . . . . . . . . . 140 gnus-score-edit-file . . . . . . . . . . . . . . . . . . . . . . . . . . 131 gnus-score-edit-insert-date . . . . . . . . . . . . . . . . . . . 140 gnus-score-exact-adapt-limit . . . . . . . . . . . . . . . . . 141 gnus-score-expiry-days . . . . . . . . . . . . . . . . . . . . . . 135 gnus-score-file-suffix. . . . . . . . . . . . . . . . . . . . . . . . . 134 gnus-score-find-bnews . . . . . . . . . . . . . . . . . . . . . . . 135 gnus-score-find-hierarchical . . . . . . . . . . . . . . . . . . 135 gnus-score-find-score-files-function. . . . . . . . . . . . 134 gnus-score-find-single. . . . . . . . . . . . . . . . . . . . . . . . 134 gnus-score-find-trace . . . . . . . . . . . . . . . . . . . . . . . . 131 gnus-score-flush-cache . . . . . . . . . . . . . . . . . . 131, 133 gnus-score-followup-article . . . . . . . . . . . . . . . . . . . 143 gnus-score-followup-thread . . . . . . . . . . . . . . . . . . 144 gnus-score-interactive-default-score . . . . . . . . . . . 134 gnus-score-menu-hook . . . . . . . . . . . . . . . . . . . . . . . 162 gnus-score-mimic-keymap . . . . . . . . . . . . . . . . . . . 133 gnus-score-mode-hook . . . . . . . . . . . . . . . . . . . . . . . 140 gnus-score-over-mark. . . . . . . . . . . . . . . . . . . . . . . . 134 gnus-score-pretty-print . . . . . . . . . . . . . . . . . . . . . . 140 gnus-score-search-global-directories . . . . . . . . . . . 145 gnus-score-set-expunge-below . . . . . . . . . . . . . . . . 132 gnus-score-set-mark-below . . . . . . . . . . . . . . . . . . . 132 gnus-score-thread-simplify . . . . . . . . . . . . . . . . . . . 135 gnus-score-uncacheable-files. . . . . . . . . . . . . . . . . . 134 gnus-secondary-select-methods . . . . . . . . . . . . . . . . . 3 gnus-secondary-servers . . . . . . . . . . . . . . . . . . . . . . . . 3 gnus-select-article-hook. . . . . . . . . . . . . . . . . . . . . . . 38 gnus-select-group-hook . . . . . . . . . . . . . . . . . . . . . . . 15 gnus-select-method. . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 gnus-selected-tree-face. . . . . . . . . . . . . . . . . . . . . . . . 73 gnus-sent-message-ids-file . . . . . . . . . . . . . . . . . . . . . 85 gnus-sent-message-ids-length . . . . . . . . . . . . . . . . . . 85 gnus-server-add-server. . . . . . . . . . . . . . . . . . . . . . . . 92 gnus-server-close-all-servers . . . . . . . . . . . . . . . . . . . 95 gnus-server-close-server. . . . . . . . . . . . . . . . . . . . . . . 95 gnus-server-copy-server . . . . . . . . . . . . . . . . . . . . . . . 92 gnus-server-deny-server . . . . . . . . . . . . . . . . . . . . . . . 95 gnus-server-edit-server. . . . . . . . . . . . . . . . . . . . . . . . 92 gnus-server-equal . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 gnus-server-exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 gnus-server-kill-server . . . . . . . . . . . . . . . . . . . . . . . . 92 gnus-server-line-format . . . . . . . . . . . . . . . . . . . . . . . 91 gnus-server-list-servers . . . . . . . . . . . . . . . . . . . . . . . 92 gnus-server-menu-hook . . . . . . . . . . . . . . . . . . . . . . 162 gnus-server-mode-hook . . . . . . . . . . . . . . . . . . . . . . . 91 gnus-server-mode-line-format . . . . . . . . . . . . . . . . . 92 gnus-server-open-all-servers . . . . . . . . . . . . . . . . . . . 95 gnus-server-open-server. . . . . . . . . . . . . . . . . . . . . . . 95 gnus-server-read-server . . . . . . . . . . . . . . . . . . . . . . . 92 gnus-server-regenerate-server. . . . . . . . . . . . . . . . . . 92 gnus-server-remove-denials. . . . . . . . . . . . . . . . . . . . 95 gnus-server-scan-server . . . . . . . . . . . . . . . . . . . . . . . 92 gnus-server-to-method. . . . . . . . . . . . . . . . . . . . . . . 211 gnus-server-yank-server. . . . . . . . . . . . . . . . . . . . . . . 92 gnus-set-active. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 gnus-shell-command-separator . . . . . . . . . . . . . . . 174 gnus-show-all-headers . . . . . . . . . . . . . . . . . . . . . . . . 81 gnus-show-mime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 gnus-show-mime-method . . . . . . . . . . . . . . . . . . . . . 82 gnus-show-threads . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 gnus-signature-face. . . . . . . . . . . . . . . . . . . . . . . . . . . 64 gnus-signature-limit . . . . . . . . . . . . . . . . . . . . . . . . . . 69 gnus-signature-separator . . . . . . . . . . . . . . . . . . 64, 69 gnus-simplify-ignored-prefixes . . . . . . . . . . . . . . . . . 48 gnus-simplify-subject-functions. . . . . . . . . . . . . . . . 48 gnus-simplify-subject-fuzzy . . . . . . . . . . . . . . . . . . . 48 gnus-simplify-subject-fuzzy-regexp . . . . . . . . . . . . 48 gnus-simplify-subject-re . . . . . . . . . . . . . . . . . . . . . . 48 gnus-simplify-whitespace . . . . . . . . . . . . . . . . . . . . . 49 gnus-single-article-buffer. . . . . . . . . . . . . . . . . . . . . . 84 gnus-sorted-header-list . . . . . . . . . . . . . . . . . . . . . . . 81 gnus-soup-add-article . . . . . . . . . . . . . . . . . . . . . . . 118 gnus-soup-directory . . . . . . . . . . . . . . . . . . . . . . . . . 118 gnus-soup-pack-packet . . . . . . . . . . . . . . . . . . . . . . 118 gnus-soup-packer . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 gnus-soup-packet-directory . . . . . . . . . . . . . . . . . . 119 gnus-soup-packet-regexp. . . . . . . . . . . . . . . . . . . . . 119 gnus-soup-prefix-file. . . . . . . . . . . . . . . . . . . . . . . . . 119 gnus-soup-replies-directory . . . . . . . . . . . . . . . . . . 118 gnus-soup-save-areas . . . . . . . . . . . . . . . . . . . . . . . . 118 gnus-soup-send-replies. . . . . . . . . . . . . . . . . . . . . . . 118 gnus-soup-unpacker . . . . . . . . . . . . . . . . . . . . . . . . . 119 gnus-souped-mark . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 gnus-sparse-mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 gnus-split-methods . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 gnus-start-date-timer. . . . . . . . . . . . . . . . . . . . . . . . . 69 gnus-started-hook. . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 gnus-startup-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 gnus-startup-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 gnus-startup-jingle . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 gnus-stop-date-timer . . . . . . . . . . . . . . . . . . . . . . . . . 69 gnus-strict-mime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 gnus-subscribe-alphabetically . . . . . . . . . . . . . . . . . . 6 gnus-subscribe-hierarchical-interactive . . . . . . . . . . 6 gnus-subscribe-hierarchically . . . . . . . . . . . . . . . . . . . 6 gnus-subscribe-interactively . . . . . . . . . . . . . . . . . . . . 6 gnus-subscribe-killed . . . . . . . . . . . . . . . . . . . . . . . . . . 6 gnus-subscribe-newsgroup-method. . . . . . . . . . . . . . 6 gnus-subscribe-options-newsgroup-method . . . . . . 7 gnus-subscribe-randomly . . . . . . . . . . . . . . . . . . . . . . 6 gnus-subscribe-zombies . . . . . . . . . . . . . . . . . . . . . . . . 6 gnus-summary-article-posted-p. . . . . . . . . . . . . . . . 76 gnus-summary-beginning-of-article . . . . . . . . . . . . 39 gnus-summary-best-unread-article . . . . . . . . . . . . . 38 gnus-summary-bubble-group . . . . . . . . . . . . . . . . . . 17 gnus-summary-caesar-message . . . . . . . . . . . . . . . . 66 gnus-summary-cancel-article . . . . . . . . . . . . . . . . . . 41 gnus-summary-catchup . . . . . . . . . . . . . . . . . . . . . . . 44 gnus-summary-catchup-all . . . . . . . . . . . . . . . . . . . . 44 gnus-summary-catchup-all-and-exit. . . . . . . . . . . . 78 gnus-summary-catchup-and-exit. . . . . . . . . . . . . . . 78 gnus-summary-catchup-and-goto-next-group … 78 gnus-summary-catchup-to-here . . . . . . . . . . . . . . . . 44 gnus-summary-check-current . . . . . . . . . . . . . . . . . . 37 gnus-summary-clear-above . . . . . . . . . . . . . . . . . . . . 44 gnus-summary-clear-mark-forward . . . . . . . . . . . . 43 gnus-summary-copy-article . . . . . . . . . . . . . . . . . . . 75 gnus-summary-crosspost-article . . . . . . . . . . . . . . . 75 gnus-summary-current-score . . . . . . . . . . . . . . . . . 131 gnus-summary-default-score . . . . . . . . . . . . . . . . . 134 gnus-summary-delete-article . . . . . . . . . . . . . . . . . . 75 gnus-summary-describe-briefly . . . . . . . . . . . . . . . . 77 gnus-summary-describe-group. . . . . . . . . . . . . . . . . 77 gnus-summary-down-thread . . . . . . . . . . . . . . . . . . 51 gnus-summary-dummy-line-format . . . . . . . . . . . . 47 gnus-summary-edit-article . . . . . . . . . . . . . . . . . . . . 75 gnus-summary-edit-global-kill. . . . . . . . . . . . . . . . 146 gnus-summary-edit-local-kill . . . . . . . . . . . . . . . . . 146 gnus-summary-edit-parameters. . . . . . . . . . . . . . . . 78 gnus-summary-end-of-article . . . . . . . . . . . . . . . . . . 39 gnus-summary-enter-digest-group . . . . . . . . . . . . . 77 gnus-summary-execute-command. . . . . . . . . . . . . . 77 gnus-summary-exit. . . . . . . . . . . . . . . . . . . . . . . . . . . 78 gnus-summary-exit-hook . . . . . . . . . . . . . . . . . . . . . 78 gnus-summary-exit-no-update. . . . . . . . . . . . . . . . . 78 gnus-summary-expand-window . . . . . . . . . . . . . . . . 78 gnus-summary-expire-articles . . . . . . . . . . . . . . . . . 75 gnus-summary-expire-articles-now . . . . . . . . . . . . . 75 gnus-summary-expunge-below . . . . . . . . . . . . . . . 134 gnus-summary-fetch-faq . . . . . . . . . . . . . . . . . . . . . . 77 gnus-summary-first-unread-article . . . . . . . . . . . . . 37 gnus-summary-followup . . . . . . . . . . . . . . . . . . . . . . 40 gnus-summary-followup-to-mail . . . . . . . . . . . . . . . 40 gnus-summary-followup-to-mail-with-original… 40 gnus-summary-followup-with-original . . . . . . . . . . 40 gnus-summary-gather-exclude-subject . . . . . . . . . 49 gnus-summary-gather-subject-limit . . . . . . . . . . . . 47 gnus-summary-generate-hook . . . . . . . . . . . . . . . . . 76 gnus-summary-goto-article. . . . . . . . . . . . . . . . . . . . 36 gnus-summary-goto-last-article. . . . . . . . . . . . . . . . 38 gnus-summary-goto-subject . . . . . . . . . . . . . . . . . . . 36 gnus-summary-goto-unread. . . . . . . . . . . . . . . 44, 153 gnus-summary-hide-all-threads . . . . . . . . . . . . . . . . 51 gnus-summary-hide-thread . . . . . . . . . . . . . . . . . . . 51 gnus-summary-highlight . . . . . . . . . . . . . . . . . . . . . . 36 gnus-summary-ignore-duplicates . . . . . . . . . . . . . . 76 gnus-summary-import-article . . . . . . . . . . . . . . . . . 75 gnus-summary-increase-score . . . . . . . . . . . . . . . . 132 gnus-summary-insert-cached-articles . . . . . . . . . . . 77 gnus-summary-isearch-article . . . . . . . . . . . . . . . . . 39 gnus-summary-kill-below . . . . . . . . . . . . . . . . . . . . . 44 gnus-summary-kill-process-mark . . . . . . . . . . . . . . 45 gnus-summary-kill-same-subject. . . . . . . . . . . . . . . 44 gnus-summary-kill-same-subject-and-select. . . . . 44 gnus-summary-kill-thread . . . . . . . . . . . . . . . . . . . . 51 gnus-summary-limit-exclude-childless-dormant

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 gnus-summary-limit-exclude-dormant. . . . . . . . . . 46 gnus-summary-limit-include-cached. . . . . . . . . . . . 46 gnus-summary-limit-include-dormant . . . . . . . . . . 46 gnus-summary-limit-include-expunged . . . . . . . . . 46 gnus-summary-limit-include-thread . . . . . . . . . . . . 46 gnus-summary-limit-mark-excluded-as-read …. 46 gnus-summary-limit-to-age . . . . . . . . . . . . . . . . . . . 46 gnus-summary-limit-to-articles . . . . . . . . . . . . . . . . 46 gnus-summary-limit-to-author . . . . . . . . . . . . . . . . 45 gnus-summary-limit-to-marks . . . . . . . . . . . . . . . . . 45 gnus-summary-limit-to-score . . . . . . . . . . . . . . . . . . 46 gnus-summary-limit-to-subject . . . . . . . . . . . . . . . . 45 gnus-summary-limit-to-unread . . . . . . . . . . . . . . . . 45 gnus-summary-line-format . . . . . . . . . . . . . . . . . . . . 33 gnus-summary-lower-score . . . . . . . . . . . . . . . . . . . 132 gnus-summary-lower-thread. . . . . . . . . . . . . . . . . . . 51 gnus-summary-mail-crosspost-complaint . . . . . . . 40 gnus-summary-mail-forward . . . . . . . . . . . . . . . . . . 39 gnus-summary-mail-other-window . . . . . . . . . . . . . 39 gnus-summary-mail-toolbar. . . . . . . . . . . . . . . . . . 171 gnus-summary-make-false-root . . . . . . . . . . . . . . . . 47 gnus-summary-mark-above . . . . . . . . . . . . . . . . . . . 44 gnus-summary-mark-as-dormant . . . . . . . . . . . . . . 43 gnus-summary-mark-as-expirable. . . . . . . . . . . . . . 44 gnus-summary-mark-as-processable. . . . . . . . . . . . 44 gnus-summary-mark-as-read-backward. . . . . . . . . 43 gnus-summary-mark-as-read-forward . . . . . . . . . . 43 gnus-summary-mark-below . . . . . . . . . . . . . . . . . . 131 gnus-summary-mark-read-and-unread-as-read . . 38 gnus-summary-mark-region-as-read. . . . . . . . . . . . 44 gnus-summary-mark-unread-as-read . . . . . . . . . . . 38 gnus-summary-menu-hook . . . . . . . . . . . . . . . . . . . 162 gnus-summary-mode-hook . . . . . . . . . . . . . . . . . . . . 76 gnus-summary-mode-line-format . . . . . . . . . . . . . . 35 gnus-summary-move-article . . . . . . . . . . . . . . . . . . . 75 gnus-summary-next-article. . . . . . . . . . . . . . . . . . . . 37 gnus-summary-next-group . . . . . . . . . . . . . . . . . . . . 78 gnus-summary-next-page. . . . . . . . . . . . . . . . . . 37, 38 gnus-summary-next-same-subject . . . . . . . . . . . . . 37 gnus-summary-next-thread . . . . . . . . . . . . . . . . . . . 51 gnus-summary-next-unread-article . . . . . . . . . . . . 37 gnus-summary-next-unread-subject. . . . . . . . . . . . 36 gnus-summary-pick-line-format . . . . . . . . . . . . . . . 72 gnus-summary-pipe-output . . . . . . . . . . . . . . . . . . . 56 gnus-summary-pop-article . . . . . . . . . . . . . . . . . . . . 38 gnus-summary-pop-limit. . . . . . . . . . . . . . . . . . . . . . 46 gnus-summary-post-forward . . . . . . . . . . . . . . . . . . 40 gnus-summary-post-news . . . . . . . . . . . . . . . . . . . . . 40 gnus-summary-prepare . . . . . . . . . . . . . . . . . . . . . . . 77 gnus-summary-prepare-exit-hook . . . . . . . . . . . . . . 78 gnus-summary-prepare-hook . . . . . . . . . . . . . . . . . . 76 gnus-summary-prev-article. . . . . . . . . . . . . . . . . . . . 37 gnus-summary-prev-group . . . . . . . . . . . . . . . . . . . . 78 gnus-summary-prev-page . . . . . . . . . . . . . . . . . . . . . 38 gnus-summary-prev-same-subject . . . . . . . . . . . . . 37 gnus-summary-prev-thread . . . . . . . . . . . . . . . . . . . 51 gnus-summary-prev-unread-article . . . . . . . . . . . . 37 gnus-summary-prev-unread-subject. . . . . . . . . . . . 36 gnus-summary-print-article . . . . . . . . . . . . . . . . . . . 70 gnus-summary-raise-thread . . . . . . . . . . . . . . . . . . . 51 gnus-summary-read-document . . . . . . . . . . . . . . . . 77 gnus-summary-refer-article . . . . . . . . . . . . . . . . . . . 71 gnus-summary-refer-parent-article . . . . . . . . . . . . . 71 gnus-summary-refer-references . . . . . . . . . . . . . . . . 71 gnus-summary-refer-thread . . . . . . . . . . . . . . . . . . . 71 gnus-summary-remove-bookmark. . . . . . . . . . . . . . 44 gnus-summary-reparent-thread. . . . . . . . . . . . . . . . 51 gnus-summary-reply . . . . . . . . . . . . . . . . . . . . . . . . . 39 gnus-summary-reply-with-original . . . . . . . . . . . . . 39 gnus-summary-rescan-group . . . . . . . . . . . . . . . . . . 78 gnus-summary-rescore. . . . . . . . . . . . . . . . . . . . . . . 131 gnus-summary-reselect-current-group . . . . . . . . . . 78 gnus-summary-resend-bounced-mail . . . . . . . . . . . 39 gnus-summary-resend-message . . . . . . . . . . . . . . . . 39 gnus-summary-respool-article . . . . . . . . . . . . . . . . . 75 gnus-summary-respool-query. . . . . . . . . . . . . . . . . . 75 gnus-summary-respool-trace . . . . . . . . . . . . . . . . . . 76 gnus-summary-rethread-current . . . . . . . . . . . . . . . 51 gnus-summary-same-subject . . . . . . . . . . . . . . . . . . 33 gnus-summary-save-article . . . . . . . . . . . . . . . . . . . . 56 gnus-summary-save-article-body-file . . . . . . . . . . . 56 gnus-summary-save-article-file . . . . . . . . . . . . . . . . 56 gnus-summary-save-article-folder . . . . . . . . . . . . . . 56 gnus-summary-save-article-mail . . . . . . . . . . . . . . . 56 gnus-summary-save-article-rmail . . . . . . . . . . . . . . 56 gnus-summary-save-article-vm . . . . . . . . . . . . . . . . 56 gnus-summary-save-body-in-file . . . . . . . . . . . . . . . 57 gnus-summary-save-in-file . . . . . . . . . . . . . . . . . . . . 57 gnus-summary-save-in-folder . . . . . . . . . . . . . . . . . . 57 gnus-summary-save-in-mail . . . . . . . . . . . . . . . . . . . 57 gnus-summary-save-in-rmail . . . . . . . . . . . . . . . . . . 57 gnus-summary-save-in-vm . . . . . . . . . . . . . . . . . . . . 57 gnus-summary-save-newsrc . . . . . . . . . . . . . . . . . . . 78 gnus-summary-save-process-mark . . . . . . . . . . . . . 45 gnus-summary-scroll-down. . . . . . . . . . . . . . . . . . . . 38 gnus-summary-scroll-up . . . . . . . . . . . . . . . . . . . . . . 38 gnus-summary-search-article-backward . . . . . . . . 77 gnus-summary-search-article-forward . . . . . . . . . . 77 gnus-summary-select-article-buffer . . . . . . . . . . . . 39 gnus-summary-selected-face. . . . . . . . . . . . . . . . . . . 35 gnus-summary-set-bookmark . . . . . . . . . . . . . . . . . 44 gnus-summary-set-score . . . . . . . . . . . . . . . . . . . . . 131 gnus-summary-show-all-threads . . . . . . . . . . . . . . . 51 gnus-summary-show-article . . . . . . . . . . . . . . . . . . . 38 gnus-summary-show-thread . . . . . . . . . . . . . . . . . . . 51 gnus-summary-sort-by-author . . . . . . . . . . . . . . . . . 70 gnus-summary-sort-by-date . . . . . . . . . . . . . . . . . . . 70 gnus-summary-sort-by-lines . . . . . . . . . . . . . . . . . . . 70 gnus-summary-sort-by-number . . . . . . . . . . . . . . . . 70 gnus-summary-sort-by-score . . . . . . . . . . . . . . . . . . 70 gnus-summary-sort-by-subject . . . . . . . . . . . . . . . . 70 gnus-summary-stop-page-breaking. . . . . . . . . . . . . 66 gnus-summary-supersede-article . . . . . . . . . . . . . . . 41 gnus-summary-thread-gathering-function . . . . . . 49 gnus-summary-tick-above. . . . . . . . . . . . . . . . . . . . . 44 gnus-summary-tick-article-forward . . . . . . . . . . . . 43 gnus-summary-toggle-header . . . . . . . . . . . . . . . . . . 66 gnus-summary-toggle-mime . . . . . . . . . . . . . . . . . . . 66 gnus-summary-toggle-threads . . . . . . . . . . . . . . . . . 51 gnus-summary-toggle-truncation . . . . . . . . . . . . . . 78 gnus-summary-toolbar . . . . . . . . . . . . . . . . . . . . . . 171 gnus-summary-top-thread . . . . . . . . . . . . . . . . . . . . 51 gnus-summary-universal-argument . . . . . . . . . . . . 77 gnus-summary-unmark-all-processable . . . . . . . . . 45 gnus-summary-unmark-as-processable . . . . . . . . . 44 gnus-summary-up-thread . . . . . . . . . . . . . . . . . . . . . 51 gnus-summary-update-hook . . . . . . . . . . . . . . . . . . 35 gnus-summary-verbose-header . . . . . . . . . . . . . . . . 66 gnus-summary-wake-up-the-dead . . . . . . . . . . . . . . 79 gnus-summary-wide-reply . . . . . . . . . . . . . . . . . . . . 39 gnus-summary-wide-reply-with-original . . . . . . . . 39 gnus-summary-write-article-file . . . . . . . . . . . . . . . 56 gnus-summary-yank-process-mark . . . . . . . . . . . . . 45 gnus-summary-zcore-fuzz . . . . . . . . . . . . . . . . . . . . . 34 gnus-supercite-regexp . . . . . . . . . . . . . . . . . . . . . . . . 64 gnus-supercite-secondary-regexp . . . . . . . . . . . . . . 64 gnus-suppress-duplicates. . . . . . . . . . . . . . . . . . . . . . 80 gnus-suspend-gnus-hook . . . . . . . . . . . . . . . . . . . . . . 25 gnus-symbolic-argument . . . . . . . . . . . . . . . . . . . . . 154 gnus-thread-expunge-below . . . . . . . . . . . . . . . . . . . 50 gnus-thread-hide-killed . . . . . . . . . . . . . . . . . . . . . . . 50 gnus-thread-hide-subtree . . . . . . . . . . . . . . . . . . . . . 50 gnus-thread-ignore-subject. . . . . . . . . . . . . . . . . . . . 50 gnus-thread-indent-level . . . . . . . . . . . . . . . . . . . . . . 50 gnus-thread-operation-ignore-subject . . . . . . . . . . 52 gnus-thread-score-function . . . . . . . . . . . . . . . . . . . . 52 gnus-thread-sort-by-author . . . . . . . . . . . . . . . . . . . 52 gnus-thread-sort-by-date . . . . . . . . . . . . . . . . . . . . . 52 gnus-thread-sort-by-number . . . . . . . . . . . . . . . . . . 52 gnus-thread-sort-by-score . . . . . . . . . . . . . . . . . . . . . 52 gnus-thread-sort-by-subject . . . . . . . . . . . . . . . . . . . 52 gnus-thread-sort-by-total-score . . . . . . . . . . . . . . . . 52 gnus-thread-sort-functions . . . . . . . . . . . . . . . . . . . . 52 gnus-ticked-mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 gnus-topic-copy-group . . . . . . . . . . . . . . . . . . . . . . . . 27 gnus-topic-copy-matching. . . . . . . . . . . . . . . . . . . . . 27 gnus-topic-create-topic . . . . . . . . . . . . . . . . . . . . . . . 26 gnus-topic-delete. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 gnus-topic-display-empty-topics . . . . . . . . . . . . . . . 26 gnus-topic-edit-parameters. . . . . . . . . . . . . . . . . . . . 28 gnus-topic-indent . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 gnus-topic-indent-level . . . . . . . . . . . . . . . . . . . . . . . 26 gnus-topic-kill-group . . . . . . . . . . . . . . . . . . . . . . . . . 27 gnus-topic-line-format . . . . . . . . . . . . . . . . . . . . . . . . 26 gnus-topic-list-active . . . . . . . . . . . . . . . . . . . . . . . . . 28 gnus-topic-mark-topic . . . . . . . . . . . . . . . . . . . . . . . . 27 gnus-topic-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 gnus-topic-mode-hook . . . . . . . . . . . . . . . . . . . . . . . . 26 gnus-topic-move-group . . . . . . . . . . . . . . . . . . . . . . . 27 gnus-topic-move-matching . . . . . . . . . . . . . . . . . . . . 27 gnus-topic-remove-group . . . . . . . . . . . . . . . . . . . . . 27 gnus-topic-rename . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 gnus-topic-select-group . . . . . . . . . . . . . . . . . . . . . . . 27 gnus-topic-sort-groups-by-alphabet . . . . . . . . . . . . 28 gnus-topic-sort-groups-by-level . . . . . . . . . . . . . . . . 28 gnus-topic-sort-groups-by-method . . . . . . . . . . . . . 28 gnus-topic-sort-groups-by-rank . . . . . . . . . . . . . . . . 28 gnus-topic-sort-groups-by-score . . . . . . . . . . . . . . . 28 gnus-topic-sort-groups-by-unread. . . . . . . . . . . . . . 28 gnus-topic-toggle-display-empty-topics . . . . . . . . . 27 gnus-topic-topology . . . . . . . . . . . . . . . . . . . . . . . . . . 28 gnus-topic-unindent . . . . . . . . . . . . . . . . . . . . . . . . . . 27 gnus-topic-unmark-topic. . . . . . . . . . . . . . . . . . . . . . 27 gnus-topic-yank-group. . . . . . . . . . . . . . . . . . . . . . . . 27 gnus-total-expirable-newsgroups . . . . . . . . . . . . . 108 gnus-tree-brackets . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 gnus-tree-line-format . . . . . . . . . . . . . . . . . . . . . . . . . 73 gnus-tree-minimize-window . . . . . . . . . . . . . . . . . . . 74 gnus-tree-mode-hook . . . . . . . . . . . . . . . . . . . . . . . . . 73 gnus-tree-mode-line-format . . . . . . . . . . . . . . . . . . . 73 gnus-tree-parent-child-edges . . . . . . . . . . . . . . . . . . 74 gnus-uncacheable-groups . . . . . . . . . . . . . . . . . . . . . 54 gnus-undo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 gnus-undo-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 gnus-unload . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 gnus-unread-mark . . . . . . . . . . . . . . . . . . . . . . . . 38, 42 gnus-update-format . . . . . . . . . . . . . . . . . . . . . . . . . 154 gnus-update-score-entry-dates . . . . . . . . . . . . . . . 135 gnus-updated-mode-lines . . . . . . . . . . . . . . . . . . . . 160 gnus-use-adaptive-scoring. . . . . . . . . . . . . . . . . . . . 140 gnus-use-article-prefetch . . . . . . . . . . . . . . . . . . . . . . 53 gnus-use-cache. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 gnus-use-cross-reference . . . . . . . . . . . . . . . . . . . . . . 79 gnus-use-demon . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 gnus-use-dribble-file . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 gnus-use-full-window . . . . . . . . . . . . . . . . . . . . . . . . 157 gnus-use-grouplens . . . . . . . . . . . . . . . . . . . . . . . . . . 147 gnus-use-long-file-name . . . . . . . . . . . . . . . . . . . 54, 58 gnus-use-nocem. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 gnus-use-scoring . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 gnus-use-toolbar . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 gnus-use-trees . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 gnus-use-undo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 gnus-useful-groups . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 gnus-uu-be-dangerous . . . . . . . . . . . . . . . . . . . . . . . . 61 gnus-uu-correct-stripped-uucode . . . . . . . . . . . . . . 61 gnus-uu-decode-binhex . . . . . . . . . . . . . . . . . . . . . . . 60 gnus-uu-decode-postscript . . . . . . . . . . . . . . . . . . . . 60 gnus-uu-decode-postscript-and-save . . . . . . . . . . . 60 gnus-uu-decode-postscript-and-save-view. . . . . . . 60 gnus-uu-decode-postscript-view . . . . . . . . . . . . . . . 60 gnus-uu-decode-save . . . . . . . . . . . . . . . . . . . . . . . . . 60 gnus-uu-decode-unshar . . . . . . . . . . . . . . . . . . . . . . . 59 gnus-uu-decode-unshar-and-save . . . . . . . . . . . . . . 59 gnus-uu-decode-unshar-and-save-view . . . . . . . . . 60 gnus-uu-decode-unshar-view . . . . . . . . . . . . . . . . . . 59 gnus-uu-decode-uu . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 gnus-uu-decode-uu-and-save . . . . . . . . . . . . . . . . . . 59 gnus-uu-decode-uu-and-save-view . . . . . . . . . . . . . 59 gnus-uu-decode-uu-view . . . . . . . . . . . . . . . . . . . . . . 59 gnus-uu-digest-headers . . . . . . . . . . . . . . . . . . . . . . . 85 gnus-uu-digest-mail-forward . . . . . . . . . . . . . . . . . . 40 gnus-uu-digest-post-forward . . . . . . . . . . . . . . . . . . 40 gnus-uu-do-not-unpack-archives . . . . . . . . . . . . . . . 61 gnus-uu-grab-move. . . . . . . . . . . . . . . . . . . . . . . . . . . 61 gnus-uu-grab-view . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 gnus-uu-grabbed-file-functions . . . . . . . . . . . . . . . . 61 gnus-uu-ignore-default-archive-rules . . . . . . . . . . . 61 gnus-uu-ignore-default-view-rules. . . . . . . . . . . . . . 61 gnus-uu-ignore-files-by-name . . . . . . . . . . . . . . . . . . 61 gnus-uu-ignore-files-by-type. . . . . . . . . . . . . . . . . . . 61 gnus-uu-invert-processable . . . . . . . . . . . . . . . . . . . . 45 gnus-uu-kill-carriage-return . . . . . . . . . . . . . . . . . . . 61 gnus-uu-mark-all. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 gnus-uu-mark-buffer . . . . . . . . . . . . . . . . . . . . . . . . . 45 gnus-uu-mark-by-regexp . . . . . . . . . . . . . . . . . . . . . . 45 gnus-uu-mark-over . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 gnus-uu-mark-region . . . . . . . . . . . . . . . . . . . . . . . . . 45 gnus-uu-mark-series . . . . . . . . . . . . . . . . . . . . . . . . . . 45 gnus-uu-mark-sparse . . . . . . . . . . . . . . . . . . . . . . . . . 45 gnus-uu-mark-thread . . . . . . . . . . . . . . . . . . . . . 45, 51 gnus-uu-notify-files. . . . . . . . . . . . . . . . . . . . . . . . . . . 59 gnus-uu-post-include-before-composing . . . . . . . . 62 gnus-uu-post-length . . . . . . . . . . . . . . . . . . . . . . . . . . 62 gnus-uu-post-news . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 gnus-uu-post-separate-description . . . . . . . . . . . . . 62 gnus-uu-post-threaded . . . . . . . . . . . . . . . . . . . . . . . 62 gnus-uu-pre-uudecode-hook . . . . . . . . . . . . . . . . . . . 62 gnus-uu-save-in-digest . . . . . . . . . . . . . . . . . . . . . . . . 62 gnus-uu-tmp-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 gnus-uu-unmark-articles-not-decoded . . . . . . . . . . 61 gnus-uu-unmark-thread . . . . . . . . . . . . . . . . . . . 45, 51 gnus-uu-user-archive-rules . . . . . . . . . . . . . . . . . . . . 60 gnus-uu-user-view-rules. . . . . . . . . . . . . . . . . . . . . . . 60 gnus-uu-user-view-rules-end. . . . . . . . . . . . . . . . . . . 60 gnus-uu-view-and-save . . . . . . . . . . . . . . . . . . . . . . . 61 gnus-uu-view-with-metamail . . . . . . . . . . . . . . . . . . 62 gnus-valid-select-methods. . . . . . . . . . . . . . . . . . . . 220 gnus-verbose . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 gnus-verbose-backends . . . . . . . . . . . . . . . . . . . . . . 173 gnus-version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 gnus-view-pseudo-asynchronously . . . . . . . . . . . . . 62 gnus-view-pseudos . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 gnus-view-pseudos-separately . . . . . . . . . . . . . . . . . 63 gnus-visible-headers . . . . . . . . . . . . . . . . . . . . . . . . . . 81 gnus-visual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 gnus-visual-mark-article-hook . . . . . . . . . . . . . . . . . 35 gnus-window-min-height. . . . . . . . . . . . . . . . . . . . . 158 gnus-window-min-width . . . . . . . . . . . . . . . . . . . . . 158 gnus-xmas-glyph-directory. . . . . . . . . . . . . . . . . . . 171 gnus-xmas-logo-color-alist . . . . . . . . . . . . . . . . . . . 171 gnus-xmas-logo-color-style . . . . . . . . . . . . . . . . . . . 171 gnus-xmas-modeline-glyph. . . . . . . . . . . . . . . . . . . 171 group buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 group buffer format . . . . . . . . . . . . . . . . . . . . . . . . . . 11 group description . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 group highlighting . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 group information . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 group level . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 group listing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 group modeline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 group movement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 group parameters. . . . . . . . . . . . . . . . . . . . . . . . . 19, 28 group rank . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 group score. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 group score commands . . . . . . . . . . . . . . . . . . . . . . 133 group selection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 group timestamps. . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 group-group-find-parameter . . . . . . . . . . . . . . . . . 211 GroupLens . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 grouplens-best-unread-article . . . . . . . . . . . . . . . . 148 grouplens-newsgroups . . . . . . . . . . . . . . . . . . . . . . . 148 grouplens-next-unread-article . . . . . . . . . . . . . . . . 148 grouplens-prediction-display . . . . . . . . . . . . . . . . . 148 grouplens-pseudonym . . . . . . . . . . . . . . . . . . . . . . . 147 grouplens-score-thread . . . . . . . . . . . . . . . . . . . . . . 148

H

head . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 header . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 help group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 114 hiding headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 highlighting . . . . . . . . . . . . . . . . . . . . . 12, 63, 161, 178 highlights . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 hilit19 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38, 177

I

ignored groups. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 incoming mail files . . . . . . . . . . . . . . . . . . . . . . . . . . 103 incoming mail treatment . . . . . . . . . . . . . . . . . . . . 108 info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 information on groups. . . . . . . . . . . . . . . . . . . . . . . . 30 InReference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 120 interaction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 interactive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 internal variables . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 invalid characters in file names. . . . . . . . . . . . . . . 174 ISO 8601. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 ISO8601 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 ispell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 ispell-message . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86

J

Jem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165

K

kibozing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 kill files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146, 147 killed groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205

L

level . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 levels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 limiting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 links . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 LIST overview.fmt . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 list server brain damage . . . . . . . . . . . . . . . . . . . . . 108 local variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 loose threads . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

M

mail . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39, 85, 100, 204 mail folders . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 mail group commands . . . . . . . . . . . . . . . . . . . . . . . . 75 mail message . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 mail nov spool . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 mail splitting . . . . . . . . . . . . . . . . . . . . . . . . . . 100, 103 mail washing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 mail-extract-address-components . . . . . . . . . . . . . . 33 mail-to-news gateways . . . . . . . . . . . . . . . . . . . . . . 121 MAILHOST. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 mailing lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 making digests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 making groups. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 manual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31, 203 mark as unread . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 marking groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 marks. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 max-lisp-eval-depth . . . . . . . . . . . . . . . . . . . . . . . . . 209 mbox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 mbox folders . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 mc-verify. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 merging groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 message . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 Message-ID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 message-sent-hook . . . . . . . . . . . . . . . . . . . . . . . . . . 144 message-this-is-mail . . . . . . . . . . . . . . . . . . . . . . . . . . 89 message-this-is-news . . . . . . . . . . . . . . . . . . . . . . . . . 89 messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 metamail. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 metamail-buffer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 MH folders . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 mh-e mail spool . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 mime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 MIME digest . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 MMDF mail box. . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 mode lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160, 173 mode reader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 moderation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 mouse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 move mail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 movemail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 moving articles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 Mule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179

N

native . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 new features. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 new groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 new messages. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 news. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 news backends . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 news spool . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 newsgroup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 nnbabyl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 nnbabyl-active-file . . . . . . . . . . . . . . . . . . . . . . . . . . 111 nnbabyl-get-new-mail . . . . . . . . . . . . . . . . . . . 110, 111 nnbabyl-mbox-file. . . . . . . . . . . . . . . . . . . . . . . . . . . 111 nnchoke. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 nndir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 113 nndoc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 114 nndoc-article-type . . . . . . . . . . . . . . . . . . . . . . . . . . 115 nndoc-post-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 nndraft . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 nndraft-directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 nneething. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 113 nneething-exclude-files . . . . . . . . . . . . . . . . . . . . . . 114 nneething-map-file . . . . . . . . . . . . . . . . . . . . . . . . . . 114 nneething-map-file-directory . . . . . . . . . . . . . . . . . 114 nnfolder. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 nnfolder-active-file . . . . . . . . . . . . . . . . . . . . . . . . . . 112 nnfolder-directory. . . . . . . . . . . . . . . . . . . . . . . . . . . 112 nnfolder-generate-active-file. . . . . . . . . . . . . . . . . . 113 nnfolder-get-new-mail. . . . . . . . . . . . . . . . . . . 110, 113 nnfolder-newsgroups-file . . . . . . . . . . . . . . . . . . . . . 113 nnfolder-save-buffer-hook . . . . . . . . . . . . . . . . . . . . 113 nngateway-address . . . . . . . . . . . . . . . . . . . . . . . . . . 121 nngateway-header-transformation . . . . . . . . . . . . 121 nngateway-mail2news-header-transformation . . 122 nngateway-simple-header-transformation . . . . . 122 nnheader-file-name-translation-alist . . . . . . . . . . 174 nnheader-get-report . . . . . . . . . . . . . . . . . . . . . . . . . 217 nnheader-head-chop-length . . . . . . . . . . . . . . . . . . 174 nnheader-max-head-length. . . . . . . . . . . . . . . . . . . 173 nnheader-ms-strip-cr . . . . . . . . . . . . . . . . . . . . . . . . 109 nnheader-report . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 nnkiboze . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 123 nnkiboze-directory . . . . . . . . . . . . . . . . . . . . . . . . . . 123 nnkiboze-generate-groups. . . . . . . . . . . . . . . . . . . . 123 nnmail-cache-accepted-message-ids . . . . . . . . . . . 103 nnmail-crash-box . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 nnmail-crosspost. . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 nnmail-crosspost-link-function . . . . . . . . . . . . . . . 101 nnmail-delete-file-function . . . . . . . . . . . . . . . . . . . 103 nnmail-delete-incoming. . . . . . . . . . . . . . . . . . . . . . 103 nnmail-expiry-wait. . . . . . . . . . . . . . . . . . . . . . . . . . 107 nnmail-expiry-wait-function . . . . . . . . . . . . . . 21, 107 nnmail-keep-last-article . . . . . . . . . . . . . . . . . 105, 108 nnmail-message-id-cache-file . . . . . . . . . . . . . . . . . 109 nnmail-message-id-cache-length . . . . . . . . . . . . . . 109 nnmail-movemail-program . . . . . . . . . . . . . . . . . . . 103 nnmail-pop-password . . . . . . . . . . . . . . . . . . . . . . . 102 nnmail-pop-password-required . . . . . . . . . . . . . . . 102 nnmail-post-get-new-mail-hook . . . . . . . . . . . . . . 102 nnmail-pre-get-new-mail-hook . . . . . . . . . . . . . . . 102 nnmail-prepare-incoming-header-hook . . . . . . . . 109 nnmail-prepare-incoming-hook . . . . . . . . . . 102, 108 nnmail-prepare-incoming-message-hook . . . . . . . 109 nnmail-procmail-directory . . . . . . . . . . . . . . . . . . . 105 nnmail-procmail-suffix . . . . . . . . . . . . . . . . . . 102, 105 nnmail-read-incoming-hook . . . . . . . . . . . . . . . . . . 101 nnmail-remove-leading-whitespace. . . . . . . . . . . . 109 nnmail-remove-list-identifiers . . . . . . . . . . . . . . . . 109 nnmail-remove-tabs . . . . . . . . . . . . . . . . . . . . . . . . . 109 nnmail-resplit-incoming . . . . . . . . . . . . . . . . . . . . . 105 nnmail-split-abbrev-alist . . . . . . . . . . . . . . . . . . . . 104 nnmail-split-fancy . . . . . . . . . . . . . . . . . . . . . . . . . . 103 nnmail-split-fancy-syntax-table . . . . . . . . . . . . . . 104 nnmail-split-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 nnmail-split-methods. . . . . . . . . . . . . . . . . . . . . . . . 100 nnmail-spool-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 nnmail-tmp-directory . . . . . . . . . . . . . . . . . . . . . . . 103 nnmail-treat-duplicates. . . . . . . . . . . . . . . . . . . . . . 109 nnmail-use-long-file-names. . . . . . . . . . . . . . . . . . . 103 nnmail-use-procmail . . . . . . . . . . . . . . . . . . . . 102, 105 nnmbox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 nnmbox-active-file . . . . . . . . . . . . . . . . . . . . . . 110, 111 nnmbox-get-new-mail . . . . . . . . . . . . . . . . . . . 110, 111 nnmbox-mbox-file. . . . . . . . . . . . . . . . . . . . . . . . . . . 110 nnmh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 nnmh-be-safe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 nnmh-directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 nnmh-get-new-mail . . . . . . . . . . . . . . . . . . . . . 110, 112 nnml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 nnml-active-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 nnml-directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 nnml-generate-nov-databases . . . . . . . . . . . . . . . . 112 nnml-get-new-mail . . . . . . . . . . . . . . . . . . . . . 110, 112 nnml-newsgroups-file . . . . . . . . . . . . . . . . . . . . . . . . 111 nnml-nov-file-name . . . . . . . . . . . . . . . . . . . . . . . . . 112 nnml-nov-is-evil . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 nnml-prepare-save-mail-hook . . . . . . . . . . . . . . . . 112 nnsoup. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 nnsoup-active-file . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 nnsoup-always-save . . . . . . . . . . . . . . . . . . . . . . . . . 120 nnsoup-directory. . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 nnsoup-pack-replies . . . . . . . . . . . . . . . . . . . . . . . . . 118 nnsoup-packer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 nnsoup-packet-directory . . . . . . . . . . . . . . . . . . . . . 120 nnsoup-packet-regexp . . . . . . . . . . . . . . . . . . . . . . . 120 nnsoup-replies-directory . . . . . . . . . . . . . . . . . . . . . 119 nnsoup-replies-format-type . . . . . . . . . . . . . . . . . . 119 nnsoup-replies-index-type . . . . . . . . . . . . . . . . . . . 119 nnsoup-set-variables . . . . . . . . . . . . . . . . . . . . . . . . 120 nnsoup-tmp-directory . . . . . . . . . . . . . . . . . . . . . . . 119 nnsoup-unpacker. . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 nnspool . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 nnspool-active-file . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 nnspool-active-times-file . . . . . . . . . . . . . . . . . . . . . 100 nnspool-history-file . . . . . . . . . . . . . . . . . . . . . . . . . 100 nnspool-inews-program . . . . . . . . . . . . . . . . . . . . . . . 99 nnspool-inews-switches . . . . . . . . . . . . . . . . . . . . . . . 99 nnspool-lib-dir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 nnspool-newsgroups-file . . . . . . . . . . . . . . . . . . . . . . 99 nnspool-nov-directory . . . . . . . . . . . . . . . . . . . . . . . . 99 nnspool-nov-is-evil . . . . . . . . . . . . . . . . . . . . . . . . . . 100 nnspool-sift-nov-with-sed . . . . . . . . . . . . . . . . . . . . 100 nnspool-spool-directory . . . . . . . . . . . . . . . . . . . . . . 99 nntp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 nntp authentification. . . . . . . . . . . . . . . . . . . . . . . . . 95 nntp server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 nntp-address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 nntp-authinfo-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 nntp-authinfo-function . . . . . . . . . . . . . . . . . . . . . . . 95 nntp-buggy-select . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 nntp-connection-timeout . . . . . . . . . . . . . . . . . . . . . 96 nntp-end-of-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 nntp-maximum-request. . . . . . . . . . . . . . . . . . . . . . . 96 nntp-nov-gap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 nntp-nov-is-evil . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 nntp-open-connection-function . . . . . . . . . . . . . . . . 97 nntp-open-network-stream . . . . . . . . . . . . . . . . . . . . 97 nntp-open-rlogin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 nntp-open-ssl-stream . . . . . . . . . . . . . . . . . . . . . . . . . 98 nntp-open-telnet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 nntp-open-telnet-envuser . . . . . . . . . . . . . . . . . . . . . 98 nntp-port-number. . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 nntp-prepare-server-hook . . . . . . . . . . . . . . . . . . . . . 99 nntp-record-commands . . . . . . . . . . . . . . . . . . . . . . . 99 nntp-rlogin-parameters . . . . . . . . . . . . . . . . . . . . . . . 97 nntp-rlogin-program . . . . . . . . . . . . . . . . . . . . . . . . . 97 nntp-rlogin-user-name . . . . . . . . . . . . . . . . . . . . 97, 98 nntp-send-authinfo. . . . . . . . . . . . . . . . . . . . . . . . . . . 95 nntp-send-mode-reader . . . . . . . . . . . . . . . . . . . . . . . 95 nntp-server-action-alist . . . . . . . . . . . . . . . . . . . . . . . 96 nntp-server-hook. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 nntp-server-opened-hook . . . . . . . . . . . . . . . . . . . . . 95 nntp-telnet-command . . . . . . . . . . . . . . . . . . . . . . . . 97 nntp-telnet-parameters . . . . . . . . . . . . . . . . . . . . . . . 97 nntp-telnet-passwd. . . . . . . . . . . . . . . . . . . . . . . . . . . 97 nntp-telnet-shell-prompt. . . . . . . . . . . . . . . . . . . . . . 98 nntp-telnet-switches. . . . . . . . . . . . . . . . . . . . . . . . . . 97 nntp-telnet-user-name . . . . . . . . . . . . . . . . . . . . . . . . 97 nntp-warn-about-losing-connection . . . . . . . . . . . . 99 nntp-xover-commands . . . . . . . . . . . . . . . . . . . . . . . . 98 NNTPSERVER. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 nnvirtual. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 nnvirtual-always-rescan . . . . . . . . . . . . . . . . . . . . . 123 nnweb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 120 nnweb-max-hits . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 nnweb-search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 nnweb-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 nnweb-type-definition . . . . . . . . . . . . . . . . . . . . . . . 121 nocem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 nov . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 nov . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98, 204

O

offline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117, 124 overview.fmt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

P

parameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 parent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 parent articles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 persistent articles . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 pick and read. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 POP mail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 post . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40, 85 posting styles. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 PostScript . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60, 70 pre-fetch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 process mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 process/prefix convention. . . . . . . . . . . . . . . . . . . . 153 procmail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 pseudo-articles. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

R

rank . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 rcvstore . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 reading init file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 reading mail. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 reading news . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 referring articles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 rejected articles. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 renaming groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 reply . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85, 204 reporting bugs . . . . . . . . . . . . . . . . . . . . . . . . . 178, 209 restarting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 reverse scoring. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 RFC 1036 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 RFC 1153 digest. . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 RFC 341 digest . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 RFC 822 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 RFC1522 decoding. . . . . . . . . . . . . . . . . . . . . . . . . . 102 rmail mbox . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111, 114 rnews batch files . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 root . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 rule variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60

S

saving .newsrc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 saving articles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 scanning new news. . . . . . . . . . . . . . . . . . . . . . . . . . . 30 score cache . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 score commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 score decays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 Score File Atoms . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 score file format . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 score file group parameter . . . . . . . . . . . . . . . . . . . . 21 score variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 scoring. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 scoring crossposts. . . . . . . . . . . . . . . . . . . . . . . . . . . 144 scoring tips. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 searching the Usenet . . . . . . . . . . . . . . . . . . . . . . . . 120 secondary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 sed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 select method . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 select methods. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 selecting articles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 sending mail. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 sent messages. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 series . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 server buffer format . . . . . . . . . . . . . . . . . . . . . . . . . . 91 server commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 server errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 setting marks. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 setting process marks . . . . . . . . . . . . . . . . . . . . . . . . 44 shared articles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 shell archives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 signatures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 slave . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 slocal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 slow machine . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 smiley-circle-color. . . . . . . . . . . . . . . . . . . . . . . . . . . 170 smiley-data-directory . . . . . . . . . . . . . . . . . . . . . . . 170 smiley-deformed-regexp-alist . . . . . . . . . . . . . . . . . 170 smiley-features-color . . . . . . . . . . . . . . . . . . . . . . . . 170 smiley-flesh-color . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 smiley-mouse-face. . . . . . . . . . . . . . . . . . . . . . . . . . . 171 smiley-nosey-regexp-alist . . . . . . . . . . . . . . . . . . . . 170 smiley-tongue-color . . . . . . . . . . . . . . . . . . . . . . . . . 170 smileys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 solid groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 Son-of-RFC 1036 . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 sorting groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 SOUP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 sox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 spam . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164, 172 spamming . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 sparse articles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 splitting mail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 starting up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 startup files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 styles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 subscription. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5, 15 summary buffer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 summary buffer format . . . . . . . . . . . . . . . . . . . . . . . 33 summary exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 summary movement. . . . . . . . . . . . . . . . . . . . . . . . . . 36 summary sorting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 superseding articles . . . . . . . . . . . . . . . . . . . . . . . . . . 41 symbolic prefixes . . . . . . . . . . . . . . . . . . . . . . . . . . . 154

T

terminology . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 thread commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 thread root. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 threading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46, 205 timestamps. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 to-address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 to-group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 to-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 todo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 topic commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 topic parameters . . . . . . . . . . . . . . . . . . . . . . . . . 28, 29 topic sorting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 topic topology . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 topic variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 topics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 topology . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 total-expire. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 transient-mark-mode . . . . . . . . . . . . . . . . . . . . . . . . 153 trees . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 troubleshooting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209

U

UCE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 undo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 unix mail box . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 Unix mbox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 unloading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 unplugged. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 unshar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 unsolicited commercial email . . . . . . . . . . . . . . . . 172 Usenet searches. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 utility functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 uudecode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 uuencoded articles . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

V

velveeta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 viewing files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 virtual groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 virtual server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 visible . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 visible group parameter . . . . . . . . . . . . . . . . . . . . . . 23 visual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161

W

washing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66, 205 window height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 window width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 windows configuration . . . . . . . . . . . . . . . . . . . . . . 157

X

x-face . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 XEmacs. . . . . . . . . . . . . . . . . . . . . . . . . . . 167, 179, 226 XOVER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 Xref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

Z

zombie groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205

Chapter 12: Key Index 257

12 Key Index

!

! (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

#

# (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 # (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

&

& (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

*

* (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

,

, (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 , (GroupLens) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 , (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

.

. (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 . (Pick) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 . (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

/

/ * (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 / / (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 / a (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 / c (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 / C (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 / d (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 / D (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 / E (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 / m (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 / n (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 / t (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 / T (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 / u (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 / v (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 / w (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

=

= (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

?

? (Article). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 ? (Browse) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 ? (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 ? (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

@

@ (Agent Summary). . . . . . . . . . . . . . . . . . . . . . . . . . 128

>

> (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

^

^(Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 ^(Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

<

< (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

A

a (Category). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 a (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 a (Server) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 a (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 A > (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 A < (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 A a (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 A A (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 A d (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 A g (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 A k (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 A l (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 A m (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 A M (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 A P (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 A R (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 A s (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 A s (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 A T (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 A T (Topic) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 A u (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 A z (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

B

b (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 B (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 24 B B (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 B c (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 B DEL (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 B e (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 B i (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 B m (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 B M-C-e (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . 75 B p (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 B q (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 B r (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 B t (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 B w (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

C

c (Category). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 c (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 C (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 c (Server) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 C (Server) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 c (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 C (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 C-c ^(Article) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 C-c C-c (Article). . . . . . . . . . . . . . . . . . . . . . . . . . . 75 C-c C-c (Post) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 C-c C-c (Score). . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 C-c C-d (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 C-c C-d (Score). . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 C-c C-i (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 C-c C-m (Article). . . . . . . . . . . . . . . . . . . . . . . . . . . 83 C-c C-p (Score). . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 C-c C-s (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 C-c C-s C-a (Summary). . . . . . . . . . . . . . . . . . . . . . . 70 C-c C-s C-d (Summary). . . . . . . . . . . . . . . . . . . . . . . 70 C-c C-s C-i (Summary). . . . . . . . . . . . . . . . . . . . . . . 70 C-c C-s C-l (Summary). . . . . . . . . . . . . . . . . . . . . . . 70 C-c C-s C-n (Summary). . . . . . . . . . . . . . . . . . . . . . . 70 C-c C-s C-s (Summary). . . . . . . . . . . . . . . . . . . . . . . 70 C-c C-x (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 C-c M-C-x (Group). . . . . . . . . . . . . . . . . . . . . . . . . . . 24 C-c M-g (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 C-d (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 C-k (Group). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 C-k (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 C-k (Topic). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 C-t (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 C-w (Group). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 C-w (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 C-x C-t (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 C-y (Group). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 C-y (Topic). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

D

D (Server) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 d (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 D (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 D e (Draft) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 DEL (Article). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 DEL (Group). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 DEL (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

E

e (Server) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 e (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 E (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

F

F (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 f (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 F (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

G

g (Binary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 g (Category). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 g (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 g (Server) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 g (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 G a (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 G b (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 G c (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 G C-n (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 G C-p (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 G d (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 G D (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 G DEL (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 G e (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 G E (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 G f (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 G f (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 G g (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 G h (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 G j (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 G k (Group). . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 123 G l (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 G m (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 G M-n (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 G M-p (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 G n (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 G N (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 G o (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 G p (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 G P (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 G p (Topic) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 G P a (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 G P l (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 G P m (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 G P r (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 G P u (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 G P v (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 G r (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 G S a (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 G s b (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 G S l (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 G S m (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 G s p (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 G s r (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 G S r (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 G s s (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 G S u (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 G S v (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 G s w (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 G u (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 G v (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 G V (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 G w (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

H

h (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 H d (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 H d (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 H f (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 H f (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 H h (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 H i (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 H v (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

J

J # (Agent Summary) . . . . . . . . . . . . . . . . . . . . . . . . 128 j (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 j (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 J a (Agent Group) . . . . . . . . . . . . . . . . . . . . . . . . . . 128 J a (Agent Server) . . . . . . . . . . . . . . . . . . . . . . . . . 128 J c (Agent Group) . . . . . . . . . . . . . . . . . . . . . . . . . . 127 J c (Agent Summary) . . . . . . . . . . . . . . . . . . . . . . . . 128 J M-# (Agent Summary). . . . . . . . . . . . . . . . . . . . . . 128 J r (Agent Server) . . . . . . . . . . . . . . . . . . . . . . . . . 128 J s (Agent Group) . . . . . . . . . . . . . . . . . . . . . . . . . . 128 J S (Agent Group) . . . . . . . . . . . . . . . . . . . . . . . . . . 128 J u (Agent Group) . . . . . . . . . . . . . . . . . . . . . . . . . . 127

K

k (Category). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 k (GroupLens) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 k (Server) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 k (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

L

l (Browse) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 l (Category). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 l (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 L (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 l (Server) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 l (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

M

m (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 m (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 M ? (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 M b (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 M b (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 M B (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 M c (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 M C (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 M C-c (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 M d (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 M e (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 M H (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 M k (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 M K (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 M m (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 M P a (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 M P b (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 M P i (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 M P k (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 M P p (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 M P r (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 M P R (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 M P s (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 M P S (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 M P t (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 M P T (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 M P u (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 M P U (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 M P v (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 M P w (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 M P y (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 M r (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 M S (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 M t (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 M u (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 M U (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 M V c (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 M V k (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 M V m (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 M V u (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 M w (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 M-# (Group). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 M-# (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 M-& (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 M-* (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 M-^(Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 M-c (Group). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 M-c (Server). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 M-C-d (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 M-C-e (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 M-C-k (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 M-C-l (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 M-C-RET (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 M-d (Group). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 M-g (Group). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 M-g (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 M-i (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 M-k (Group). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 M-K (Group). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 M-k (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 M-K (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 M-n (Group). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 M-n (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 M-o (Server). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 M-p (Group). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 M-p (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 M-r (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 M-RET (Group). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 M-RET (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 M-s (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 M-SPACE (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 M-TAB (Article) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 M-TAB (Topic). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 M-u (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 M-x gnus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 M-x gnus-agent-expire . . . . . . . . . . . . . . . . . . . . . 128 M-x gnus-binary-mode . . . . . . . . . . . . . . . . . . . . . . . 73 M-x gnus-bug . . . . . . . . . . . . . . . . . . . . . . . . . . 178, 209 M-x gnus-change-server . . . . . . . . . . . . . . . . . . . . . . 7

                                                M-x gnus-group-clear-data-on-native-groups

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 16

                                                M-x gnus-group-move-group-to-server . . . . . . . . 7

M-x gnus-no-server . . . . . . . . . . . . . . . . . . . . . . . . . . 4 M-x gnus-other-frame . . . . . . . . . . . . . . . . . . . . . . . . 3 M-x gnus-pick-mode . . . . . . . . . . . . . . . . . . . . . . . . . 72 M-x gnus-update-format . . . . . . . . . . . . . . . . . . . . 154

                                                M-x nnfolder-generate-active-file . . . . . . . . 113
                                                M-x nnkiboze-generate-groups . . . . . . . . . . . . . 123
                                                M-x nnmail-split-history. . . . . . . . . . . . . . . . . . 101

N

n (Browse) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 n (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 N (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 n (GroupLens) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 n (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 N (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 nnmail-split-history. . . . . . . . . . . . . . . . . . . . . . 101

O

O (Server) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 o (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 O b (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 O f (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 O F (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 O h (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 O m (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 O o (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 O p (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 O r (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 O s (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 O v (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

P

p (Browse) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 p (Category). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 p (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 P (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 p (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 P (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

Q

q (Browse) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 q (Category). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 q (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Q (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 q (Server) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 q (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 Q (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

R

r (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 R (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 r (GroupLens) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 R (Server) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 r (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 R (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 RET (Browse). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 RET (Group). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 RET (Pick) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 RET (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 RET (Topic). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

S

s (Article). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 s (Category). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 s (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 s (Server) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 s (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 S (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 S C-k (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 S D b (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 S D r (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 S f (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 S F (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 S k (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 S l (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 S m (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 S M-c (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 S n (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 S N (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 S o m (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 S O m (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 S o p (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 S O p (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 S p (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 S r (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 S R (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 S s (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 S t (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 S u (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 S w (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 S w (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 S W (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 S y (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 S z (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 SPACE (Article) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 SPACE (Browse) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 SPACE (Group). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 SPACE (Pick). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 SPACE (Server) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 SPACE (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . 37, 38

T

T # (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 T # (Topic) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 t (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 T ^(Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 T c (Topic) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 T C (Topic) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 T d (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 T D (Topic) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 T DEL (Topic) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 T h (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 T H (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 T H (Topic) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 T i (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 T k (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 T l (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 T m (Topic) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 T M (Topic) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 T M-# (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 T M-# (Topic) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 T n (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 T n (Topic) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 T o (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 T p (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 T r (Topic) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 T s (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 T S (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 T S a (Topic) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 T S l (Topic) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 T S m (Topic) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 T S r (Topic) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 T S u (Topic) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 T S v (Topic) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 T t (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 T T (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 T TAB (Topic) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 T u (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 TAB (Article). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 TAB (Topic). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

U

u (Browse) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 u (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 U (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 u (Pick) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

V

V (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 V c (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 V C (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 V e (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 V f (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 V F (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 V m (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 V R (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 V s (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 V S (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 V t (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 V x (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132

W

W b (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 W B (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 W c (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 W d (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 W e (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 W E a (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 W E A (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 W E l (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 W E m (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 W E s (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 W E t (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 W f (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 W f (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 W H a (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 W H c (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 W H h (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 W H s (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 W l (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 W m (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 W o (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 W q (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 W r (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 W t (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 W T e (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 W T i (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 W T l (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 W T o (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 W T s (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 W T u (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 W v (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 W w (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 W W a (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 W W b (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 W W c (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 W W C (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 W W C-c (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 W W h (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 W W p (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 W W P (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 W W s (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

X

x (Summary). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 X b (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 X o (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 X p (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 X P (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 X s (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 X S (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 X u (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 X U (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 X v p (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 X v P (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 X v s (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 X v S (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 X v u (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 X v U (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

Y

y (Server) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 Y c (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 Y g (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

Z

z (Group) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Z c (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 Z C (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 Z E (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 Z G (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 Z n (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 Z N (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 Z P (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 Z R (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 Z s (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 Z Z (Summary) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

264 Gnus Manual

Short Contents

The Gnus Newsreader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1 Starting Gnus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2 The Group Buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3 The Summary Buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 4 The Article Buffer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 5 Composing Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 6 Select Methods . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 7 Scoring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 8 Various . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 9 The End . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 10 Appendices. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 11 Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 12 Key Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 ii Gnus Manual

Table of Contents

The Gnus Newsreader . . . . . . . . . . . . . . . . . . . . . . . . . 1

1 Starting Gnus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

1.1 Finding the News . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2 The First Time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.3 The Server is Down . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.4 Slave Gnusae . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.5 Fetching a Group. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.6 New Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

1.6.1 Checking New Groups . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.6.2 Subscription Methods . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.6.3 Filtering New Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

1.7 Changing Servers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 1.8 Startup Files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 1.9 Auto Save . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 1.10 The Active File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 1.11 Startup Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

2 The Group Buffer . . . . . . . . . . . . . . . . . . . . . . . . 11

2.1 Group Buffer Format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.1.1 Group Line Specification . . . . . . . . . . . . . . . . . . . . . . . 11 2.1.2 Group Modeline Specification. . . . . . . . . . . . . . . . . . . 12 2.1.3 Group Highlighting . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

2.2 Group Maneuvering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.3 Selecting a Group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.4 Subscription Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.5 Group Data. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 2.6 Group Levels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 2.7 Group Score . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.8 Marking Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.9 Foreign Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 2.10 Group Parameters. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 2.11 Listing Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 2.12 Sorting Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 2.13 Group Maintenance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 2.14 Browse Foreign Server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 2.15 Exiting Gnus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 2.16 Group Topics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

2.16.1 Topic Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 2.16.2 Topic Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 2.16.3 Topic Sorting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 2.16.4 Topic Topology. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

2.16.5 Topic Parameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 2.17 Misc Group Stuff. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

2.17.1 Scanning New Messages . . . . . . . . . . . . . . . . . . . . . . . 30 2.17.2 Group Information . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 2.17.3 Group Timestamp . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 2.17.4 File Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

3 The Summary Buffer . . . . . . . . . . . . . . . . . . . . . 33

3.1 Summary Buffer Format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 3.1.1 Summary Buffer Lines . . . . . . . . . . . . . . . . . . . . . . . . . 33 3.1.2 Summary Buffer Mode Line . . . . . . . . . . . . . . . . . . . . 35 3.1.3 Summary Highlighting . . . . . . . . . . . . . . . . . . . . . . . . . 35

3.2 Summary Maneuvering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 3.3 Choosing Articles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

3.3.1 Choosing Commands . . . . . . . . . . . . . . . . . . . . . . . . . . 37 3.3.2 Choosing Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

3.4 Scrolling the Article . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 3.5 Reply, Followup and Post . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

3.5.1 Summary Mail Commands . . . . . . . . . . . . . . . . . . . . . 39 3.5.2 Summary Post Commands . . . . . . . . . . . . . . . . . . . . . 40

3.6 Canceling Articles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 3.7 Marking Articles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

3.7.1 Unread Articles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 3.7.2 Read Articles. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 3.7.3 Other Marks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 3.7.4 Setting Marks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 3.7.5 Setting Process Marks . . . . . . . . . . . . . . . . . . . . . . . . . 44

3.8 Limiting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 3.9 Threading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

3.9.1 Customizing Threading . . . . . . . . . . . . . . . . . . . . . . . . 47 3.9.1.1 Loose Threads . . . . . . . . . . . . . . . . . . . . . . . . 47 3.9.1.2 Filling In Threads. . . . . . . . . . . . . . . . . . . . . 49 3.9.1.3 More Threading . . . . . . . . . . . . . . . . . . . . . . 50 3.9.1.4 Low-Level Threading . . . . . . . . . . . . . . . . . . 50

3.9.2 Thread Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 3.10 Sorting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 3.11 Asynchronous Article Fetching . . . . . . . . . . . . . . . . . . . . . . . . . 53 3.12 Article Caching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 3.13 Persistent Articles. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 3.14 Article Backlog . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 3.15 Saving Articles. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 3.16 Decoding Articles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

3.16.1 Uuencoded Articles . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 3.16.2 Shell Archives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 3.16.3 PostScript Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 3.16.4 Other Files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 3.16.5 Decoding Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . 60

3.16.5.1 Rule Variables . . . . . . . . . . . . . . . . . . . . . . . 60 3.16.5.2 Other Decode Variables . . . . . . . . . . . . . . 61 3.16.5.3 Uuencoding and Posting . . . . . . . . . . . . . . 62

3.16.6 Viewing Files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 3.17 Article Treatment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

3.17.1 Article Highlighting. . . . . . . . . . . . . . . . . . . . . . . . . . . 63 3.17.2 Article Fontisizing . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 3.17.3 Article Hiding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 3.17.4 Article Washing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 3.17.5 Article Buttons. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 3.17.6 Article Date. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 3.17.7 Article Signature . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

3.18 Article Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 3.19 Summary Sorting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 3.20 Finding the Parent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 3.21 Alternative Approaches . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

3.21.1 Pick and Read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 3.21.2 Binary Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

3.22 Tree Display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 3.23 Mail Group Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 3.24 Various Summary Stuff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76

3.24.1 Summary Group Information . . . . . . . . . . . . . . . . . . 76 3.24.2 Searching for Articles . . . . . . . . . . . . . . . . . . . . . . . . . 77 3.24.3 Summary Generation Commands . . . . . . . . . . . . . . 77 3.24.4 Really Various Summary Commands . . . . . . . . . . . 77

3.25 Exiting the Summary Buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 3.26 Crosspost Handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 3.27 Duplicate Suppression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

4 The Article Buffer . . . . . . . . . . . . . . . . . . . . . . . . 81

4.1 Hiding Headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 4.2 Using mime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 4.3 Customizing Articles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 4.4 Article Keymap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 4.5 Misc Article . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

5 Composing Messages. . . . . . . . . . . . . . . . . . . . . . 85

5.1 Mail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 5.2 Post. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 5.3 Posting Server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 5.4 Mail and Post. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 5.5 Archived Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 5.6 Posting Styles. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 5.7 Drafts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 5.8 Rejected Articles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90

6 Select Methods . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

6.1 The Server Buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 6.1.1 Server Buffer Format . . . . . . . . . . . . . . . . . . . . . . . . . . 91 6.1.2 Server Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 6.1.3 Example Methods . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 6.1.4 Creating a Virtual Server . . . . . . . . . . . . . . . . . . . . . . 93 6.1.5 Server Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 6.1.6 Servers and Methods. . . . . . . . . . . . . . . . . . . . . . . . . . . 94 6.1.7 Unavailable Servers . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

6.2 Getting News . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 6.2.1 nntp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 6.2.2 News Spool . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

6.3 Getting Mail. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 6.3.1 Getting Started Reading Mail . . . . . . . . . . . . . . . . . 100 6.3.2 Splitting Mail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 6.3.3 Mail Backend Variables . . . . . . . . . . . . . . . . . . . . . . . 101 6.3.4 Fancy Mail Splitting . . . . . . . . . . . . . . . . . . . . . . . . . . 103 6.3.5 Mail and Procmail . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 6.3.6 Incorporating Old Mail . . . . . . . . . . . . . . . . . . . . . . . 106 6.3.7 Expiring Mail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 6.3.8 Washing Mail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 6.3.9 Duplicates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 6.3.10 Not Reading Mail . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 6.3.11 Choosing a Mail Backend . . . . . . . . . . . . . . . . . . . . 110

6.3.11.1 Unix Mail Box . . . . . . . . . . . . . . . . . . . . . . 110 6.3.11.2 Rmail Babyl . . . . . . . . . . . . . . . . . . . . . . . . 111 6.3.11.3 Mail Spool . . . . . . . . . . . . . . . . . . . . . . . . . 111 6.3.11.4 MH Spool . . . . . . . . . . . . . . . . . . . . . . . . . . 112 6.3.11.5 Mail Folders . . . . . . . . . . . . . . . . . . . . . . . . 112

6.4 Other Sources. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 6.4.1 Directory Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 6.4.2 Anything Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 6.4.3 Document Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114

6.4.3.1 Document Server Internals. . . . . . . . . . . . 115 6.4.4 SOUP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117

6.4.4.1 SOUP Commands. . . . . . . . . . . . . . . . . . . . 118 6.4.4.2 soup Groups . . . . . . . . . . . . . . . . . . . . . . . . 119 6.4.4.3 SOUP Replies . . . . . . . . . . . . . . . . . . . . . . . 120

6.4.5 Web Searches . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 6.4.6 Mail-To-News Gateways. . . . . . . . . . . . . . . . . . . . . . . 121

6.5 Combined Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 6.5.1 Virtual Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 6.5.2 Kibozed Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123

6.6 Gnus Unplugged . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 6.6.1 Agent Basics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 6.6.2 Agent Categories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125

6.6.2.1 Category Syntax . . . . . . . . . . . . . . . . . . . . . 125 6.6.2.2 The Category Buffer . . . . . . . . . . . . . . . . . 126

6.6.2.3 Category Variables . . . . . . . . . . . . . . . . . . . 127 6.6.3 Agent Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127

6.6.3.1 Group Agent Commands . . . . . . . . . . . . . 127 6.6.3.2 Summary Agent Commands . . . . . . . . . . 128 6.6.3.3 Server Agent Commands . . . . . . . . . . . . . 128

6.6.4 Agent Expiry. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 6.6.5 Outgoing Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 6.6.6 Agent Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 6.6.7 Example Setup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 6.6.8 Batching Agents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130

7 Scoring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131

7.1 Summary Score Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 7.2 Group Score Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 7.3 Score Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 7.4 Score File Format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 7.5 Score File Editing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 7.6 Adaptive Scoring. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 7.7 Home Score File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 7.8 Followups To Yourself . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 7.9 Scoring Tips . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 7.10 Reverse Scoring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 7.11 Global Score Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 7.12 Kill Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 7.13 Converting Kill Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 7.14 GroupLens . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147

7.14.1 Using GroupLens . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 7.14.2 Rating Articles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 7.14.3 Displaying Predictions . . . . . . . . . . . . . . . . . . . . . . . 148 7.14.4 GroupLens Variables. . . . . . . . . . . . . . . . . . . . . . . . . 149

7.15 Advanced Scoring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 7.15.1 Advanced Scoring Syntax . . . . . . . . . . . . . . . . . . . . 150 7.15.2 Advanced Scoring Examples . . . . . . . . . . . . . . . . . . 150 7.15.3 Advanced Scoring Tips. . . . . . . . . . . . . . . . . . . . . . . 151

7.16 Score Decays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152

8 Various . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153

8.1 Process/Prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 8.2 Interactive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 8.3 Symbolic Prefixes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 8.4 Formatting Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154

8.4.1 Formatting Basics . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 8.4.2 Mode Line Formatting . . . . . . . . . . . . . . . . . . . . . . . . 155 8.4.3 Advanced Formatting . . . . . . . . . . . . . . . . . . . . . . . . . 155 8.4.4 User-Defined Specs . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 8.4.5 Formatting Fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156

8.5 Windows Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 8.6 Faces and Fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 8.7 Compilation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 8.8 Mode Lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 8.9 Highlighting and Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 8.10 Buttons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 8.11 Daemons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 8.12 NoCeM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 8.13 Undo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 8.14 Moderation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 8.15 XEmacs Enhancements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167

8.15.1 Picons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 8.15.1.1 Picon Basics . . . . . . . . . . . . . . . . . . . . . . . . 167 8.15.1.2 Picon Requirements . . . . . . . . . . . . . . . . . 167 8.15.1.3 Easy Picons . . . . . . . . . . . . . . . . . . . . . . . . 168 8.15.1.4 Hard Picons . . . . . . . . . . . . . . . . . . . . . . . . 168 8.15.1.5 Picon Useless Configuration. . . . . . . . . . 169

8.15.2 Smileys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 8.15.3 Toolbar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 8.15.4 Various XEmacs Variables. . . . . . . . . . . . . . . . . . . . 171

8.16 Fuzzy Matching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 8.17 Thwarting Email Spam . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 8.18 Various Various . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173

9 The End . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 10 Appendices . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177

10.1 History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 10.1.1 Why?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 10.1.2 Compatibility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 10.1.3 Conformity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 10.1.4 Emacsen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 10.1.5 Contributors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 10.1.6 New Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181

10.1.6.1 (ding) Gnus . . . . . . . . . . . . . . . . . . . . . . . . 181 10.1.6.2 September Gnus . . . . . . . . . . . . . . . . . . . . 183 10.1.6.3 Red Gnus . . . . . . . . . . . . . . . . . . . . . . . . . . 185 10.1.6.4 Quassia Gnus. . . . . . . . . . . . . . . . . . . . . . . 186

10.1.7 Newest Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 10.2 The Manual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 10.3 Terminology . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 10.4 Customization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207

10.4.1 Slow/Expensive nntp Connection . . . . . . . . . . . . 207 10.4.2 Slow Terminal Connection. . . . . . . . . . . . . . . . . . . . 207 10.4.3 Little Disk Space . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 10.4.4 Slow Machine . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208

10.5 Troubleshooting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 10.6 A Programmerś Guide to Gnus. . . . . . . . . . . . . . . . . . . . . . . . 210

10.6.1 Gnus Utility Functions . . . . . . . . . . . . . . . . . . . . . . . 210 10.6.2 Backend Interface . . . . . . . . . . . . . . . . . . . . . . . . . . . 211

10.6.2.1 Required Backend Functions . . . . . . . . . 212 10.6.2.2 Optional Backend Functions . . . . . . . . . 215 10.6.2.3 Error Messaging . . . . . . . . . . . . . . . . . . . . 217 10.6.2.4 Writing New Backends . . . . . . . . . . . . . . 218 10.6.2.5 Hooking New Backends Into Gnus. … 220 10.6.2.6 Mail-like Backends . . . . . . . . . . . . . . . . . . 221

10.6.3 Score File Syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 10.6.4 Headers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 10.6.5 Ranges . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 10.6.6 Group Info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 10.6.7 Extended Interactive. . . . . . . . . . . . . . . . . . . . . . . . . 226 10.6.8 Emacs/XEmacs Code . . . . . . . . . . . . . . . . . . . . . . . . 226 10.6.9 Various File Formats. . . . . . . . . . . . . . . . . . . . . . . . . 227

10.6.9.1 Active File Format . . . . . . . . . . . . . . . . . . 227 10.6.9.2 Newsgroups File Format . . . . . . . . . . . . . 227

10.7 Emacs for Heathens . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 10.7.1 Keystrokes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 10.7.2 Emacs Lisp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228

10.8 Frequently Asked Questions. . . . . . . . . . . . . . . . . . . . . . . . . . . 230 10.8.1 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 10.8.2 Customization. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 10.8.3 Reading News . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 10.8.4 Reading Mail. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 x Gnus Manual

11 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239

12 Key Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257